### v1

# Deep Discrete Encoders: Identifiable Deep Generative Models for Rich Data with Discrete Latent Layers

Abstract
In the era of generative AI, deep generative models (DGMs) with latent representations have gained tremendous popularity. Despite their impressive empirical performance, the statistical properties of these models remain underexplored. DGMs are often overparametrized, non-identifiable, and uninterpretable black boxes, raising serious concerns when deploying them in high-stakes applications. Motivated by this, we propose an interpretable deep generative modeling framework for rich data types with discrete latent layers, called Deep Discrete Encoders (DDEs). A DDE is a directed graphical model with multiple binary latent layers. Theoretically, we propose transparent identifiability conditions for DDEs, which imply progressively smaller sizes of the latent layers as they go deeper. Identifiability ensures consistent parameter estimation and inspires an interpretable design of the deep architecture. Computationally, we propose a scalable estimation pipeline of a layerwise nonlinear spectral initialization followed by a penalized stochastic approximation EM algorithm. This procedure can efficiently estimate models with exponentially many latent components. Extensive simulation studies validate our theoretical results and demonstrate the proposed algorithms’ excellent performance. We apply DDEs to three diverse real datasets for hierarchical topic modeling, image representation learning, response time modeling in educational testing, and obtain interpretable findings.
Corresponding Author: Yuqi Gu. Email: yuqi.gu@columbia.edu. Research is partially supported by NSF Grant DMS-2210796.
Keywords: Identifiability; Interpretable Artificial Intelligence; Representation Learning; Deep Belief Network; Directed Graphical Model.
1 Introduction
In the era of generative AI, deep generative models (DGMs) with latent representations have gained tremendous popularity across various domains. DGMs achieve impressive empirical success due to their rich modeling power and predictive capacity, and are useful tools to generate images, text, and audio (Hinton et al.,, 2006; Lee et al.,, 2009; Kingma and Welling,, 2014; Salakhutdinov,, 2015). However, these models are often subject to statistical issues regarding the model identifiability, interpretability, and parameter estimation reliability. Indeed, most deep learning models are heavily overparametrized black boxes, with more parameters than the number of samples, and are fundamentally non-identifiable. The lack of identifiability means that there may be very different parameter values that give the same marginal distribution of the observed data, leading to inconsistent parameter estimation. In such cases, it is impossible to guarantee the reproducibility of the learned latent representations across different training instances and the validity of the downstream inference. Additionally, the deep layers often merely serve as tools for inducing flexible data distributions but without a meaningful substantive interpretation. These issues raise serious concerns when deploying DGMs in high-stakes applications such as education, medicine, and health care.
We address the above problems from statisticians’ perspectives by proposing a broad family of interpretable and identifiable deep generative models for rich types of data, called Deep Discrete Encoders (DDEs). DDEs are directed graphical models with potentially deep discrete latent layers to generate the bottom-layer multivariate observed data. A key feature of DDEs is that the latent layers are discrete and organized in progressively smaller sizes as they go deeper; see Figure 1. This architecture induces very expressive models via the exponentially many mixture components (since each configuration of the discrete latent vector gives a mixture component), and also has the nice interpretation of increasingly more general latent features in deeper layers (Bengio et al.,, 2013). DDEs are motivated by both popular generative models in deep learning and latent variable models in educational and psychological measurement. While these two areas rarely intersect in the past, we leverage the insights from both fields to inspire the theory and methodology of DDEs.
Figure 1 displays the graphical model representations of a typical DDE alongside a DDE estimated from real data. The right panel of Figure 1 shows that fitting DDEs to a dataset of text documents uncovers interesting hierarchical latent topics as well as an interpretable word generating mechanism; see more details in Section 6. We emphasize that there is no restriction on the types of observed data; for example, the bottom data layer can be modeled by any exponential family distributions. In the text data example, Poisson distribution is used to model the word counts in documents. Our real data examples in Section 6 range from word counts in text documents to binary pixel values in images, to continuous response times of students in digital educational assessments. Such flexibility makes DDEs attractive for many practical applications ranging from machine learning to domain sciences.
[…]
rectechsportscargraphicsspaceleagueplayoffpartsopinioncolornasaorbit
Figure 1: Left: Example graphical model representation of DDEs. Right: Simplified estimated DDE structure of the latent topics for the 20 newsgroups dataset. Only the shaded nodes are observed. In the right panel, edge widths are proportional to coefficients’ absolute values. The black/red edge colors imply positive/negative coefficients, respectively.
The main contributions in this paper include rigorous identifiability theory and scalable computational pipelines for DDEs. For identifiability, we propose general identifiability conditions in terms of the dependence graph structure between adjacent layers in the graphical model (corresponding to the directed arrows in Figure 1). We work under the minimal possible assumptions in order to flexibly cover the various examples mentioned above. Next, we provide an informal statement of the identifiability conditions. Under an identifiable DDE, we also prove that a penalized-likelihood based estimator is consistent for estimating both the continuous parameters as well as the discrete graph structure.
Theorem (Informal version of Theorems 1 and 2).
The DDE with a known number of latent variables is identifiable up to latent variable permutation in each layer, as long as each latent variable has at least three exclusive children. Under a weaker notion of “generic identifiability”, this condition can be relaxed to that each latent variable has at least three children that are not necessarily exclusive.
For computation, we propose a scalable estimation pipeline for DDEs. The multiple layers of nonlinear latent variables in DDEs lead to a highly nonconvex optimization landscape with potentially exponentially many local optima. To address this challenging setting, our computational pipeline features a nuanced layerwise nonlinear spectral initialization followed by a penalized stochastic approximation EM algorithm. This procedure can efficiently handle models with a large number of latent variables. We achieve favorable simulation results for as many as $K^{(1)}=30$ binary latent variables in the shallowest latent layer, which amount to $2^{K^{(1)}}$ mixture components and define a very expressive model. Extensive simulation studies not only validate the identifiability results, but also demonstrate the excellent performance of the proposed algorithms. We apply DDEs to real data in three diverse tasks, including hierarchical topic modeling, image representation learning, and response time modeling in digital educational testing. Across these applications, DDEs extract highly interpretable latent structures and learn useful representations for downstream analyses.
We make brief remarks to place DDEs in the rapidly emerging field of generative AI. For the considered unsupervised learning setting, we use “generative model” to refer to probabilistic models where observed data are generated conditional on hidden variables. Classical examples with one hidden layer include mixture models and mixed membership models. Recently, powerful multilayer models have been proposed for more complex tasks, which either consist of multiple latent layers or a single latent layer transformed by deep neural networks. Popular models in machine learning include deep belief networks (Hinton et al.,, 2006), deep Boltzmann machines (Salakhutdinov and Hinton,, 2009), variational autoencoders (Ranzato and Szummer,, 2008), generative adversarial networks (Goodfellow et al.,, 2014), diffusion models (Sohl-Dickstein et al.,, 2015), and transformer-based models (Vaswani et al.,, 2017) such as large language models. See Section 2.2 for more discussions.
Organization. Section 2 formally defines DDEs and elaborates on its connection with existing models. Section 3 provides identifiability results of DDEs and proves the consistency of a penalized-likelihood estimator. Section 4 presents scalable computational algorithms for estimating DDEs. Section 5 and Section 6 present simulation studies and real-data applications under various data types. Section 7 concludes the paper. All technical proofs, and additional details about computation and data analysis are in the Supplementary Material.
2 Deep Discrete Encoders
Notation. For a positive integer $K$, denote $[K]=\{1,\ldots,K\}$. For a matrix $\mathbf{G}$ with $J$ rows, let $\mathbf{g}_{1},\ldots,\mathbf{g}_{J}$ denote its row vectors. Let $\mathbf{0}_{K},\mathbf{1}_{K}$ be the all-zero vector and all-one vector of length $K$, respectively. Let $\mathbf{e}_{k}$ be the $K$-dimensional canonical basis vector. Let $\operatorname{\emph{g}_{\text{logistic}}}(x):={e^{x}}/({1+e^{x}})$ denote the logistic/sigmoid function. For a finite set $I$, let $S_{I}$ denote the collection of all permutation maps of $I$, and let $\text{id}_{I}$ be the identity permutation on $I$. For nonnegative sequences $\{(a_{N},b_{N})\}_{N\geq 1}$, write $a_{N}\ll b_{N}$ when ${a_{N}}/{b_{N}}\to 0$ as $N\to\infty$.
2.1 Model Definition
We formally define the $D$-latent-layer DDE as a generative model with $D$ discrete latent layers. For each $d\in[D]$, assume that the $(d-1)$th layer is generated conditional on the $d$th. Here, only the bottom layer (indexed by $d=0$) is observed, and all other layers are latent. The bottom layer data can take arbitrary values, but all latent variables are binary, similar to the celebrated deep belief networks (Hinton et al.,, 2006). Let $\mathbf{Y}=(Y_{1},\ldots,Y_{J})\in\times_{j=1}^{J}\mathcal{Y}_{j}$ denote the $J$-dimensional observed responses, where $\mathcal{Y}_{j}$ is the sample space for the $j$th response; see the end of this subsection for concrete examples. We work under the minimal assumption that each $\mathcal{Y}_{j}$ is a separable metric space. Denote the $d$th latent layer as $\mathbf{A}^{(d)}=\big{(}A_{1}^{(d)},\ldots,A_{K^{(d)}}^{(d)}\big{)}\in\{0,1\}^{K^{(d)}}$, which is a $K^{(d)}$-dimensional binary vector. A DDE has a shrinking-ladder-shaped architecture, with the dimension of each layer decreasing as $d$ increases: $K^{(D)}<\ldots<K^{(1)}<J$. See Figure 1 for a graphical model representation.
We further specify the distribution of the directed graphical model in a top-down manner. The directed edges in Figure 1 are pointing downward, meaning that the deepest latent variables in the top layer $d=D$ are generated first. The top layer latent variables are assumed to be independent Bernoullis with parameter ${\boldsymbol{p}}=(p_{1},\ldots,p_{K^{(D)}})$:
$\mathbb{P}(\mathbf{A}^{(D)}=\boldsymbol{\alpha}^{(D)})=\prod_{k=1}^{K^{(D)}}\mathbb{P}(A_{k}^{(D)}=\alpha_{k}^{(D)})=\prod_{k=1}^{K_{D}}p_{k}^{\alpha_{k}^{(D)}}(1-p_{k})^{1-\alpha_{k}^{(D)}},\quad\forall\boldsymbol{\alpha}^{(D)}\in\{0,1\}^{K^{(D)}}.$ (1)
Next, define the middle latent layers inductively as follows. For each $d>1$, suppose that $\mathbb{P}(\mathbf{A}^{(d)})$, the distribution of the $d$th layer, is given. Then, we define the $(d-1)$th layer distribution by assuming the conditional independence of $A_{1}^{(d-1)},\ldots,A_{K^{(d-1)}}^{(d-1)}$ given $\mathbf{A}^{(d)}$:
$\mathbb{P}(\mathbf{A}^{(d-1)}=\boldsymbol{\alpha}^{(d-1)}\mid\mathbf{A}^{(d)})=\prod_{k=1}^{K^{(d-1)}}\mathbb{P}(A^{(d-1)}_{k}=\alpha^{(d-1)}_{k}\mid\mathbf{A}^{(d)}),\quad\forall\boldsymbol{\alpha}^{(d-1)}\in\{0,1\}^{K^{(d-1)}}.$ (2)
Here, we additionally model each conditional distribution in (2) as
$A_{k}^{(d-1)}\mid\mathbf{A}^{(d)}\sim\text{Bernoulli}\Big{(}\operatorname{\emph{g}_{\text{logistic}}}\big{(}\beta_{k,0}^{(d)}+\sum_{l=1}^{K^{(d)}}\beta_{k,l}^{(d)}A_{l}^{(d)}\big{)}\Big{)},$ (3)
where $\operatorname{\emph{g}_{\text{logistic}}}$ is the logistic function that maps the real-valued linear combinations to the $[0,1]$-valued Bernoulli parameters (one may use alternative link functions $g:\mathbb{R}\to[0,1]$, such as the probit link). Collect the $\beta_{k,l}$-parameters in a $K^{(d-1)}\times(K^{(d)}+1)$ matrix $\mathbf{B}^{(d)}$, whose first column is the intercepts $\big{(}\beta_{k,0}^{(d)}\big{)}_{k\in[K^{(d-1)}]}$ and remaining parts are $\big{(}\beta_{k,l}^{(d)}\big{)}_{k\in[K^{(d-1)}],l\in[K^{(d)}]}$.
Finally, we model the bottom layer for the observed data. The observed $\mathbf{Y}=(Y_{1},\ldots,Y_{J})$ are modeled by assuming the conditional independence of $Y_{1},\ldots,Y_{J}$ given $\mathbf{A}^{(1)}$. As the observations $\mathbf{Y}$ are not necessarily binary, we replace the Bernoulli conditional distributions in (3) by a general parametric family of the form
$Y_{j}\mid\mathbf{A}^{(1)}\sim\text{ParFam}_{j}\Big{(}g_{j}\Big{(}\beta_{j,0}^{(1)}+\sum_{k=1}^{K^{(1)}}\beta_{j,k}^{(1)}A_{k}^{(1)},\gamma_{j}\Big{)}\Big{)}.$ (4)
Here, $\text{ParFam}_{j}$ denotes a specific parametric family, and let $H_{j}$ denote its parameter space. For convenience, let $p_{j}$ be the probability mass/density function of $\text{ParFam}_{j}$. The $g_{j}:\mathbb{R}\times[0,\infty)\rightarrow H_{j}$ is a known link function that maps the linear combinations to the parameters of the given parametric family. Here, $\gamma_{j}>0$ denotes the dispersion parameter, when there exists one. Throughout the paper, we will state all results under the more general assumption that there exists a dispersion parameter in the parametric family in (4). If not, one may ignore the notation $\gamma$. We further elaborate on the specific parameterizations in (4) for various response types $\mathcal{Y}_{j}$ at the end of this section.
Following the definition of directed graphical models (Koller and Friedman,, 2009), we can write the joint distribution of all observed and latent variables based on (1)–(4):
$\mathbb{P}\big{(}\mathbf{Y},\{\mathbf{A}^{(d)}\}_{d\in[D]}\big{)}=\mathbb{P}(\mathbf{Y}\mid\mathbf{A}^{(1)})\prod_{d=2}^{D}\mathbb{P}(\mathbf{A}^{(d-1)}\mid\mathbf{A}^{(d)})\mathbb{P}(\mathbf{A}^{(D)}).$
The marginal distribution of $\mathbf{Y}$ is obtained by marginalizing out all of the $D$ latent layers:
$~{}\mathbb{P}(\mathbf{Y})=$ $~{}\sum_{\begin{subarray}{c}\boldsymbol{\alpha}^{(d)}\in\{0,1\}^{K^{(d)}}\\ \forall d\in[D]\end{subarray}}\mathbb{P}(\mathbf{Y}|\mathbf{A}^{(1)}=\boldsymbol{\alpha}^{(1)})\prod_{d=2}^{D}\mathbb{P}(\mathbf{A}^{(d-1)}=\boldsymbol{\alpha}^{(d-1)}\mid\mathbf{A}^{(d)}=\boldsymbol{\alpha}^{(d)})\mathbb{P}(\mathbf{A}^{(D)}=\boldsymbol{\alpha}^{(D)}).$ (5)
The $D$-latent-layer DDE is parametrized by $({\boldsymbol{p}},\mathcal{B},\boldsymbol{\gamma})$, where :=$\mathcal{B}:=\{\mathbf{B}^{(d)}\}_{d\in[D]}$ and $\boldsymbol{\gamma}:=(\gamma_{1},\ldots,\gamma_{J})$. Upon the above marginalization, the induced DDE is a highly expressive model with exponentially many latent mixture components. However, this expressivity also introduces identifiability and computation challenges, which we address in Sections 3 and 4.
In many scenarios, it is desirable for the coefficients $\mathcal{B}$ to be sparse, as this leads to a more interpretable and parsimonious data-generating mechanism. The interpretability stems from that if a latent variable is connected to only a few, rather than all, variables in the layer below, then these children variables can help pinpoint the interpretation of the latent parent. Similar sparse deep generative architectures have been considered in deep exponential families (Ranganath et al.,, 2015), Bayesian pyramids (Gu and Dunson,, 2023), and deep cognitive diagnostic models (Gu,, 2024). Additionally, as will be shown in Section 3.1, the sparsity of the coefficients play a key role in facilitating identifiability. To encode the sparsity of $\mathbf{B}^{(d)}$, for each $d\in[D]$, define a $K^{(d-1)}\times K^{(d)}$ binary matrix $\mathbf{G}^{(d)}=(g_{k,l})$, where $g_{k,l}=1$ if the corresponding coefficient $\beta_{k,l}^{(d)}$ is nonzero, and 0 otherwise. Define $K^{(0)}:=J$ for notational convenience. By (2) and (3), $\mathbf{G}^{(d)}$ can also be viewed as the adjacency matrix or “graphical matrix” between the $(d-1)$th layer and the $d$th layer in the graphical model representation (see Figure 1). Collect all graphical matrices $\mathbf{G}^{(d)}$s by defining $\mathcal{G}:=\{\mathbf{G}^{(d)}\}_{d\in[D]}$. Now, we formally define $D$-latent-layer DDEs, which also incorporate $\mathcal{G}$ as unknown parameters.
Definition 1 (DDE).
A $D$-latent-layer DDE with parameters $({\boldsymbol{p}},\mathcal{B},\mathcal{G},\boldsymbol{\gamma})$ is a statistical model with marginal distribution of the observed data as given in (5). When $\mathcal{K}$ is known and fixed, DDEs can be viewed as parametric families with parameters $(\boldsymbol{\Theta},\mathcal{G})$ and probability mass/density functions $\mathbb{P}_{\boldsymbol{\Theta},{\mathcal{G}}}$, where $\boldsymbol{\Theta}:=({\boldsymbol{p}},\mathcal{B},\boldsymbol{\gamma})$ denotes all continuous parameters.
Next, we give some examples of the various response types $\mathcal{Y}_{j}$ allowed in the DDE framework, along with the corresponding link functions $g$ and parametrizations for (4). As mentioned in the Introduction, the later numerical experiments consider three types of responses: (i) binary, (ii) count, and (iii) continuous. We model each of these responses using (i) Bernoulli with $g=g_{\text{logistic}}$, (ii) Poisson with an exponential link $g(x)=e^{x}$, and (iii) Normal with an identity link $g(x,y)=(x,y)$, respectively. Modeling other data types is also straightforward. While not required, we typically consider that the data types are the same across the $p$ features. In such cases, we omit the subscript in $\mathcal{Y}_{j},g_{j}$, and write $\mathcal{Y}$ and $g$.
2.2 Connections with Existing Models
Deep Probabilistic Graphical Models in Machine Learning.
There are many deep latent variable models in the machine learning literature, including the popular deep Boltzmann machine (DBM, Salakhutdinov and Hinton,, 2009), deep belief networks (DBNs, Hinton et al.,, 2006), and deep exponential families (DEFs, Ranganath et al.,, 2015). The DBM and DBN both contain multiple binary latent layers and differ in the directions of the edges; see Figure 2. DBM and DBN have been originally proposed to model binary data and have been later extended to handle continuous or count responses (Cho et al.,, 2013; Gan et al.,, 2015; Li et al.,, 2019). We recommend the review Salakhutdinov, (2015) for more details and references on DBM and DBNs. DEFs are an unsupervised modeling framework that uses exponential families to model conditional distributions for each layer.
[…]
DBM
[…]
DBN
[…]
DEF
[…]
DDE
Figure 2: Comparison of the graphical structure of DDEs to relevant deep generative models. DBM: binary data and binary latent. DBN: binary data and binary latent. DEF: exponential family conditional distributions. DDE: general-response data and binary latent.
Despite their numerous empirical successes, the aforementioned machine learning models are usually fundamentally non-identifiable due to an enormous number of latent variables and parameters organized in a complex nonlinear architecture. These models are typically heavily overparametrized, making it challenging to understand and interpret the latent representations. Moreover, popular existing estimation procedures for these models are developed to maximize a tractable but less theoretically understood alternative to the likelihood, such as contrastive divergence or layer-wise variational approximation.
As to be shown later, DDEs resolve all these issues by assuming a shrinking-ladder architecture of entirely discrete latent variables and potentially sparse layerwise connections (see Figure 2). This allows us to establish identifiability and consistency as well as effectively reduce the model dimension and interpret the latent structure. Recently, Gu and Dunson, (2023) proposed Bayesian pyramids, which are identifiable multilayer discrete latent variable models with a pyramid structure. But the methodology and identifiability theory therein are restricted to multivariate categorical data. DDEs significantly broaden the applicability of Bayesian pyramids by modeling arbitrarily flexible data distributions while still remaining identifiable. For computation, we propose scalable frequentist estimation approaches for DDEs, rather than adopting Bayesian MCMC inference as in Gu and Dunson, (2023).
Psychometrics.
Multidimensional binary latent variables are also popular for modeling students’ item response data in educational measurement (Junker and Sijtsma,, 2001; von Davier,, 2008; de la Torre,, 2011). These psychometric models are known as cognitive diagnostic models (CDMs). A CDM uses a single latent layer of cognitive skills to model a student’s binary responses to many questions in a test, with a sparse loading graph between the observed and latent layers. Here, each binary latent variable represents a student’s mastery or deficiency of a cognitive skill, and the sparse graph between the observed item responses and latent skills encodes which skills each item is designed to measure in the test. The proposed DDEs substantially generalize this modeling idea by allowing (a) multilayer latent variables, which can model one’s knowledge structure at multiple resolutions ranging from fine-grained details to general concepts (Gu,, 2024), and (b) modeling a rich class of general responses, going beyond the typical binary correct/incorrect responses in educational tests.
3 Theoretical Guarantees for General DDEs
3.1 Model Identifiability
In this section, we establish the identifiability of DDEs. Assuming known numbers of latent variables, we prove that the continuous model parameters and the discrete graph structures between layers can be uniquely identified under certain conditions on the true graphical matrices $\mathbf{G}^{(d)}$’s. We first make an assumption to address some trivial ambiguities.
Assumption 1.
Assume that the graphical matrices $\mathcal{G}=\{\mathbf{G}^{(d)}\}_{d\in[D]}$ and proportion parameters $\boldsymbol{p}$ satisfy the following conditions. Here, for notational convenience, let $K^{(0)}:=J$.
  1. (a)
    For all $k\in[K^{(D)}]$, $p_{k}\in(0,1)$.
  2. (b)
    For all $d\in[D]$, $\mathbf{G}^{(d)}$ does not have all-zero columns and is faithful in the sense that for any $k\in[K^{(d-1)}],l\in[K^{(d)}]$, $g_{k,l}^{(d)}=0$ if and only if $\beta_{k,l+1}^{(d)}=0$.
  3. (c)
    For any $d\in[D]$, all column sums of $\mathbf{B}^{(d)}$ except for the first column are strictly positive.
Condition (a) and the first part of (b) is required for the latent dimension $\mathcal{K}$ to be well-defined, in the sense that removing or adding a latent variable must change the marginal distribution (5). Condition (b) is a standard faithfulness assumption in graphical models (e.g. see Definition 3.8 in Koller and Friedman,, 2009) that follows from our definition of $\mathbf{G}^{(d)}$. Condition (c) is introduced to avoid trivial sign-flipping of latent variables. This condition ensures that for each latent variable $A_{k}^{(d)}$, the value $A_{k}^{(d)}=1$ implies a larger coefficient of the $(k+1)$th row of $\mathbf{B}^{(d)}$, and may be replaced with other monotonicity assumptions. For example, one can alternatively assume that the first nonzero coefficient in each column of $\mathbf{B}^{(d)}$ is positive. We emphasize that condition (c) is much weaker compared to the nonnegative coefficient assumption $\beta_{j,k}\geq 0$, which is a popular assumption for various identifiable latent variable models (Donoho and Stodden,, 2003; Chen et al.,, 2020; Lee and Gu,, 2024).
Now, we formally define the parameter space and the notion of identifiability. For multilayer latent variable models, there are inevitable latent variable permutation issues within each latent layer. Hence, we introduce the notion of identifiability up to equivalence.
Definition 2 (Parameter space).
Consider a $D$-latent-layer DDE with $\mathcal{K}=\{K^{(d)}\}_{d\in[D]}$ latent variables. We define the parameter space of the continuous parameters $\boldsymbol{\Theta}$ given $\mathcal{G}$ as $\Omega_{\mathcal{K}}(\boldsymbol{\Theta};\mathcal{G}):=\{\boldsymbol{\Theta}:\beta_{l,k}^{(d)}\neq 0\text{ if and only if }g_{l,k}^{(d)}=1,\gamma_{j}>0\}.$ Define the joint parameter space for all parameters to be $\Omega_{\mathcal{K}}(\boldsymbol{\Theta},\mathcal{G}):=\{(\boldsymbol{\Theta},\mathcal{G}):\boldsymbol{\Theta}\in\Omega_{\mathcal{K}}(\boldsymbol{\Theta};\mathcal{G})\}.$
Definition 3 (Identifiability up to equivalence).
For a D-layer DDE with $\mathcal{K}$ latent variables, we define an equivalence relationship “$\sim_{\mathcal{K}}$” by setting $(\boldsymbol{\Theta},\mathcal{G})\sim_{\mathcal{K}}(\widetilde{\boldsymbol{\Theta}},\widetilde{\mathcal{G}})$ if and only if $\boldsymbol{\gamma}=\widetilde{\boldsymbol{\gamma}}$ and there exists permutations $\sigma^{(d)}\in S_{[K^{(d)}]}$ for all $d\in[D]$ such that the following hold:
  • $p_{k}=\widetilde{p}_{\sigma^{(D)}(k)}$
  • ${g}_{l,k}^{(d)}=\widetilde{{g}}_{\sigma^{(d-1)}(l),\sigma^{(d)}(k)}^{(d)}$ and $\beta_{l,k}^{(d)}=\widetilde{\beta}_{\sigma^{(d-1)}(l),\sigma^{(d)}(k)}^{(d)}$ for all $d\in[D]$, $k\in[K^{(d)}],l\in[K^{(d-1)}]$
Here, we set $\sigma^{(0)}=\textnormal{id}_{[J]}$. We say that the DDE with true parameters $(\boldsymbol{\Theta}^{\star},\mathcal{G}^{\star})$ is identifiable up to $\sim_{\mathcal{K}}$ when for any alternate parameter value $(\boldsymbol{\Theta},\mathcal{G})\in\Omega_{\mathcal{K}}(\boldsymbol{\Theta},\mathcal{G})$ with $\mathbb{P}_{\boldsymbol{\Theta},\mathcal{G}}=\mathbb{P}_{\boldsymbol{\Theta}^{\star},\mathcal{G}^{\star}}$, it holds that $(\boldsymbol{\Theta},\mathcal{G})\sim_{\mathcal{K}}(\boldsymbol{\Theta}^{\star},\mathcal{G}^{\star})$. Here, $\mathbb{P}_{\boldsymbol{\Theta},\mathcal{G}}$ is the marginal distribution of $\mathbf{Y}$ defined in (5).
Despite the seemingly heavy notation, the equivalence relationship is quite natural. Here, each permutation $\sigma^{(d)}$ addresses the label-switching issue for indexing the latent variables in the $d$th latent layer. Below is a main result on the identifiability of $D$-latent-layer DDEs.
Theorem 1 (Identifiability of DDEs).
Consider the $D$-latent-layer DDE with $\mathcal{K}$ latent variables and true parameters $(\boldsymbol{\Theta}^{\star},\mathcal{G}^{\star})$. Suppose that for all $d\leq D-1$, the true graphical matrices $\mathbf{G}^{(d)\star}$ and parameters $\mathbf{B}^{(d)\star}$ satisfy the following conditions:
  1. A.
    $\mathbf{G}^{(d)\star}$ contains two identity matrices after permuting the rows. Without the loss of generality, suppose that the first $2K^{(d)}$ rows of $\mathbf{G}^{(d)\star}$ are $\big{[}\mathbf{I}_{K^{(d)}},\mathbf{I}_{K^{(d)}}\big{]}^{\top}.$
  2. B.
    For any $\boldsymbol{\alpha}^{(d)}\neq\boldsymbol{\alpha}^{\prime(d)}\in\{0,1\}^{K^{(d)}}$, there exists $j>2K^{(d)}$ such that $\sum_{k=1}^{K^{(d)}}\beta_{j,k}^{(d)\star}(\alpha_{k}^{(d)}-\alpha_{k}^{(d)^{\prime}})\neq 0$.
Then, the model components $(\boldsymbol{\Theta},\mathcal{G})$ are identifiable.
Our key observation behind proving the identifiability of the complex deep generative structures in DDEs is that, with discrete latent layers, it suffices to establish identifiability for a one-latent-layer model and proceed in a layer-wise manner inductively. To elaborate, consider a “collapsed” DDE where the latent layers indexed by $d=2,\ldots,D$ are marginalized out to give a probability mass function for the first latent layer: $\{\mathbb{P}(\mathbf{A}^{(1)}=\boldsymbol{\alpha}^{(1)}):\boldsymbol{\alpha}^{(1)}\in\{0,1\}^{K^{(1)}}\}$. If the collapsed model is proven to be identifiable, that means the conditional distributions $\mathbb{P}(Y_{j}\mid\mathbf{A}^{(1)})$ and the marginal distribution $\mathbb{P}(\mathbf{A}^{(1)})$ can be uniquely identified from the data distribution $\mathbb{P}(\mathbf{Y})$. In this case, we can determine all values of $\mathbb{P}(\mathbf{A}^{(1)}=\boldsymbol{\alpha}^{(1)})$ for every $\boldsymbol{\alpha}^{(1)}\in\{0,1\}^{K^{(1)}}$. Then, we can theoretically treat the shallowest latent layer $\mathbf{A}^{(1)}$ as if it were observed, because its probability mass function (pmf) is now identified and known. Then, viewing this pmf as the marginal distribution of $K^{(1)}$-dimensional observed variables of a $(D-1)$-latent-layer Bernoulli-DDE, we can inductively identify all parameters via a layerwise argument. Please see Supplementary Material S.1 for the detailed proof.
We interpret the conditions in Theorem 1. Condition A requires each latent variable to have at least two exclusive children in the layer below. Condition B is more technical and is introduced to distinguish the different binary latent configurations $\boldsymbol{\alpha}\neq\boldsymbol{\alpha}^{\prime}$. Condition B holds when $\mathbf{G}^{(d)\star}$ also contains exclusive children for all latent variables.
The exclusive children requirements in condition A can be further relaxed under a weaker notion of generic identifiability. Generic identifiability allows a measure-zero subset of the parameter space to be non-identifiable, and often holds under weaker conditions than that in Definition 3. As the concept was originally proposed under a continuous parameter space (Allman et al.,, 2009), we consider the following modified definition that considers a smaller parameter space for the coefficients $\mathcal{B}$ given the true graphical matrices.
Definition 4 (Generic identifiability).
Consider a $D$-latent-layer DDE with $\mathcal{K}$ latent variables, graphical matrices $\mathcal{G}^{\star}$, and true parameters belonging to $\Omega_{\mathcal{K}}(\boldsymbol{\Theta};\mathcal{G}^{\star})$. The model is generically identifiable up to $\sim_{\mathcal{K}}$ when
[…]
is a measure-zero set with respect to $\Omega_{\mathcal{K}}(\boldsymbol{\Theta};\mathcal{G}^{\star})$.
We introduce an additional assumption on the parametric families used to model the conditional distribution $Y_{j}\mid\mathbf{A}^{(1)}$ in (4). This is a technical assumption that arises from our proof technique for dealing with measure-zero sets. This assumption holds for all example parametric families described in Section 2.1, and more generally for exponential families with an analytic log-partition function.
Assumption 2 (Analytic family).
Let $p(\cdot;\eta,\gamma)$ be the pmf/pdf of a parametric family, indexed by $\eta,\gamma$ and equipped with a sample space $\mathcal{Y}$. We say that $p$ is analytic when the pmf/pdf $p(Y;\eta,\gamma)$ is analytic in both $\eta,\gamma$, for all $Y\in\mathcal{Y}$.
Now, we are ready to state the generic identifiability result for $D$-latent-layer DDEs.
Theorem 2 (Generic identifiability of DDEs).
Consider the $D$-latent-layer DDE with $\mathcal{K}$ latent variables and true graphical matrices $\mathcal{G}^{\star}$, where all parametric families and link function $g_{j}$s are analytic, and the true parameter lives in $\Omega_{\mathcal{K}}(\boldsymbol{\Theta};\mathcal{G}^{\star})$. Suppose that for all $d\leq D-1$, the true graphical matrices $\mathbf{G}^{(d)}$ satisfy the following condition C:
  1. C.
    After permuting the rows, we can write $\mathbf{G}^{(d)\star}=\left[\mathbf{G}_{1}^{(d)\top},\mathbf{G}_{2}^{(d)\top},\mathbf{G}_{3}^{(d)\top}\right]^{\top}$, where $\mathbf{G}_{1}^{(d)}$ and $\mathbf{G}_{2}^{(d)}$ are $K^{(d)}\times K^{(d)}$ matrices whose diagonal entries are ones:
    $\mathbf{G}_{1}^{(d)},~{}\mathbf{G}_{2}^{(d)}=$ $\begin{pmatrix}1&*&\cdots&*\\ *&1&\cdots&*\\ \vdots&\vdots&\ddots&\vdots\\ *&*&\cdots&1\end{pmatrix}$, where each * denotes an arbitrary value in $\{0,1\}$, and $\mathbf{G}_{3}^{(d)}$ does not have any all-zero columns.
Then, the model components $(\boldsymbol{\Theta},\mathcal{G})$ are generically identifiable.
Condition C relaxes conditions A and B in Theorem 1. While condition A requires two exclusive children for each latent variable, condition C does not require any exclusive child and allows more complicated dependence structures. Additionally, condition B for the remaining $J-2K^{(d)}$ variables is relaxed to a simple non-zero column condition on $\mathbf{G}^{(d)}_{3}$. Thus, condition C does not concern any continuous parameter values and just depends on the graph structure $\mathbf{G}^{(d)}$, and can be used as a practical criterion to assess identifiability.
We place our identifiability results in the literature on the identifiability of generative models and latent variable models. While the identifiability of generative models has attracted increasing attention in machine learning (Hyvarinen et al.,, 2019; Khemakhem et al.,, 2020; Moran et al.,, 2022; Kivva et al.,, 2022), many of these results require additional information such as auxiliary covariates. More importantly, almost all of these results build on nonlinear independent component analysis (Oja and Hyvarinen,, 2000) or variational autoencoders (Ranzato and Szummer,, 2008), both of which essentially have only one latent layer of random variables transformed by deterministic deep neural networks. Consequently, these results cannot be applied to DDEs with multiple latent layers organized in a probabilistic graphical model. Since uncertainty occurs and accumulates in each layer of a DDE, addressing identifiability in such cases requires fundamentally different techniques due to a complicated marginal likelihood. At the high level, our proof techniques are based on transforming the marginal distribution of data into a tensor and invoking the uniqueness of tensor decompositions to establish identifiability. In statistics, the study of identifiability has a long history but also mainly concerns relatively simple latent structures with only one latent layer (Anderson and Rubin,, 1956; Koopmans and Reiersol,, 1950; Allman et al.,, 2009; Xu and Shang,, 2018). It has been largely unknown whether DGMs with many nonlinear latent layers can be identifiable for general data as considered in this paper.
3.2 Estimation Consistency
In this section, we propose a penalized maximum likelihood estimation method for DDEs, and show that the estimator is consistent for both the continuous parameters and the discrete graph structures. Suppose that the numbers of latent variables in all layers are known. We maximize the following objective function to estimate parameters ${\boldsymbol{\Theta}}=({\boldsymbol{p}},\mathcal{B},\boldsymbol{\gamma})$:
$\widehat{\boldsymbol{\Theta}}\in\operatorname*{argmax}_{\boldsymbol{\Theta}}\left[\ell(\boldsymbol{\Theta}\mid\mathbf{Y})-\sum_{d=1}^{D}p_{\lambda_{N},\tau_{N}}(\mathbf{B}^{(d)})\right],$ (6)
where $\ell(\boldsymbol{\Theta}\mid\mathbf{Y})=\sum_{i=1}^{N}\log\mathbb{P}(\mathbf{Y}_{i}\mid\boldsymbol{\Theta})$ denotes the marginal log-likelihood function given a sample $\mathbf{Y}_{1},\ldots,\mathbf{Y}_{N}$ of size $N$, defined based on the marginal distribution of $\mathbf{Y}$ in (5). From now on, we slightly abuse notation and let $\mathbf{Y}$ denote the $N\times J$ data matrix including $\mathbf{Y}_{1},\ldots,\mathbf{Y}_{N}$ as rows. Using the estimated coefficients in ^$\widehat{\mathcal{B}}=(\widehat{\beta}_{l,k}^{(d)})$, the layer-wise graphical matrices in $\mathcal{G}$ can be estimated by reading off the sparsity pattern of $\widehat{\mathcal{B}}$:
$\widehat{g}_{l,k}^{(d)}:=\mathbbm{1}(\widehat{\beta}_{l,k}^{(d)}\neq 0)\quad\text{for all}\quad d\in[D],~{}k\in[K^{(d)}],~{}l\in[K^{(d-1)}].$ (7)
For some tuning parameters $\lambda_{N},\tau_{N}>0$, $p_{\lambda_{N},\tau_{N}}:\mathbb{R}\to[0,\infty)$ is a sparsity-inducing symmetric penalty that is nondecreasing on $[0,\infty)$, nondifferentiable at 00, differentiable at $(0,\tau_{N})$, $p_{\lambda_{N},\tau_{N}}\propto{\lambda_{N}}/{\tau_{N}}$ around 00 and satisfy
$p_{\lambda_{N},\tau_{N}}(b)=0,~{}\text{if}~{}b=0;\quad p_{\lambda_{N},\tau_{N}}^{\prime}(b)\leq\frac{C\lambda_{N}}{\tau_{N}},~{}\text{if}~{}|b|\leq\tau_{N};\quad p_{\lambda_{N},\tau_{N}}(b)=\lambda_{N},\quad\text{if}~{}|b|\geq\tau_{N}.$
Note that $\lambda_{N}$ is the magnitude of the penalty, and $\tau_{N}$ is the point of truncation. Our assumption for the penalty $p_{\lambda_{N},\tau_{N}}$ includes common truncated sparsity-inducing penalties such as the TLP (Truncated Lasso Penalty; Shen et al.,, 2012) and SCAD (Smoothly Clipped Absolute Deviation; Fan and Li,, 2001). With a slight abuse of notation, in (6), we view the penalty $p_{\lambda_{N},\tau_{N}}$ as a function of matrices by letting it be the sum of the entrywise penalties: $p_{\lambda_{N},\tau_{N}}(\mathbf{B}^{(d)})=\sum_{k\in[K^{(d-1)}],\ell\in[K^{(d)}]}p_{\lambda_{N},\tau_{N}}(\beta_{k,\ell}^{(d)}).$
Assuming a compact parameter space with bounded coefficients $\mathbf{B}^{(d)}$, we prove that the estimator defined in (6) and (7) results in consistent estimation.
Theorem 3.
[…]
Remark 1.
One natural question is to whether our estimator would still be consistent when the number of latent variables are unknown. This extension is not straightforward since the number of the top-layer latent variable, $K^{(D)}$, determines the number of deepest mixture components of DDEs. Estimating the number of mixture components is a challenging problem even in simple parametric models, and often leads to a slower (than ${1}/{\sqrt{N}}$) rate of convergence in parameter estimation (Goffinet et al.,, 1992; Ho and Nguyen,, 2016). Note that for such cases, the Fisher information becomes singular, and Theorem 3 cannot be applied. We will give a practical method to choose $\mathcal{K}$ at the end of Section 4.
4 Scalable Computation Pipeline
4.1 Warm-up Example: Penalized EM Algorithm
This subsection first presents a penalized EM algorithm as a warm-up example, before presenting the scalable computational pipelines in Sections 4.2 (stage one) and 4.3 (stage two).
We describe the methods for two-layer DDEs; they are conceptually easy to generalize to DDEs with more layers. Since the coefficients $\mathbf{B}^{(1)}$ and $\mathbf{B}^{(2)}$ are potentially sparse, we propose a penalized EM algorithm (PEM) to compute the penalized maximum likelihood estimator in (6). The PEM is an iterative procedure that consists of an expectation step followed by a penalized maximization step (Green,, 1990; Chen et al.,, 2015). In the $(t+1)$th iteration, the E-step computes the expectation of the complete data penalized-log-likelihood
$\ell_{c}(\mathbf{Y},\mathbf{A}^{(1)},\mathbf{A}^{(2)};\boldsymbol{\Theta})-$ $\sum_{d=1}^{2}p_{\lambda_{N},\tau_{N}}(\mathbf{B}^{(d)})=~{}\sum_{i=1}^{N}\Bigg{[}\log\mathbb{P}(\mathbf{A}_{i}^{(2)};{\boldsymbol{p}})+\log\mathbb{P}(\mathbf{A}_{i}^{(1)}\mid\mathbf{A}_{i}^{(2)};\mathbf{B}^{(2)})$
$+\log\mathbb{P}(\mathbf{Y}_{i}\mid\mathbf{A}_{i}^{(1)},\mathbf{A}_{i}^{(2)};\mathbf{B}^{(1)},\boldsymbol{\gamma})\Bigg{]}-\sum_{d=1}^{2}p_{\lambda_{N},\tau_{N}}(\mathbf{B}^{(d)}).$
This requires calculating the conditional probability for each latent configuration using the previous parameter estimates; that is, calculating $\mathbb{P}(\mathbf{A}_{i}^{(1)}=\boldsymbol{\alpha}^{(1)},\mathbf{A}_{i}^{(2)}=\boldsymbol{\alpha}^{(2)}\mid\mathbf{Y};\boldsymbol{\Theta}^{[t]})$ for all $i\in[N]$, $\boldsymbol{\alpha}^{(1)}\in\{0,1\}^{K_{1}}$, and $\boldsymbol{\alpha}^{(2)}\in\{0,1\}^{K_{2}}$. Here, the superscript $[t]$ denotes the $t$th iteration estimates. In the M-step, we update the parameters by maximizing the expectation computed in the E step, which boils down to solving the following three maximizations:
${\boldsymbol{p}}^{[t+1]}:=$ $\operatorname*{argmax}_{{\boldsymbol{p}}}\sum_{i=1}^{N}\mathbb{E}\left[\log\mathbb{P}(\mathbf{A}_{i}^{(2)};{\boldsymbol{p}});{\boldsymbol{p}}^{[t]}\right],$ (8)
$\mathbf{B}^{(2),[t+1]}:=$ $\operatorname*{argmax}_{\mathbf{B}^{(2)}}\sum_{i=1}^{N}\mathbb{E}\left[\log\mathbb{P}(\mathbf{A}_{i}^{(1)}\mid\mathbf{A}_{i}^{(2)};\mathbf{B}^{(2)});\mathbf{B}^{(2),[t]}\right]-p_{\lambda_{N},\tau_{N}}(\mathbf{B}^{(2)}),$ (9)
$(\mathbf{B}^{(1),[t+1]},\boldsymbol{\gamma}^{[t+1]}):=$ $\operatorname*{argmax}_{\mathbf{B}^{(1)},\boldsymbol{\gamma}}\sum_{i=1}^{N}\mathbb{E}\left[\log\mathbb{P}(\mathbf{Y}_{i}\mid\mathbf{A}_{i}^{(1)},\mathbf{A}_{i}^{(2)});\mathbf{B}^{(1),[t]},\boldsymbol{\gamma}^{[t]})\right]-p_{\lambda_{N},\tau_{N}}(\mathbf{B}^{(1)}).$ (10)
Here, the optimizations for each layer are separated; we update the top-layer proportion parameters ${\boldsymbol{p}}$ in (8), the middle latent layer coefficients $\mathbf{B}^{(2)}$ in (9), and the bottom layer coefficients $(\mathbf{B}^{(1)},\boldsymbol{\gamma})$ in (10). Additionally, due to the conditional independence assumption in each layer, the maximizations can be further simplified into low-dimensional optimizations over each row of $\mathbf{B}$. Algorithm 1 summarizes the PEM algorithm. The details of the simplified M-step and the implementation are in Supplementary Material S.3.1 and S.3.3.
Data: $\mathbf{Y},\mathcal{K}$, tuning parameters $\lambda_{N},\tau_{N}$.
Initialize $\boldsymbol{\Theta}^{[0]}$ using the layerwise initialization in Algorithm 2.
while log-likelihood has not converged do
       In the $[t+1]$th iteration,
      // E-step
      for $(i,\boldsymbol{\alpha}^{(1)},\operatorname{\boldsymbol{\alpha}^{(2)}}2)\in[N]\times\{0,1\}^{K^{(1)}}\times\{0,1\}^{K^{(2)}}$ do
$\varphi^{[t+1]}_{i,\operatorname{\boldsymbol{\alpha}^{(1)}},\operatorname{\boldsymbol{\alpha}^{(2)}}2}=\mathbb{P}(\mathbf{A}_{i}^{(1)}=\boldsymbol{\alpha}^{(1)},\mathbf{A}_{i}^{(2)}=\boldsymbol{\alpha}^{(2)}\mid\mathbf{Y};\boldsymbol{\Theta}^{[t]})$
       end for
      // M-step
       update $\boldsymbol{\Theta}^{[t+1]}=({\boldsymbol{p}}^{[t+1]},\mathbf{B}^{(1),[t+1]},\mathbf{B}^{(2),[t+1]},\boldsymbol{\gamma}^{[t+1]})$ by solving (8)-(10).
end while
Estimate $\mathbf{G}$ based on the sparsity structure of $\widehat{\mathbf{B}}$ according to (7).
Output : Estimated continuous parameters $\widehat{\boldsymbol{\Theta}}$ and graphical matrices $\widehat{\mathbf{G}}^{(1)}$, $\widehat{\mathbf{G}}^{(2)}$.
Algorithm 1 Penalized EM (PEM) algorithm for the two-latent-layer DDE
4.2 Stage One: Layerwise Double-SVD Initialization
The multiple layers of nonlinearity in DDEs lead to a highly nonconvex optimization landscape with potentially exponentially many local optima (Sutskever et al.,, 2013). If the penalized EM algorithm starts with an initialization close to a local optima, it can get stuck and fail to converge to the global maximizer of the penalized log-likelihood function. Hence, for highly complex latent variable models such as DDEs, it is critical to initialize the optimization algorithm wisely. We propose a novel layerwise nonlinear spectral initialization strategy, which enjoys low computational complexity and reasonably high accuracy. This spectral initialization serves as the first stage in the proposed computational pipeline.
Spectral methods have mostly been used for linear low-rank models, and existing approaches are not directly applicable for DDEs with a deep nonlinear structure. To address this, we propose a nuanced layerwise procedure utilizing the double SVD procedure for denoising low-rank generalized linear factor models (Zhang et al.,, 2020) and the SVD-based Varimax to find sparse rotations of the factor loadings (Rohe and Zeng,, 2023).
Data: Data matrix $\mathbf{Y}_{N\times J}$, latent dimensions $\mathcal{K}$.
  1. 1.
    De-noise the data matrix $\mathbf{Y}$ using a first SVD, and linearize this matrix by applying the inverse-link function $(\mu\circ g)^{-1}$ elementwisely. Let $\widehat{\mathbf{Z}}$ denote the inverted matrix.
  2. 2.
    Let $\widehat{\mathbf{Z}}_{0}$ be the column-centered version of $\widehat{\mathbf{Z}}$, and compute its rank-$K^{(1)}$ approximation by a second SVD: $\widehat{\mathbf{Z}}_{0}\approx\mathbf{U}_{N\times K^{(1)}}\mathbf{\Sigma}_{K^{(1)}\times K^{(1)}}\mathbf{V}_{J\times K^{(1)}}^{\top}$.
  3. 3.
    Rotate $\mathbf{V}$ according to the Varimax criteria, and denote it as $\widehat{\mathbf{B}}^{(1)}$. Modify the sign ($\pm$) of each column so that 2(c) is satisfied.
  4. 4.
    Use the relationship $\widehat{\mathbf{Z}}\approx[\mathbf{1}_{N},{\mathbf{A}}^{(1)}]{\mathbf{B}}^{(1)\top}$ to estimate $\mathbf{A}^{(1)}$.
  5. 5.
    Now, suppose that the estimated $\widehat{\mathbf{A}}^{(1)}$ is the “observed data” of a one-latent-layer DDE. Repeat steps 1-4 to estimate ${\mathbf{B}}^{(2)}$ and $\mathbf{A}^{(2)}$.
Algorithm 2 Outline of the Layerwise Double-SVD Initialization
Our main idea is to view the responses $\mathbf{Y}$ as a perturbation of a population expectation $\mathbb{E}(\mathbf{Y}\mid\mathbf{A}^{(1)},\mathbf{B}^{(1)})=(\mu\circ g)\left([\mathbf{1}_{N},\mathbf{A}^{(1)}]\mathbf{B}^{(1)\top}\right)$, which is an elementwise (nonlinear) transformation of a low-rank matrix. Here, $\mu:H\to\mathcal{Y}$ is the known mean function of the observed-layer parametric family in (4) and $g:\mathbb{R}\to H$ is the link function. The function $(\mu\circ g)$ is equal to $\operatorname{\emph{g}_{\text{logistic}}}$ for Bernoulli responses, the exponential function for Poisson, and the identity function for Normal. When $(\mu\circ g)$ is nonlinear, we use the aforementioned double SVD procedure (Zhang et al.,, 2020). This procedure applies a first SVD to de-noise the data, and then linearizes the data through inverting the link function $(\mu\circ g)$, and finally performs a second SVD to find the low-rank matrix $\widehat{\mathbf{Z}}\approx[\mathbf{1}_{N},\mathbf{A}^{(1)}]\mathbf{B}^{(1)\top}$. Next, noting that the true coefficient matrix $\mathbf{B}^{(1)}$ is sparse, we estimate it by seeking a sparse rotation of the right singular subspace of $\widehat{\mathbf{Z}}$, using the popular Varimax criterion (Kaiser,, 1958; Rohe and Zeng,, 2023). This procedure also provides an estimate of the latent variables $\mathbf{A}^{(1)}$, which can be treated as the “observed data” to initialize the deeper layer’s $\mathbf{B}^{(2)}$ and $\mathbf{A}^{(2)}$ in a similar fashion as described above. This layer-by-layer algorithm readily generalizes to deeper models and is reminiscent of the greedy learning strategy for DBNs (Hinton et al.,, 2006; Salakhutdinov,, 2015).
We summarize the overall procedure in Algorithm 2, and postpone further details of each step to Supplementary Material S.2. In the Supplement, we also illustrate the effectiveness of the spectral initialization by comparing the estimation accuracy of our two-stage computational pipeline to that of the EM algorithm with a random initialization.
4.3 Stage Two: Penalized SAEM Algorithm
For DDEs with a large number of latent variables, the E-step in Algorithm 1 requires computing all conditional probabilities $\mathbb{P}(\mathbf{A}^{(1)}_{i}=\boldsymbol{\alpha}^{(1)},\mathbf{A}^{(2)}_{i}=\boldsymbol{\alpha}^{(2)}\mid\mathbf{Y};\boldsymbol{\Theta}^{[t]})$ for all $\boldsymbol{\alpha}^{(1)}\in\{0,1\}^{K^{(1)}}$ and $\boldsymbol{\alpha}^{(2)}\in\{0,1\}^{K^{(2)}}$. This requires computing and storing $O(N\times 2^{\sum_{d=1}^{D}K^{(d)}})$ terms. The exponential dependency in $K^{(d)}$ is concerning even for moderately large latent dimensions, say $K^{(d)}=10$, and quickly becomes prohibitive for larger $K^{(d)}$. Therefore, we propose a penalized Stochastic Approximate EM (SAEM; see Delyon et al.,, 1999; Kuhn and Lavielle,, 2004) by modifying both the E-step and M-step to more scalable versions using approximate sampling. As we illustrate below, this is a method with linear dependence of $\sum_{d=1}^{D}K^{(d)}$ on the computing time as well as memory, and is a scalable alternative to Algorithm 1.
We elaborate on the details on how to modify the PEM to derive the SAEM. First, we replace the E-step to a simulation step, which consists of simulating only a small number (denoted as $C$) of posterior samples of the latent variables. As exact sampling from the joint distribution $\mathbb{P}(\mathbf{A}_{i}^{(1)},\mathbf{A}_{i}^{(2)}\mid\mathbf{Y};\boldsymbol{\Theta}^{[t]})$ is expensive, we sample each latent variable from their complete conditionals. That is, we sample each $A_{i,k}^{(1),[t+1]}$ from $\mathbb{P}(A_{i,k}^{(1)}\mid(-),\boldsymbol{\Theta}^{[t]})$, where ()(-)( - ) denotes the estimates of all other latent variables from the $t$th iteration. Since the latent variables are binary, the conditional distribution is Bernoulli and easy to evaluate. Consequently, the computationally expensive E-step is replaced by the simulation step, which computes and stores only $O(N\times\sum_{d=1}^{D}K^{(d)})$ terms. In terms of choosing the number of samples $C$ in each iteration, we empirically find that taking $C=1$ is computationally efficient without sacrificing much accuracy. Additional simulation studies in Supplementary Material S.4.2 demonstrate that taking a larger $C$ results in only little improvement in accuracy, but becomes significantly slower. This choice of $C=1$ was also suggested in the original paper that proposed the SAEM (Delyon et al.,, 1999).
Second, we modify the M-step objective function by replacing the exact conditional probability values (denoted as $\varphi^{[t+1]}$ in Algorithm 1) to sample-based quantities, and also stochastically averaging the objective functions. For example, we modify (9) as follows:
$Q^{(2),[t+1]}(\mathbf{B}^{(2)})$
$:=(1-\theta_{t+1})Q^{(2),[t]}(\mathbf{B}^{(2)})+\theta_{t+1}\sum_{i=1}^{N}\log\mathbb{P}(\mathbf{A}_{i}^{(1)}=\mathbf{A}_{i}^{(1),[t+1]}\mid\mathbf{A}_{i}^{(2)}=\mathbf{A}_{i}^{(2),[t+1]};\mathbf{B}^{(2)}),$ (11)
$\mathbf{B}^{(2),[t+1]}:=\operatorname*{argmax}_{\mathbf{B}^{(2)}}\left[Q^{(2),[t+1]}(\mathbf{B}^{(2)})-p_{\lambda_{N},\tau_{N}}(\mathbf{B}^{(2)})\right],$ (12)
where $Q^{(2),[0]}=0$ and $\theta_{t+1}$ is a pre-determined step size that decreases in $t$. Here, in the log probability term in (11), $\mathbf{A}_{i}^{(d)}$ denotes the latent random variable and $\mathbf{A}_{i}^{(d),[t+1]}$ denotes the realized sample from the simulation step in the current $(t+1)$th iteration. In the $(t+1)$th iteration, we update the objective function $Q^{(2),[t+1]}$ by taking a weighted average of the previous objective function $Q^{(2),[t]}$, and the conditional probabilities computed using the current iteration’s simulated samples $\mathbf{A}^{(1),[t+1]}$. Then, in (12) we compute the parameters that maximize the penalized objective function.
Data: $\mathbf{Y},\mathcal{K}$, tuning parameters $\lambda_{N},\tau_{N}$.
Initialize $\mathbf{A}^{(1),[0]},\mathbf{A}^{(2),[0]}$ and $\boldsymbol{\Theta}^{[0]}$ based on Algorithm 2.
while $\|{\boldsymbol{\Theta}}^{[t]}-{\boldsymbol{\Theta}}^{[t-1]}\|$ is larger than a threshold do
       In the $t$th iteration,
      // Simulation-step
      for $(i,k)\in[N]\times[K^{(1)}]$ do
             Sample each $A_{i,k}^{(1),[t+1]}$ from the complete conditional $\mathbb{P}(A_{i,k}^{(1)}\mid(-))$ using the previous parameter estimates $\boldsymbol{\Theta}^{[t]},\mathbf{A}^{[t]}$
      for $(i,l)\in[N]\times[K^{(2)}]$ do
             Sample each $A_{i,l}^{(2),[t+1]}$ from the complete conditional $\mathbb{P}(A_{i,l}^{(2)}\mid(-))$ using the previous parameter estimates $\boldsymbol{\Theta}^{[t]},\mathbf{A}^{[t]}$
      // Stochastic approximation M-step
      update the parameters $\boldsymbol{\Theta}^{[t+1]}$ by maximizing the stochastic averaged objectives (e.g. see (12))
Estimate $\mathbf{G}$ based on the sparsity structure of $\widehat{\mathbf{B}}$ according to (7).
Output : Estimated continuous parameters $\widehat{\boldsymbol{\Theta}}$, estimated graphical matrices $\mathcal{G}$.
Algorithm 3 Penalized SAEM algorithm for the two-latent-layer DDE
We similarly modify the optimizations in (8)–(10). Algorithm 3 summarizes our final SAEM algorithm with $C=1$, where exact formulas are deferred to Supplement S.3.1. Similar to the PEM algorithm, the M-step updates can be written in terms of low-dimensional maximizations for each row of the coefficient matrices.
Selecting the latent dimension $\mathcal{K}$.
To apply the above computation pipeline to real data, one additionally needs to specify the number of latent variables, $\mathcal{K}$. For this purpose, we propose a layer-wise estimation strategy, which can be incorporated into our initialization procedure in Algorithm 2. Recall the denoised data matrix $\widehat{\mathbf{Z}}$ from Step 1 of Algorithm 2 and let $\sigma_{1},\sigma_{2},\ldots$ be its singular values in the descending order. Given a candidate grid $\mathfrak{K}$, we estimate the number of latent variables in the first layer based on the largest spectral ratio: $\widehat{K}^{(1)}:=\operatorname*{argmax}_{k\in\mathfrak{K}}({\sigma_{k}}/{\sigma_{k+1}})-1$. Now, given $\widehat{K}^{(1)}$, we proceed with the remaining steps of Algorithm 2 to estimate the first-layer latent variables $\mathbf{A}^{(1)}$. By treating the estimated $\mathbf{A}^{(1)}$ as the observed variables of a one-latent-layer DDE, we can repeat the above procedure to estimate $\widehat{K}^{(2)}$. See Supplement S.3.2 for alternative selection criteria for $\mathcal{K}$.
5 Simulation Studies
We conduct extensive simulation studies in various settings to assess the performance of the proposed computation pipeline (Algorithms 2 and 3) and validate our identifiability conditions (in Section 3.1). We generate data from two-latent-layer DDEs, exploring a total of 90 true settings by varying the following:
  1. (a)
    three parametric families: Bernoulli, Poission, Normal,
  2. (b)
    three paramter dimensions: $(J,K^{(1)},K^{(2)})=(18,6,2),(54,18,6),(90,30,10)$,
  3. (c)
    two parameter values: see (13) below,
  4. (d)
    five varying sample sizes: $N=500,1000,2000,4000,8000$.
Given a value of $K^{(2)}$, set $K^{(1)}=3K^{(2)}$ and $J=9K^{(2)}$, which allows a large latent dimension with as many as $K^{(1)}=30$ binary latent variables in the shallowest latent layer.
In terms of the true parameter values, we consider two sets of values based on the strict and generic identifiability conditions in Theorems 1 and 2, respectively. In particular, we define $\mathcal{B}_{s}=\{\mathbf{B}_{s}^{(d)}\}_{d=1,2}$ that satisfy Theorem 1 as follows:
$\mathbf{B}_{s}^{(d)}=\begin{pmatrix}-2\mathbf{1}_{K^{(d)}}&4\mathbf{I}_{K^{(d)}}\\ -4\mathbf{1}_{K^{(d)}}&4\mathbf{I}_{K^{(d)}}\\ -2\mathbf{1}_{K^{(d)}}&\mathbf{B}_{1}^{(d)}\end{pmatrix},\quad\text{where}\quad\beta_{1;j,k}^{(d)}:=\begin{cases}4&\text{if }k=j,\\ 4/3&\text{ else if }|k-j|=K^{(d)}/2,\\ 0&\text{otherwise}.\end{cases}$ (13)
We also consider $\mathcal{B}_{g}=\{\mathbf{B}_{g}^{(d)}\}_{d=1,2}$ that satisfy Theorem 2, whose exact values are displayed in Supplement S.4.1. We set the top-layer proportion parameters as $p_{k}=0.5$ for all $k\in[K^{(2)}]$, and set the dispersion parameters for the Normal distribution as $\gamma_{j}=\sigma_{j}^{2}=1$ for all $j\in[J]$.
Figure 3: Estimation error for $\mathcal{G}$ and $\boldsymbol{\Theta}$ under the two-latent-layer DDE with parameters $\mathcal{B}_{s}$ and various observed-layer parametric families.
For each scenario, we run $100$ independent simulations, and display the estimation results in Figure 3. The results under the generic identifiable parameters $\mathcal{B}_{g}$ and computation times are included in Supplementary Material S.4.1. We measure the estimation accuracy of the graphical matrices $\mathcal{G}$ by computing the average entrywise accuracy. For the continuous parameters $\boldsymbol{\Theta}$, we report the root mean squared error (RMSE). Under all response types and parameter values, the estimation errors for both $\mathcal{G}$ and $\boldsymbol{\Theta}$ decrease as the sample size $N$ increases. This empirically justifies the identifiability and consistency results. Additionally, by comparing the estimation accuracy across different parametric families, we observe that the Bernoulli is the most challenging to estimate, and the Normal is the easiest. Regarding the true parameters, the sparse model with coefficients $\mathcal{B}_{s}$ is easier to estimate. This is because sparsity helps the initialization and reduces the effective parameter dimension. Finally, under coefficients $\mathcal{B}_{s}$, estimation accuracy improves as the latent and observed dimensions increase, which is a consequence of the decreasing percentage of nonzero entries in $\mathcal{B}$.
In Supplementary Material S.4.3, we present simulation results where the latent dimension $\mathcal{K}$ is unknown. We illustrate that the spectral-gap estimator has near-perfect selection accuracy for a large $N$ (e.g., larger than $4000$), and is superior in terms of both accuracy and computation time compared to alternative estimation strategies.
6 Real Data Applications
We illustrate DDEs’ interpretability, representation power, and downstream prediction accuracy on three diverse real-world datasets. Supplementary Material S.5.1 gives the preprocessing details of all datasets.
6.1 Binary Data: Bernoulli-DDE for MNIST Handwritten Digits
The MNIST dataset for handwritten digits is very popular for classification as well as unsupervised learning (Deng,, 2012). We fit the two-latent-layer DDE with binary responses (Bernoulli-DDE), where the observed layer distributions in (4) are $Y_{j}\mid(\mathbf{A}^{(1)}=\boldsymbol{\alpha}^{(1)})\sim\text{Ber}\big{(}\operatorname{\emph{g}_{\text{logistic}}}(\beta_{j,0}^{(1)}+\sum_{k\in[K^{(1)}]}\beta_{j,k}^{(1)}\alpha_{k}^{(1)})\big{)}$. This resembles existing generative models for images such as DBN and DBM, but we instead consider a much low-dimensional shrinking-ladder shaped latent structure that is identifiable and interpretable. For easier presentation, we consider the subset of images whose true digit labels are 0, 1, 2, and 3. After preprocessing, our training set consists of $N=20,679$ images each with $J=264$ binary pixels. Compared to many existing works that analyzed MNIST, we are considering a more challenging fully-unsupervised setting by holding out all other information about the images, such as the true labels, number of classes, and the spatial location among the pixels.
Basis image [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Positive part [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Negative part [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 1: From left to right: Estimated basis images ${\boldsymbol{\beta}}_{0,0}^{(1)}$, ${\boldsymbol{\beta}}_{0,1}^{(1)}$, ${\boldsymbol{\beta}}_{0,2}^{(1)}$, ${\boldsymbol{\beta}}_{0,3}^{(1)}$, ${\boldsymbol{\beta}}_{0,4}^{(1)}$, ${\boldsymbol{\beta}}_{0,5}^{(1)}$ from the MNIST data, which define sparse subregions in the $28\times 28$ image. The second and third row shows the most significant portion of the basis images by thresholding at the value ±$\pm 1.5$.
We fit the two-latent-layer DDE using Algorithm 1 with the latent dimensions set to $K^{(1)}=5$ and $K^{(2)}=2$ (see the Supplementary Material S.5.2 for the rationale for this choice). Table 1 displays the first-layer coefficients $\mathbf{B}^{(1)}$ by rearranging the each column into the original $28\times 28$ grid. Note that the value of $\mathbf{B}^{(1)}$ is in the logit-scale, so the negative coefficients make the corresponding pixel more likely to be zero. From the reshaped columns ${\boldsymbol{\beta}}_{0,0}^{(1)}$, ${\boldsymbol{\beta}}_{0,1}^{(1)}$, …, ${\boldsymbol{\beta}}_{0,5}^{(1)}$ of $\mathbf{B}^{(1)}$, we can interpret the meaning of each latent variable: $A_{1}^{(1)}=1$ indicates a zero-like shape, $A_{5}^{(1)}=1$ indicates symmetric curves on the upper-left and bottom-right corners, and the other latent variables represent different rotations. Additionally, using the deeper graphical matrix $\mathbf{G}^{(2)}$ displayed in the Supplementary Material S.5.4, we can also interpret the top layer latent variables as broader information about the images. For example, $A_{1}^{(2)}$ indicates large pixel density and $A_{2}^{(2)}$ indicates symmetry with respect to the x-axis.
Refer to captionRefer to caption
Figure 4: Left: Decision tree to estimate the digit $L=0,1,2,3$. Center: Estimated latent representations $\widehat{\mathbf{A}}^{(1)},\widehat{\mathbf{A}}^{(2)}$. Right: Estimated and true digits in the train set.
The learned shallower and deeper latent representations $(\widehat{\mathbf{A}}^{(1)},\widehat{\mathbf{A}}^{(2)})$ are evaluated under two measures of performance: classification accuracy and reconstruction accuracy. We first estimate the latent variables using the $\varphi$ matrix in the EM algorithm:
$(\widehat{\mathbf{A}}^{(1)}_{i},\widehat{\mathbf{A}}^{(2)}_{i})=\operatorname*{argmax}_{\boldsymbol{\alpha}^{(1)}\in\{0,1\}^{K_{1}},~{}\boldsymbol{\alpha}^{(2)}\in\{0,1\}^{K_{2}}}\varphi_{i,\boldsymbol{\alpha}^{(1)},\boldsymbol{\alpha}^{(2)}};\quad\boldsymbol{\alpha}^{(\ell)}\in\{0,1\}^{K_{\ell}},~{}\ell=1,2.$ (14)
Then, a decision tree that classifies the binary latent representations to the categorical label is built by using the misclassification error as the splitting criteria; see the left panel of Figure 4. The center and right panels display the estimated latent traits and estimated labels. Our classifier leads to a high train/test accuracy of $92.0\%/92.6\%$, even though our model is not fine-tuned for image classification. Although the state-of-art machine learning methods can achieve an accuracy as high as $97\%$ (Monnier et al.,, 2020), we point out that the main goal here is not classification, but on interpretability and parsimony; indeed, the DDE uses more limited information and is a less complex but more interpretable model that provides the generative process for the image. In Table 3, we compare the classification accuracy and pixel-wise reconstruction accuracy of the two-latent-layer DDE with alternative interpretable models, such as the latent class model (LCM) and the one-latent-layer DDE. The results show that the two-latent-layer DDE performs the best. We also display example generated images alongside their latent configurations in Table 2, which illustrates various handwriting styles for each digit. Supplementary Material S.5.4 additionally displays sample reconstructed images based on these three methods, and illustrates that the LCM and the one-latent-layer DDE generate noisier and less flexible images than the two-latent-layer DDE.
$[1,0,1,0,1]$ $[0,0,0,0,0]$ $[0,1,0,1,1]$ $[0,0,1,1,1]$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] $[1,0,1,1,1]$ $[0,0,0,1,0]$ $[0,1,1,0,1]$ $[0,0,1,1,0]$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Table 2: Example generated images from the two-latent-layer DDE and their latent representations. Accuracy LCM 1-DDE 2-DDE Train classif. (%) 89.5 84.8 92.0 Test classif. (%) 90.9 86.2 92.6 Train recon. (%) 48.2 79.5 79.6 Test recon. (%) 48.5 79.1 79.9 Table 3: Classification accuracy and pixel-wise reconstruction accuracy for the MNIST dataset.
6.2 Count Data: Poisson-DDE Hierarchical Topic Modeling
Next, we apply DDEs to learn hierarchical latent topics from text documents. Within the field of topic modeling, it is natural for the topics to be correlated with each other (Blei and Lafferty,, 2006), and hierarchical topic modeling is often adopted (Griffiths et al.,, 2003; Paisley et al.,, 2014; Chakraborty et al.,, 2024). While many of the existing works assume a tree-structured hierarchy, DDEs flexibly allow multiple parents for each variable.
We analyze the text corpus from the 20 newsgroups dataset (Lang,, 1995), which was previously analyzed by other topic models with binary latent variables (Srivastava et al.,, 2013; Gan et al.,, 2015). The training set of the corpus consists of approximately 11,000 documents and 2,000 words, and is based on the email messages across 20 newsgroups. After preprocessing by focusing on 12 newsgroups and removing stop words and infrequent words, the dataset consists of $N=5,883$ documents and $J=653$ words. While not used for fitting the DDE, the label (newsgroup) of each document is also provided in the dataset, alongside three larger categories: recreation, computer, and science.
Y1.0e0]ptrecreationtechnologycarssportsmotorcyclesoftwarelocation/nameshardwaregraphicsspacegeb bike banks riding pitt dod bmw gordonhockeynhlplayoffsteamsplayersleagueseasonbaseballenginecheappartsturnedmodelsdealermarketcarfilesdirectoryfilemscopydeletedsizecharlespeter ny chicagodevicesperiodyorkthomasusamotherboard dx CPU mhz controller board ram cardsmouse motif lines color display screen window imageorbit earth research science nasa development project institute
Figure 5: DDE estimated from the 20 newsgroups dataset. For each shallow layer latent variable, we display the top eight representative words. The width of the upper layer arrows is proportional to the corresponding coefficients and the red arrow indicates negative values.
We fit the two-latent-layer DDE with a Poisson-distributed data layer (Poisson-DDE) $Y_{j}\mid(\mathbf{A}^{(1)}=\boldsymbol{\alpha}^{(1)})\sim\text{Poi}\big{(}\exp\big{[}\beta_{j,0}^{(1)}+\sum_{k\in[K^{(1)}]}\beta_{j,k}^{(1)}\alpha_{k}^{(1)}\big{]}\big{)}$, and display the estimated latent structure in Figure 5. We set $K^{(1)}=8$ based on the spectral ratio estimator and $K^{(2)}=2$ based on the identifiability conditions and better interpretability. Additional details behind this choice are given in Supplementary Material S.5.2. Our choice of $K^{(2)}=2$ is also consistent with the correlation structure discovered in Gan et al., (2015). To better interpret individual latent variables, we define representative words for each topic $k$ based on the discrepancy between $\beta_{j,k}$ and all other coefficients, $(\beta_{j,l})_{l\neq k}$. That is, for each index $k$, we choose the words $j$ with the largest values of $\max\{\min\{\beta_{j,k}-\beta_{j,l}:l\neq k\},0\}$. We display the representative words for each latent variable in the bottom row of Figure 5. Here, each latent variable is named based on the representative words and the held-out newsgroup categories.
Compared to the held-out tree structure of the 12 newsgroup labels (see Figure S.7 in the Supplement), one can see that the DDE has discovered a lower-dimensional structure in Figure 5. The latent structure in DDEs allow multiple parents for each topic and effectively model the complex label dependence. For example, ‘cars’, ‘software’, ‘location/names’ have both ‘recreation’ and ‘technology’ as parents, and many bottom-layer words are assigned to multiple topics. We also observe that similar true labels are combined into a single latent variable, for example ‘computer’ and ‘science’ are combined into ‘technology’ in the second latent layer, and ‘baseball’ and ‘hockey’ are combined as ‘sports’ in the bottom latent layer.
We also compare our model fit to existing directed graphical models with matching latent dimensions: LDA (Blei et al.,, 2003) and DPFA-SBN (Gan et al.,, 2015). LDA has a single latent layer with mixed membership scores as continuous latent variables, and DPFA-SBN is a multilayer model with binary latent variables similar to DDEs. We consider the following three metrics widely used in topic modeling to measure different aspects of fit (Chen et al.,, 2023). The first measure is the perplexity, measuring the predictive likelihood of the words in the held-out set. While there are multiple notions of perplexities, we choose the definition that was adopted in similar Poisson-based models (Zhou et al.,, 2012; Gan et al.,, 2015; Ranganath et al.,, 2015, see Supplementary Material S.5.3 for more details). In addition to the train perplexity, we also display the test perplexity. The second measure is the average negative coherence, measuring the quality within each topic by computing $-\frac{1}{K^{(1)}}\sum_{k=1}^{K^{(1)}}\sum_{v_{1},v_{2}\in V_{k}}\log\left({(\text{freq}(v_{1},v_{2})+1)}/{\text{freq}(v_{2})}\right),$ where $V_{k}$ is the top $15$ representative words for the $k$th topic, and the function “freq” counts the number of documents containing the words specified in the input argument. The third measure is the similarity, computing the number of overlapping words within representative words across different topics: $\sum_{1\leq k_{1}<k_{2}\leq K^{(1)}}\sum_{v_{1}\in V_{k_{1}},v_{2}\in V_{k_{2}}}I(v_{1}=v_{2}).$ For all three measures, smaller values are better.
Table 4 summarizes the results and shows the promising fit of DDE. Compared to other models with the same latent dimensions, DDEs have better test perplexity and similarity. The similarity measure shows that while the other model fits exhibit common representative words among different topics, the representative words learned from DDE are entirely disjoint and effectively represent different topics. In terms of coherence, the DDE fit is better than LDA but worse than DPFA. We have also fit models with larger dimensions for comparison, by considering LDA with $256$ latent variables, and DPFA with $K^{(1)}=128,K^{(2)}=64$ latent variables. The dimension for LDA is motivated by that the DDE has $2562^{8}=$ 256 mixture components; while for DPFA, this is the same latent dimension specified in the original paper (Gan et al.,, 2015). We can see that while considering a larger latent dimension may help in terms of perplexity, this leads to a loss of the within-topic coherence as well as dilutes the boundary of each topic, and hence gives less interpretable results.
Model Dimension $\mathcal{K}$ Train perplexity Test perplexity Neg. coherence Similarity
LDA $8$ 499 512 276 43
LDA $256$ 269 515 321 (41450)
DPFA $8-2$ 289 499 211 5
DPFA $128-64$ 175 232 280 (3378)
DDE $8-2$ 322 398 270 0
DDE $8-3$ 322 399 275 0
Table 4: Train and test perplexity scores of different models on the 20 Newsgroups dataset. For all measures, smaller values are better. We parenthesize similarity scores for the models with different dimensions, as the measure is not normalized.
6.3 Continuous Data: Lognormal-DDE for Response Times
We apply the DDE to a response time dataset from the Trends in International Mathematics and Science Study (TIMSS) (Fishbein et al.,, 2021). We analyze the eighth-grade students’ response time for an internet-based mathematics assessment. After preprocessing, the dataset consists of $N=526$ students’ response times for $J=29$ questions. Here, we fit the two-latent-layer DDE where a lognormal distribution is used to model the continuous positive response times (Lognormal-DDE): $Y_{j}\mid(\mathbf{A}^{(1)}=\boldsymbol{\alpha}^{(1)})\sim\text{lognormal}\big{(}\beta_{j,0}^{(1)}+\sum_{k\in[K^{(1)}]}\beta_{j,k}^{(1)}\alpha_{k}^{(1)},\gamma_{j}\big{)}$. This dataset includes additional information regarding the latent structure $\mathbf{G}^{(1)}$, which is the so-called $Q$-matrix in the cognitive diagnostic modeling literature (von Davier,, 2008; Lee and Gu,, 2024). The provisional $\mathbf{G}^{(1)}$ matrix specifies $K^{(1)}=7$ latent cognitive skills (Number, Algebra, Geometry, Data and Probability, Knowing, Applying, and Reasoning) as well as the skills that are required to solve each item. That is, $g^{(1)}_{j,k}=1$ if item $j$ requires latent skill $k$ to solve it. Hence, following the confirmatory latent variable modeling convention in psychometrics, we estimate the DDE parameters by fixing $\mathbf{G}^{(1)}$ to this given structure.
Response \ Latent skill $A_{1}^{(2)}$ $A_{1}^{(1)}$: Number $A_{2}^{(1)}$: Algebra $A_{3}^{(1)}$: Geometry $A_{4}^{(1)}$: Data and Prob
Agree a lot 0.83 0.72 0.88 0.86 0.87
Agree a little 0.80 0.72 0.85 0.84 0.85
Disagree a little 0.76 0.74 0.85 0.84 0.84
Disagree a lot 0.69 0.67 0.76 0.75 0.78
Table 5: Average latent variable estimate for each response category for the question “Mathematics is one of my favorite subjects”.
We fit the two-latent-layer DDE with $K^{(1)}=7,~{}K^{(2)}=1$, and estimate the latent skills based on the posterior probability using (14). We compare the estimated latent variables with the held-out information of each student’s categorical response to a survey question: “Mathematics is one of my favorite subjects”, and display the results in Table 5. The first column shows that the higher-order latent variable, $A_{1}^{(2)}$, is highly correlated with the magnitude that students like math. This suggests that $A_{1}^{(2)}$ can be interpreted as a general indicator for the students’ interest in math, while the fine-grained latent variables $\mathbf{A}^{(1)}$ represent more specific skill mastery for each domain. In addition, we observe that the students who enjoy math tend to have a higher probability of mastering specific skills as well. This is coherent with the fact that the estimated $\mathbf{B}^{(1)}$-coefficients are all positive.
7 Discussion
This paper makes contributions to core AI problems from statisticians’ perspectives by proposing a broad family of interpretable DGMs with solid identifiability guarantees, scalable computation pipelines, and promising application potential. It also opens up interesting directions for future research. Theoretically, it would be desirable to extend our identifiability and consistency results to the challenging setting with an unknown number of latent variables. We conjecture that one may be able to identify $\mathcal{K}$ under additional restrictions on the parameter space or considering specific parametric families. Second, it would be interesting to theoretically justify the good performance of the spectral-gap estimator for estimating the latent dimensions under general nonlinear link functions. Another interesting problem pertains to high-dimensional settings, where the number of observed responses, $J$, may grow with the sample size $N$. Our current notion of identifiability focuses on identifying the population model parameters under the traditional asymptotics with a fixed $J$, where the latent variables are marginalized out in the likelihood. As modern datasets often comes with a large number of observed features, it would be interesting to explore whether our identifiability and estimability results can be generalized to such settings.
In terms of the methodology and applications, an immediate future direction is working to effectively estimate and deploy deeper DDEs with $D\geq 3$ latent layers. The current theory and computational pipeline conceptually readily generalizes to deeper architectures than considered in this initial paper. It remains to test and refine these deeper models’ performance in larger-scale real-world data. Also, it would be interesting to extend DDEs to datasets with additional covariates. For example, the MNIST dataset comes with the actual digit labels as well as the spatial structure of the pixels in the image. In addition, many real datasets contain multiple modalities of data. For example, one may consider a dataset consisting of images uploaded in social media alongside text data such as tags. It is worthwhile to generalize DDEs to model such modern multimodal data. Finally, it would be interesting to extend DDEs for identifiable causal representation learning (Schölkopf et al.,, 2021) to uncover causal structures among the higher-order latent variables.
Supplementary Material.
The Supplement contains all technical proofs, additional theoretical results, and additional details about algorithms, simulations, and real data analyses.
Supplementary Material
This Supplementary Material is organized as follows. Section S.1 proves all main theorems and provides additional identifiability results under one-layer saturated models and variants. Section S.2 provides details regarding the spectral initialization algorithm. Section S.3 gives details regarding the EM algorithms (Algorithms 1 and 3) such as M-step update formulas, implementation details, and selection of the number of latent variables. Section S.4 provides various additional simulation results under (a) generically identifiable true parameters, (b) varying numbers of Monte Carlo samples in the SAEM algorithm, (c) unknown latent dimensions. Finally, Section S.5 gives additional data analysis details such as preprocessing, latent dimension selection, and additional visualizations.
S.1 Proof of Theorems
Recall from Section 3.1 that our identifiability results for general DDEs with multiple latent layers build upon the identifiability of a model with only one latent layer (the shallowest latent layer), where the deeper latent layers have been marginalized out. Here, we formally state identifiability conditions for such one-latent-layer saturated models in Section S.1.1, before proving the identifiability results for general DDEs in Section S.1.2. We prove the claims stated for the one-latent-layer models in Section S.1.3, and Theorem 3 in Section S.1.4. Additional identifiability results under related models with interaction effects (instead of the main-effect DDEs introduced in the main paper) are presented in Section S.1.5.
S.1.1 Identifiability Under One-latent-layer Saturated Models
The one-latent-layer saturated model is defined as follows.
Definition S.1 (One-latent-layer saturated model).
The one-latent-layer saturated model with $K^{(1)}$ latent variables, responses $\mathbf{Y}\in\prod_{j\in[J]}\mathcal{Y}_{j}$, and parameters $(\boldsymbol{\Theta}^{(1)},\mathbf{G}^{(1)})$ is defined by the distribution of $\mathbf{Y}\mid\mathbf{A}^{(1)}$ in (4), and the saturated latent distribution
$\mathbb{P}(\mathbf{A}^{(1)}=\boldsymbol{\alpha})=\pi_{\boldsymbol{\alpha}},\quad\text{for all}\quad\boldsymbol{\alpha}\in\{0,1\}^{K^{(1)}}.$ (S.1)
Here, the parameter $\boldsymbol{\pi}:=(\pi_{\boldsymbol{\alpha}})_{\boldsymbol{\alpha}}$ satisfies $\pi_{\boldsymbol{\alpha}}\in(0,1)$ and $\sum_{\boldsymbol{\alpha}}\pi_{\boldsymbol{\alpha}}=1$. The notation $\boldsymbol{\Theta}^{(1)}:=(\boldsymbol{\pi},\mathbf{B}^{(1)},\boldsymbol{\gamma})$ collects all continuous parameters. We define the parameter spaces $\Omega_{K^{(1)}}(\boldsymbol{\Theta}^{(1)};\mathbf{G}^{(1)})$ and $\Omega_{K^{(1)}}(\boldsymbol{\Theta}^{(1)},\mathbf{G}^{(1)})$ similar to Definition 2 in the main paper.
Here, the term “saturated” indicates that no additional distributional assumptions are imposed on the latent variables, except that they are discrete. Similar to Definition 3, we define an equivalence relationship $\sim_{K^{(1)}}$ when the parameters are identical up to label switching, and use it to define identifiability.
Definition S.2 (Equivalence relation).
For the one-latent-layer saturated model, define an equivalence relationship “$\sim_{K^{(1)}}$” by setting $(\boldsymbol{\Theta}^{(1)},\mathbf{G}^{(1)})\sim_{K^{(1)}}(\widetilde{\boldsymbol{\Theta}}^{(1)},\widetilde{\mathbf{G}}^{(1)})$ if and only if $\boldsymbol{\gamma}=\widetilde{\boldsymbol{\gamma}}$ and there exist a permutation $\sigma^{(1)}\in S_{[K^{(1)}]}$ such that the following conditions hold:
  • $\boldsymbol{\pi}_{(\alpha_{\sigma(1)},\ldots,\alpha_{\sigma(K^{(1)})})}=\widetilde{\boldsymbol{\pi}}_{\boldsymbol{\alpha}}$ for all $\boldsymbol{\alpha}\in\{0,1\}^{K^{(1)}}$
  • ${g}_{j,k}^{(1)}=\widetilde{{g}}_{j,\sigma^{(1)}(k)}^{(1)}$ and $\beta_{j,k}^{(1)}=\widetilde{\beta}_{j,\sigma^{(1)}(k)}^{(1)}$ for all $k\in[K^{(1)}],j\in[J]$.
We say that the one-latent-layer saturated model with true parameters $(\boldsymbol{\Theta}^{(1)\star},\mathbf{G}^{(1)\star})$ is identifiable up to $\sim_{K^{(1)}}$, if for any alternate parameter value $(\boldsymbol{\Theta}^{(1)},\mathbf{G}^{(1)})\in\Omega_{K^{(1)}}(\boldsymbol{\Theta}^{(1)},\mathbf{G}^{(1)})$ with $\mathbb{P}_{\boldsymbol{\Theta}^{(1)},\mathbf{G}^{(1)}}=\mathbb{P}_{\boldsymbol{\Theta}^{(1)\star},\mathbf{G}^{(1)\star}}$, it holds that $(\boldsymbol{\Theta}^{(1)},\mathbf{G}^{(1)})\sim_{\mathcal{K}}(\boldsymbol{\Theta}^{(1)\star},\mathbf{G}^{(1)\star})$. Here, $\mathbb{P}_{\boldsymbol{\Theta}^{(1)},\mathbf{G}^{(1)}}$ is the marginal distribution of $\mathbf{Y}$, which follows from (4) and (S.1).
Under these definitions, we state identifiability results for the one-latent-layer saturated model. Proposition 1 and Proposition 2 are one-layer analogues of Theorem 1 and Theorem 2, respectively. We postpone the proofs of these results to Section S.1.3.
Proposition 1.
Given the knowledge of $K^{(1)}$, the one-latent-layer saturated model with parameters $(\boldsymbol{\Theta}^{\star},\mathbf{G}^{(1)\star})\in\Omega_{K^{(1)}}(\boldsymbol{\Theta}^{(1)},\mathbf{G}^{(1)})$ is identifiable up to $\sim_{K^{(1)}}$ when the true parameters $\mathbf{B}^{(1)\star},\mathbf{G}^{(1)\star}$ satisfy conditions A, B from Theorem 1. In particular, condition B holds when $\mathbf{G}^{(1)\star}$ contains another identity matrix.
Proposition 2.
Consider the one-latent-layer saturated model where all parametric families and link functions $g_{j}$s in (4) are analytic, and the true parameter lives in $\Omega_{K^{(1)}}(\boldsymbol{\Theta}^{(1)};\mathbf{G}^{(1)\star})$. Then, the model is generically identifiable when $\mathbf{G}^{(1)\star}$ satisfies condition C from Theorem 2.
S.1.2 Proof of Theorems 1 and 2
We prove the identifiability results for DDEs using Propositions 1 and 2.
Proof of Theorem 1.
Our argument is based on applying Proposition 1 in a layer-wise manner. First, consider the bottom two layers with
$\pi^{(1)}_{\boldsymbol{\alpha}^{(1)}}=\mathbb{P}(\mathbf{A}^{(1)}=\boldsymbol{\alpha}^{(1)}),\quad\forall\boldsymbol{\alpha}^{(1)}\in\{0,1\}^{K_{1}}$ (S.2)
defined by marginalizing out the deeper latent layers. Now, we can consider this as the one-layer model with proportion parameters $\boldsymbol{\pi}^{(1)}=(\pi^{(1)}_{\boldsymbol{\alpha}^{(1)}},\boldsymbol{\alpha}^{(1)}\in\{0,1\}^{K_{1}})$. Then, Proposition 1 gives the identifiability of $\mathbf{B}^{(1)},\mathbf{G}^{(1)},\boldsymbol{\pi}^{(1)}$ up to $\sim_{K^{(1)}}$.
Having identified $\boldsymbol{\pi}^{(1)}$, the marginal distribution of the shallowest latent layer $\mathbf{A}^{(1)}$ is uniquely identified. We generalize the notation in (S.2) and define $\boldsymbol{\pi}^{(d)}$ similarly. Inductively, for $1\leq d<D$, we apply Proposition 1 by considering $\mathbf{A}^{(d)}$ and $\boldsymbol{\pi}^{(d)}$ as the “observed” binary response vector and its proportion parameters that characterize its marginal probability mass function:
$\pi^{(d)}_{\boldsymbol{\alpha}^{(d)}}=\mathbb{P}(\mathbf{A}^{(d)}=\boldsymbol{\alpha}^{(d)}),\quad\forall\boldsymbol{\alpha}^{(d)}\in\{0,1\}^{K_{d}}.$
Consequently, the parameters $\mathbf{B}^{(d+1)},\mathbf{G}^{(d+1)},\boldsymbol{\pi}^{(d+1)}$ between the $d$th and $(d+1)$th layers are identified up to $\sim_{K^{(d+1)}}$. In particular, when $d=D-1$, it remains to determine ${\boldsymbol{p}}$ from the unstructured proportion parameter vector $\boldsymbol{\pi}^{(D)}$. Since we already have identified the $D$th layer labels up to $\sim_{K^{(D)}}$, this simply follows by marginalizing out the irrelevant coordinates:
$p_{k}=\mathbb{P}(A_{k}^{(D)}=1)=\sum_{\boldsymbol{\alpha}^{(D)}:\alpha^{(D)}_{k}=1}\pi_{\boldsymbol{\alpha}^{(D)}}^{(D)}.$
The proof is complete. ∎
Proof of Theorem 2.
Proposition 2 shows that the non-identifiable measure-zero set of the one-latent-layer saturated model only depends on the coefficients $\mathbf{B}^{(1)}$ and $\boldsymbol{\gamma}$. By marginalizing out all layers except the bottom two layers, the DDE becomes a one-latent-layer saturated model with parameters $\boldsymbol{\Theta}^{(1)}:=(\pi^{(1)},\mathbf{B}^{(1)},\mathbf{G}^{(1)},\boldsymbol{\gamma})$. Since we assume that $\mathbf{G}^{(1)}$ satisfies condition C, the parameters $\boldsymbol{\Theta}^{(1)}$ are identifiable (upto a permutation $\sigma^{(1)}\in S_{[K^{(1)}]}$) as long as $(\mathbf{B}^{(1)},\boldsymbol{\gamma})\not\in N^{(1)}$. Here, $N^{(1)}$ is a measure-zero subset of the coefficient space $\Omega(\mathbf{B}^{(1)},\boldsymbol{\gamma};\mathbf{G}^{(1)})$.
Now, assuming $\mathbf{B}^{(1)}\not\in N^{(1)}$, we can use a similar argument for deeper layers inductively. For $12\leq d\leq D-$ 1, let $N^{(d)}$ be the non-identifiable measure-zero subset of the $d$th layer coefficient space $\Omega(\mathbf{B}^{(d)};\mathbf{G}^{(d)})$. Note that we can still apply Proposition 2 since the conditional distribution of $\mathbf{A}^{(d)}\mid\mathbf{A}^{(d+1)}$ is modeled as a Bernoulli distribution with a logistic link $g_{\text{logistic}}$, which is indeed analytic and satisfy Assumption 2. Consequently, as long as $\mathbf{B}^{(d)}\not\in N^{(d)}$ for all $11\leq d\leq D-$ 1, the $D$-layer DDE is identifiable up to permutations of the latent variables within each layer. The proof is complete since $\cup_{d=1}^{D}(N^{(d)})^{c}$ is a union of a finite number of measure-zero sets, and hence again measure-zero. ∎
S.1.3 Proof of Propositions 1 and 2
S.1.3.1 Additional Notations
We introduce additional notations that will be used to prove identifiability results for the one-latent-layer saturated model. Most of these notations are consistent with Lee and Gu, (2024). First, we omit the superscript “${(1)}$” that indicates the first latent layer when dealing with the one-latent-layer saturated model. Let $j,k,\boldsymbol{\alpha}$ denote typical indices for $j\in[J],k\in[K],\boldsymbol{\alpha}\in\{0,1\}^{K}$. Write $\mathbf{G}=\mathbf{G}^{(1)}=\{\mathbf{g}_{1}^{\top},\ldots,\mathbf{g}_{J}^{\top}\}^{\top}$ and let $H_{j}:=\{k\in[K]:G_{j,k}=1\}$ be the index of the parent latent variables for $Y_{j}$. Recall that for each $j$, the sample space $\mathcal{Y}_{j}$ is a separable metric space. Let $m_{j}$ be a base measure on $\mathcal{Y}_{j}$, this will be the counting measure for discrete sample spaces and the Lebesgue measure for continuous cases. Given $j$ and $\boldsymbol{\alpha}$, define the measure $\mathbb{P}_{j,\boldsymbol{\alpha}}$ on $\mathcal{Y}_{j}$ by setting
[…]
In other words, $\mathbb{P}_{j,\boldsymbol{\alpha}}$ denotes the conditional distribution of $Y_{j}\mid\mathbf{A}=\boldsymbol{\alpha}$ in (4).
For each $j\in[J]$, construct measurable subsets $S_{1,j},\ldots,S_{\kappa_{j},j}\subseteq\mathcal{Y}_{j}$ with $\kappa_{j}\geq 2$, where the collection of vectors $\mathbf{s}_{j}(\boldsymbol{\alpha}):=\big{(}\mathbb{P}_{j,\boldsymbol{\alpha}}(S_{1,j}),\ldots,\mathbb{P}_{j,\boldsymbol{\alpha}}(S_{\kappa_{j},j})\big{)}_{\boldsymbol{\alpha}\in\{0,1\}^{K}}$ is “faithful” in the following sense:
  1. (a)
    for $\boldsymbol{\alpha},\boldsymbol{\alpha}^{\prime}$ with $\boldsymbol{\alpha}_{H_{j}}=\boldsymbol{\alpha}^{\prime}_{H_{j}}$, it holds that $\mathbf{s}_{j}(\boldsymbol{\alpha})=\mathbf{s}_{j}(\boldsymbol{\alpha}^{\prime})$,
  2. (b)
    there exists $\boldsymbol{\alpha},\boldsymbol{\alpha}^{\prime}$ with $\boldsymbol{\alpha}_{H_{j}}\neq\boldsymbol{\alpha}^{\prime}_{H_{j}}$ such that $\mathbf{s}_{j}(\boldsymbol{\alpha})\neq\mathbf{s}_{j}(\boldsymbol{\alpha}^{\prime})$.
This construction is possible since Assumption 1(a) on the parameter space lead to a faithful graphical model. Without loss of generality, suppose that $S_{\kappa_{j},j}=\mathcal{Y}_{j}$ for all $j$. Also, define the following (unordered) set
$\mathcal{S}_{j}:=\Big{\{}\mathbf{s}_{j}(\boldsymbol{\alpha}):\boldsymbol{\alpha}\in\{0,1\}^{K}\Big{\}}.$ (S.4)
Define $\mathbf{N}_{1}$ to be a $\kappa_{1}...\kappa_{K}\times 2^{K}$ matrix by setting
$\mathbf{N}_{1}((l_{1},...,l_{K}),\boldsymbol{\alpha}):=\mathbb{P}(Y_{1}\in S_{l_{1},1},...,Y_{K}\in S_{l_{K},K}\mid\boldsymbol{\alpha}).$
Here, we index the $2^{K}$ columns of $\mathbf{N}_{1}$ using the binary vector $\boldsymbol{\alpha}\in\{0,1\}^{K}$, and the rows by $\xi_{1}=(l_{1},\ldots,l_{K})$, where $l_{j}\in[\kappa_{j}]$. Similarly, let $\mathbf{N}_{2}$ be a $\kappa_{K+1}\ldots\kappa_{2K}\times 2^{K}$ matrix whose $((l_{K+1},\ldots,l_{2K}),\boldsymbol{\alpha})$-th entry is $\mathbb{P}(Y_{K+1}\in S_{l_{K+1},K+1},...,Y_{2K}\in S_{l_{2K},2K}\mid\boldsymbol{\alpha})$, and $\mathbf{N}_{3}$ be a $\kappa_{2k+1}\ldots\kappa_{J}\times 2^{K}$ matrix whose $((l_{2K+1},\ldots,l_{J}),\boldsymbol{\alpha})$-th entry is $\mathbb{P}(Y_{2K+1}\in S_{l_{2K+1},1},...,Y_{J}\in S_{l_{J},J}\mid\boldsymbol{\alpha})$. Similar to $\mathbf{N}_{1}$, we index the rows of $\mathbf{N}_{2}$ and $\mathbf{N}_{3}$ by $\xi_{2}=(l_{K+1},\ldots,l_{2K})$ and $\xi_{3}=(l_{2K+1},\ldots,l_{J})$, respectively. For notational simplicity, let $\upsilon_{1}=\prod_{k=1}^{K}\kappa_{k},~{}\upsilon_{2}=\prod_{k=K+1}^{2K}\kappa_{k},~{}\upsilon_{3}=\prod_{k=2K+1}^{J}\kappa_{k}$. Note that the assumption $S_{\kappa_{j},j}=\mathcal{Y}_{j}$ implies forces the last row in all $\mathbf{N}_{a}$s to be $\mathbf{1}_{2^{K}}^{\top}$.
Next, let $\mathbf{P}_{0}$ be a 3-way marginal probability tensor with size $\upsilon_{1}\times\upsilon_{2}\times\upsilon_{3}$, defined as
$\mathbf{P}_{0}(\xi_{1},\xi_{2},\xi_{3})$ $=\mathbb{P}(Y_{1}\in S_{l_{1}},\ldots,Y_{J}\in S_{l_{J}})$
$=\sum_{\boldsymbol{\alpha}}\pi_{\boldsymbol{\alpha}}\mathbf{N}_{1}((l_{1},\ldots,l_{K}),\boldsymbol{\alpha})\mathbf{N}_{2}((l_{K+1},\ldots,l_{2K}),\boldsymbol{\alpha})\mathbf{N}_{3}((l_{2K+1},\ldots,l_{J}),\boldsymbol{\alpha}).$
We introduce an additional notation for tensor products as follows. For $a=1,2,3$, consider $\nu_{a}\times r$ matrices $\mathbf{M}_{a}$ whose $l$th column is indexed as $\mathbf{m}_{a,l}$. Also, let $\circ$ denote the outer product between vectors. Then, we define the tensor product of $\mathbf{M}_{1},\mathbf{M}_{2},\mathbf{M}_{3}$ as
$[\mathbf{M}_{1},\mathbf{M}_{2},\mathbf{M}_{3}]:=\sum_{l=1}^{r}\mathbf{m}_{1,l}\circ\mathbf{m}_{2,l}\circ\mathbf{m}_{3,l}.$
Using this notation, we can write $\mathbf{P}_{0}$ as follows:
$\mathbf{P}_{0}=[\mathbf{N}_{1}\text{Diag}(\boldsymbol{\pi}),\mathbf{N}_{2},\mathbf{N}_{3}].$ (S.5)
Now, our notation is almost identical to that in the proof of Theorem 1 in Lee and Gu, (2024). Under conditions A and B from Theorem 1, we can apply step 3 and the first two paragraphs of step 4 there to argue that the decomposition (S.5) is unique up to a column permutation. We summarize this in the below Lemma S.1.
Lemma S.1 (Theorem 1 in Lee and Gu, (2024)).
Consider the one-latent-layer saturated model, where the true parameters satisfy conditions A and B. Let $\mathbf{P}_{0}$ be the 3-way marginal probability tensor under these parameters. Then, the tensor decomposition $\mathbf{P}_{0}=[\mathbf{M}_{1},\mathbf{M}_{2},\mathbf{M}_{3}]$ is unique up to a common column permutation. Here, $\mathbf{M}_{a}$s are $\upsilon_{a}\times 2^{K}$ matrices, whose last rows are the all-one vector $\mathbf{1}_{2^{K}}^{\top}$ for $a=2,3$. Additionally, assuming that the graphical matrix $\mathbf{G}$ is known, the model parameters $(\boldsymbol{\Theta},\mathbf{G})$ are identifiable up to sign-flipping.
S.1.3.2 Proof of Proposition 1
The following Lemma will be crucially utilized to identify and recover $\widetilde{\mathbf{G}}$. We present its proof after proving the main Proposition.
Lemma S.2.
Suppose that the parameters $(\mathbf{B},\mathbf{G})$ satisfy part (a) in Assumption 1. For any $j$, define $H_{j},\mathcal{S}_{j},\mathbf{s}_{j}(\boldsymbol{\alpha})$ as above.
  1. (a)
    $|H_{j}|$ and $|\mathcal{S}_{j}|$ satisfies:
    • $|H_{j}|=0$ if and only if $|\mathcal{S}_{j}|=1$
    • $|H_{j}|=1$ if and only if $|\mathcal{S}_{j}|=2$
    • $|H_{j}|\geq 2$ if and only if $|\mathcal{S}_{j}|\geq 3$.
  2. (b)
    $\Big{\{}\mathbf{s}_{j}(\boldsymbol{\alpha}):\alpha_{k}=1\Big{\}}=\Big{\{}\mathbf{s}_{j}(\boldsymbol{\alpha}):\alpha_{k}=0\Big{\}}$ if and only if $g_{j,k}$ = 0.
Proof of Proposition 1.
Our proof builds upon Lemma S.1, which proved a more general notion of nonparametric identifiability of the one-layer model, but under a given graphical matrix $\mathbf{G}$. The cited Theorem 1 proves that the continuous parameters are identifiable up to sign flipping, given $\mathbf{G}$. In our setting, Assumption 1(c) resolves the sign flipping issue and the continuous parameters can be uniquely determined. Consequently, it suffices to show that $\mathbf{G}$ is identifiable up to the equivalence relation $\underset{K^{(1)}}{\sim}$. We separate this proof into two steps.
Step 1: Tensor decomposition and setup. Consider a one-latent-layer saturated model with true parameters $(\boldsymbol{\Theta},\mathbf{G})$ that satisfies conditions A, B. Suppose there exists an alternative set of parameters $(\widetilde{\boldsymbol{\Theta}},\widetilde{\mathbf{G}})$ that define a same marginal distribution of $\mathbf{Y}$. Define the notation ~$\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}$ similar as ${\mathbb{P}}_{j,{\boldsymbol{\alpha}}}$ in (S.3), and also define $\widetilde{\mathbf{N}}_{1},\widetilde{\mathbf{N}}_{2},\widetilde{\mathbf{N}}_{3}$ as the conditional probability matrices that specify the value of
$\mathbb{P}\big{(}(Y_{1},\ldots,Y_{K})\mid\mathbf{A}=\widetilde{\boldsymbol{\alpha}}\big{)},\quad\mathbb{P}\big{(}(Y_{K+1}\ldots,Y_{2K})\mid\mathbf{A}=\widetilde{\boldsymbol{\alpha}}\big{)},\quad\mathbb{P}\big{(}(Y_{2K+1},\ldots,Y_{J})\mid\mathbf{A}=\widetilde{\boldsymbol{\alpha}}\big{)}$
under the alternative parameters. Here, each column of $\widetilde{\mathbf{N}}_{a}$ is denoted by $\widetilde{\boldsymbol{\alpha}}\in\{0,1\}^{K}$, and the last row of each $\widetilde{\mathbf{N}}_{a}$ is the all-one vector $\mathbf{1}_{2^{K}}^{\top}$. Then, the marginal probability tensor $\mathbf{P}_{0}$ defined in (S.5) can be written as
$\mathbf{P}_{0}=[{\mathbf{N}}_{1}\text{{diag}}({\boldsymbol{\pi}}),{\mathbf{N}}_{2},{\mathbf{N}}_{3}]=[\widetilde{\mathbf{N}}_{1}\text{{diag}}(\widetilde{\boldsymbol{\pi}}),\widetilde{\mathbf{N}}_{2},\widetilde{\mathbf{N}}_{3}].$ (S.6)
By applying Lemma S.1, the tensor decomposition in (S.6) is unique, so $\widetilde{\mathbf{N}}_{1}\text{{diag}}({\boldsymbol{\pi}}),\widetilde{\mathbf{N}}_{2},\widetilde{\mathbf{N}}_{3}$ and $\mathbf{N}_{1}\text{{diag}}({\boldsymbol{\pi}}),\mathbf{N}_{2},\mathbf{N}_{3}$ are identical up to a common column permutation, say $\mathfrak{S}\in S_{\{0,1\}^{K}}$. In particular, as the last row of $\widetilde{\mathbf{N}}_{1}\text{{diag}}(\widetilde{\boldsymbol{\pi}})$ and $\mathbf{N}_{1}\text{{diag}}({\boldsymbol{\pi}})$ is exactly $\widetilde{\boldsymbol{\pi}}^{\top}$ and $\boldsymbol{\pi}^{\top}$, $\big{(}\widetilde{\boldsymbol{\pi}},\widetilde{\mathbf{N}}_{1}\big{)}$ and $\big{(}\boldsymbol{\pi},\mathbf{N}_{1}\big{)}$ are also identical up to the same permutation $\mathfrak{S}$.
We use this observation to prove that $\widetilde{\mathbf{G}}$ is equivalent to $\mathbf{G}$ under $\underset{K^{(1)}}{\sim}$. One subtlety for identifying $\mathbf{G}$ is that there is no information about the $2^{K}$ column indices of $\widetilde{\mathbf{N}}_{1},\widetilde{\mathbf{N}}_{2},\widetilde{\mathbf{N}}_{3}$, so we cannot read off the conditional dependence structure directly. We tackle this problem based on the key observation that the unordered set $\mathcal{S}_{j}$ (defined in (S.4)) can be written as
$\mathcal{S}_{j}=$$\Big{\{}\big{(}\mathbb{P}_{j,\boldsymbol{\alpha}}(S_{1,j}),\ldots,\mathbb{P}_{j,\boldsymbol{\alpha}}(S_{\kappa_{j},j})\big{)}:{\boldsymbol{\alpha}}\in\{0,1\}^{K}\Big{\}}$ (S.7)
$=$ $\Big{\{}\big{(}\mathbf{N}_{a_{j}}((\kappa_{1},\ldots,\kappa_{j-1},{1},\kappa_{j+1},\ldots,\kappa_{J}),\boldsymbol{\alpha}),\cdots,\mathbf{N}_{a_{j}}((\kappa_{1},\ldots,\kappa_{j-1},{\kappa_{j}},\kappa_{j+1},\ldots,\kappa_{J}),\boldsymbol{\alpha})\big{)}:\boldsymbol{\alpha}\Big{\}}$
$=$ $\Big{\{}\big{(}\widetilde{\mathbf{N}}_{a_{j}}((\kappa_{1},\ldots,\kappa_{j-1},{1},\kappa_{j+1},\ldots,\kappa_{J}),\widetilde{\boldsymbol{\alpha}}),\cdots,\widetilde{\mathbf{N}}_{a_{j}}((\kappa_{1},\ldots,\kappa_{j-1},{\kappa_{j}},\kappa_{j+1},\ldots,\kappa_{J}),\widetilde{\boldsymbol{\alpha}})\big{)}:\widetilde{\boldsymbol{\alpha}}\Big{\}}$
[…]
and provides enough information about $\widetilde{\mathbf{g}}_{j}$. Here, the index $a_{j}=1,2,3$ can be understood from the context, for example $a_{j}=1$ when $j\leq K$, $a_{j}=2$ when $K<j\leq 2K$.
Step 2: Proving the equivalence by constructing a column permutation. Now, we show that there exists a permutation $\sigma\in S_{[K]}$ such that $\widetilde{g}_{j,k}=g_{j,\sigma(k)}$ for all $j\in[J],k\in[K]$. We first construct such a permutation $\sigma$ by observing $\mathcal{S}_{1},\ldots,\mathcal{S}_{K}$. For $k\in[K]$, (S.7) implies $|\mathcal{S}_{k}|=2$, and part (a) of Lemma S.2 constraints $\widetilde{\mathbf{g}}_{k}$ to be a standard basis vector. Hence, we can define $\sigma(k)$ such that $\widetilde{\mathbf{g}}_{k}=\mathbf{e}_{\sigma(k)}$. To see that $\sigma$ is indeed a permutation, we have to show that $\sigma(k)\neq\sigma(l)$ for $k\neq l$. But this is immediate by noting that the cardinality of the pmf vector of $\{\mathbb{P}(Y_{k}\in S_{1,k},Y_{l}\in S_{1,l}\mid\mathbf{A}=\boldsymbol{\alpha}):\boldsymbol{\alpha}\}$ is 2 if and only if $k=l$. For future purposes, let us partition the $2^{K}$ row indices $\widetilde{\boldsymbol{\alpha}}$ in (S.8) into two groups $T_{k}$ and $T_{k}^{c}$ based on the value of (~$\big{(}\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{1,j}),\ldots,\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{\kappa_{j},j})\big{)}$. Then, the columns of $\widetilde{\mathbf{N}}$ that correspond to the $\widetilde{\boldsymbol{\alpha}}\in T_{k}$ must be identical to the columns of ${\mathbf{N}}$ indexed by $\boldsymbol{\alpha}$s such that $\alpha_{\sigma(k)}=0$ (or 1).
Now, for each $j>K$ and $k\in[K]$, we show that $\widetilde{g}_{j,k}=g_{j,\sigma(k)}$ and complete the proof. By part (b) of Lemma S.2, we have $\widetilde{g}_{j,k}$ = 0 if and only if
{(~$\big{\{}\big{(}\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{1,j}),\ldots,\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{\kappa_{j},j})\big{)}:\widetilde{\boldsymbol{\alpha}}\in T_{k}\big{\}}=\big{\{}\big{(}\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{1,j}),\ldots,\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{\kappa_{j},j})\big{)}:\widetilde{\boldsymbol{\alpha}}\not\in T_{k}\big{\}}.$ (S.9)
For notational simplicity, let $\boldsymbol{\alpha}_{\sigma}:=(\alpha_{\sigma(1)},\ldots,\alpha_{\sigma(k)})$. Then, by the construction of $T_{k}$, (S.9) simplifies to
$\Big{\{}\big{(}\mathbb{P}_{j,\boldsymbol{\alpha}_{\sigma}}(S_{1,j}),\ldots,\mathbb{P}_{j,\boldsymbol{\alpha}_{\sigma}}(S_{\kappa_{j},j})\big{)}:\alpha_{\sigma(k)}=1\Big{\}}=\Big{\{}\big{(}\mathbb{P}_{j,\boldsymbol{\alpha}_{\sigma}}(S_{1,j}),\ldots,\mathbb{P}_{j,\boldsymbol{\alpha}_{\sigma}}(S_{\kappa_{j},j})\big{)}:\alpha_{\sigma(k)}=0\Big{\}}.$
Another application of part (b) of Lemma S.2 shows that this is equivalent to $g_{j,\sigma(k)}=0$. Hence, $\widetilde{g}_{j,k}=g_{j,\sigma(k)}$ and the proof is complete. ∎
Below, we provide a short proof of Lemma S.2.
Proof of Lemma S.2.
  1. (a)
    The proof is immediate from the faithfulness of the graphical matrix $\mathbf{G}$. For example, the third bullet point follows from noting that $|H_{j}|\geq 2$ implies |$|\{\sum_{k\in H_{j}}\beta_{j,k}\alpha_{k}:\boldsymbol{\alpha}\}|\geq 3$, and that |$|\{\sum_{k\in H_{j}}\beta_{j,k}\alpha_{k}:\boldsymbol{\alpha}\}|\leq 2$ implies $|H_{j}|\leq 1$.
  2. (b)
    The “if” part immediately follows by conditional independence. For the “only if” part, by considering the contrapositive statement, it suffices to show that $g_{j,k}=1$ implies
    $\{\mathbf{s}_{j}(\boldsymbol{\alpha}):{\alpha_{k}=1}\}\neq\{\mathbf{s}_{j}(\boldsymbol{\alpha}):{\alpha_{k}=0}\}.$
    By writing out the parametrization and using the identifiability of the parametric family in (4), it suffices to show that $\{\sum_{l\neq k}\beta_{j,l}\alpha_{l}+\beta_{j,k}\}\neq\{\sum_{l\neq k}\beta_{j,l}\alpha_{l}\}$. But this is immediate since $\beta_{j,k}\neq 0$.
S.1.3.3 Proof of Proposition 2
For the sake of notational simplicity, for a given coefficient matrix $\mathbf{B}$, define
$\eta_{j,\boldsymbol{\alpha}}:=\beta_{j,0}+\sum_{k=1}^{K}\beta_{j,k}\alpha_{k}$ (S.10)
for each $j$ and $\boldsymbol{\alpha}\in\{0,1\}^{K}$. For an alternative coefficient matrix $\widetilde{\mathbf{B}}$, we similarly define
$\widetilde{\eta}_{j,\widetilde{\boldsymbol{\alpha}}}:=\widetilde{\beta}_{j,0}+\sum_{k=1}^{K}\widetilde{\beta}_{j,k}\widetilde{\alpha}_{k}$
for $\widetilde{\boldsymbol{\alpha}}\in\{0,1\}^{K}$.
Before presenting the proof of Proposition 2, we state two lemmas whose proof is postponed to the end of the subsection. Our first lemma is a relaxation of Lemma S.1, and guarantees an almost sure unique tensor decomposition of $\mathbf{P}_{0}$. Recall the definition of $\mathbf{P}_{0}$ from (S.5).
Lemma S.3 (Modification of Theorem 2 in Lee and Gu, (2024)).
Consider the one-latent-layer saturated model with a true graphical matrix $\mathbf{G}^{(1)\star}$ that satisfies condition C, and analytic parametric families $p(\cdot;\eta,\gamma)$ and link functions $g_{j}$. Then, the rank $2^{K}$ tensor decomposition $\mathbf{P}_{0}=[\mathbf{M}_{1},\mathbf{M}_{2},\mathbf{M}_{3}]$ is unique up to column permutations for $\Omega_{K}(\boldsymbol{\Theta}^{(1)};\mathbf{G}^{(1)\star})\setminus\mathcal{N}_{1}$. Here, $\mathbf{M}_{a}$s are $\upsilon_{a}\times 2^{K}$ matrices, whose last rows are the all-one vector $\mathbf{1}_{2^{K}}^{\top}$ for $a=2,3$. The set $\mathcal{N}_{1}$ is a measure-zero subset of $\Omega_{K}(\boldsymbol{\Theta}^{(1)};\mathbf{G}^{(1)\star})$ that only imposes restrictions on $\mathbf{B},\boldsymbol{\gamma}$ and not on the mixture proportion parameters $\boldsymbol{\pi}$.
Our next lemma provides additional structure regarding the $2^{K}$ column indices in the tensor decomposition in Lemma S.3. This is a nontrivial problem as we wish to identify the parameters up to a label switching of $K$ binary latent variables. Here, to formally state the label switching, recall the notation of $\sim_{K}$ Definition S.2 and write $\mathbf{B}\sim_{K}\widetilde{\mathbf{B}},\mathbf{G}\sim_{K}\widetilde{\mathbf{G}}$, $\boldsymbol{\alpha}\sim_{K}\widetilde{\boldsymbol{\alpha}}$ when there exists a permutation $\sigma\in S_{[K]}$ such that $\beta_{j,l}=\widetilde{\beta}_{j,\sigma(l)}$, $g_{j,l}=\widetilde{g}_{j,\sigma(l)}$, and $\boldsymbol{\alpha}=(\widetilde{\alpha}_{\sigma(1)},\ldots,\widetilde{\alpha}_{\sigma(K)}).$
Lemma S.4.
Suppose that there exist two sets of parameters $(\mathbf{B},\mathbf{G})$, $(\widetilde{\mathbf{B}},\widetilde{\mathbf{G}})$ that satisfies Assumption 1 and defines an identical $\eta_{j,\boldsymbol{\alpha}}$ in the following sense: there exists a permutation $\mathfrak{S}\in S_{\{0,1\}^{K}}$ such that
$\eta_{j,\boldsymbol{\alpha}}=\widetilde{\eta}_{j,\mathfrak{S}(\boldsymbol{\alpha})},$ (S.11)
for all $j$, $\boldsymbol{\alpha}$. Then, we have $(\mathbf{B},\mathbf{G},\boldsymbol{\alpha})\sim_{K}(\widetilde{\mathbf{B}},\widetilde{\mathbf{G}},\mathfrak{S}(\boldsymbol{\alpha}))$ for “generic” parameters $\mathbf{B}\not\in\Omega(\mathbf{B};\mathbf{G})\setminus\mathcal{N}_{2}$, where $\mathcal{N}_{2}$ is a measure-zero subset of $\Omega(\mathbf{B};\mathbf{G})$.
Proof of Proposition 2.
We work under the same notations introduced at the beginning of the section. We make one additional assumption regarding the finite subsets $\mathcal{D}_{j}:=(S_{1,j},\ldots,S_{\kappa_{j},j})$ of the sample space $\mathcal{Y}_{j}$ as follows. We assume that $\mathcal{D}_{j}\subset\mathcal{C}_{j}$, where $\mathcal{C}_{j}$ is a countable separating class whose values determine probability measure on $\mathcal{Y}_{j}$. The existence of such a separating class is a consequence of $\mathcal{Y}_{j}$ being a separating metric space; see Step 1 in the proof of Theorem 1 in Lee and Gu, (2024) for a proof.
Suppose that $\mathbf{G}$ satisfy condition C, $\boldsymbol{\Theta}\in\Omega_{\mathcal{K}}(\boldsymbol{\Theta};\mathbf{G})$, and that there exists an alternate parameter $\widetilde{\boldsymbol{\Theta}},\widetilde{\mathbf{G}}$ that defines the same marginal likelihood. Here, we are dropping all superscripts in $\mathbf{G}=\mathbf{G}^{(1)\star}$ for simplicity. We show that $(\boldsymbol{\Theta},\mathbf{G})\sim_{\mathcal{K}}(\widetilde{\boldsymbol{\Theta}},\widetilde{\mathbf{G}})$ for $\boldsymbol{\Theta}\in\Omega_{\mathcal{K}}(\boldsymbol{\Theta};\mathbf{G})\setminus(\mathcal{N}_{1}\cup\mathcal{N}_{2})$, where $\mathcal{N}_{1},\mathcal{N}_{2}$ are measure-zero sets that will be defined later. Recall that for $a=1,2,3$, $\mathbf{N}_{a}$ are the $\upsilon_{a}\times 2^{K}$ conditional probability matrices that describe $\mathbf{Y}\mid(\mathbf{A}=\boldsymbol{\alpha})$ under the true parameters $\boldsymbol{\Theta},\mathbf{G}$. Similarly, define $\widetilde{\mathbf{N}}_{a}$ as the corresponding matrices under the alternative parameters $\widetilde{\boldsymbol{\Theta}},\widetilde{\mathbf{G}}$. Similar to the proof of Proposition 1, we start from the rank $2^{K}$ decomposition of the marginal probability tensor in (S.5). By applying Lemma S.3, the decomposition
$\mathbf{P}_{0}=[\mathbf{N}_{1}\text{{diag}}(\boldsymbol{\pi}),\mathbf{N}_{2},\mathbf{N}_{3}]=[\widetilde{\mathbf{N}}_{1}\text{{diag}}(\widetilde{\boldsymbol{\pi}}),\widetilde{\mathbf{N}}_{2},\widetilde{\mathbf{N}}_{3}]$ (S.12)
is unique up to (the $2^{K}$) column permutations, for true parameters $\boldsymbol{\Theta}\in\Omega_{K}(\boldsymbol{\Theta};\mathbf{G})\setminus\mathcal{N}_{1}$. Here, $\mathcal{N}_{1}$ is a measure-zero subset of $\Omega_{K}(\boldsymbol{\Theta};\mathbf{G})$ that is defined in Lemma S.3. This implies that $(\boldsymbol{\pi},\mathbf{N}_{1},\mathbf{N}_{2},\mathbf{N}_{3})$ and $(\widetilde{\boldsymbol{\pi}},\widetilde{\mathbf{N}}_{1},\widetilde{\mathbf{N}}_{2},\widetilde{\mathbf{N}}_{3})$ are identical up to a common column permutation $\mathfrak{S}\in S_{\{0,1\}^{K}}$. In other words, for any $\boldsymbol{\alpha}\in\{0,1\}^{K}$, the $\boldsymbol{\alpha}$th column in $\mathbf{N}_{a}$ is identical to the $\mathfrak{S}(\boldsymbol{\alpha})$th column of $\widetilde{\mathbf{N}}_{a}$ and $\pi_{\boldsymbol{\alpha}}=\widetilde{\pi}_{\mathfrak{S}(\boldsymbol{\alpha})}$.
By considering the row of $\mathbf{N}_{a}$ that corresponds to the set $S_{l_{j},j}$, we have
$\mathbb{P}_{j,\boldsymbol{\alpha}}(S_{l_{j},j})$ $=\mathbf{N}_{a}(\kappa_{\cdot},\ldots,\kappa_{j-1},l_{j},\kappa_{j+1},\ldots,\kappa_{\cdot},\boldsymbol{\alpha})$ (S.13)
$=\widetilde{\mathbf{N}}_{a}(\kappa_{\cdot},\ldots,\kappa_{j-1},l_{j},\kappa_{j+1},\ldots,\kappa_{\cdot},\mathfrak{S}(\boldsymbol{\alpha}))=\widetilde{\mathbb{P}}_{j,\mathfrak{S}(\boldsymbol{\alpha})}(S_{l_{j},j})$
for all $j,l_{j}\in[\kappa_{j}]$, and $\boldsymbol{\alpha}$. Furthermore, note that this identity holds for any finite subset $\mathcal{D}_{j}$ of $\mathcal{C}_{j}$, as the column indices of $\widetilde{\mathbf{N}}_{a}$ are determined by the alternative model and do not depend on the discretization $\mathcal{D}_{j}$. Hence, (S.13) holds for all $S_{l_{j},j}\in\mathcal{C}_{j}$. Since $\mathcal{C}_{j}$ is a separating class, we have $\mathbb{P}_{j,\boldsymbol{\alpha}}=\widetilde{\mathbb{P}}_{j,\mathfrak{S}(\boldsymbol{\alpha})}$. Recalling that $\mathbb{P}_{j,\boldsymbol{\alpha}}=\text{ParFam}_{j}(\eta_{j,\boldsymbol{\alpha}},\gamma_{j})$ for some identifiable parametric family (see (S.3)), we must have $\eta_{j,\boldsymbol{\alpha}}=\widetilde{\eta}_{j,\mathfrak{S}(\boldsymbol{\alpha})}$ and $\gamma_{j}=\widetilde{\gamma}_{j}$ for all $j,\boldsymbol{\alpha}$.
Now, additionally assuming that $\boldsymbol{\Theta}\not\in\mathcal{N}_{2}$, we can apply Lemma S.4. Here, $\mathcal{N}_{2}$ is the null set defined in Lemma S.4. Then, we have $(\mathbf{B},\mathbf{G})\sim_{K}(\widetilde{\mathbf{B}},\widetilde{\mathbf{G}})$. Also, because $\boldsymbol{\alpha}\sim_{K}\mathfrak{S}(\boldsymbol{\alpha})$, $\widetilde{\pi}_{\mathfrak{S}(\boldsymbol{\alpha})}=\pi_{\boldsymbol{\alpha}}$ implies $\widetilde{\boldsymbol{\pi}}\sim_{K}\boldsymbol{\pi}$ and the proof is complete. ∎
We finally present the postponed proof of Lemmas S.3 and S.4. While the main proof idea of Lemma S.3 is similar to that of Theorem 2 in Lee and Gu, (2024), we provide a detailed proof for the sake of completeness.
Proof of Lemma S.3.
For a matrix $\mathbf{N}$, let $rk_{k}(\mathbf{N})$ be the Kruskal column-rank of $\mathbf{N}$, that is, the largest integer $r$ such that any $r$ columns of $\mathbf{N}$ are linearly independent. We claim that it suffices to show that
$rk_{k}(\mathbf{N}_{1})=2^{K},~{}rk_{k}(\mathbf{N}_{2})=2^{K},~{}rk_{k}(\mathbf{N}_{3})\geq 2$
for generic parameters in $\Omega(\boldsymbol{\Theta};\mathbf{G})\setminus\mathcal{N}$. Assuming this, one can apply Kruskal’s Theorem (Kruskal,, 1977), which guarantees the uniqueness of the three-way tensor decomposition of $\mathbf{P}_{0}$ up to a universal column permutation and gives the desired result. Along the way, we show that the candidate set of non-identifiable parameters, $\mathcal{N}$, only imposes restrictions on $\mathbf{B}$ and $\boldsymbol{\gamma}$ but not on $\boldsymbol{\pi}$.
For the remainder of the proof, let ${\boldsymbol{\beta}}_{I_{1},I_{2}}$ denote the sub-matrix of the $J\times(K+1)$ coefficient matrix ${\boldsymbol{\beta}}$ whose rows and columns are indexed by $I_{1}$ and $I_{2}$, respectively. Similarly, $\boldsymbol{\gamma}_{I_{1}}$ denotes the sub-vector of $\boldsymbol{\gamma}$ by collecting the entries indexed by $I_{1}$.
Proof of $rk_{k}(\mathbf{N}_{1})=2^{K}$.
First, write the parameter space
$\Omega({\boldsymbol{\beta}}_{1:K,0:K},\boldsymbol{\gamma}_{1:K};\mathbf{G}_{1})=\{{\boldsymbol{\beta}}_{1:K,0:K},\boldsymbol{\gamma}_{1:K}:\beta_{j,k}\neq 0\text{ for }g_{j,k}=1\}$
as a finite union of open, connected subsets of Euclidean space $\mathbb{R}^{\sum_{j,k\leq K}g_{j,k}}\times{\mathbb{R}}^{K}$. Without loss of generality, let
$\Omega_{\text{positive}}({\boldsymbol{\beta}}_{1:K,0:K},\boldsymbol{\gamma}_{1:K};\mathbf{G}_{1}):=\{{\boldsymbol{\beta}}_{1:K,0:K},\boldsymbol{\gamma}_{1:K}:\beta_{j,k}>0\text{ for }g_{j,k}=1\}$
be our domain. Here, we consider $\mathbf{N}_{1}=\mathbf{N}_{1}(\boldsymbol{\beta}_{1:K,0:K},\boldsymbol{\gamma}_{1:K})$ to be a matrix-valued function of $(\boldsymbol{\beta}_{1:K,0:K},\boldsymbol{\gamma}_{1:K})$. Because $\mathbf{N}_{1}$ has full column rank if and only if $\det(\mathbf{N}_{1}^{\top}\mathbf{N}_{1})\neq 0$, it suffices to show that
$\{{\boldsymbol{\beta}}_{1:K,0:K}\in\Omega_{\text{positive}}({\boldsymbol{\beta}}_{1:K,0:K},\boldsymbol{\gamma}_{1:K};\mathbf{G}_{1}):\det(\mathbf{N}_{1}^{\top}\mathbf{N}_{1})=0\}$
is a measure-zero set in $\Omega({\boldsymbol{\beta}}_{1:K,0:K},\boldsymbol{\gamma}_{1:K};\mathbf{G}_{1})$. Note that the following argument also holds for other connected sub-domains of $\Omega({\boldsymbol{\beta}}_{1:K,0:K},\boldsymbol{\gamma}_{1:K};\mathbf{G}_{1})$, and the union of a finite number of measure-zero sets are still measure-zero.
In particular, when $\mathbf{G}_{1}=\mathbf{I}_{K}$, the proof of Theorem 1 in Lee and Gu, (2024) showed that $\mathbf{N}_{1}(\boldsymbol{\beta}_{1:K,0:K},\boldsymbol{\gamma}_{1:K})$ always have full column rank. We use technical real analysis arguments to claim that this statement can be generalized to an arbitrary $\mathbf{G}_{1}$ that satisfies $\text{{diag}}(\mathbf{G}_{1})=\mathbf{I}_{K}$. Consider the mapping
$(\det(\mathbf{N}_{1}^{\top}\mathbf{N}_{1}))(\boldsymbol{\beta}_{1:K,0:K},\boldsymbol{\gamma}_{1:K}):\Omega(\boldsymbol{\beta}_{1:K,0:K},\boldsymbol{\gamma}_{1:K};\mathbf{G}_{1})\rightarrow\mathbb{R}.$ (S.14)
Observe that $\det(\mathbf{N}_{1}^{\top}\mathbf{N}_{1})$ defined in (S.14) is a polynomial of entries of $\mathbf{N}_{1}$. Each entry in $\mathbf{N}_{1}$ can be written in the form of
$\mathbf{N}_{1}((l_{1},\ldots,l_{K}),\boldsymbol{\alpha})=\prod_{j=1}^{K}\int_{S_{l_{j},j}}p_{j}(y_{j};g_{j}(\beta_{j,0}+\sum_{k}\beta_{j,k}\alpha_{k}),\gamma_{j})m(dy_{j}).$
Since we assume that the density $p_{j}$ and link function $g_{j}$ are analytic, each entry of $\mathbf{N}_{1}$ is analytic. Consequently, $\det(\mathbf{N}_{1}^{\top}\mathbf{N}_{1})$ is also analytic.
Next, note that $\det(\mathbf{N}_{1}^{\top}\mathbf{N}_{1})$ cannot be identically zero, in other words, there exists $(\boldsymbol{\beta}^{\star}_{1:K,0:K},\boldsymbol{\gamma}^{\star}_{1:K})$ such that $\det(\mathbf{N}_{1}^{\top}\mathbf{N}_{1})(\boldsymbol{\beta}^{\star}_{1:K,0:K},\boldsymbol{\gamma}^{\star}_{1:K})\neq 0$. This is because one can make small perturbations from a parameter value in $\Omega(\boldsymbol{\beta}_{1:K,0:K},\boldsymbol{\gamma}_{1:K};\mathbf{I}_{K})$ by setting $\beta_{j,k}=\epsilon$ for indices with $g_{j,k}=1$ without making the determinant become zero. Hence, by the following technical Lemma, we conclude that
$\{\boldsymbol{\beta}_{1:K,0:K},\boldsymbol{\gamma}_{1:K}\in\Omega(\boldsymbol{\beta}_{1:K,0:K},\boldsymbol{\gamma}_{1:K};\mathbf{G}_{1}):(\det(\mathbf{N}_{1}^{\top}\mathbf{N}_{1}))(\boldsymbol{\beta}_{1:K,0:K},\boldsymbol{\gamma}_{1:K})=0\}$
is a null set in $\Omega(\boldsymbol{\beta}_{1:K,0:K},\boldsymbol{\gamma}_{1:K},\mathbf{G}_{1})$. The conclusion $rk_{k}(\mathbf{N}_{2})=2^{K}$ automatically follows.
Lemma S.5 (Mityagin, (2020)).
Let $f:\Omega\rightarrow\mathbb{R}$ be a real analytic function defined on an open, connected domain $\Omega\in\mathbb{R}^{d}$ that is not identically zero. Then, $\{\omega\in\Omega:f(\omega)=0\}$ is a measure-zero set in $\Omega$.
Proof of $rk_{k}(\mathbf{N}_{3})\geq 2$.
Theorem 1 in Lee and Gu, (2024) showed that $rk_{k}(\mathbf{N}_{3})\geq 2$ under condition B. Hence, it suffices to show that condition B holds for generic parameters in $\Omega({\boldsymbol{\beta}}_{2K+1:J,1:K};\mathbf{G}_{3})$. Fix any $\boldsymbol{\alpha}\neq\boldsymbol{\alpha}^{\prime}$, and let $l=l(\boldsymbol{\alpha},\boldsymbol{\alpha}^{\prime})$ be an index among $[K]$ such that $\alpha_{l}\neq\alpha_{l}^{\prime}$. Since we assume that all columns of $\mathbf{G}_{3}$ are nonzero, there exists $j=j(\boldsymbol{\alpha},\boldsymbol{\alpha}^{\prime})>2K$ such that $g_{j,l}=1$. As $\beta_{j,l}(\alpha_{l}-\alpha_{l}^{\prime})\neq 0$,
$\{{\boldsymbol{\beta}}_{j,1:K}\in\Omega({\boldsymbol{\beta}}_{j,1:K};\mathbf{g}_{j}):\sum_{k=1}^{K}\beta_{j,k}g_{j,k}(\alpha_{k}-\alpha_{k}^{\prime})=0\}$
is a measure-zero set in $\Omega({\boldsymbol{\beta}}_{j,1:K};\mathbf{g}_{j})$. Consequently,
$\Phi_{\boldsymbol{\alpha},\boldsymbol{\alpha}^{\prime}}:=\{{\boldsymbol{\beta}}_{2K+1:J,1:K}\in\Omega({\boldsymbol{\beta}}_{2K+1:J,1:K};\mathbf{G}_{3}):\sum_{k=1}^{K}\beta_{j,k}g_{j(\boldsymbol{\alpha},\boldsymbol{\alpha}^{\prime}),k}(\alpha_{k}-\alpha_{k}^{\prime})=0\}$
is a measure-zero set in $\Omega({\boldsymbol{\beta}}_{2K+1:J,1:K};\mathbf{G}_{3})$. The proof is complete by taking a union over all $\boldsymbol{\alpha}\neq\boldsymbol{\alpha}^{\prime}$. ∎
Proof of Lemma S.4.
We first assume that the true parameter $\mathbf{B}$ belongs in the coefficient space
$\Omega^{g}(\mathbf{B};\mathbf{G}):=\{\mathbf{B}:\text{for each $j$, all elements of the form }\sum_{k\in H_{j}}\beta_{j,k}a_{k},a_{k}=0,1,\text{ are distinct}\}.$
It is clear that the complement $\Omega(\mathbf{B};\mathbf{G})\setminus\Omega^{g}(\mathbf{B};\mathbf{G})$ can be spelled out as a finite union of lower dimensional subspaces, and has measure-zero with respect to $\Omega(\mathbf{B};\mathbf{G})$. The following proof is somewhat technical, but the main idea is to show that one can entirely recover $\mathbf{B},\mathbf{G}$, $\mathfrak{S}$ up to label switching, based on the values of $\eta_{j,\boldsymbol{\alpha}}$ while not using the indices $\boldsymbol{\alpha}$. Steps 1 and 2 characterizes the coefficients $\mathbf{B}$ based on the values $\{\eta_{j,\boldsymbol{\alpha}}\}_{\boldsymbol{\alpha}}$, which allows us to prove equivalence of $\mathbf{B},\mathbf{G}$ and $\widetilde{\mathbf{B}},\widetilde{\mathbf{G}}$ in step 3.
Step 1: recovering $|H_{j}|$ and $|\mathbf{B}|$.
We first claim that for each $j$, there uniquely exists an integer $I_{j}$, coefficients $c_{j,0}$, $c_{j,1}>\ldots>c_{j,I_{j}}>0$ that satisfy
$\eta_{j,\boldsymbol{\alpha}}=\beta_{j,0}+\sum_{l\in H_{j}}\beta_{j,l}\alpha_{l}=c_{j,0}+\sum_{k\leq I_{j}}c_{j,k}(\mathfrak{T}_{j}(\boldsymbol{\alpha}))_{k}$ (S.15)
for all $\boldsymbol{\alpha}$ and some permutation $\mathfrak{T}_{j}\in S_{\{0,1\}^{K}}$. Here, we also prove that that $I_{j}=|H_{j}|$.
The existence directly follows as we can take $I_{j}=|H_{j}|$ and define $\{c_{j,k}:k\leq I_{j}\}$ to be the coefficients $\{|\beta_{j,l}|:l\in H_{j}\}$ in decreasing order. Then, for each $k\leq I_{j}$, there must be an $l_{k}\in H_{j}$ such that $c_{j,k}=|\beta_{j,l_{k}}|>0$. Consequently, we can construct the permutation $\mathfrak{T}_{j}$ by setting
$(\mathfrak{T}_{j}(\boldsymbol{\alpha}))_{k}:=\begin{cases}\alpha_{l_{k}}&\text{if }\beta_{j,l_{k}}>0\\ 1-\alpha_{l_{k}}&\text{if }\beta_{j,l_{k}}<0,\end{cases}$
for $k\leq I_{j}$ and arbitrarily defining the remaining coordinates. By plugging in these definitions, we get
$\sum_{k\leq I_{j}}c_{j,k}(\mathfrak{T}_{j}(\boldsymbol{\alpha}))_{k}$ $=\sum_{l\in H_{j}}\text{sgn}(\beta_{j,l})|\beta_{j,l}|\alpha_{l}+\sum_{l\in H_{j},\beta_{j,l}<0}|\beta_{j,l}|$
$=\sum_{l\in H_{j}}\beta_{j,l}\alpha_{l}-\sum_{l\in H_{j},\beta_{j,l}<0}\beta_{j,l}.$
Hence, (S.15) holds for $c_{j,0}:=\beta_{j,0}+\sum_{l\in H_{j},\beta_{j,l}<0}\beta_{j,l}$.
For uniqueness, we provide an inductive characterization of the $c_{j,k}$s based on the values $U_{j}:=\{\eta_{j,\boldsymbol{\alpha}}:\boldsymbol{\alpha}\in\{0,1\}^{K}\}$. First, $c_{j,0}$ must be the minimum of the set $U_{j}$. Next, we define $c_{j,I_{j}}$ by noting that $c_{j,0}+c_{j,I_{j}}$ should be the smallest element in $\{\eta_{j,\boldsymbol{\alpha}}:\eta_{j,\boldsymbol{\alpha}}>c_{j,0}\}$. Inductively for each $k<I_{j}$, given $c_{j,0}$ and $c_{j,k+1},\ldots,c_{j,I_{j}}$, $c_{j,0}+c_{j,k}$ must be the smallest element in $U_{j}\setminus\{c_{j,0}+\sum_{l>k}c_{j,l}a_{l}:a_{l}=0,1\}$. We continue the induction until the set $U_{j}\setminus\{c_{j,0}+\sum_{l>k}c_{j,l}a_{l}:a_{l}=0,1\}$ is empty. By considering the true parameterization and the fact that $\mathbf{B}\in\Omega^{g}(\mathbf{B};\mathbf{G})$, we have $|U_{j}|=2^{|H_{j}|}$, and hence $|H_{j}|=I_{j}$. Additionally, this observation gives $c_{j,1}>\ldots>c_{j,|H_{j}|}>0$.
Step 2: Recovering the indexing $\boldsymbol{\alpha}$.
While we have recovered the absolute values of the coefficients $|\beta_{j,k}|$ in Step 1, they are ordered based on their absolute values. Thus, the ordering is different across different $j$. In this step, we fully recover $\beta_{j,k}$ up to a column permutation $\tau\in S_{[K]}$, by representing each $\beta_{j,k}$ in terms of $\eta$ and $c$-values. To this end, we inductively construct sets $T_{j,k},V_{j,k}\subseteq[2^{K}]$ for all $j\in[J]$ and $k\leq|H_{j}|$. We claim that they correspond to ${\boldsymbol{\alpha}}$s with identical ${\alpha}_{\sigma_{j}(k)}$ values, for some injective mapping $\sigma_{j}:[|H_{j}|]\to[K]$ (see (S.18) for a precise statement). This claim will be crucially utilized to recover the coefficients $\beta_{j,k}$.
First, fix $j\in[J]$, $k\leq|H_{j}|$, and define $T_{j,k}^{(1)}=\{\boldsymbol{\alpha}:\eta_{j,\boldsymbol{\alpha}}=c_{j,0}\}$ and $V_{j,k}^{(1)}=\{{\boldsymbol{\alpha}}:\eta_{j,{\boldsymbol{\alpha}}}=c_{j,0}+c_{j,k}\}$. Since there exists a unique index $l_{k}$ such that $|\beta_{j,l_{k}}|=c_{j,k}>0$, these two sets have distinct $\alpha_{l}$ values if and only if $l=l_{k}$. Define $\sigma_{j}(k)$ as this $l_{k}$. Since $l_{k}\neq l_{k^{\prime}}$ for $k\neq k^{\prime}$, $\sigma_{j}$ is injective. Without the loss of generality, suppose $\beta_{j,l_{k}}>0$. This is only for the sake of explicitly characterizing the sets $T_{j,k}$ and $V_{j,k}$, and does not affect their construction. Under this assumption, $\boldsymbol{\alpha}\in T_{j,k}^{(1)}$ must satisfy $\alpha_{\sigma_{j}(k)}=0$, $\boldsymbol{\alpha}\in V_{j,k}^{(1)}$ must satisfy $\alpha_{\sigma_{j}(k)}=1$. Also, as the values of $\alpha_{l}$ for $l\not\in H_{j}$ does not change the value of $\eta_{j,\boldsymbol{\alpha}}$, we have $|T_{j,k}^{(1)}|=|V_{j,k}^{(1)}|=2^{K-|H_{j}|}$.
Inductively for $t>1$, let
$\mathcal{T}_{j,k}^{[t]}$ :=$:=\{\boldsymbol{\alpha}\not\in T_{j,k}^{(t-1)},V_{j,k}^{(t-1)}:\eta_{j,\boldsymbol{\alpha}}=\min_{\boldsymbol{\alpha}^{\prime}\not\in T_{j,k}^{(t-1)},V_{j,k}^{(t-1)}}\eta_{j,\boldsymbol{\alpha}^{\prime}}\},$ (S.16)
$\mathcal{V}_{j,k}^{[t]}$ :=$:=\{\boldsymbol{\alpha}\not\in T_{j,k}^{(t-1)},V_{j,k}^{(t-1)}:\eta_{j,\boldsymbol{\alpha}}=\min_{\boldsymbol{\alpha}^{\prime}\not\in T_{j,k}^{(t-1)},V_{j,k}^{(t-1)}}\eta_{j,\boldsymbol{\alpha}^{\prime}}+c_{j,k}\},$ (S.17)
and define $T_{j,k}^{[t]}=T_{j,k}^{(t-1)}\cup\mathcal{T}_{j,k}^{[t]}$ and $V_{j,k}^{[t]}=V_{j,k}^{(t-1)}\cup\mathcal{V}_{j,k}^{[t]}$. By the construction in (S.16) and (S.17), again because $\beta_{j,\sigma_{j}(k)}>0$, we must have $\alpha_{\sigma_{j}(k)}=0$ for $\boldsymbol{\alpha}\in\mathcal{T}_{j,k}^{[t]}$, and $\alpha_{\sigma_{j}(k)}=1$ for $\boldsymbol{\alpha}\in\mathcal{V}_{j,k}^{[t]}$. We also have $|\mathcal{T}_{j,k}^{[t]}|=|\mathcal{V}_{j,k}^{[t]}|=2^{K-|H_{j}|}$, which inductively gives $|T_{j,k}^{[t]}|=|V_{j,k}^{[t]}|=2^{K-|H_{j}|}t$. We continue the construction until $T_{j,k}^{[t]}\cup V_{j,k}^{[t]}=\{0,1\}^{K}$, that is when $t=2^{|H_{j}|-1}$.
Finally, we define $T_{j,k}:=T_{j,k}^{\left(2^{|H_{j}|-1}\right)}$ and $V_{j,k}:=V_{j,k}^{\left(2^{|H_{j}|-1}\right)}$ as the final induction outputs. Then, $\{T_{j,k},V_{j,k}\}$ is a partition of $\{0,1\}^{K}$ with equal cardinality, where $T_{j,k}=\{\boldsymbol{\alpha}:\alpha_{\sigma_{j}(k)}=0\}$ and $V_{j,k}=\{\boldsymbol{\alpha}:\alpha_{\sigma_{j}(k)}=1\}$. In general, without the positivity assumption $\beta_{j,\sigma_{j}(k)}>0$, we can conclude that
$\{T_{j,k},V_{j,k}\}=\big{\{}\{\boldsymbol{\alpha}:\alpha_{\sigma_{j}(k)}=0\},\{\boldsymbol{\alpha}:\alpha_{\sigma_{j}(k)}=1\}\big{\}},$ (S.18)
as an unordered set. As all columns of $\mathbf{G}$ are not empty (see Assumption 1 (b)), for each $l$, we must have at least one $j$ such that $g_{j,l}=1$. Hence, the set of all possible partitions $\{\{T_{j,k},V_{j,k}\}:j\in[J],k\in[|H_{j}|]\}$ must take exactly $K$ distinct values. Let us index each element by $\{T_{l},V_{l}\}$ for ${l\in[K]}$, and let $\tau\in S_{[K]}$ be a permutation such that
$\{T_{l},V_{l}\}=\big{\{}\{\boldsymbol{\alpha}:\alpha_{\tau(l)}=0\},\{\boldsymbol{\alpha}:\alpha_{\tau(l)}=1\}\big{\}}.$ (S.19)
Without loss of generality, suppose that $T_{l}$ and $V_{l}$ are defined so that the mean of the vector $(\eta_{j,\boldsymbol{\alpha}}:\boldsymbol{\alpha}\in T_{l})$ is strictly smaller than the mean of $(\eta_{j,\boldsymbol{\alpha}}:\boldsymbol{\alpha}\in V_{l})$. Then, the monotonicity assumption Assumption 1(c) implies that
$T_{l}=\{\boldsymbol{\alpha}:\alpha_{\tau(l)}=0\},\quad V_{l}=\{\boldsymbol{\alpha}:\alpha_{\tau(l)}=1\},$ (S.20)
and resolves the sign-flipping for the latent variable $A_{\tau(l)}$. Next, to recover the parameter $\beta_{j,\tau(l)}$ for each $j$ and $l$, we claim that
  1. (a)
    if $\{T_{l},V_{l}\}\not\in\{\{T_{j,k},V_{j,k}\}:k\leq|H_{j}|\}$, $\beta_{j,\tau(l)}=0$.
  2. (b)
    if $\{T_{l},V_{l}\}=\{T_{j,k},V_{j,k}\}$ for some $k\leq|H_{j}|$, $|\beta_{j,\tau(l)}|=c_{j,k}$ and
    $\text{sgn}(\beta_{j,\tau(l)})=\begin{cases}1&\text{ if }T_{l}=T_{j,k},\\ -1&\text{ if }T_{l}=V_{j,k}.\end{cases}$
Here, to show (a), suppose $\{T_{l},V_{l}\}\not\in\{\{T_{j,k},V_{j,k}\}:k\leq|H_{j}|\}$. Then, by the characterization in (S.18) and (S.19), we must have $\sigma_{j}(k)\neq\tau(l)$ for all $k\leq|H_{j}|$. Recalling that $\sigma_{j}$ collects all indices such that $|\beta_{j,\sigma_{j}(k)}|>0$, we must have $\beta_{j,\tau(l)}=0$. To prove part (b), note that the assumption implies $\sigma_{j}(k)=\tau(l)$, so $|\beta_{j,\tau(l)}|=|\beta_{j,\sigma_{j}(k)}|=c_{j,k}$. The claim regarding the sign holds because $T_{j,k}$ has a smaller average of $\eta_{j,\cdot}$ values than $V_{j,k}$ by construction. Thus, $T_{l}=T_{j,k}$ if and only if $\beta_{j,\tau(l)}>0$.
Step 3: explicit representation of all parameters.
Finally, we prove the desired result by applying the same characterization of the parameters under the alternative values $\widetilde{\eta}_{j,\widetilde{\boldsymbol{\alpha}}}$. Here, note that $\mathbf{B}\in\Omega^{g}(\mathbf{B};\mathbf{G})$ implies that $\widetilde{\mathbf{B}}\in\Omega^{g}(\mathbf{B};\mathbf{G})$. We apply steps 1-2 using $\widetilde{\eta}_{j,\widetilde{\boldsymbol{\alpha}}}$ instead of $\eta_{j,\boldsymbol{\alpha}}$. Step 1 holds without any change, and we obtain the same coefficients $c_{j,k}$s as well as $|H_{j}|=|\widetilde{H}_{j}|$. Next, in step 2, the definition of $T_{j,k},V_{j,k}$ in (S.16), (S.17) needs to be modified to be the collection of $\widetilde{\boldsymbol{\alpha}}$s instead of $\boldsymbol{\alpha}$s. Let $\widetilde{T}_{j,k}$ and $\widetilde{V}_{j,k}$ be the corresponding sets under the alternative parametrization. As the coefficients $c_{j,k}$ are identical, $({T}_{j,k},V_{j,k})$ and $(\widetilde{T}_{j,k},\widetilde{V}_{j,k})$ are consistent in the sense that
$\widetilde{T}_{j,k}=\mathfrak{S}({T}_{j,k}),\quad\widetilde{V}_{j,k}=\mathfrak{S}({V}_{j,k}).$ (S.21)
In other words, by viewing $T_{j,k}$ as the collection of indices $\boldsymbol{\alpha}$ in $\eta_{j,\boldsymbol{\alpha}}$, $\widetilde{T}_{j,k}$ and $T_{j,k}$ are the same collection of indices. Thus, in terms of defining $\widetilde{T}_{l}$ and $\widetilde{V}_{l}$, we can take the same indexing for $l\in[K]$ as in (S.19) by letting $\widetilde{T}_{l}:=\mathfrak{S}(T_{l})$ and $\widetilde{V}_{l}:=\mathfrak{S}(V_{l})$. Hence, for some permutation $\widetilde{\tau}\in S_{[K]}$, we must have
$\widetilde{T}_{l}$ =$=\{\widetilde{\boldsymbol{\alpha}}:\widetilde{\alpha}_{\widetilde{\tau}(l)}=0\}=\mathfrak{S}(\{\boldsymbol{\alpha}:\alpha_{\tau(l)}=0\})=\mathfrak{S}(T_{l}),$ (S.22)
$\quad\widetilde{V}_{l}$ =$=\{\widetilde{\boldsymbol{\alpha}}:\widetilde{\alpha}_{\widetilde{\tau}(l)}=1\}=\mathfrak{S}(\{\boldsymbol{\alpha}:\alpha_{{\tau}(l)}=1\})=\mathfrak{S}(V_{l}).$ (S.23)
Now, we finish the proof by proving $\mathbf{B}\sim\widetilde{\mathbf{B}}$ and $\boldsymbol{\alpha}\sim\mathfrak{S}(\boldsymbol{\alpha})$. By taking $\sigma:=\widetilde{\tau}\cdot\tau^{-1}$ in the definition of the equivalence relations, it suffices to show $\beta_{j,\tau(l)}=\widetilde{\beta}_{j,\widetilde{\tau}(l)}$ for all $j,l$, and $\alpha_{\tau(l)}=(\mathfrak{S}(\boldsymbol{\alpha}))_{\widetilde{\tau}(l)}$ for all $\boldsymbol{\alpha},l$. The conclusion for $\beta_{j,\tau(l)}$ follows from its characterization in Step 2. For any $j,l$, the relationship (S.21) and (S.22) shows that $\{T_{l},V_{l}\}\in\{\{T_{j,k},V_{j,k}\}:k\leq|H_{j}|\}$ if and only if $\{\widetilde{T}_{l},\widetilde{V}_{l}\}\in\{\{\widetilde{T}_{j,k},\widetilde{V}_{j,k}\}:k\leq|H_{j}|\}$. For case (a), we have $\beta_{j,\tau(l)}=\widetilde{\beta}_{j,\widetilde{\tau}(l)}=0$. For case (b), we have $|\beta_{j,\tau(l)}|=c_{j,k}=|\widetilde{\beta}_{j,\widetilde{\tau}(l)}|$ as well as $\text{sgn}(\beta_{j,\tau(l)})=\text{sgn}(\widetilde{\beta}_{j,\widetilde{\tau}(l)})$, thus $\beta_{j,\tau(l)}=\widetilde{\beta}_{j,\widetilde{\tau}(l)}$. We next show the claim for $\alpha$. Without loss of generality, suppose $\alpha_{\tau(l)}=0$. Then, by applying (S.20) and (S.22), we have $\mathfrak{S}(\boldsymbol{\alpha})\in\widetilde{T}_{l}$ and $(\mathfrak{S}(\boldsymbol{\alpha}))_{\widetilde{\tau}(l)}=0$. Similarly, for $\alpha_{\tau(l)}=1$, we get $(\mathfrak{S}(\boldsymbol{\alpha}))_{\widetilde{\tau}(l)}=1$ and the proof is complete. ∎
S.1.4 Proof of Theorem 3
Proof.
For simplicity, we omit displaying the dependence of data in the likelihood $\ell(\boldsymbol{\Theta})=\ell(\boldsymbol{\Theta}\mid\mathbf{Y})$ and omit the subscript in the equivalence relation $\sim_{\mathcal{K}}$. We also define the objective function in (6) as
$Q_{N}(\boldsymbol{\Theta}):=\frac{1}{N}\Big{[}-\ell(\boldsymbol{\Theta})+\sum_{d=1}^{D}p_{\lambda_{N},\tau_{N}}(\mathbf{B}^{(d)})\Big{]}.$
Below, we separately prove the consistency of the continuous parameters $\boldsymbol{\Theta}$ and discrete parameters $\mathcal{G}$. In the proof, we use the usual Bachmann-Landau asymptotic notations for deterministic sequences as well as its analog for random sequences.
We first show that $\widehat{\boldsymbol{\Theta}}$ is consistent up to label permutations. This follows by modifying the usual consistency proof of M-estimators (Van der Vaart,, 2000) under our identifiability notion. Fix $\epsilon>0$, and define the pointwise limit of $Q_{N}(\boldsymbol{\Theta})$ as
$Q_{\infty}(\boldsymbol{\Theta}):=-\mathbb{E}_{\boldsymbol{\Theta}^{\star}}\log\mathbb{P}(\mathbf{Y}\mid\boldsymbol{\Theta}).$
Define the $\epsilon$-ball around $\boldsymbol{\Theta}^{\star}$ under the equivalence relation $\sim$ as
$B_{\sim}(\epsilon,\boldsymbol{\Theta}^{\star}):=\{\boldsymbol{\Theta}:\lVert\boldsymbol{\Theta}^{\prime}-\boldsymbol{\Theta}^{\star}\rVert\leq\epsilon\text{ for some }\boldsymbol{\Theta}^{\prime}\sim\boldsymbol{\Theta}\}.$
Also, define the constant $\eta:=\inf_{\boldsymbol{\Theta}\not\in B_{\sim}(\epsilon,\boldsymbol{\Theta}^{\star})}Q_{\infty}(\boldsymbol{\Theta})-Q_{\infty}(\boldsymbol{\Theta}^{\star})$. Since we assume the model identifiability up to $\sim$, we must have $Q_{\infty}(\boldsymbol{\Theta})-Q_{\infty}(\boldsymbol{\Theta}^{\star})=\text{KL}\big{(}\mathbb{P}(\cdot\mid\boldsymbol{\Theta})\|\mathbb{P}(\cdot\mid\boldsymbol{\Theta}^{\star})\big{)}>0$ for all $\boldsymbol{\Theta}\not\sim\boldsymbol{\Theta}^{\star}$. Here, KL denotes the Kullback–Leibler divergence. Then, we get $\eta>0$, as we are considering a compact parameter space. By a standard argument, we have
$\mathbb{P}(\min_{\boldsymbol{\Theta}^{\prime}\sim\widehat{\boldsymbol{\Theta}}}$ $\lVert\boldsymbol{\Theta}^{\prime}-\boldsymbol{\Theta}^{\star}\rVert>\epsilon)\leq\mathbb{P}\big{(}\min_{\boldsymbol{\Theta}\not\in B_{\sim}(\epsilon,\boldsymbol{\Theta}^{\star})}Q_{N}(\boldsymbol{\Theta})\leq Q_{N}(\boldsymbol{\Theta}^{\star})\big{)}$ (S.24)
$\leq\mathbb{P}\big{(}\min_{\boldsymbol{\Theta}\not\in B_{\sim}(\epsilon,\boldsymbol{\Theta}^{\star})}Q_{N}(\boldsymbol{\Theta})\leq Q_{N}(\boldsymbol{\Theta}^{\star}),~{}\sup_{\boldsymbol{\Theta}}\big{|}Q_{N}(\boldsymbol{\Theta})-Q_{\infty}(\boldsymbol{\Theta})\big{|}<\frac{\eta}{2}\big{)}+o(1).$ (S.25)
Here, (S.24) uses the definition that $\widehat{\boldsymbol{\Theta}}$ is a minimizer of $Q_{N}(\boldsymbol{\Theta})$. The inequality in (S.25) follows from the uniform law of large numbers, which holds under a compact parameter space and a vanishing penalty with $\lambda_{N}=o(N)$. Now, the proof is complete by noting that the first term in (S.25) is zero, since $\boldsymbol{\Theta}\not\in B_{\sim}(\epsilon,\boldsymbol{\Theta}^{\star})$ and $\sup_{\boldsymbol{\Theta}}|Q_{N}(\boldsymbol{\Theta})-Q_{\infty}(\boldsymbol{\Theta})|<\frac{\eta}{2}$ implies
$Q_{N}(\boldsymbol{\Theta})>Q_{\infty}(\boldsymbol{\Theta})-\frac{\eta}{2}\geq Q_{\infty}(\boldsymbol{\Theta}^{\star})+\frac{\eta}{2}>Q_{N}(\boldsymbol{\Theta}^{\star}).$
Hence, there exists some $\widetilde{\boldsymbol{\Theta}}\sim\widehat{\boldsymbol{\Theta}}$ that is consistent for $\boldsymbol{\Theta}^{\star}$.
Now, we prove that $\widetilde{\boldsymbol{\Theta}}$ is $\sqrt{N}$-consistent, additionally using the assumption on the Fisher information. We first re-write the inequality $Q_{N}(\widetilde{\boldsymbol{\Theta}})\leq Q_{N}(\boldsymbol{\Theta}^{\star})$ as
$-\ell(\widetilde{\boldsymbol{\Theta}})+\ell(\boldsymbol{\Theta}^{\star})\leq p_{\lambda_{N},\tau_{N}}(\mathbf{B}_{0})-p_{\lambda_{N},\tau_{N}}(\widetilde{\mathbf{B}}).$ (S.26)
By a Taylor expansion, we can bound the LHS of (S.26) as
$-\ell(\widetilde{\boldsymbol{\Theta}})+\ell(\boldsymbol{\Theta}^{\star})$ $=-(\widetilde{\boldsymbol{\Theta}}-\boldsymbol{\Theta}^{\star})^{\top}\ell^{\prime}(\boldsymbol{\Theta}^{\star})+\frac{1}{2}(\widetilde{\boldsymbol{\Theta}}-\boldsymbol{\Theta}^{\star})^{\top}(NI(\boldsymbol{\Theta}^{\star})+o_{p}(N))(\widetilde{\boldsymbol{\Theta}}-\boldsymbol{\Theta}^{\star})$
$\geq\|\widetilde{\boldsymbol{\Theta}}-\boldsymbol{\Theta}^{\star}\|_{2}O_{p}(\sqrt{N})+N\|\widetilde{\boldsymbol{\Theta}}-\boldsymbol{\Theta}^{\star}\|_{2}^{2}\left(\frac{\lambda_{\min}(I(\boldsymbol{\Theta}^{\star}))}{2}+o_{p}(1)\right).$
Here, $\lambda_{\min}(I(\boldsymbol{\Theta}^{\star}))>0$ denotes the smallest eigenvalue of the positive definite Fisher information $I(\boldsymbol{\Theta}^{\star})$. On the other hand, the RHS of (S.26) must be negative since $\tau_{N}\rightarrow 0$. Indeed, for $\beta_{l,k}^{(d)\star}\neq 0$ and a large enough $N$, consistency gives $|\widetilde{\beta}_{l,k}^{(d)}|=|\beta_{l,k}^{(d)\star}|+o_{p}(1)>\tau_{N}$, and the bound
$p_{\lambda_{N},\tau_{N}}(\mathbf{B}_{0}^{(d)})-p_{\lambda_{N},\tau_{N}}(\widetilde{\mathbf{B}}^{(d)})$ $\leq\sum_{l,k:\beta_{l,k}^{(d)\star}\neq 0}\left[p_{\lambda_{N},\tau_{N}}(\beta_{l,k}^{(d)\star})-p_{\lambda_{N},\tau_{N}}(\widetilde{\beta}_{l,k}^{(d)})\right]=0$
holds with high probability. Hence, we have
$\|\widetilde{\boldsymbol{\Theta}}-\boldsymbol{\Theta}^{\star}\|_{2}O_{p}(\sqrt{N})+\frac{\lambda_{\min}(I(\boldsymbol{\Theta}^{\star}))}{2}N\|\widetilde{\boldsymbol{\Theta}}-\boldsymbol{\Theta}^{\star}\|_{2}^{2}\leq 0$
with high probability, which is impossible when $\|\widetilde{\boldsymbol{\Theta}}-\boldsymbol{\Theta}^{\star}\|_{2}\gg\frac{1}{\sqrt{N}}$. Thus, $\|\widetilde{\boldsymbol{\Theta}}-\boldsymbol{\Theta}^{\star}\|_{2}=O_{p}\left(\frac{1}{\sqrt{N}}\right)$.
[…]
$\partial_{\beta_{l,k}^{(d)}}\ell(\widetilde{\boldsymbol{\Theta}}):=\frac{\partial\ell(\boldsymbol{\Theta})}{\partial\beta_{l,k}^{(d)}}\mid_{\boldsymbol{\Theta}=\widetilde{\boldsymbol{\Theta}}}=p^{\prime}_{\lambda_{N},\tau_{N}}(\widetilde{\beta}_{l,k}^{(d)})=\Theta_{p}\left(\frac{\lambda_{N}}{\tau_{N}}\right).$
But we have a contradiction because a Taylor expansion of the partial derivative gives
$\partial_{\beta_{l,k}^{(d)}}\ell(\widetilde{\boldsymbol{\Theta}})=\partial_{\beta_{l,k}^{(d)}}\ell(\boldsymbol{\Theta}^{\star})+NO_{p}(\widetilde{\boldsymbol{\Theta}}-\boldsymbol{\Theta}^{\star})=O_{p}\left(\sqrt{N}\right),$
and $\sqrt{N}\ll{\lambda_{N}}/{\tau_{N}}.$ Hence, we must have $\widetilde{g}_{l,k}^{(d)}=0$, and the proof is complete. ∎
S.1.5 Detour: Identifiability of One-latent-layer Saturated Models With Interaction Effects of Binary Latent Variables
One may ask whether the linear/additive parametrization $\beta_{j,0}+\sum_{k\in[K]\beta_{j,k}}A_{k}$ in the one-latent-layer saturated model is necessary for its identifiability. We next show that this is not the case, and prove that identifiability of $\mathbf{G}$ can be established under two more flexible nonlinear (in terms of the dependence on $\mathbf{A}$) parametric models commonly used in psychometrics, using the exact same conditions A and B. Here, to handle different parametrizations, define $\eta_{j,\boldsymbol{\alpha}}$ to be the nonlinear parameter for $Y_{j}\mid\mathbf{A}$:
$Y_{j}\mid(\mathbf{A}=\boldsymbol{\alpha})\sim\text{ParFam}_{j}(\eta_{j,\boldsymbol{\alpha}}),$
and rewrite condition B as follows:
  1. B’.
    For any $\boldsymbol{\alpha}\neq\boldsymbol{\alpha}^{\prime}$, there exists $j>2K$ such that $\eta_{j,\boldsymbol{\alpha}}\neq\eta_{j,\boldsymbol{\alpha}^{\prime}}$.
We first consider the ExpDINA model (see Definition 4 in Lee and Gu,, 2024) [^1] While this model is originally named “Exponential-family based”, this is not required for our identifiability conclusion. In other words, the family Parfamj in (S.27) can be any paramteric family., which considers the following conjunctive form of $\eta_{j,\boldsymbol{\alpha}}$:
$Y_{j}\mid(\mathbf{A}=\boldsymbol{\alpha})\sim\textnormal{ParFam}_{j}\Big{(}g_{j}\big{(}\beta_{j,0}+\beta_{j,1}\prod_{k=1}^{K}\alpha_{k}^{q_{j,k}},\gamma_{j}\big{)}\Big{)}.$ (S.27)
In other words, the conditional distribution has two possible parameter values based on whether $\prod_{k=1}^{K}\alpha_{k}^{q_{j,k}}=1$ (or equivalently, whether $\mathbf{A}\succeq\mathbf{q}_{j}$). To resolve the sign flipping ambiguity, let as assume $\beta_{j,1}>0$ for all $j$ instead of Assumption 1(c). Under this model, both parts of Lemma S.2 does not hold since $|\mathcal{S}_{j}|=2$ whenever $|H_{j}|\geq 1$. Hence, we focus on partitioning $\{0,1\}^{K}$ by grouping the binary patterns $\boldsymbol{\alpha}$ that takes the same values of (S.7), instead of just looking at the cardinality of $\mathcal{S}_{j}$. We formally state this claim in the following Lemma. Consequently, this Lemma can be used in place of Lemma S.2 to prove identifiability of the exploratory ExpDINA model. The proof is a direct modification of Step 2 above (first, use part (a) of Lemma S.6 to construct a permutation $\sigma\in S_{[K]}$ based on the first $K$ rows, and use part (b) to prove $\widetilde{g}_{j,k}=g_{j,\sigma(k)}$ for the other rows), and we omit the details.
Lemma S.6.
Consider an ExpDINA model. Fix any $j$ such that $|H_{j}|\geq 1$, and partition $\{0,1\}^{K}$ into two sets $T_{j}$ and $T_{j}^{c}$ based on the value of $\mathcal{S}_{j}=\{\mathbf{s}_{j}(\boldsymbol{\alpha})\}_{\boldsymbol{\alpha}}$, so that $|T_{j}|\leq|T_{j}^{c}|$. If $|T_{j}|=|T_{j}^{c}|$, we break the symmetry by additionally assuming that $\mathbf{1}_{K}\in T_{j}$. Then, the following holds.
  1. (a)
    $T_{j}=\{\boldsymbol{\alpha}:\boldsymbol{\alpha}\succeq\mathbf{g}_{j}\}$ and $|T_{j}|=2^{K-|H_{j}|}.$
  2. (b)
    Suppose that the first $K$ rows of $\mathbf{G}$ is a row-permutation of $\mathbf{I}_{K}$, in other words, there exists $\sigma\in S_{[K]}$ such that $\mathbf{g}_{k}=\mathbf{e}_{\sigma(k)}$ for all $k$. Then, for any $j$, $T_{j}\subseteq T_{k}$ if and only if $g_{j,\sigma(k)}=1$.
Proof.
  1. (a)
    Note that the parameter of the ExpDINA model takes two values, depending on the value of $\prod_{k}\alpha_{k}^{g_{j,k}}=\mathbf{1}(\boldsymbol{\alpha}\succeq\mathbf{g}_{j})$. This value is equal to one for the $2^{K-|H_{j}|}$ configurations of $\boldsymbol{\alpha}$ such that $\boldsymbol{\alpha}\succeq\mathbf{g}_{j}$, and these are exactly the elements of $T_{j}$.
  2. (b)
    The “if” part is immediate since $g_{j,\sigma(k)}=1$ implies $\mathbf{g}_{j}\succeq\mathbf{g}_{k}$.
    For the “only if” part, we prove the contrapositive. Suppose $g_{j,\sigma(k)}=0$. Then, $\boldsymbol{\alpha}=\mathbf{1}_{K}$ and $\boldsymbol{\alpha}^{\prime}:=(\alpha_{1},\ldots,\alpha_{\sigma(k)-1},0,\alpha_{\sigma(k)+1},\ldots,\alpha_{K})$ have the same conditional distribution as in (S.27). Hence, $\boldsymbol{\alpha}$ and $\boldsymbol{\alpha}^{\prime}$ must both belong in $T_{j}$ but $\boldsymbol{\alpha}^{\prime}\not\in T_{k}$, so $T_{j}\not\subseteq T_{k}$.
Next, as a second example, we consider the flexible ExpGDM (see Definition 1 in the Supplementary Material of Lee and Gu,, 2024) and prove its identifiability under the same conditions A and B’. This model considers all possible linear and interaction effects between the latent variables:
$Y_{j}\mid\mathbf{A}\sim$ $\textnormal{ParFam}\Big{(}\eta_{j,\boldsymbol{\alpha}},~{}\gamma_{j}\Big{)},$
$\text{where}\quad\eta_{j,\boldsymbol{\alpha}}=$ $\beta_{j,\varnothing}+{\sum}_{k=1}^{K}\beta_{j,k}\left\{q_{j,k}\alpha_{k}\right\}+{\sum}_{1\leq k_{1}<k_{2}\leq K}\beta_{j,k_{1}k_{2}}\left\{q_{j,k_{1}}\alpha_{k_{1}}\right\}\left\{q_{j,k_{2}}\alpha_{k_{2}}\right\}$
$\qquad+\cdots+\beta_{j,H_{j}}{\prod}_{k\in H_{j}}\left\{q_{j,k}\alpha_{k}\right\}.$
The above model incorporates all possible main and interaction effects of the parent latent variables in the conditional distribution of $Y_{j}\mid\mathbf{A}$. It is clear that this model generalizes both the one-latent-layer saturated model in Definition S.1 and the ExpDINA model. To address the sign-flipping issue, we assume that
$\eta_{j,\boldsymbol{\alpha}}>\eta_{j,\boldsymbol{\alpha}^{\prime}}\quad\text{for}\quad\boldsymbol{\alpha}\succeq\mathbf{g}_{j},~{}\boldsymbol{\alpha}^{\prime}\not\succeq\mathbf{g}_{j}.$ (S.28)
This is a stronger assumption compared to Assumption 1(c). We impose this modified monotonicity condition as Assumption 1(c) cannot resolve the sign-flipping ambiguity, as we illustrate this in the following example.
Example S.1.
Consider a toy setting of $J=K=2$ and $\mathbf{G}=\mathbf{I}_{2}$ with $\beta_{1,1}=\beta_{2,2}=1$. Suppose that there all intercepts and interaction effects are zero for each $j=1,2$: $\beta_{j,0}=\beta_{j,12}=0$. Consider a stronger monotonicity assumption that all main-effects are nonnegative, that is $\beta_{j,k}\geq 0$. Define an alternative model with $\widetilde{\mathbf{G}}=\begin{pmatrix}1&0\\ 1&1\end{pmatrix}$ and positive main-effects: $\widetilde{\beta}_{1,1}=1,\widetilde{\beta}_{2,1}=\widetilde{\beta}_{2,2}=1$ and no intercepts, but with a negative interaction effect $\widetilde{\beta}_{2,12}=-2$. Then, the matrix $\{\eta_{j,\boldsymbol{\alpha}}\}_{j,\boldsymbol{\alpha}}$ and $\{\widetilde{\eta}_{j,\boldsymbol{\alpha}}\}_{j,\boldsymbol{\alpha}}$ are identical up to column permutation, so these two distinct parameters are non-distinguishable.
Also assuming a monotone condition on the parametric family (see below), we can modify Lemma S.6 as follows. Consequently, under all these assumptions, the ExpGDM is also identifiable under conditions A and B’.
Definition S.3 (Monotone family).
We say that a parametric family $p(\cdot\mid\eta)$ is a monotone family with a monotone set $U$ when there exists a measurable set $U\subset\mathcal{Y}$ not depending on $\eta$ such that $\mathbb{P}(Y\in U\mid\eta)$ (or $\mathbb{P}(Y\in U\mid\eta,\gamma)$) is a strictly increasing function in $\eta$.
Note that all parametric families considered in this paper are monotone families. For example, we can take $U=\{1\},\{1,2,\ldots\},(0,\infty)$ for the Bernoulli/Poisson/Normal distribution with mean $\eta$, respectively.
Lemma S.7.
Consider an ExpGDM. Suppose all $p_{j}$s are monotone families with monotone sets $U_{j}$, and the monotonicity condition (S.28) holds. Let $T_{j}:=\{\boldsymbol{\alpha}:\mathbb{P}_{j,\boldsymbol{\alpha}}(U_{j})=\max_{\boldsymbol{\alpha}^{\prime}}\mathbb{P}_{j,\boldsymbol{\alpha}^{\prime}}(U_{j})\}$. Then, the same conclusions as in Lemma S.6 hold.
Proof.
  1. (a)
    Since we consider a monotone family, we can write $T_{j}=\{\boldsymbol{\alpha}:\eta_{j,\boldsymbol{\alpha}}=\max_{\boldsymbol{\alpha}^{\prime}}\eta_{j,\boldsymbol{\alpha}^{\prime}}\}$. Then, by assumption (S.28), $\boldsymbol{\alpha}\not\in T_{j}$ for $\boldsymbol{\alpha}\not\succeq\mathbf{g}_{j}$. Also, by the faithfulness assumption, we have $\boldsymbol{\alpha}\in T_{j}$ for all $\boldsymbol{\alpha}\succeq\mathbf{g}_{j}$.
  2. (b)
    Assuming part (a), the argument in Lemma S.6 can be applied.
S.2 Details of the Layerwise Double-SVD Initialization
S.2.1 Details of Algorithm 2
We describe full details of the layerwise double-SVD initialization in Algorithm 2. We first consider the noiseless scenario to motivate the general procedure. The setting is that we are given a $N\times J$ matrix $\mathbb{E}[\mathbf{Y}]$, and wish to recover $\mathbf{A},\mathbf{B}$. For simplicity, we consider the one-latent-layer saturated model, and omit the layer-wise superscript to simplify the notation. We also assume identical parametric families in the observed layer and a nonlinear function $\mu\circ g$.
The first step involves denoising the nonlinearity and rewriting the data matrix as a low-rank approximation. Recall that $\mu:H\to\mathbb{R}$ computes the mean of the observed-layer parametric family, $g:\mathbb{R}\to H$ is the link function, and define a function $\widetilde{g}:=\mu\circ g$. Since
$\mathbb{E}(Y_{i,j}\mid\mathbf{A}_{i})=\mu(g(\beta_{j,0}+\sum_{k}\beta_{j,k}A_{i,k},\gamma_{j}))=\widetilde{g}(\beta_{j,0}+\sum_{k}\beta_{j,k}A_{i,k}),$
we have
$\mathbf{Z}:=\widetilde{g}^{-1}(\mathbb{E}(\mathbf{Y}\mid\mathbf{A}))=[\mathbf{1}_{N},~{}\mathbf{A}]\mathbf{B}_{1}^{\top}.$
Let $\mathbf{Z}_{0}$ be the centered version of $\mathbf{Z}$ so that the column sums are zero, in other words, ${z}_{0}(i,j):={z}(i,j)-\frac{1}{N}\sum_{i^{\prime}=1}^{N}{z}(i^{\prime},j)$. Similarly, let $\mathbf{A}_{0}$ be the column-centered version of $\mathbf{A}$ with $A_{0}(i,k):=A_{(}i,k)-\frac{1}{N}\sum_{i^{\prime}=1}^{N}A(i^{\prime},l)$. Then, we have
$\mathbf{Z}_{0}=\mathbf{A}_{0}\mathbf{B}^{\top}.$ (S.29)
This is a rank $K$ decomposition, and we can write the SVD of $\mathbf{Z}_{0}$ as $\mathbf{Z}_{0}=\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{\top}$. Here, $\mathbf{U},\boldsymbol{\Sigma},\mathbf{V}$ are $N\times K$, $K\times K$, $J\times K$ matrices, respectively.
The second step addresses the rotation invariance of the SVD by finding the sparse representation in (S.29). Motivated by the sparsity of $\mathbf{B}$, we perform the Varimax rotation on $\mathbf{V}$. As Varimax finds a sparse rotation (see Rohe and Zeng,, 2023, for a theoretical justification), we expect it to find a rotation matrix $\mathbf{R}$ such that $\widehat{\mathbf{B}}:=\mathbf{V}\mathbf{R}$ has the same sparsity pattern as $\mathbf{B}$ (and the graphical matrix $\mathbf{G}$). Consequently, we can use $\widehat{\mathbf{B}}$ as a rough estimate for $\mathbf{B}$, and the sparsity pattern of $\widehat{\mathbf{B}}$ to estimate $\mathbf{G}$. The estimate for $\mathbf{A}$ follows from solving (S.29) for $\mathbf{A}_{0}$, using the estimated $\widehat{\mathbf{B}}$.
There are several subtleties to address when extending this procedure to the sample-based setting. One immediate issue is applying the inverse-link function $\widetilde{g}^{-1}$. For discrete samples, observed data may take values in the boundary of the sample space, such as $Y_{j}=0$ when $\mathcal{Y}=\mathbb{N}\cup\{0\}$. This makes the inverse function not well-defined. To resolve this, we apply the double SVD-based procedure in Zhang et al., (2020) as mentioned in the main text, which denoises and truncates the data into a subset of the sample space. The final output of this procedure is the sample-version SVD for the matrix $\mathbf{Z}_{0}$ (see step 6 in Algorithm 4).
Another subtlety arises while estimating $\mathbf{A}$ and $\mathbf{B}$ after rotating $\mathbf{V}$, since Varimax does not account for the scaling of the row/columns. Here, we exploit the discreteness of the latent variables in $\mathbf{A}$ to rescale $\mathbf{B}$. Using the decomposition (S.29), we can crudely estimate $\mathbf{A}_{0}$ using the sparse Varimax output $\widehat{\mathbf{B}}$. While the estimates $\widehat{\mathbf{A}}_{0}$ also suffer from the same scaling issue, we can still estimate the binary $\mathbf{A}$ by noting that $A_{0}(i,k)<0$ if and only if $A(i,k)=0$. Finally, using the estimated $\widehat{\mathbf{A}}$, we re-estimate $\mathbf{B}$ via (S.29), now with correct scaling; see steps 9-10 in Algorithm 4.
The entire procedure is summarized in Algorithm 4, where we also specify the choice of tuning parameters.
Data: $\mathbf{Y},\{K^{(d)}\}_{d},D$, function $\widetilde{g}$, truncation parameters $\epsilon=10^{-4},\delta=\frac{1}{2.5\sqrt{J}}$
  1. 1.
    Apply SVD to $\mathbf{Y}$ and write $\mathbf{Y}=\mathbf{U}\Sigma\mathbf{V}^{\top}$, where $\Sigma=\text{{diag}}(\sigma_{i})$ and $\sigma_{1}\geq\ldots\geq\sigma_{J}$.
  2. 2.
    Let ${\mathbf{Y}}_{\widetilde{K^{(1)}}}=\sum_{k=1}^{\widetilde{K^{(1)}}}\sigma_{k}\mathbf{u}_{k}\mathbf{v}_{k}^{\top}$, where $\widetilde{K^{(1)}}:=\max\{K^{(1)}+1,\arg\max_{k}\{\sigma_{k}\geq 1.01\sqrt{N}\}\}$.
  3. 3.
    Define $\widehat{\mathbf{Y}}_{\widetilde{K^{(1)}}}$ by truncating $\mathbf{Y}_{\widetilde{K^{(1)}}}$ to the range of responses, at level $\epsilon$. See Remark S.1 for details.
  4. 4.
    Define $\widehat{\mathbf{Z}}$ by letting $\widehat{z}(i,j)=\widetilde{g}^{-1}(\widehat{y}_{\widetilde{K^{(1)}}}(i,j))$.
  5. 5.
    Let $\widehat{\mathbf{Z}}_{0}$ be the centered version of $\widehat{\mathbf{Z}}$, that is, $\widehat{z}_{0}(i,j)=\widehat{z}(i,j)-\frac{1}{N}\sum_{k=1}^{N}\widehat{z}(k,j)$.
  6. 6.
    Apply SVD to $\widehat{\mathbf{Z}}_{0}$ and write its rank-$K^{(1)}$ approximation as $\widehat{\mathbf{Z}}_{0}\approx\widehat{\mathbf{U}}\widehat{\Sigma}\widehat{\mathbf{V}}$.
  7. 7.
    Let $\widetilde{\mathbf{V}}$ be the rotated version of $\widehat{\mathbf{V}}$ according to the Varimax criteria.
  8. 8.
    Entrywise threshold $\widetilde{\mathbf{V}}$ at $\delta$ to induce sparsity, and flip the sign of each column so that all columns have positive mean. Let $\widehat{\mathbf{G}^{1}}$ be the estimated sparsity pattern.
  9. 9.
    Estimate the centered $\mathbf{A}_{0}$ by $\widehat{\mathbf{A}}_{0}:=\widehat{\mathbf{Z}}_{0}\widetilde{\mathbf{V}}(\widetilde{\mathbf{V}}^{\top}\widetilde{\mathbf{V}})^{-1}$, and estimate $\mathbf{A}$ by reading off the signs: $\widehat{A}(i,k)=\mathbbm{1}(A_{0}(i,k)>0).$
  10. 10.
    Let $\widehat{\mathbf{A}}_{\text{long}}:=[\mathbf{1},\widehat{\mathbf{A}}]$. Estimate $\mathbf{B}_{1}$ by $\widehat{\mathbf{B}}_{1}:=C_{g}((\widehat{\mathbf{A}}^{\top}_{\text{long}}\widehat{\mathbf{A}}_{\text{long}})^{-1}\widehat{\mathbf{A}}_{\text{long}}^{\top}\widehat{\mathbf{Z}}_{0})\cdot\widehat{\mathbf{G}}^{1}$, where $\cdot$ is the element-wise product and $C_{g}$ is a positive constant that is defined in Remark S.2
  11. 11.
    Let $\mathbf{Y}=\widehat{\mathbf{A}}$ and $g$ be the logistic function. Go back to Step 1 to estimate the next layer coefficient matrix. Continue until reaching the deepest layer.
  12. 12.
    For the deepest layer, estimate ${\boldsymbol{p}}$ by setting $\widehat{{\boldsymbol{p}}}_{k}:=\frac{1}{N}\sum_{i=1}^{N}\widehat{A}(i,k)$.
Output : $\widehat{{\boldsymbol{p}}},\{\widehat{\mathbf{B}}^{(d)}\}_{d}$.
Algorithm 4 Spectral initialization for Algorithm 1
Remark S.1 (Truncating $\widehat{\mathbf{Y}}$).
We explain more on the truncation details in Step 3 by considering specific response types. For Normal responses, the original sample space is $\mathbb{R}$ and the truncation (Steps 1-4 in Algorithm 4) may be omitted. For Binary responses, we set
$\widehat{y}_{K^{(1)}}(i,j)=\begin{cases}\epsilon,&\text{ if }y_{K^{(1)}}(i,j)=0,\\ 1-\epsilon,&\text{ if }y_{K^{(1)}}(i,j)=1.\end{cases}$
For Poisson responses, we set
$\widehat{y}_{K^{(1)}}(i,j)=\begin{cases}\epsilon,&\text{ if }y_{K^{(1)}}(i,j)<\epsilon,\\ y_{K^{(1)}}(i,j),&\text{ otherwise.}\end{cases}$
In terms of implementing the method, we follow the suggestions of Zhang et al., (2020) with $\epsilon=10^{-4}$.
Remark S.2 (Constant $C_{g}$).
In Step 10 of Algorithm 4, $C_{g}>0$ is an artificial scaling constant that depends on the link function $\widetilde{g}$. This is introduced to better adjust the scaling of $\mathbf{B}$, as the nonlinear transform $\widetilde{g}^{-1}$ leads to a potentially biased estimate for $\widehat{\mathbf{Z}}_{0}$. This adjustment is unnecessary for Normal-based DDEs, where $g$ is the identity link, and in such cases, one can simply set $C_{g}=1$. For the Bernoulli or Poisson-based DDE where $\widetilde{g}$ is the logistic or exponential function, we choose $C_{g}=\frac{1}{2}$ as the scaling factor based on simulation results.
S.2.2 Estimation Accuracy Without the Spectral Initialization
We illustrate the effectiveness of our spectral initialization for by comparing it with EM parameter estimates obtained under random initialization. We show that even for the low dimensional setting $(J,K^{(1)},K^{(2)})=(18,6,2)$, the overall model complexity may be too large for an EM algorithm with random initialization to converge to the global optimum. In contrast, the spectral initialization provides a reliable starting point. In Table S.1, we compare the accuracy of the PEM estimates under (a) random initialization and (b) spectral initialization. Here, we consider the same three parametric families as the main paper, the identifiable true parameter values $\mathcal{B}_{s}$ (see eq. (13)), and two sample sizes $N=1000,4000$. We set the random initialization as follows:
$p_{k},B^{(1)}_{j,k},B^{(2)}_{k,l}$ $\sim\text{Unif}(0,1),\quad\mbox{for all }k\in[K^{(1)}],~{}j\in[J],~{}l\in[K^{(2)}],$
$B^{(1)}_{j,0},B^{(2)}_{k,0}$ $\sim\text{Unif}(-1,0),\quad\gamma_{j}\sim\text{Unif}(0.5,1.5),$
where Unif$(a,b)$ denotes the uniform distribution on the interval $(a,b)$.
ParFam Accuracy($\mathcal{G}$) RMSE$(\boldsymbol{\Theta})$ Time (s) # iterations
Initialization \$N$ 1000 4000 1000 4000 1000 4000 1000 4000
Bernoulli Random 0.617 0.547 1.37 1.32 23.6 48.4 20.1 27.11
Spectral 0.966 0.992 0.30 0.20 6.7 37.5 4.1 4.2
Poisson Random 0.743 0.725 1.47 1.49 20.4 26.8 21.9 23.2
Spectral 0.999 1 0.16 0.08 3.6 6.4 4.4 4.0
Normal Random 0.595 0.581 1.71 1.83 36.6 347.7 14.7 16.9
Spectral 0.996 1 0.13 0.06 1.2 3.0 4.1 4.4
Table S.1: Accuracy measures for $\mathcal{G}$ and $\boldsymbol{\Theta}$, computation time and iterations for 2-layer DDE estimates under different initializations. For the Accuracy$(\mathcal{G})$ column, larger is better. For the other columns, smaller is better.
The results in Table S.1 clearly illustrates that random initialization does not effectively converge to the true parameter values. In contrast, the spectral initialization results in a significantly smaller estimator error as well as shorter time and smaller numbers of iterations, demonstrating its superior performance.
It may be worth mentioning that a significant proportion of local optimizers arise from boundary cases of the identifiability condition. For example, in the Normal case, the local optimizers emptied out one or more columns in the $\mathbf{B}^{(1)}$ and $\mathbf{B}^{(2)}$ matrices, or exhibit two linearly dependent columns. This corresponds to the set of parameters excluded by Assumption 1.
S.3 Additional Algorithm Details
S.3.1 Detailed Update Formulas
Simplified M-step for the PEM Algorithm.
We display more explicit expressions for the M-step in Algorithm 1. Recall that we need to solve three optimizations detailed in equations (8)-(10). By plugging-in the distributional formulas from Section 2, we can simplify each equation as follows. The computations are straightforward and we just display the simplified formulas.
First, the maximization in (8) simplifies to updating
$p_{l}^{[t+1]}=\frac{\sum_{i,\operatorname{\boldsymbol{\alpha}^{(1)}},\operatorname{\boldsymbol{\alpha}^{(2)}}2}\alpha_{l}^{(2)}\varphi^{[t+1]}_{i,\operatorname{\boldsymbol{\alpha}^{(1)}},\operatorname{\boldsymbol{\alpha}^{(2)}}2}}{N},$
for $l\in[K^{(2)}]$. The maximization in (9) simplifies to maximizing the $k$th row of $\mathbf{B}^{(2)}$ for each $k\in[K^{(1)}]$:
${\boldsymbol{\beta}}_{k}^{(2),[t+1]}$
$=$ $\underset{{\boldsymbol{\beta}}^{(2)}\in\mathbb{R}^{K^{(2)}+1}}{\operatorname*{argmax}}\sum_{i,\operatorname{\boldsymbol{\alpha}^{(1)}},\operatorname{\boldsymbol{\alpha}^{(2)}}2}\log\mathbb{P}(\mathbf{A}_{i,k}=\alpha_{k}^{(1)};g_{\text{logistic}}(\beta_{0}^{(2)}+\sum_{l=1}^{K^{(2)}}\alpha_{l}^{(2)}\beta_{l}^{(2)}))\varphi_{i,\operatorname{\boldsymbol{\alpha}^{(1)}},\operatorname{\boldsymbol{\alpha}^{(2)}}2}^{[t+1]}-p_{\lambda_{2},\tau}({\boldsymbol{\beta}}^{(2)}).$
Finally, the maximization in (10) boils down to maximizing the $j$th row of $\mathbf{B}^{(1)}$ and $\gamma_{j}$ for each $j\in[J]$:
$({\boldsymbol{\beta}}_{j}^{(1),[t+1]},\gamma_{j}^{[t+1]})$
$=$ $\underset{{\boldsymbol{\beta}}^{(1)}\in\mathbb{R}^{K^{(1)}+1},\gamma>0}{\operatorname*{argmax}}~{}\sum_{i,\operatorname{\boldsymbol{\alpha}^{(1)}},\operatorname{\boldsymbol{\alpha}^{(2)}}2}\log\mathbb{P}(Y_{i,j};\beta_{0}^{(1)}+\sum_{k=1}^{K^{(1)}}\alpha_{k}^{(1)}\beta_{k}^{(1)},\gamma)\varphi_{i,\operatorname{\boldsymbol{\alpha}^{(1)}},\operatorname{\boldsymbol{\alpha}^{(2)}}2}^{[t+1]}-p_{\lambda_{1,N},\tau}({\boldsymbol{\beta}}^{(1)}),$
Simplified Simulation Step for the SAEM Algorithm.
We display the complete conditionals for the simulation step in the $(t+1)$th iteration of the SAEM (see Algorithm 3). First, we sample each $A_{i,l}^{(2),[t+1]}$ from the following distribution:
$\mathbb{P}(A_{i,l}^{(2)}=\alpha_{l}^{(2)}\mid(-))$ $\propto p_{l}^{[t]\alpha_{l}^{(2)}}(1-p_{l}^{[t]})^{1-\alpha_{l}^{(2)}}\prod_{k=1}^{K^{(1)}}g_{\text{logistic}}(A_{i,k}^{(1),[t]};e^{\eta_{k,\mathbf{A}_{i}^{(2),[t]}}})$
$\propto p_{l}^{[t]\alpha_{l}^{(2)}}(1-p_{l}^{[t]})^{1-\alpha_{l}^{(2)}}\prod_{k=1}^{K^{(1)}}\frac{e^{A_{i,k}^{(1),[t]}\beta^{(2),[t]}_{k,l}\alpha_{l}^{(2)}}}{1+e^{\eta_{k,\mathbf{A}_{i}^{(2),[t]}}}}.$
Here, (-) denotes the samples/parameter values computed in the previous ($t$th) iteration, excluding the random variable of interest, $A_{i,l}^{(2),[t]}$. The notation
$\eta_{k,\mathbf{A}_{i}^{(2),[t]}}:=\beta_{k,0}^{(2),[t]}+\sum_{l^{\prime}\neq l}\beta_{k,l^{\prime}}^{(2),[t]}A_{i,l^{\prime}}^{(2),[t]}+\beta_{k,l}^{(2),[t]}\alpha_{l}^{(2)}$
denotes the linear combinations computed under $\boldsymbol{\Theta}^{[t]},\mathbf{A}^{(2),[t]}$. As $\alpha_{l}^{(2)}=0/1$, sampling from this distribution is straightforward by computing the above expression.
Next, we sample each $A_{i,k}^{(1),[t+1]}$ similarly from the complete conditionals:
$\mathbb{P}(A_{i,k}^{(1)}=\alpha_{k}^{(1)}\mid(-))$ $\propto e^{\alpha_{k}^{(1)}\eta_{k,\mathbf{A}_{i}^{(2),[t]}}}\prod_{j=1}^{J}\mathbb{P}\left(Y_{i,j};g(\eta_{j,\mathbf{A}_{i}^{(1),[t]}},\gamma_{j}^{[t]})\right).$
Here, $\eta_{j,\mathbf{A}_{i}^{(1),[t]}}$ is similarly defined as $\beta_{j,0}^{(1),[t]}+\sum_{k^{\prime}\neq k}\beta_{j,k^{\prime}}^{(1),[t]}A_{i,k^{\prime}}^{(1),[t]}+\beta_{j,k}^{(1),[t]}\alpha_{k}^{(1)}$.
Simplified M-step for the SAEM Algorithm.
In the main paper, we have motivated the SAEM M-step in terms of the parameter $\mathbf{B}^{(2)}$ (see (12)). Here, for completeness, we present the fully expanded formulas for updating each parameter.
First, for each $l\in[K^{(2)}]$, we update $p_{l}$ in closed form as follows:
$p_{l}^{[t+1]}:=\frac{\sum_{i}A_{i,l}^{(2),[t+1]}}{N}.$
Next, for $k\in[K^{(1)}]$, we update the $k$th row of $\mathbf{B}^{(2)}$ (denoted as ${\boldsymbol{\beta}}_{k}^{(2)}$) as follows:
$Q_{k}^{(2),[t+1]}({\boldsymbol{\beta}}_{k}^{(2)})$ $:=(1-\theta_{t+1})Q_{k}^{(2),[t]}({\boldsymbol{\beta}}_{k}^{(2)})+\theta_{t+1}\sum_{i=1}^{N}\log\mathbb{P}(A_{i,k}^{(1)}=A_{i,k}^{(1),[t+1]}\mid\mathbf{A}_{i}^{(2)}=\mathbf{A}_{i}^{(2),[t+1]};{\boldsymbol{\beta}}_{k}^{(2)}),$
${\boldsymbol{\beta}}_{k}^{(2),[t+1]}$ $:=\operatorname*{argmax}_{{\boldsymbol{\beta}}_{k}^{(2)}}\left[Q_{k}^{(2),[t+1]}({\boldsymbol{\beta}}_{k}^{(2)})-p_{\lambda_{N},\tau_{N}}({\boldsymbol{\beta}}_{k}^{(2)})\right],$
Finally, for each $j\in[J]$, we update the $j$th row of $\mathbf{B}^{(1)}$ (denoted as ${\boldsymbol{\beta}}_{j}^{(1)}$) and $\gamma_{j}$ (if it exists) as follows:
$Q_{j}^{(1),[t+1]}({\boldsymbol{\beta}}_{j}^{(1)},\gamma_{j})$ $:=(1-\theta_{t+1})Q_{j}^{(1),[t]}({\boldsymbol{\beta}}_{j}^{(1)},\gamma_{j})+\theta_{t+1}\sum_{i=1}^{N}\log\mathbb{P}\left(Y_{i,j}\mid\mathbf{A}_{i}^{(1)}=\mathbf{A}_{i}^{(1),[t+1]};{\boldsymbol{\beta}}_{j}^{(1)},\boldsymbol{\gamma}_{j}\right),$
$({\boldsymbol{\beta}}_{j}^{(1),[t+1]},\gamma_{j}^{[t+1]})$ $:=\operatorname*{argmax}_{{\boldsymbol{\beta}}_{j}^{(1)},\gamma_{j}}\left[Q_{j}^{(1),[t+1]}({\boldsymbol{\beta}}_{j}^{(1)},\gamma_{j})-p_{\lambda_{N},\tau_{N}}({\boldsymbol{\beta}}_{j}^{(1)})\right].$ (S.30)
S.3.2 Estimating the Number of Latent Variables
Recall that we have proposed a spectral-ratio estimator to select the latent dimension $\mathcal{K}$ in Section 4.3. In this supplement, we propose two alternative estimators motivated by popular methods for selecting the number of latent variables in other statistical problems (Shen and Wong,, 1994; Melnykov and Melnykov,, 2012; Chen and Li,, 2022). The performance of these will be later assessed in a simulation study in Section S.4.3.
Continuing from the setup in Section 4.3, we address the estimation of $\mathcal{K}$ under the two-latent-layer DDE. We first focus on the scenario where the number of deepest latent variables, $K^{(2)}$, is known, and our goal is to select $K^{(1)}$ from a candidate grid $\mathfrak{K}$. This assumption is often justified in real-world applications where prior knowledge of the number of latent labels or classes in the dataset is available. The objective is to uncover finer-grained latent structures that capture additional generative information beyond the known labels.
  • EBIC: We compute the extended BIC (EBIC; Chen and Chen,, 2008) for each $k\in\mathfrak{K}$, and select the model with the smallest EBIC:
    $K^{(1)}:=\operatorname*{argmin}_{k\in\mathfrak{K}}\text{EBIC}(k).$
    We postpone the details of the formula of EBIC$(k)$ to Section S.3.4. Here, we consider EBIC instead of more traditional information criteria such as AIC and BIC since it is designed to handle large parameter spaces.
  • LRT: The class of models with $k\in\mathfrak{K}$ can be viewed as a nested set of regular parametric models, and one can apply the method of sieves to select $K^{(1)}$ (Shen and Wong,, 1994). In other words, we start from the smallest $k$ and sequentially conduct level-$\alpha$ likelihood ratio tests for $H_{0}:K^{(1)}=k$ v.s. $H_{1}:K^{(1)}=k+1$ using the $\chi^{2}$ limiting distributions, until when we cannot reject the alternative; define $\widehat{K}^{(1)}$ as this $k$.
Below is a discussion regarding the theoretical and computational properties of the three proposed estimators (spectral-ratio, EBIC, LRT). Theoretically, consistency of all three methods can be justified under varying assumptions. Additionally assuming that $K^{(1)}$ is identifiable, the EBIC is known to be consistent even under a diverging $J$ (Chen and Chen,, 2008). The LRT estimator can be justified by standard arguments invoking Wilk’s theorem (Wilks,, 1938). Finally, under the asymptotic regime $N,J\to\infty$ and assuming that $\mu\circ g$ is linear, $\sigma_{1},\ldots,\sigma_{r}$ is close to the top $r$ singular values of $\mathbf{Y}$ and standard eigenvalue perturbation arguments (Weyl’s theorem) guarantee that the spectral ratio estimator for $K^{(1)}$ is consistent.
In terms of computation, the spectral-ratio estimator is the most desirable as it just requires computing the SVD of the denoised data matrix once. The other two methods (EBIC, LRT) require re-fitting the model (using Algorithm 3) for each candidate value of $K^{(1)}$ as well as computing the likelihood. This limits their usage when the upper bound for $K^{(1)}$ is large. Furthermore, the model re-fitting issue for EBIC and LRT is amplified when the number of the top layer latent variables $K^{(2)}$ is also unknown, as the cardinality of the candidate set for $\mathcal{K}=(K^{(1)},K^{(2)})$ increases. On the other hand, the spectral-ratio estimates the number of latent variables in a layerwise manner, and is computationally efficient even under an unknown $K^{(2)}$.
For analyzing real data, we recommend making the final selection of $\mathcal{K}$ by also incorporating qualitative aspects of the data such as domain knowledge and interpretability.
S.3.3 Implementation Details
We elaborate on the choices made to implement the EM Algorithms 1 and 3.
Tuning parameter selection.
For practical implementation, the penalty function and the tuning parameters $\lambda_{N},\tau_{N}$ must be specified. In this work, we consider the truncated Lasso penalty function (TLP) proposed by Shen et al., (2012), that is $p_{\lambda,\tau}(b):=\lambda\min(|b|,\tau)$. Preliminary simulations revealed that the results are very similar under other penalties such as SCAD.
Based on the consistency result in Theorem 3, we consider $\lambda_{N}=N^{1/4},\tau_{N}=3N^{-0.3}$ for the simulation studies. For real data analysis, tuning parameters were selected from the following grid:
$\lambda_{1,N},\lambda_{2,N}\in\{N^{1/8},N^{2/8},N^{3/8}\},\quad\tau_{N}\in 2\{N^{-1/8},N^{-2/8},N^{-3/8}\}.$
Here, $\lambda_{a,N}$ ($a=1,2$) denotes the penalty size for the $a$th latent layer coefficient $\mathbf{B}^{(a)}$. Note that distinct $\lambda_{N}$-values are used for each layer, which is to better accommodate the larger uncertainty in the deeper layer.
While there are many ways to select tuning parameters such as exact or approximate cross-validation and information criteria-based methods, we choose to use the latter in order to encourage sparsity. We select the model with the smallest extended Bayesian information criterion (EBIC). This approach is preferred over cross-validation in unsupervised settings, where likelihood-based loss functions often yield non-sparse, overfitted models (Chetverikov et al.,, 2021).
Additionally, to implement the SAEM, the step size $\theta_{t}$ that determines the weight of the stochastic averaging needs to be specified. Here, following the standard Robbins-Monro condition (Robbins and Monro,, 1951), we set $\theta_{t}=1/t$.
Convergence criteria.
For the penalized EM algorithm, we set the convergence criteria such that we terminate the algorithm when the difference between the log-likelihood function values of two consecutive iterations is less than $N/500$, as the log-likelihood is proportional to $N$. For the SAEM, since we do not directly compute the log-likelihood, convergence is declared when the difference of the vectorized $L^{2}$ norm of the continuous parameters is smaller than $K^{(2)}/2$. Under the spectral initialization, the PEM and SAEM generally took less than 10 and 5 iterations until convergence, respectively.
In terms of implementing the PEM algorithm to select the number of latent variables, $K^{(1)}$ or $K^{(2)}$, we consider a more generous threshold of $5003\times N/$ 500. This is for the sake of faster implementation, as we only use the resulting likelihood to implement the EBIC and LRT method.
M-step implementation.
As our M-step in both algorithms (PEM and SAEM) boils down to solving multiple lower-dimensional maximization of dimensions less or equal to $K^{(1)}+1$, we choose to simply apply a built-in optimizing package that directly computes the global maximizers. This is because there are already other approximations being made in the SAEM, and we did not want to increase the randomness for our simulation studies.
For practical implementation for a larger dataset, we recommend the user to modify the M-step to a faster but approximate optimization procedure. For example, one may choose to do a one-step gradient ascent in each iteration of the M-step.
Additional latent variable permutation for simulations.
In our simulations, resolving the latent variable permutation is necessary to accurately compute the errors. A naive approach involves computing the error across all possible permutations, but this quickly becomes computationally infeasible, even for moderate values of $K^{(1)}=18,K^{(2)}=6$. To address this, we formulate the optimal latent variable permutation as an assignment problem and solve it efficiently using the following bottom-up approach as follows.
First, we construct a $K^{(1)}\times K^{(1)}$ cost matrix, where each entry is the squared $L^{2}$ norm between the corresponding column vectors of $\mathbf{B}^{(1)}$ and $\widehat{\mathbf{B}}^{(1)}$. Next, use the Hungarian algorithm (Kuhn,, 1955) to find a column permutation that minimizes the total assignment cost. As the column indices of $\mathbf{B}^{(1)}$ correspond to the row indices of $\mathbf{B}^{(2)}$, we permute the rows of $\mathbf{B}^{(2)}$ accordingly. Finally, we apply the same procedure to find an optimal permutation of the $\mathbf{B}^{(2)}$ columns. This method ensures computational efficiency while accurately resolving the latent variable permutation.
S.3.4 Definition of EBIC
Consider a 2-latent layer DDE with $\mathcal{K}=(K^{(1)},K^{(2)})$ latent variables. This is a parametric model with $|\boldsymbol{\Theta}|=J(K^{(1)}+1)+K^{(1)}(K^{(2)}+1)+K^{(2)}+J$ (if there exists a dispersion parameter) or $J(K^{(1)}+1)+K^{(1)}(K^{(2)}+1)+K^{(2)}$ (otherwise) parameters. Here, we do not count the discrete parameters $\mathcal{G}$, as they are implied by the coefficients $\mathcal{B}$. Let $\widehat{\boldsymbol{\Theta}}_{\mathcal{K}}$ and $\ell_{\mathcal{K}}(\widehat{\boldsymbol{\Theta}}_{\mathcal{K}};\mathbf{Y})$ respectively denote the MLE and the log-likelihood under the 2-layer DDE with $\mathcal{K}$ parameters. Also, let df$(\mathcal{K})$ be the number of non-zero parameters in $\widehat{\boldsymbol{\Theta}}_{\mathcal{K}}$. Then, following Chen and Chen, (2008), the EBIC objective in Section S.3.2 is defined as follows:
$EBIC(\mathcal{K}):=-2\ell_{\mathcal{K}}(\widehat{\boldsymbol{\Theta}}_{\mathcal{K}};\mathbf{Y})+\text{df}(\mathcal{K})\log N+2{|\boldsymbol{\Theta}|\choose\text{df}(\mathcal{K})}.$
S.4 Additional Simulation Results
S.4.1 Simulation Results Under the Generic Identifiable Parameters $\mathcal{B}_{g}$ and Computation Time
Here, we present the omitted details from Section 5 in the main paper, regarding (a) the estimation accuracy under generically identifiable parameters, (b) performance of the PEM algorithm, and (c) computation time.
Estimation accuracy under generically identifiable parameters.
We present the analogs of Figure 3 under the generically identifiable true parameters $\mathcal{B}_{g}$ defined as:
$\mathbf{B}_{g}^{(d)}=\begin{pmatrix}-2\mathbf{1}_{K^{(d)}}&\mathbf{B}_{2}^{(d)}\\ -4\mathbf{1}_{K^{(d)}}&\mathbf{B}_{2}^{(d)\top}\\ -2\mathbf{1}_{K^{(d)}}&\mathbf{B}_{1}^{(d)}\end{pmatrix},\quad\text{where}\quad\beta_{2;j,k}^{(d)}$ :=$:=\begin{cases}4&\text{if }k=j,\\ 4/3&\text{ else if }0<k-j\leq\lceil K^{(d)}/3\rceil,\\ 0&\text{otherwise},\end{cases}$ (S.31)
and $\mathbf{B}_{1}^{(d)}$ is the matrix defined in Equation 13. Note that $\mathbf{B}_{g}^{(d)}$ is a less sparse version of $\mathbf{B}_{s}^{(d)}$, where the identity matrices are modified to $\mathbf{B}_{2}^{(d)}$. While $\mathbf{B}_{2}^{(d)}$ and $\mathcal{B}_{g}$ have the same value of 4 on the main diagonals, $\mathcal{B}_{s}$ is sparser. In particular, $\mathcal{B}_{s}$ has two exclusive children per latent variable whereas $\mathcal{B}_{g}$ has none.
While all other simulation settings are identical to that described in the main text, we make the following two changes. First, we do not consider the largest parameter dimension of $(J,K^{(1)},K^{(2)})=(90,30,10)$ in this generically identifiable setting. This is because the layerwise initialization turned out to be less effective in this high-dimensional but less-sparse scenario. Second, for Poisson-based-DDEs, we modify the intercept values to be smaller compared to the values in (13) as follows. When $K^{(2)}=2$, we consider a smaller intercept parameter for $\mathbf{B}_{g}^{(1)}$ as follows:
$\beta_{2;j,0}^{(1)}:=\begin{cases}-3&\text{ if }k\leq K^{(1)},\\ -5&\text{ else if }K^{(1)}<k\leq 2K^{(1)},\\ -2&\text{ otherwise}.\end{cases}$
Similarly for $K^{(2)}=6$, we work under the smaller intercept values of
$\beta_{2;j,0}^{(1)}:=\begin{cases}-10&\text{ if }\sum_{k=1}^{K^{(1)}}\beta_{j,k}\geq 8,\\ -5&\text{ otherwise}.\end{cases}$
This adjustment is to make the Poisson parameters not explode, as using the original intercept values in (S.31) makes some Poisson parameters for $Y_{j}\mid\mathbf{A}$ very large and generates unrealistic data.
Figure S.1: Estimation error for $\mathcal{G}$ and $\boldsymbol{\Theta}$ under the 2-layer DDE with parameters $\mathcal{B}_{g}$ and various observed-layer parametric families. Note that the y-axis varies across plots.
The estimation accuracy for $\mathcal{G}$ and $\boldsymbol{\Theta}$ is displayed in Figure S.1. While we see a similar trend as in the results under the true parameters $\mathcal{B}_{s}$ (Figure 3), the overall error values are larger and indicate that estimation under the less sparse model is more challenging. Recall that we are considering a different Poisson intercept compared to that under $\mathcal{B}_{s}$, so we cannot directly compare the Poisson results.
Estimation accuracy of the PEM algorithm.
We compare the performance of PEM (Algorithm 1) and SAEM (Algorithm 3) for each parametric family. We implement the PEM under the same simulation setup as SAEM, described in Section 5. However, the PEM implementation failed for higher latent dimensions, specifically under the latent dimension of $(J,K^{(1)},K^{(2)})=(54,18,6)$ or higher, due to memory allocation issues. As a result, PEM was only applied in the low-dimensional scenario with with $(J,K^{(1)},K^{(2)})=(18,6,2)$.
The results, presented in Tables S.3-S.5 (error of estimating $\mathcal{G}$) and S.6-S.8 (error of estimating $\boldsymbol{\Theta}$) in Section S.4.4, show that the PEM estimates exhibit a faster rate of convergence as $N$ increases. In particular, the RMSE values decrease at the parametric rate ${1}/{\sqrt{N}}$, as established in Theorem 3. We believe that the lower estimation accuracy under SAEM is due to multiple approximations within the SAEM algorithm such as approximate sampling and the objective function updates. We recommend using PEM instead of SAEM to estimate DDEs when the latent dimension is small.
Computation time.
Figure S.2 reports the computation times for all simulations in Section 5.1 and Section S.4.1. The results show that computation time varies across parametric families. A common trend is that both SAEM and PEM slow down as the sample size $N$ increases. Additionally, SAEM becomes slower as the parameter dimension increases. Comparing SAEM and PEM is somewhat subtle, as their relative computation times depend on the response type and different convergence criteria. However, preliminary simulations under the dimension $(J,K^{(1)},K^{(2)})=(45,15,5)$ indicate that PEM is slower than SAEM across all three responses types. Furthermore, PEM fails to operate under larger dimensions due to high memory requirements. These observations support the conclusion in the main paper that SAEM is desirable for scenarios involving large latent dimensions.
Figure S.2: Log computation time (in seconds) for the simulation results in Section 5 and Section S.4.1, under true parameters $\mathcal{B}_{s}$ (top row) and $\mathcal{B}_{g}$ (bottom row). The legends indicate the value of $K^{(1)}$ and estimation algorithm.
S.4.2 Experiments With Varying Numbers of Gibbs Samples in Algorithm 3
Here, we conduct additional experiments to assess the effect of the number of Gibbs samples, $C$, on both estimation accuracy and computation time. These experiments follow the setup in Section S.2.2, but we implement the SAEM algorithm (Algorithm 3) varying $C=1,5,25$.
The simulation results in Table S.2 indicate that smaller values of $C$ result in faster computation across all parametric families, with minimal loss in estimation accuracy. Here, the dependence of computation time on $C$ arises at the M-step, where the maximization objective is a sum of $O(C)$ functions. A larger $C$ complicates the objective function and makes the optimization slower. Based on these findings, we selected $C=1$ as a baseline for the SAEM procedure.
ParFam Accuracy($\mathcal{G}$) RMSE$(\boldsymbol{\Theta})$ time (s) # iterations
# Gibbs $C$$N$ 1000 4000 1000 4000 1000 4000 1000 4000
Bernoulli 1 0.916 0.962 0.43 0.36 3.1 10.0 3.0 3.6
5 0.920 0.961 0.43 0.38 5.3 41.4 2.6 3.2
25 0.923 0.962 0.43 0.37 26.0 198.5 2.5 3.1
Poisson 1 0.993 0.999 0.26 0.24 3.5 7.0 3.2 3.2
5 0.999 1 0.24 0.24 6.5 25.4 3.3 3.2
25 0.998 1 0.24 0.24 34.7 120.3 3.4 3.2
Normal 1 0.994 1 0.16 0.14 0.5 1.1 2.0 2.0
5 0.995 1 0.15 0.13 1.0 2.2 2.0 2.0
25 0.994 1 0.16 0.13 2.2 8.8 2.0 2.0
Table S.2: Accuracy measures for $\mathcal{G}$ and $\boldsymbol{\Theta}$, computation time and iterations for 2-layer DDE estimates under varying number of Gibbs samples $C$. For the first column, larger is better. For the other columns, smaller is better.
S.4.3 Simulations for Selecting the Number of Latent Variables
We first evaluate the performance of the three estimators (denoted as EBIC, LRT, Spectral) for selecting $K^{(1)}$, as introduced in Section S.3.2, assuming the knowledge of $K^{(2)}$. Since the EBIC and LRT estimators require likelihood computation, we restrict the simulations to the configuration $(J,K^{(1)},K^{(2)})=(18,6,2)$. For each candidate model with $k\in\mathfrak{K}$ first-latent-layer variables, the likelihood is computed using parameter estimates from the PEM algorithm. For the LRT estimator, we have set the significance level to be $\alpha=0.01$ and used the $\chi^{2}$ limiting distribution for each sequential test. Continuing from the simulation settings in Section 5, we consider two sets of true parameters: $\mathcal{B}_{s},\mathcal{B}_{g}$ and three sets of observed-layer parametric families: Bernoulli, Poisson, and Normal. In addition to the sample sizes $N$ in the previous section, $N=6000$ is also considered to better assess the large-sample accuracy. Here, we assume that the number of the top layer latent variables $K^{(2)}=2$ is known, and select $K^{(1)}$ from the candidate set $\mathfrak{K}=[2K^{(2)},J/2)\cap\mathbb{N}=\{4,5,6,7,8\}$. The equality in $2K^{(2)}\leq K^{(1)}$ is allowed to consider an equal number of underfitted/overfitted models.
Figure S.3: Selection accuracy of $K^{(1)}$ under the two-latent-layer DDE with true parameters $\mathcal{B}_{s}$. The spectral ratio estimator shows near-perfect accuracy for large $N$.
Figure S.4: Simulation results for selecting $K^{(1)}$ under the 2-layer DDE with true parameters $\mathcal{B}_{g}$.
We fit the three estimators $400$ times for each scenario, and display the correct selection percentage for the Bernoulli, Poisson, and Normal-based DDEs with true parameters $\mathcal{B}_{s}$/$\mathcal{B}_{g}$ in Figure S.3/Figure S.4. Comparing the three estimators, the spectral ratio-based estimator has near-perfect accuracy when $N$ is large enough, say $4000$. This demonstrates the empirical consistency of the spectral ratio estimator. In contrast, the consistency of the EBIC and LRT estimators is not clear for Bernoulli and Poisson responses [^2] To be more precise, as we are using approximate level $\alpha=0.01$ tests, the correct selection percentage of the LRT estimator should converge to 0.99.. Thus, we conclude that for a large enough $N$, it is desirable to select $K^{(1)}$ using the spectral ratio estimator. Among the three response types, the Bernoulli case with its nonlinear link function and limited response values is the most challenging. In contrast, the Normal case with a linear observed layer and continuous responses achieves near-perfect selection accuracy, which is consistent with earlier observations regarding parameter estimation accuracy.
Figure S.5: Selection accuracy of $K^{(2)}$ under two-latent-layer DDEs with true parameters $\mathcal{B}_{s}$.
Figure S.6: Simulation results for selecting $K^{(2)}$ under the 2-layer DDE with true parameters $\mathcal{B}_{g}$.
We also apply these three estimators to select $K^{(2)}$, assuming that $K^{(1)}=6$ is correctly estimated or known. The candidate set for $K^{(2)}$ is $\mathfrak{K}:=\{1,2,3\}$. Figure S.5/Figure S.6 displays the correct selection percentage under the true parameters $\mathcal{B}_{s}$/$\mathcal{B}_{g}$. Here, we implement the LRT estimator by naively assuming that Wilk’s theorem holds. The overall trends are similar to those observed in the previous figures for selecting $K^{(1)}$. Under the sparse true parameter $\mathcal{B}_{s}$, the spectral ratio estimator performs well across all response types and the LRT estimator has the lowest, but still decent accuracy. Interestingly, the EBIC estimator outperforms the spectral ratio estimator in more challenging scenarios, such as Bernoulli responses with small $N$ or under the Poisson responses with less sparse true parameters $\mathcal{B}_{g}$. Unlike the case for selecting $K^{(1)}$, it is not clear that the spectral ratio estimator outperforms the EBIC estimator for selecting $K^{(2)}$.
Based on these experiments, we conclude that in the most general setting of a two-latent-layer DDE where both $K^{(1)}$ and $K^{(2)}$ are unknown, a two-step approach is effective. First, the spectral ratio estimator can be used to select $K^{(1)}$. Then, $K^{(2)}$ can be determined using either the EBIC or the spectral ratio estimator, incorporating domain knowledge if needed.
S.4.4 Exact Values of Estimation Errors
We display the complete results of the simulation outputs corresponding to plots in Section 5, Section S.4.1, and Section S.4.3.
Accuracy of Estimating $\mathcal{G}$.
Tables S.3-S.5 display the average entrywise estimation accuracy for $\mathcal{G}$ under each parametric family. These values correspond to the upper rows of Figures 3 and S.1.
[…]
Bernoulli $\mathcal{B}_{s}$ (18,6,2) SAEM 0.899 0.931 0.956 0.974 0.985
PEM 0.926 0.966 0.986 0.992 0.992
(54,18,6) SAEM 0.910 0.967 0.990 0.996 0.998
(90,30,10) SAEM 0.914 0.971 0.991 0.996 0.998
$\mathcal{B}_{g}$ (18,6,2) SAEM 0.872 0.909 0.937 0.962 0.974
PEM 0.891 0.937 0.970 0.986 0.992
(54,18,6) SAEM 0.792 0.854 0.902 0.926 0.968
Table S.3: $\text{Acc}(\mathcal{G})$ under the Bernoulli 2-layer DDE.
[…]
Poisson $\mathcal{B}_{s}$ (18,6,2) SAEM 0.985 0.993 0.999 0.999 0.999
PEM 0.994 0.999 1 1 1
(54,18,6) SAEM 0.987 0.996 0.999 1 1
(90,30,10) SAEM 0.986 0.996 1 1 1
$\mathcal{B}_{g}$ (18,6,2) SAEM 0.981 0.988 0.997 0.992 0.997
PEM 0.994 0.995 0.998 0.998 1
(54,18,6) SAEM 0.801 0.845 0.861 0.867 0.891
Table S.4: $\text{Acc}(\mathcal{G})$ under the Poisson 2-layer DDE.
Parfam True parameter $(J,K^{(1)},K^{(2)})$ Algorithm 500 1000 2000 4000 8000
Normal $\mathcal{B}_{s}$ (18,6,2) SAEM 0.985 0.993 0.999 0.999 0.999
PEM 0.992 0.996 1 1 1
(54,18,6) SAEM 0.996 0.998 0.999 1 1
(90,30,10) SAEM 0.995 0.998 1 1 1
$\mathcal{B}_{g}$ (18,6,2) SAEM 0.993 0.994 0.996 0.998 0.998
PEM 0.992 0.997 0.999 1 1
(54,18,6) SAEM 0.993 0.995 0.998 0.999 1
Table S.5: $\text{Acc}(\mathcal{G})$ under the Normal 2-layer DDE.
Accuracy of Estimating $\boldsymbol{\Theta}$.
Tables S.6S.8 reports the RMSE values for estimating continuous parameters $\boldsymbol{\Theta}$. These values correspond to the bottom rows of Figures 3 and S.1.
[…]
Bernoulli $\mathcal{B}_{s}$ (18,6,2) SAEM 0.498 0.431 0.386 0.359 0.341
PEM 0.404 0.304 0.231 0.203 0.184
(54,18,6) SAEM 0.392 0.289 0.231 0.213 0.208
(90,30,10) SAEM 0.356 0.240 0.204 0.185 0.177
$\mathcal{B}_{g}$ (18,6,2) SAEM 0.553 0.474 0.420 0.400 0.387
PEM 0.520 0.398 0.288 0.206 0.165
(54,18,6) SAEM 0.740 0.650 0.596 0.577 0.412
Table S.6: $\text{RMSE}(\boldsymbol{\Theta})$ under the Bernoulli 2-layer DDE.
[…]
Poisson $\mathcal{B}_{s}$ (18,6,2) SAEM 0.281 0.258 0.245 0.243 0.242
PEM 0.219 0.158 0.108 0.080 0.061
(54,18,6) SAEM 0.289 0.168 0.153 0.147 0.145
(90,30,10) SAEM 0.281 0.195 0.125 0.117 0.116
$\mathcal{B}_{g}$ (18,6,2) SAEM 0.309 0.260 0.247 0.231 0.229
PEM 0.206 0.145 0.101 0.072 0.053
(54,18,6) SAEM 1.210 1.147 1.100 0.988 0.918
Table S.7: $\text{RMSE}(\boldsymbol{\Theta})$ under the Poisson 2-layer DDE.
[…]
Normal $\mathcal{B}_{s}$ (18,6,2) SAEM 0.189 0.156 0.147 0.137 0.133
PEM 0.170 0.128 0.080 0.063 0.054
(54,18,6) SAEM 0.117 0.090 0.072 0.063 0.057
(90,30,10) SAEM 0.108 0.074 0.054 0.045 0.041
$\mathcal{B}_{g}$ (18,6,2) SAEM 0.177 0.148 0.133 0.115 0.094
PEM 0.179 0.137 0.094 0.069 0.055
(54,18,6) SAEM 0.174 0.141 0.098 0.090 0.086
Table S.8: $\text{RMSE}(\boldsymbol{\Theta})$ under the Normal 2-layer DDE.
Accuracy of selecting $\mathcal{K}$.
Tables S.9-S.10 reports the accuracy of selecting $\mathcal{K}$ under Bernoulli and Poisson-based DDEs, which correspond to Figure S.3 and Figure S.4. Accuracy values for Normal-based DDEs are omitted, as all methods demonstrated near-perfect accuracy in this setting. Similarly, Tables S.11-S.12 display the results for selecting $K^{(2)}$, which corresponds to Figures S.5 and S.6.
These tables provide additional insights beyond those presented in figures, which only display the correct selection probability for the events $\widehat{K}^{(1)}=6$ and $\widehat{K}^{(2)}=2$. The percentage of incorrect estimates in the tables reveal systematic tendencies of the estimators: the EBIC and LRT estimators frequently overselect ($\widehat{K}^{(1)}\geq K^{(1)}$) whereas the spectral estimator sometimes underselects ($\widehat{K}^{(1)}\leq K^{(1)}$).
ParFam, value $N$ Method \ $\widehat{K}^{(1)}$ 4 5 6 7 8
Bernoulli, $\mathcal{B}_{s}$ 500 EBIC 0.25 96.25 3.5
LRT 0.5 98.75 0.75
Spectral 1.75 30.0 67.5 0.75
1000 EBIC 94.5 5.5
LRT 96.5 3.5
Spectral 0.25 18.25 81.5
2000 EBIC 88.25 11.75
LRT 93.5 6.25 0.25
Spectral 13.0 87.0
4000 EBIC 86.0 12.25 1.75
LRT 90.25 9.25 0.5
Spectral 3.75 96.25
6000 EBIC 86.0 13.5 1.5
LRT 87.75 12.25
Spectral 0.5 99.5
8000 EBIC 87.75 11.0 1.25
LRT 88.25 10.75 1.00
Spectral 0.75 99.25
Bernoulli, $\mathcal{B}_{g}$ 500 EBIC 84.5 13.75 1.75
LRT 0.75 1.0 86.25 12.0
Spectral 8.5 4.25 87.25
1000 EBIC 87.25 10.5 2.5
LRT 80.5 18.5 1.0
Spectral 0.5 99.5
2000 EBIC 85.25 11.5 3.25
LRT 82.75 16.25 1.0
Spectral 100.0
4000 EBIC 97.0 3.0
LRT 92.25 7.75
Spectral 100.0
6000 EBIC 97.5 2.5
LRT 91.5 8.5
Spectral 100.0
8000 EBIC 98.0 2.0
LRT 90.75 9.25
Spectral 100.0
Table S.9: Empirical distribution of the estimated $\widehat{K}^{(1)}$ values under the Bernoulli DDE with true $K^{(1)}=6$ and parameters $\mathcal{B}_{s},\mathcal{B}_{g}$. For each sample size, we present the method with the highest accuracy in bold. “ParFam” is short for “parametric family”.
ParFam, value $N$ Method \ $\widehat{K}^{(1)}$ 4 5 6 7 8
Poisson, $\mathcal{B}_{s}$ 500 EBIC 99.5 0.5
LRT 99.5 0.5
Spectral 100.0
1000 EBIC 99.0 1.0
LRT 99.0 1.0
Spectral 100.0
2000 EBIC 97.5 2.5
LRT 96.25 3.75
Spectral 100.0
4000 EBIC 98.25 1.5 0.25
LRT 97.25 2.75
Spectral 100.0
6000 EBIC 99.0 1.0
LRT 97.25 2.75
Spectral 100.0
8000 EBIC 98.75 0.75 0.5
LRT 97.75 1.75 0.5
Spectral 100.0
Poisson, $\mathcal{B}_{g}$ 500 EBIC 92.0 7.75 0.25
LRT 1.0 91.0 8.0
Spectral 8.5 91.5
1000 EBIC 90.0 9.5 0.5
LRT 88.5 10.75 0.75
Spectral 2.75 97.25
2000 EBIC 90.5 8.5 1.0
LRT 0.25 87.25 11.5 1.0
Spectral 0.5 99.5
4000 EBIC 91.25 7.25 1.5
LRT 89.0 9.25 1.75
Spectral 100.0
6000 EBIC 95.0 4.5 0.5
LRT 91.25 7.75 1.0
Spectral 100.0
8000 EBIC 94.0 5.0 1.0
LRT 90.75 8.25 1.0
Spectral 100.0
Table S.10: Empirical distribution of the estimated $\widehat{K}^{(1)}$ values under the Poisson DDE with true $K^{(1)}=6$ and parameters $\mathcal{B}_{s},\mathcal{B}_{g}$. For each sample size, we present the method with the highest accuracy in bold. “ParFam” is short for “parametric family”.
ParFam, value $N$ Method \ $\widehat{K}^{(2)}$ 1 2 3
Bernoulli, $\mathcal{B}_{s}$ 500 EBIC 0.5 99.5
LRT 16.75 83.0 0.25
Spectral 12.0 85.5 2.5
1000 EBIC 100.0
LRT 0.5 98.75 0.75
Spectral 6.5 91.5 2.0
2000 EBIC 100.0
LRT 98.0 2.0
Spectral 3.25 96.5 0.25
4000 EBIC 99.5 0.5
LRT 96.0 4.0
Spectral 2.75 97.25
6000 EBIC 98.0 2.0
LRT 93.0 7.0
Spectral 2.0 98.0
8000 EBIC 97.25 2.75
LRT 94.25 5.75
Spectral 2.0 98.0
Bernoulli, $\mathcal{B}_{g}$ 500 EBIC 0.75 99.25
LRT 13.5 86.0 0.5
Spectral 7.0 91.75 1.25
1000 EBIC 99.25 0.75
LRT 1.5 96.25 2.25
Spectral 3.5 95.25 1.25
2000 EBIC 99.25 0.75
LRT 0.5 94.0 5.5
Spectral 8.75 90.75 0.5
4000 EBIC 98.25 1.75
LRT 98.25 1.75
Spectral 2.75 96.25 1.0
6000 EBIC 97.5 2.5
LRT 96.5 3.5
Spectral 1.75 97.75 0.5
8000 EBIC 99.25 0.75
LRT 99.25 0.75
Spectral 0.5 99.0 0.5
Table S.11: Empirical distribution of the estimated $\widehat{K}^{(2)}$ values under the Bernoulli DDE with true $K^{(2)}=2$ and parameters $\mathcal{B}_{s},\mathcal{B}_{g}$. For each sample size, we present the method with the highest accuracy in bold. “ParFam” is short for “parametric family”.
Parfam, value $N$ Method \ $\widehat{K}^{(2)}$ 1 2 3
Poisson, $\mathcal{B}_{s}$ 500 EBIC 99.75 0.25
LRT 99.5 0.5
Spectral 0.25 99.75
1000 EBIC 99.75 0.25
LRT 99.75 0.25
Spectral 100.0
2000 EBIC 99.25 0.75
LRT 99.0 1.0
Spectral 100.0
4000 EBIC 98.25 1.75
LRT 98.25 1.75
Spectral 100.0
6000 EBIC 99.0 1.0
LRT 98.25 1.75
Spectral 100.0
8000 EBIC 98.5 1.5
LRT 98.5 1.5
Spectral 100.0
Poisson, $\mathcal{B}_{g}$ 500 EBIC 1.0 97.5 1.5
LRT 1.75 96.75 1.5
Spectral 16.5 81.75 1.75
1000 EBIC 0.25 96.25 3.5
LRT 0.75 95.5 3.75
Spectral 13.75 84.75 1.5
2000 EBIC 95.75 4.25
LRT 0.5 94.0 5.5
Spectral 8.75 90.75 0.5
4000 EBIC 0.25 96.5 3.25
LRT 0.75 94.75 4.5
Spectral 5.5 94.0 0.5
6000 EBIC 97.25 2.75
LRT 97.0 3.0
Spectral 3.5 96.0 0.5
8000 EBIC 0.25 97.5 2.25
LRT 0.25 97.0 2.75
Spectral 3.5 96.0 0.5
Table S.12: Empirical distribution of the estimated $\widehat{K}^{(2)}$ values under the Poisson DDE with true $K^{(2)}=2$ and parameters $\mathcal{B}_{s},\mathcal{B}_{g}$. For each sample size, we present the method with the highest accuracy in bold. “ParFam” is short for “parametric family”.
S.5 Data Analysis Details and Additional Results
S.5.1 Preprocessing
MNIST Data.
We work with the preprocessed version of the default training set, which consists of $60,000$ images, each containing information on the $28^{2}$ pixel values. Initially, each pixel takes integer values between 00 and $255$. As the data values are highly concentrated around values near 0 or larger than 200, we transform the data into binary responses by thresholding at a value of 128. In other words, pixels with values exceeding 128 are assigned a binary value of 1, while the rest are set to 0. For the sake of easier presentation and computational efficiency, we work on the subset of the dataset whose true digit labels are equal to $0,1,2,3$, and also discard some pixels with uniformly small values by selecting $J=264$ pixels whose average pixel values are larger than 40. After preprocessing, the training set consists of $N=20,679$ unlabeled images. We identically preprocess the test set, which leads to a total of $N=4,157$ images.
20 Newsgroups Data.
The dataset provides a default partition of the train and test sets based on chronological order, and we use this partition. The preprocessing of the training data was carried out in three main steps. First, to reduce the signal-to-noise ratio and enhance interpretability, we focus on a subset of labels. To elaborate, we consider the newsgroup articles that belong to the large class of computer, recreation, and science (see the top-latent layer labels in Figure S.7). A manual inspection revealed that the newsgroup with label sci.crypt has a wide range of topics such as government and politics and was often cross-referenced to those newsgroups, so we did not include these documents in our dataset. Second, documents of extreme lengths were filtered out by removing the shortest $5\%$ and longest $1\%$ of all documents. This procedure is standard in the literature, and has been shown to increase the signal-to-noise ratio (Ke and Wang,, 2024). Third, we construct our dictionary by excluding infrequent words (with less than 100 occurrences) and screening out stop-words (such as the, he, in) and topic-irrelevant words (such as like, must, since) using the R package tm. As the original dataset consists of email texts, we performed a manual screening to remove uninformative email-related vocabulary, such as edu, com, net, and cmu. Finally, we conducted a secondary filtering of the short documents, as the removal of stop-words led to some documents being uninformative. The resulting processed train dataset is a sparse $5883\times 653$ count matrix.
Y1.0e0]ptrecreationcomputersciencesports-baseballcarsmotorcyclesports-hockeyos-windowswindowsgraphicshardware-ibmhardware-macmedicinespaceelectronics
Figure S.7: Nested structure of the true held-out labels for the 20 newsgroups dataset.
To process the test dataset, we go through the same first and second steps described above. We continue using the dictionary constructed from the train set, with $J=653$ words. After filtering out the short/long documents, the resulting data matrix consists of $N=3320$ documents.
TIMSS Response Time Data.
We use the preprocessed dataset provided in Lee and Gu, (2024). Additionally, we remove all rows of the data matrix with missing values. While our estimation procedure can be modified to handle missing data (missing at random) by modifying the likelihood function, we take this additional preprocessing step for the sake of consistency with other parts of the paper.
S.5.2 Selecting the Number of Latent Variables
Figure S.8: Based on the spectral estimator for $K^{(1)}$, (left) we select $K^{(1)}=5$ for MNIST, and (right) $K^{(1)}=8$ for 20 newsgroups. The peak, highlighted in red, correspond to the selected values. We omit displaying the first few eigenvalue ratios for better illustration.
MNIST Data.
We select $K^{(1)}=5$ based on the spectral-gap estimator, as illustrated in the left panel of Figure S.8. Also, we set $K^{(2)}=2$ motivated by our identifiability requirement $2K^{(2)}<K^{(1)}$. This choice is also supported by the fact that there are four true labels. Each latent configuration $\mathbf{A}^{(2)}=(A_{1}^{(2)},A_{2}^{(2)})\in\{0,1\}^{2}$ uniquely corresponds to each digit, as illustrated in the center panel of Figure 4.
20 Newsgroups Data.
We select $K^{(1)}=8$ based on the spectral-gap estimator, as illustrated in the right panel of Figure S.8. For $K^{(2)}$, the EBIC values are quite similar for $K^{(2)}=2,3$ with values of $1.3273\times 10^{6}$ and $1.3277\times 10^{6}$, respectively. Also, the spectral-gap is similar for both values $K^{(2)}=2,3$. Hence, we select $K^{(2)}$ based on the interpretability of the inferred latent structure. The estimated latent structure with $K^{(2)}=3$ is displayed in Figure S.9. Compared to Figure 5 in the main text, the ‘technology’ topic has split into two groups whose interpretation is somewhat blurry. Notably, these two ‘technology’ latent variables do not match the labels of ‘computer’ and ‘science’ provided by the dataset (see Figure S.7). For instance, the ‘technology 2’ variable has arrows to the finer topics of motorcycle, software, and space. Hence, we select $K^{(2)}=2$ based on this lack of semantic distinction.
Y1.0e0]ptrecreationtechnology 1technology 2carssportsmotorcyclesoftwarelocation/nameshardwaregraphicsspace
Figure S.9: Graphical structure of the latent topics, where we fit the 2-layer DDE with $K^{(2)}=3,K^{(1)}=8$. The width of the upper layer arrows is proportional to the corresponding coefficients and the red arrow indicates negative values.
TIMSS Response Time Data.
The dataset comes with a rough classification of the seven latent skills into $K^{(2)}=2$ categories: content skills and cognitive skills. However, we observed that fitting $K^{(2)}=2$ results in one of the proportion parameters exhibiting an unusually large value of $p_{2}=0.96$. This indicates that $A_{2}^{(2)}=1$ for $96\%$ of the students, and this skill is redundant. Hence, we have adjusted the number of higher-order latent variables to $K^{(2)}=1$, which partitions the students into two groups based on the value of $A^{(2)}_{1}$. The EBIC for the two cases were also comparable and justified this choice.
S.5.3 Performance Evaluation for the 20 Newsgroups Dataset
Perplexity.
Perplexity is a very popular measure to evaluate topic models and many other machine learning models. Following the original definition from Blei et al., (2003) that considers each word as an individual sample, we define
$\text{perplexity}(\mathbf{Y}\mid\mathbf{A},\boldsymbol{\Theta}):=\exp\left[-\frac{\sum_{i,j}Y_{i,j}\log\left(\frac{\lambda_{i,j}}{\sum_{j^{\prime}}\lambda{i,j^{\prime}}}\right)}{\sum_{i,j}Y_{i,j}}\right].$ (S.32)
Here, $\lambda_{i,j}=\exp\big{(}\beta_{j,0}^{(1)}+\sum_{k\in[K^{(1)}]}\beta_{j,k}^{(1)}A_{i,k}^{(1)}\big{)}$ is the Poisson parameter for the conditional distribution $Y_{i,j}\mid\mathbf{A}_{i}^{(1)}$. The motivation for (S.32) is that under the Poisson likelihood for (4), the joint distribution of $(Y_{i,1},\ldots,Y_{i,J})\mid\mathbf{A}_{i}$ follows a Multinomial distribution
$\text{Multi}\left(\sum_{j}Y_{i,j},\left(\frac{\lambda_{i,1}}{\sum_{j^{\prime}}\lambda{i,j^{\prime}}},\ldots,\frac{\lambda_{i,J}}{\sum_{j^{\prime}}\lambda{i,j^{\prime}}}\right)\right).$
This definition is consistent with that used for Poisson factor analysis (Zhou et al.,, 2012; Gan et al.,, 2015) as well as LDA (Blei et al.,, 2003).
To evaluate (S.32), we use the parameter estimates $\widehat{\boldsymbol{\Theta}}$ from our proposed method. As (S.32) also requires the knowledge of each latent variable $\mathbf{A}_{i}^{(1)}$, latent variable estimates needs to be computed. To compute train perplexity, we use the estimator (14) from the main text. For test perplexity, we adopt the approach of Gan et al., (2015), first computing the posterior distribution of $\mathbf{A}_{i}^{(1)}$ by using $80\%$ of the words in each document, and evaluating perplexity based on the remaining $20\%$ words.
Implementations of existing topic modeling algorithms.
We describe the implementation details of alternative topic modeling algorithms that were used for model comparison in Table 4. LDA is implemented by the function LDA in the R package topicmodels (Grün and Hornik,, 2011), using the variational EM algorithm. For DPFA, we have used the original MATLAB codes that were publicly available on the first author’s GitHub page [^3] https://github.com/zhegan27/dpfa_icml2015/tree/master (Gan et al.,, 2015). Based on the empirical findings in Gan et al., (2015), which indicate that the DPFA-SBN model optimized using the SGNHT method performs the best among their proposed methods, we chose this implementation. The method was run using the default tuning parameters. Note while Gan et al., (2015) also analyzed the 20 Newsgroups dataset (see Table 1 in the cited paper), their reported perplexity values are not directly comparable to our results due to preprocessing. Notably, our preprocessing steps enhanced the signal-to-noise ratio and leads to a smaller perplexity.
S.5.4 Additional Visualization of MNIST Dataset
[…]
Figure S.10: Graphical model representation of the learned latent structure. The edge widths are proportional to coefficients’ absolute values. The edge colors (black/red) imply positive/negative coefficients, respectively.
In Figure S.10, we visualize the learned latent structure of the MNIST dataset. The corresponding graphical matrix $\mathbf{G}^{(2)}$ was used to interpret the top layer latent variables in the main text. For example, $A_{1}^{(2)}$ is connected to all lower-layer variables, and we interpret this as an indicator for the pixel density of the image.
Label True LCM 1-layer DDE Spectral init 2-layer DDE
0 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
1 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
2 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
3 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table S.13: True (first column) and reconstructed (other columns) images under various estimators. Here, each pixel takes values in $[0,1]$, where a darker shade indicates a larger value.
We additionally visualize examples of reconstructed images under various models in Table S.13, which have been used to compute the pixel-wise reconstruction accuracy in Table 3 in the main paper. The reconstructions are based on the corresponding conditional mean $\mathbb{E}\left(\mathbf{Y}\mid\mathbf{A}^{(1)}\right)$ for each latent representation $\mathbf{A}^{(1)}$, which is rearranged in the original $28\times 28$ grid. The visualization illustrates that the reconstructions from the 2-layer DDE exhibit greater clarity compared to those from alternative models.

### v2

@@
 DGMs are often overparametrized, non-identifiable, and uninterpretable black boxes, raising serious concerns when deploying them in high-stakes applications.
-Motivated by this, we propose an interpretable deep generative modeling framework for rich data types with discrete latent layers, called Deep Discrete Encoders (DDEs).
+Motivated by this, we propose interpretable deep generative models for rich data types with discrete latent layers, called Deep Discrete Encoders (DDEs).
 A DDE is a directed graphical model with multiple binary latent layers.
@@
 This procedure can efficiently estimate models with exponentially many latent components.
-Extensive simulation studies validate our theoretical results and demonstrate the proposed algorithms’ excellent performance.
-We apply DDEs to three diverse real datasets for hierarchical topic modeling, image representation learning, response time modeling in educational testing, and obtain interpretable findings.
-Corresponding Author: Yuqi Gu. Email: yuqi.gu@columbia.edu. Research is partially supported by NSF Grant DMS-2210796.
+Extensive simulation studies for high-dimensional data and deep architectures validate our theoretical results and demonstrate the excellent performance of our algorithms.
+We apply DDEs to three diverse real datasets with different data types to perform hierarchical topic modeling, image representation learning, and response time modeling in educational testing.
+Corresponding Author: Yuqi Gu. Email: yuqi.gu@columbia.edu.
 Keywords: Identifiability; Interpretable Artificial Intelligence; Representation Learning; Deep Belief Network; Directed Graphical Model.
@@
 The black/red edge colors imply positive/negative coefficients, respectively.
-The main contributions in this paper include rigorous identifiability theory and scalable computational pipelines for DDEs.
-For identifiability, we propose general identifiability conditions in terms of the dependence graph structure between adjacent layers in the graphical model (corresponding to the directed arrows in Figure 1).
-We work under the minimal possible assumptions in order to flexibly cover the various examples mentioned above.
+The main contributions of this paper include rigorous identifiability theory and scalable computational pipelines for DDEs.
+For identifiability, we propose general identifiability conditions in terms of the probabilistic graph structures between layers in the graphical model (corresponding to the directed arrows in Figure 1).
+We work under the minimal possible assumptions in order to flexibly cover the various examples and data types mentioned above.
 Next, we provide an informal statement of the identifiability conditions.
@@
 Theorem (Informal version of Theorems 1 and 2).
-The DDE with a known number of latent variables is identifiable up to latent variable permutation in each layer, as long as each latent variable has at least three exclusive children.
-Under a weaker notion of “generic identifiability”, this condition can be relaxed to that each latent variable has at least three children that are not necessarily exclusive.
+The DDE is identifiable up to latent variable permutation in each layer, as long as each latent variable has at least three pure children (which only has one parent variable).
+Under a weaker notion of “generic identifiability”, this condition can be relaxed to that each latent variable has at least three children that are not necessarily pure.
 For computation, we propose a scalable estimation pipeline for DDEs.
 The multiple layers of nonlinear latent variables in DDEs lead to a highly nonconvex optimization landscape with potentially exponentially many local optima.
-To address this challenging setting, our computational pipeline features a nuanced layerwise nonlinear spectral initialization followed by a penalized stochastic approximation EM algorithm.
+To address this challenging issue, our computational pipeline features a nuanced layerwise nonlinear spectral initialization followed by a penalized stochastic approximation EM algorithm.
 This procedure can efficiently handle models with a large number of latent variables.
@@
 Extensive simulation studies not only validate the identifiability results, but also demonstrate the excellent performance of the proposed algorithms.
-We apply DDEs to real data in three diverse tasks, including hierarchical topic modeling, image representation learning, and response time modeling in digital educational testing.
+We apply DDEs to real data in three diverse tasks, including hierarchical topic modeling, image representation learning, and multimodal modeling in digital educational testing.
 Across these applications, DDEs extract highly interpretable latent structures and learn useful representations for downstream analyses.
@@
 For the considered unsupervised learning setting, we use “generative model” to refer to probabilistic models where observed data are generated conditional on hidden variables.
-Classical examples with one hidden layer include mixture models and mixed membership models.
 Recently, powerful multilayer models have been proposed for more complex tasks, which either consist of multiple latent layers or a single latent layer transformed by deep neural networks.
-Popular models in machine learning include deep belief networks (Hinton et al.,, 2006), deep Boltzmann machines (Salakhutdinov and Hinton,, 2009), variational autoencoders (Ranzato and Szummer,, 2008), generative adversarial networks (Goodfellow et al.,, 2014), diffusion models (Sohl-Dickstein et al.,, 2015), and transformer-based models (Vaswani et al.,, 2017) such as large language models.
+Popular models in machine learning include deep belief networks (DBNs, Hinton et al., , 2006), deep Boltzmann machines (DBMs, Salakhutdinov and Hinton, , 2009), variational autoencoders (Ranzato and Szummer, , 2008), generative adversarial networks (Goodfellow et al., , 2014), diffusion models (Sohl-Dickstein et al., , 2015), and transformer-based models (Vaswani et al., , 2017) such as large language models.
+Our proposed models are most closely related to DBNs and DBMs.
 See Section 2.2 for more discussions.
@@
 Notation.
-For a positive integer $K$, denote $[K]=\{1,\ldots,K\}$.
+For a positive integer $K$, denote $K]=\{1,\ldots,K\}$.
 For a matrix $\mathbf{G}$ with $J$ rows, let $\mathbf{g}_{1},\ldots,\mathbf{g}_{J}$ denote its row vectors.
@@
 For a finite set $I$, let $S_{I}$ denote the collection of all permutation maps of $I$, and let $\text{id}_{I}$ be the identity permutation on $I$.
-For nonnegative sequences $\{(a_{N},b_{N})\}_{N\geq 1}$, write $a_{N}\ll b_{N}$ when ${a_{N}}/{b_{N}}\to 0$ as $N\to\infty$.
 2.1 Model Definition
-We formally define the $D$-latent-layer DDE as a generative model with $D$ discrete latent layers.
-For each $d\in[D]$, assume that the $(d-1)$th layer is generated conditional on the $d$th.
+The $D$-latent-layer DDE is a generative model with $D$ discrete latent layers.
+For each $d\in[D]$, assume that the (d1)(d-1)th layer is generated conditional on the $d$th.
 Here, only the bottom layer (indexed by $d=0$) is observed, and all other layers are latent.
@@
 Denote the $d$th latent layer as $\mathbf{A}^{(d)}=\big{(}A_{1}^{(d)},\ldots,A_{K^{(d)}}^{(d)}\big{)}\in\{0,1\}^{K^{(d)}}$, which is a $K^{(d)}$-dimensional binary vector.
-A DDE has a shrinking-ladder-shaped architecture, with the dimension of each layer decreasing as $d$ increases: $K^{(D)}<\ldots<K^{(1)}<J$.
+A DDE has a shrinking-ladder-shaped deep architecture, with the dimension of each layer decreasing as $d$ increases: $K^{(D)}<\ldots<K^{(1)}<J$.
 See Figure 1 for a graphical model representation.
@@
 For each $d>1$, suppose that $\mathbb{P}(\mathbf{A}^{(d)})$, the distribution of the $d$th layer, is given.
-Then, we define the $(d-1)$th layer distribution by assuming the conditional independence of $A_{1}^{(d-1)},\ldots,A_{K^{(d-1)}}^{(d-1)}$ given $\mathbf{A}^{(d)}$:
+Then, we define the (d1)(d-1)th layer distribution by assuming the conditional independence of $A_{1}^{(d-1)},\ldots,A_{K^{(d-1)}}^{(d-1)}$ given $\mathbf{A}^{(d)}$:
 $\mathbb{P}(\mathbf{A}^{(d-1)}=\boldsymbol{\alpha}^{(d-1)}\mid\mathbf{A}^{(d)})=\prod_{k=1}^{K^{(d-1)}}\mathbb{P}(A^{(d-1)}_{k}=\alpha^{(d-1)}_{k}\mid\mathbf{A}^{(d)}),\quad\forall\boldsymbol{\alpha}^{(d-1)}\in\{0,1\}^{K^{(d-1)}}.$ (2)
@@
 $Y_{j}\mid\mathbf{A}^{(1)}\sim\text{ParFam}_{j}\Big{(}g_{j}\Big{(}\beta_{j,0}^{(1)}+\sum_{k=1}^{K^{(1)}}\beta_{j,k}^{(1)}A_{k}^{(1)},\gamma_{j}\Big{)}\Big{)}.$ (4)
-Here, $\text{ParFam}_{j}$ denotes a specific parametric family, and let $H_{j}$ denote its parameter space.
+Here, $\text{ParFam}_{j}$ denotes a specific identifiable parametric family, and let $H_{j}$ denote its parameter space.
 For convenience, let $p_{j}$ be the probability mass/density function of $\text{ParFam}_{j}$.
@@
 $~\mathbb{P}(\mathbf{Y})=$ $~\sum_{\begin{subarray}{c}\boldsymbol{\alpha}^{(d)}\in\{0,1\}^{K^{(d)}}\\ \forall d\in[D]\end{subarray}}\mathbb{P}(\mathbf{Y}|\mathbf{A}^{(1)}=\boldsymbol{\alpha}^{(1)})\prod_{d=2}^{D}\mathbb{P}(\mathbf{A}^{(d-1)}=\boldsymbol{\alpha}^{(d-1)}\mid\mathbf{A}^{(d)}=\boldsymbol{\alpha}^{(d)})\mathbb{P}(\mathbf{A}^{(D)}=\boldsymbol{\alpha}^{(D)}).$ (5)
-The $D$-latent-layer DDE is parametrized by $({\boldsymbol{p}},\mathcal{B},\boldsymbol{\gamma})$, where :=$\mathcal{B}:=\{\mathbf{B}^{(d)}\}_{d\in[D]}$ and $\boldsymbol{\gamma}:=(\gamma_{1},\ldots,\gamma_{J})$.
+The $D$-latent-layer DDE is parametrized by $({\boldsymbol{p}},\mathcal{B},\boldsymbol{\gamma})$, where $\mathcal{B}:=\{\mathbf{B}^{(d)}\}_{d\in[D]}$ and $\boldsymbol{\gamma}:=(\gamma_{1},\ldots,\gamma_{J})$.
 Upon the above marginalization, the induced DDE is a highly expressive model with exponentially many latent mixture components.
@@
 Define $K^{(0)}:=J$ for notational convenience.
-By (2) and (3), $\mathbf{G}^{(d)}$ can also be viewed as the adjacency matrix or “graphical matrix” between the $(d-1)$th layer and the $d$th layer in the graphical model representation (see Figure 1).
-Collect all graphical matrices $\mathbf{G}^{(d)}$s by defining $\mathcal{G}:=\{\mathbf{G}^{(d)}\}_{d\in[D]}$.
-Now, we formally define $D$-latent-layer DDEs, which also incorporate $\mathcal{G}$ as unknown parameters.
+By (2) and (3), $\mathbf{G}^{(d)}$ can also be viewed as the adjacency matrix or “graphical matrix” between the (d1)(d-1)th layer and the $d$th layer in the graphical model representation (see Figure 1).
+Collect all $\mathbf{G}^{(d)}$s by defining $\mathcal{G}:=\{\mathbf{G}^{(d)}\}_{d\in[D]}$, and collect the number of latent variables in each layer by defining $\mathcal{K}:=\{K^{(d)}\}_{d\in[D]}.$ Now, we formally define $D$-latent-layer DDEs, which also incorporate $\mathcal{G}$ as unknown parameters.
 Definition 1 (DDE).
@@
 2.2 Connections with Existing Models
-Deep Probabilistic Graphical Models in Machine Learning.
-There are many deep latent variable models in the machine learning literature, including the popular deep Boltzmann machine (DBM, Salakhutdinov and Hinton,, 2009), deep belief networks (DBNs, Hinton et al.,, 2006), and deep exponential families (DEFs, Ranganath et al.,, 2015).
+Related deep latent variable models in the machine learning literature include the deep Boltzmann machine (DBM, Salakhutdinov and Hinton, , 2009), deep belief networks (DBNs, Hinton et al., , 2006), and deep exponential families (DEFs, Ranganath et al., , 2015).
 The DBM and DBN both contain multiple binary latent layers and differ in the directions of the edges; see Figure 2.
@@
 DEFs are an unsupervised modeling framework that uses exponential families to model conditional distributions for each layer.
-[…]
+$A_{1}^{(2)}$$A_{2}^{(2)}$$A_{3}^{(2)}$$A_{1}^{(1)}$$A_{2}^{(1)}$$A_{3}^{(1)}$$Y_{1}$$Y_{2}$$Y_{3}$
 DBM
-[…]
+$A_{1}^{(2)}$$A_{2}^{(2)}$$A_{3}^{(2)}$$A_{1}^{(1)}$$A_{2}^{(1)}$$A_{3}^{(1)}$$Y_{1}$$Y_{2}$$Y_{3}$
 DBN
-[…]
+$A_{1}^{(2)}$$A_{2}^{(2)}$$A_{3}^{(2)}$$A_{1}^{(1)}$$A_{2}^{(1)}$$A_{3}^{(1)}$$Y_{1}$$Y_{2}$$Y_{3}$
 DEF
-[…]
+$A_{1}^{(2)}$$A_{1}^{(1)}$$A_{2}^{(1)}$$Y_{1}$$Y_{2}$$Y_{3}$
 DDE
@@
 This allows us to establish identifiability and consistency as well as effectively reduce the model dimension and interpret the latent structure.
-Recently, Gu and Dunson, (2023) proposed Bayesian pyramids, which are identifiable multilayer discrete latent variable models with a pyramid structure.
+On a related note, several identifiable deep models have recently been proposed.
+Gu and Dunson, (2023) proposed Bayesian pyramids, which are identifiable multilayer discrete latent variable models with a pyramid structure.
 But the methodology and identifiability theory therein are restricted to multivariate categorical data.
-DDEs significantly broaden the applicability of Bayesian pyramids by modeling arbitrarily flexible data distributions while still remaining identifiable.
-For computation, we propose scalable frequentist estimation approaches for DDEs, rather than adopting Bayesian MCMC inference as in Gu and Dunson, (2023).
-Psychometrics.
-Multidimensional binary latent variables are also popular for modeling students’ item response data in educational measurement (Junker and Sijtsma,, 2001; von Davier,, 2008; de la Torre,, 2011).
-These psychometric models are known as cognitive diagnostic models (CDMs).
-A CDM uses a single latent layer of cognitive skills to model a student’s binary responses to many questions in a test, with a sparse loading graph between the observed and latent layers.
-Here, each binary latent variable represents a student’s mastery or deficiency of a cognitive skill, and the sparse graph between the observed item responses and latent skills encodes which skills each item is designed to measure in the test.
-The proposed DDEs substantially generalize this modeling idea by allowing (a) multilayer latent variables, which can model one’s knowledge structure at multiple resolutions ranging from fine-grained details to general concepts (Gu,, 2024), and (b) modeling a rich class of general responses, going beyond the typical binary correct/incorrect responses in educational tests.
-3 Theoretical Guarantees for General DDEs
+DDEs significantly broaden the applicability of Bayesian pyramids by modeling arbitrarily flexible data types while still remaining identifiable.
+Kong et al., (2024) considered a modeling framework allowing both discrete and continuous latent variables, where the discrete part of the model can have more flexible graphical structures than the multi-layer structure.
+However, as a tradeoff, they consider a weaker notion of identifiability (up to atomic cover structures) and propose less explicit conditions for identifiability.
+Finally, Anandkumar et al., (2013) and Xie et al., (2024) consider continuous latent variable models and prove identifiability under potentially more general graph structures, but assume that all conditional distributions are linear.
+In contrast, DDEs allow arbitrary nonlinear link functions, which improves representation power.
+We remark that DDEs are identifiable even when higher-order interaction effects among latent variables are in the model (as detailed in Supplement S.1.6).
+We provide additional literature review regarding identifiable VAEs and psychometric models in Supplement S.6.
+3 Theoretical Guarantees for DDEs
 3.1 Model Identifiability
 In this section, we establish the identifiability of DDEs.
-Assuming known numbers of latent variables, we prove that the continuous model parameters and the discrete graph structures between layers can be uniquely identified under certain conditions on the true graphical matrices $\mathbf{G}^{(d)}$’s.
+Assuming known numbers of latent variables, we prove that both the continuous model parameters and the discrete graph structures between layers can be uniquely identified under certain conditions on the true graphical matrices $\mathbf{G}^{(d)}$’s.
 We first make an assumption to address some trivial ambiguities.
+For notational convenience, denote $K^{(0)}:=J$ and $A_{j}^{(0)}:=Y_{j}$.
 Assumption 1.
 Assume that the graphical matrices $\mathcal{G}=\{\mathbf{G}^{(d)}\}_{d\in[D]}$ and proportion parameters $\boldsymbol{p}$ satisfy the following conditions.
-Here, for notational convenience, let $K^{(0)}:=J$.
   1. (a)
@@
 Definition 3 (Identifiability up to equivalence).
-For a D-layer DDE with $\mathcal{K}$ latent variables, we define an equivalence relationship “$\sim_{\mathcal{K}}$” by setting $(\boldsymbol{\Theta},\mathcal{G})\sim_{\mathcal{K}}(\widetilde{\boldsymbol{\Theta}},\widetilde{\mathcal{G}})$ if and only if $\boldsymbol{\gamma}=\widetilde{\boldsymbol{\gamma}}$ and there exists permutations $\sigma^{(d)}\in S_{[K^{(d)}]}$ for all $d\in[D]$ such that the following hold:
+[…]
   • $p_{k}=\widetilde{p}_{\sigma^{(D)}(k)}$
@@
 Despite the seemingly heavy notation, the equivalence relationship is quite natural.
-Here, each permutation $\sigma^{(d)}$ addresses the label-switching issue for indexing the latent variables in the $d$th latent layer.
-Below is a main result on the identifiability of $D$-latent-layer DDEs.
+For example, in the right panel of Figure 1, the location of the latent variables in each layer can be arbitrarily permuted without changing the likelihood.
+In other words, the latent variable “sports” and “car” can be equivalently indexed by $(1,2)$ or $(2,1)$ as long as the associated arrows are permuted accordingly.
+For the $d$th latent layer, the permutation $\sigma^{(d)}$ corresponds to this fundamental yet relatively trivial label-switching map.
+The following theorem is our first main result on the identifiability of DDEs.
+Here, we say that a variable $X$ is a “pure child” of $Y$ if $Y$ is the only parent of $X$.
 Theorem 1 (Identifiability of DDEs).
-Consider the $D$-latent-layer DDE with $\mathcal{K}$ latent variables and true parameters $(\boldsymbol{\Theta}^{\star},\mathcal{G}^{\star})$.
-Suppose that for all $d\leq D-1$, the true graphical matrices $\mathbf{G}^{(d)\star}$ and parameters $\mathbf{B}^{(d)\star}$ satisfy the following conditions:
+Consider a $D$-latent-layer DDE with true parameters $(\boldsymbol{\Theta}^{\star},\mathcal{G}^{\star})$, where the number of latent variables, $\mathcal{K}$, is given.
+Suppose that for all $d\leq D-1$, the true graphical structures $\mathbf{G}^{(d)\star}$ and parameters $\mathbf{B}^{(d)\star}$ satisfy the following conditions:
   1. A.
-    $\mathbf{G}^{(d)\star}$ contains two identity matrices after permuting the rows.
-Without the loss of generality, suppose that the first $2K^{(d)}$ rows of $\mathbf{G}^{(d)\star}$ are $\big{[}\mathbf{I}_{K^{(d)}},\mathbf{I}_{K^{(d)}}\big{]}^{\top}.$
+    Each latent variable $A_{k}^{(d)}$ has at least two pure children $A_{j_{k,1}}^{(d-1)},A_{j_{k,2}}^{(d-1)}$ according to $\mathbf{G}^{(d)\star}$.
   2. B.
-    For any $\boldsymbol{\alpha}^{(d)}\neq\boldsymbol{\alpha}^{\prime(d)}\in\{0,1\}^{K^{(d)}}$, there exists $j>2K^{(d)}$ such that $\sum_{k=1}^{K^{(d)}}\beta_{j,k}^{(d)\star}(\alpha_{k}^{(d)}-\alpha_{k}^{(d)^{\prime}})\neq 0$.
+    For any $\boldsymbol{\alpha}^{(d)}\neq\boldsymbol{\alpha}^{\prime(d)}\in\{0,1\}^{K^{(d)}}$, there exists $j\in[K^{(d-1)}]\setminus\cup_{k=1}^{K^{(d)}}\{j_{k,1},j_{k,2}\}$ such that $\sum_{k=1}^{K^{(d)}}\beta_{j,k}^{(d)\star}(\alpha_{k}^{(d)}-\alpha_{k}^{{}^{\prime}(d)})\neq 0.$
 Then, the model components $(\boldsymbol{\Theta},\mathcal{G})$ are identifiable.
 Our key observation behind proving the identifiability of the complex deep generative structures in DDEs is that, with discrete latent layers, it suffices to establish identifiability for a one-latent-layer model and proceed in a layer-wise manner inductively.
-To elaborate, consider a “collapsed” DDE where the latent layers indexed by $d=2,\ldots,D$ are marginalized out to give a probability mass function for the first latent layer: $\{\mathbb{P}(\mathbf{A}^{(1)}=\boldsymbol{\alpha}^{(1)}):\boldsymbol{\alpha}^{(1)}\in\{0,1\}^{K^{(1)}}\}$.
+To elaborate, consider a “collapsed” DDE where the latent layers indexed by $d=2,\ldots,D$ are marginalized out to give a probability mass function (pmf) for the first latent layer: $\{\mathbb{P}(\mathbf{A}^{(1)}=\boldsymbol{\alpha}^{(1)}):\boldsymbol{\alpha}^{(1)}\in\{0,1\}^{K^{(1)}}\}$.
 If the collapsed model is proven to be identifiable, that means the conditional distributions $\mathbb{P}(Y_{j}\mid\mathbf{A}^{(1)})$ and the marginal distribution $\mathbb{P}(\mathbf{A}^{(1)})$ can be uniquely identified from the data distribution $\mathbb{P}(\mathbf{Y})$.
-In this case, we can determine all values of $\mathbb{P}(\mathbf{A}^{(1)}=\boldsymbol{\alpha}^{(1)})$ for every $\boldsymbol{\alpha}^{(1)}\in\{0,1\}^{K^{(1)}}$.
-Then, we can theoretically treat the shallowest latent layer $\mathbf{A}^{(1)}$ as if it were observed, because its probability mass function (pmf) is now identified and known.
-Then, viewing this pmf as the marginal distribution of $K^{(1)}$-dimensional observed variables of a $(D-1)$-latent-layer Bernoulli-DDE, we can inductively identify all parameters via a layerwise argument.
+In this case, we can determine the structured pmf up to the inevitable permutation of the latent variables (indexed by $\sigma^{(1)}\in S_{[K^{(1)}]}$ in Definition 3).
+Then, we can theoretically treat the shallowest latent layer $\mathbf{A}^{(1)}$ as if it was observed, because its probability mass function is now identified and known.
+Then, viewing this pmf as the marginal distribution of $K^{(1)}$-dimensional observed variables of a (D1)(D-1)-latent-layer Bernoulli-DDE, we can inductively identify all parameters via a layerwise argument.
 Please see Supplementary Material S.1 for the detailed proof.
-We interpret the conditions in Theorem 1.
-Condition A requires each latent variable to have at least two exclusive children in the layer below.
+We next give an interpretation of the conditions in Theorem 1.
+Condition A requires each latent variable $A_{k}^{(d)}$ to have at least two pure children in the layer below.
+For example, the latent variable “tech” in the right panel of Figure 1 has two pure children “graphics” and “space”.
 Condition B is more technical and is introduced to distinguish the different binary latent configurations $\boldsymbol{\alpha}\neq\boldsymbol{\alpha}^{\prime}$.
-Condition B holds when $\mathbf{G}^{(d)\star}$ also contains exclusive children for all latent variables.
-The exclusive children requirements in condition A can be further relaxed under a weaker notion of generic identifiability.
+For example, condition B holds when each latent variable $A_{k}^{(d)}$ has a third pure child that is distinct from those in condition A.
+We also provide identifiability guarantees for the number of latent variables in Supplement S.1.6.
+The pure children requirements in condition A can be further relaxed under a weaker notion of generic identifiability.
 Generic identifiability allows a measure-zero subset of the parameter space to be non-identifiable, and often holds under weaker conditions than that in Definition 3.
@@
 Consider a $D$-latent-layer DDE with $\mathcal{K}$ latent variables, graphical matrices $\mathcal{G}^{\star}$, and true parameters belonging to $\Omega_{\mathcal{K}}(\boldsymbol{\Theta};\mathcal{G}^{\star})$.
-The model is generically identifiable up to $\sim_{\mathcal{K}}$ when
-[…]
-is a measure-zero set with respect to $\Omega_{\mathcal{K}}(\boldsymbol{\Theta};\mathcal{G}^{\star})$.
-We introduce an additional assumption on the parametric families used to model the conditional distribution $Y_{j}\mid\mathbf{A}^{(1)}$ in (4).
+The model is generically identifiable up to $\sim_{\mathcal{K}}$ when $\{\boldsymbol{\Theta}\in\Omega_{\mathcal{K}}(\boldsymbol{\Theta};\mathcal{G}^{\star})$: there exists $(\widetilde{\boldsymbol{\Theta}},\widetilde{\mathcal{G}})\not\sim_{\mathcal{K}}(\boldsymbol{\Theta},\mathcal{G}^{\star})$ such that $\mathbb{P}_{\widetilde{\boldsymbol{\Theta}},\widetilde{\mathcal{G}}}=\mathbb{P}_{\boldsymbol{\Theta},\mathcal{G}^{\star}}\}$ is a measure-zero set with respect to $\Omega_{\mathcal{K}}(\boldsymbol{\Theta};\mathcal{G}^{\star})$.
+For generic identifiability, we introduce an additional assumption on the parametric families used to model the conditional distribution $Y_{j}\mid\mathbf{A}^{(1)}$ in (4).
 This is a technical assumption that arises from our proof technique for dealing with measure-zero sets.
@@
 We say that $p$ is analytic when the pmf/pdf $p(Y;\eta,\gamma)$ is analytic in both $\eta,\gamma$, for all $Y\in\mathcal{Y}$.
-Now, we are ready to state the generic identifiability result for $D$-latent-layer DDEs.
+We next state the generic identifiability result for DDEs.
+In graph theory, a bipartite graph is said to have a “perfect matching” if it contains a set of edges without common vertices that covers every vertex of the graph (Hall, , 2011); see Example 1 for an illustration.
 Theorem 2 (Generic identifiability of DDEs).
-Consider the $D$-latent-layer DDE with $\mathcal{K}$ latent variables and true graphical matrices $\mathcal{G}^{\star}$, where all parametric families and link function $g_{j}$s are analytic, and the true parameter lives in $\Omega_{\mathcal{K}}(\boldsymbol{\Theta};\mathcal{G}^{\star})$.
-Suppose that for all $d\leq D-1$, the true graphical matrices $\mathbf{G}^{(d)}$ satisfy the following condition C:
+Consider the $D$-latent-layer DDE where the number of latent variables $\mathcal{K}$ is given, and all parametric families and link function $g_{j}$s are analytic.
+Let $\mathcal{G}^{\star}$ denote the true graphical matrices and suppose the true parameter lives in $\Omega_{\mathcal{K}}(\boldsymbol{\Theta};\mathcal{G}^{\star})$.
+Suppose that for all $d\leq D-1$, the true graphical structure $\mathbf{G}^{(d)}$ satisfy the following condition C:
   1. C.
-    After permuting the rows, we can write $\mathbf{G}^{(d)\star}=\left[\mathbf{G}_{1}^{(d)\top},\mathbf{G}_{2}^{(d)\top},\mathbf{G}_{3}^{(d)\top}\right]^{\top}$, where $\mathbf{G}_{1}^{(d)}$ and $\mathbf{G}_{2}^{(d)}$ are $K^{(d)}\times K^{(d)}$ matrices whose diagonal entries are ones:
-    $\mathbf{G}_{1}^{(d)},~{}\mathbf{G}_{2}^{(d)}=$ $\begin{pmatrix}1&*&\cdots&*\\ *&1&\cdots&*\\ \vdots&\vdots&\ddots&\vdots\\ *&*&\cdots&1\end{pmatrix}$, where each * denotes an arbitrary value in $\{0,1\}$, and $\mathbf{G}_{3}^{(d)}$ does not have any all-zero columns.
+    There exists a partition of the (d1)(d-1)th layer variable indices $K^{(d-1)}]=\mathcal{I}_{1}\cup\mathcal{I}_{2}\cup\mathcal{I}_{3}$ that satisfies the following properties: (i) for $a=1,2$, there is a perfect matching in the bipartite graph between $\mathbf{A}^{(d-1)}_{\mathcal{I}_{a}}$ and $\mathbf{A}^{(d)}$, (ii) each latent variable $A_{k}^{(d)}$ has at least one child among $\mathbf{A}^{(d-1)}_{\mathcal{I}_{3}}$.
 Then, the model components $(\boldsymbol{\Theta},\mathcal{G})$ are generically identifiable.
 Condition C relaxes conditions A and B in Theorem 1.
-While condition A requires two exclusive children for each latent variable, condition C does not require any exclusive child and allows more complicated dependence structures.
+While condition A requires two pure children for each latent variable, condition C does not require any pure child and allows more complex dependence structures.
 Additionally, condition B for the remaining $J-2K^{(d)}$ variables is relaxed to a simple non-zero column condition on $\mathbf{G}^{(d)}_{3}$.
 Thus, condition C does not concern any continuous parameter values and just depends on the graph structure $\mathbf{G}^{(d)}$, and can be used as a practical criterion to assess identifiability.
+Example 1 (Illustration of identifiability conditions).
+We provide a toy example to illustrate and compare conditions in Theorem 1 and Theorem 2.
+In Figure 3, condition A holds when only the solid edges exist; in this case, the latent variable $A_{1}^{(d)}$ has two pure children $A_{1}^{(d-1)},A_{3}^{(d-1)}$, and latent variable $A_{2}^{(d)}$ also has two pure children.
+In contrast, condition C allows arbitrary additional dashed arrows in the graph structure; in this case, by taking $\mathcal{I}_{1}=\{A_{1}^{(d-1)},A_{2}^{(d-1)}\}$ and $\mathcal{I}_{2}=\{A_{3}^{(d-1)},A_{4}^{(d-1)}\}$, the red and blue solid edges each form a perfect matching.
+Thus, condition C significantly relaxes the pure child condition in condition A, by allowing many more additional edges in the bipartite graph.
+$A_{1}^{(d-1)}$$A_{2}^{(d-1)}$$A_{3}^{(d-1)}$$A_{4}^{(d-1)}$$A_{5}^{(d-1)}$$A_{1}^{(d)}$$A_{2}^{(d)}$
+[…]
+Figure 3: Graphical illustration of conditions A and C.
+While condition A holds when all dashed arrows are ignored (or equivalently, =0*=0 in the matrix representation), condition C can allow arbitrarily many dashed arrows (where each * can be either zero or one).
 We place our identifiability results in the literature on the identifiability of generative models and latent variable models.
@@
 In statistics, the study of identifiability has a long history but also mainly concerns relatively simple latent structures with only one latent layer (Anderson and Rubin, , 1956; Koopmans and Reiersol, , 1950; Allman et al., , 2009; Xu and Shang, , 2018).
-It has been largely unknown whether DGMs with many nonlinear latent layers can be identifiable for general data as considered in this paper.
+In particular, Gu and Dunson, (2023) establish identifiability of a multi-layer latent structure model, but their result is only applicable to categorical response data, in contrast to the general and rich types of responses considered in this work; moreover, Gu and Dunson, (2023)’s strict identifiability result requires each latent variable to have at least three pure children.
+On the other hand, Lee and Gu, (2024) considered models for general responses which are similar to ours, but their model is restricted to only one latent layer, and they prove identifiability under a strong assumption that the true graph between the latent and observed layers is known.
+Compared to these existing results for related models, our results (i) apply to arbitrary response types as opposed to only categorical responses, (ii) require weaker identifiability conditions compared to Gu and Dunson, (2023) (in terms of both strict and generic identifiability), and (iii) do not require a known graph structure as in Lee and Gu, (2024).
 3.2 Estimation Consistency
@@
 From now on, we slightly abuse notation and let $\mathbf{Y}$ denote the $N\times J$ data matrix including $\mathbf{Y}_{1},\ldots,\mathbf{Y}_{N}$ as rows.
-Using the estimated coefficients in ^$\widehat{\mathcal{B}}=(\widehat{\beta}_{l,k}^{(d)})$, the layer-wise graphical matrices in $\mathcal{G}$ can be estimated by reading off the sparsity pattern of $\widehat{\mathcal{B}}$:
+Using the estimated coefficients in $\widehat{\mathcal{B}}=(\widehat{\beta}_{l,k}^{(d)})$, the layer-wise graphical matrices in $\mathcal{G}$ can be estimated by reading off the sparsity pattern of $\widehat{\mathcal{B}}$:
 $\widehat{g}_{l,k}^{(d)}:=\mathbbm{1}(\widehat{\beta}_{l,k}^{(d)}\neq 0)\quad\text{for all}\quad d\in[D],~k\in[K^{(d)}],~l\in[K^{(d-1)}].$ (7)
-For some tuning parameters $\lambda_{N},\tau_{N}>0$, $p_{\lambda_{N},\tau_{N}}:\mathbb{R}\to[0,\infty)$ is a sparsity-inducing symmetric penalty that is nondecreasing on $[0,\infty)$, nondifferentiable at 00, differentiable at $(0,\tau_{N})$, $p_{\lambda_{N},\tau_{N}}\propto{\lambda_{N}}/{\tau_{N}}$ around 00 and satisfy
-$p_{\lambda_{N},\tau_{N}}(b)=0,~{}\text{if}~{}b=0;\quad p_{\lambda_{N},\tau_{N}}^{\prime}(b)\leq\frac{C\lambda_{N}}{\tau_{N}},~{}\text{if}~{}|b|\leq\tau_{N};\quad p_{\lambda_{N},\tau_{N}}(b)=\lambda_{N},\quad\text{if}~{}|b|\geq\tau_{N}.$
-Note that $\lambda_{N}$ is the magnitude of the penalty, and $\tau_{N}$ is the point of truncation.
+For some tuning parameters $\lambda_{N},\tau_{N}>0$, $p_{\lambda_{N},\tau_{N}}:\mathbb{R}\to[0,\infty)$ is a sparsity-inducing symmetric penalty that satisfies several technical conditions postponed to Supplement S.1.4.
 Our assumption for the penalty $p_{\lambda_{N},\tau_{N}}$ includes common truncated sparsity-inducing penalties such as the TLP (Truncated Lasso Penalty; Shen et al., , 2012) and SCAD (Smoothly Clipped Absolute Deviation; Fan and Li, , 2001).
@@
 Theorem 3.
-[…]
-Remark 1.
-One natural question is to whether our estimator would still be consistent when the number of latent variables are unknown.
-This extension is not straightforward since the number of the top-layer latent variable, $K^{(D)}$, determines the number of deepest mixture components of DDEs.
-Estimating the number of mixture components is a challenging problem even in simple parametric models, and often leads to a slower (than ${1}/{\sqrt{N}}$) rate of convergence in parameter estimation (Goffinet et al.,, 1992; Ho and Nguyen,, 2016).
-Note that for such cases, the Fisher information becomes singular, and Theorem 3 cannot be applied.
-We will give a practical method to choose $\mathcal{K}$ at the end of Section 4.
+Consider a $D$-latent-layer DDE with true parameters $\boldsymbol{\Theta}^{\star},\mathcal{G}^{\star}$ and known $\mathcal{K}$.
+Assume that the model at $\boldsymbol{\Theta}^{\star}$ is identifiable, has a non-singular Fisher information, and all entries of $\{\mathbf{B}^{(d)}\}_{d=1}^{D}$ are bounded.
+Then, the estimator $\widehat{\boldsymbol{\Theta}}$ in (6) is $\sqrt{N}$-consistent in the sense that there exists some $\widetilde{\boldsymbol{\Theta}}\sim_{\mathcal{K}}\widehat{\boldsymbol{\Theta}}$ such that $\lVert\widetilde{\boldsymbol{\Theta}}-\boldsymbol{\Theta}^{\star}\rVert=O_{p}({1}/{\sqrt{N}})$.
+Here, $\|\cdot\|$ denotes the vectorized $L^{2}$ norm.
+Additionally, the graphical matrices are consistently estimated: for $\widetilde{\mathcal{G}}$ resulting from $\widetilde{\boldsymbol{\Theta}}$ according to (7), we have $\mathbb{P}(\widetilde{\mathcal{G}}\neq\mathcal{G}^{\star})\rightarrow 0$.
 4 Scalable Computation Pipeline
-4.1 Warm-up Example: Penalized EM Algorithm
-This subsection first presents a penalized EM algorithm as a warm-up example, before presenting the scalable computational pipelines in Sections 4.2 (stage one) and 4.3 (stage two).
-We describe the methods for two-layer DDEs; they are conceptually easy to generalize to DDEs with more layers.
-Since the coefficients $\mathbf{B}^{(1)}$ and $\mathbf{B}^{(2)}$ are potentially sparse, we propose a penalized EM algorithm (PEM) to compute the penalized maximum likelihood estimator in (6).
-The PEM is an iterative procedure that consists of an expectation step followed by a penalized maximization step (Green,, 1990; Chen et al.,, 2015).
-In the $(t+1)$th iteration, the E-step computes the expectation of the complete data penalized-log-likelihood
-$\ell_{c}(\mathbf{Y},\mathbf{A}^{(1)},\mathbf{A}^{(2)};\boldsymbol{\Theta})-$ $\sum_{d=1}^{2}p_{\lambda_{N},\tau_{N}}(\mathbf{B}^{(d)})=~{}\sum_{i=1}^{N}\Bigg{[}\log\mathbb{P}(\mathbf{A}_{i}^{(2)};{\boldsymbol{p}})+\log\mathbb{P}(\mathbf{A}_{i}^{(1)}\mid\mathbf{A}_{i}^{(2)};\mathbf{B}^{(2)})$
-$+\log\mathbb{P}(\mathbf{Y}_{i}\mid\mathbf{A}_{i}^{(1)},\mathbf{A}_{i}^{(2)};\mathbf{B}^{(1)},\boldsymbol{\gamma})\Bigg{]}-\sum_{d=1}^{2}p_{\lambda_{N},\tau_{N}}(\mathbf{B}^{(d)}).$
-This requires calculating the conditional probability for each latent configuration using the previous parameter estimates; that is, calculating $\mathbb{P}(\mathbf{A}_{i}^{(1)}=\boldsymbol{\alpha}^{(1)},\mathbf{A}_{i}^{(2)}=\boldsymbol{\alpha}^{(2)}\mid\mathbf{Y};\boldsymbol{\Theta}^{[t]})$ for all $i\in[N]$, $\boldsymbol{\alpha}^{(1)}\in\{0,1\}^{K_{1}}$, and $\boldsymbol{\alpha}^{(2)}\in\{0,1\}^{K_{2}}$.
-Here, the superscript $[t]$ denotes the $t$th iteration estimates.
-In the M-step, we update the parameters by maximizing the expectation computed in the E step, which boils down to solving the following three maximizations:
-${\boldsymbol{p}}^{[t+1]}:=$ $\operatorname*{argmax}_{{\boldsymbol{p}}}\sum_{i=1}^{N}\mathbb{E}\left[\log\mathbb{P}(\mathbf{A}_{i}^{(2)};{\boldsymbol{p}});{\boldsymbol{p}}^{[t]}\right],$ (8)
-$\mathbf{B}^{(2),[t+1]}:=$ $\operatorname*{argmax}_{\mathbf{B}^{(2)}}\sum_{i=1}^{N}\mathbb{E}\left[\log\mathbb{P}(\mathbf{A}_{i}^{(1)}\mid\mathbf{A}_{i}^{(2)};\mathbf{B}^{(2)});\mathbf{B}^{(2),[t]}\right]-p_{\lambda_{N},\tau_{N}}(\mathbf{B}^{(2)}),$ (9)
-$(\mathbf{B}^{(1),[t+1]},\boldsymbol{\gamma}^{[t+1]}):=$ $\operatorname*{argmax}_{\mathbf{B}^{(1)},\boldsymbol{\gamma}}\sum_{i=1}^{N}\mathbb{E}\left[\log\mathbb{P}(\mathbf{Y}_{i}\mid\mathbf{A}_{i}^{(1)},\mathbf{A}_{i}^{(2)});\mathbf{B}^{(1),[t]},\boldsymbol{\gamma}^{[t]})\right]-p_{\lambda_{N},\tau_{N}}(\mathbf{B}^{(1)}).$ (10)
-Here, the optimizations for each layer are separated; we update the top-layer proportion parameters ${\boldsymbol{p}}$ in (8), the middle latent layer coefficients $\mathbf{B}^{(2)}$ in (9), and the bottom layer coefficients $(\mathbf{B}^{(1)},\boldsymbol{\gamma})$ in (10).
-Additionally, due to the conditional independence assumption in each layer, the maximizations can be further simplified into low-dimensional optimizations over each row of $\mathbf{B}$.
-Algorithm 1 summarizes the PEM algorithm.
-The details of the simplified M-step and the implementation are in Supplementary Material S.3.1 and S.3.3.
-Data: $\mathbf{Y},\mathcal{K}$, tuning parameters $\lambda_{N},\tau_{N}$.
-Initialize $\boldsymbol{\Theta}^{[0]}$ using the layerwise initialization in Algorithm 2.
-while log-likelihood has not converged do
-       In the $[t+1]$th iteration,
-      // E-step
-      for $(i,\boldsymbol{\alpha}^{(1)},\operatorname{\boldsymbol{\alpha}^{(2)}}2)\in[N]\times\{0,1\}^{K^{(1)}}\times\{0,1\}^{K^{(2)}}$ do
-$\varphi^{[t+1]}_{i,\operatorname{\boldsymbol{\alpha}^{(1)}},\operatorname{\boldsymbol{\alpha}^{(2)}}2}=\mathbb{P}(\mathbf{A}_{i}^{(1)}=\boldsymbol{\alpha}^{(1)},\mathbf{A}_{i}^{(2)}=\boldsymbol{\alpha}^{(2)}\mid\mathbf{Y};\boldsymbol{\Theta}^{[t]})$
-       end for
-      // M-step
-       update $\boldsymbol{\Theta}^{[t+1]}=({\boldsymbol{p}}^{[t+1]},\mathbf{B}^{(1),[t+1]},\mathbf{B}^{(2),[t+1]},\boldsymbol{\gamma}^{[t+1]})$ by solving (8)-(10).
-end while
-Estimate $\mathbf{G}$ based on the sparsity structure of $\widehat{\mathbf{B}}$ according to (7).
-Output : Estimated continuous parameters $\widehat{\boldsymbol{\Theta}}$ and graphical matrices $\widehat{\mathbf{G}}^{(1)}$, $\widehat{\mathbf{G}}^{(2)}$.
-Algorithm 1 Penalized EM (PEM) algorithm for the two-latent-layer DDE
+We present a two-stage scalable computational pipeline to compute the penalized maximum likelihood estimator in (6).
+Our proposed method builds upon the standard penalized EM algorithm (see Supplement S.3.1) by including a spectral initialization and stochastic approximation.
+We separate out each stage into Sections 4.1 (stage one) and 4.2 (stage two).
+For notational simplicity, we describe the proposed method assuming $D=2$ latent layers, which straightforwardly extends to an arbitrary number of latent layers.
 4.1 Stage One: Layerwise Double-SVD Initialization
 The multiple layers of nonlinearity in DDEs lead to a highly nonconvex optimization landscape with potentially exponentially many local optima (Sutskever et al., , 2013).
-If the penalized EM algorithm starts with an initialization close to a local optima, it can get stuck and fail to converge to the global maximizer of the penalized log-likelihood function.
+For example, if the EM algorithm starts with an initialization close to a local optima, it can get stuck and fail to converge to the global maximizer of the penalized log-likelihood function.
 Hence, for highly complex latent variable models such as DDEs, it is critical to initialize the optimization algorithm wisely.
@@
 4.2 Stage Two: Penalized SAEM Algorithm
-For DDEs with a large number of latent variables, the E-step in Algorithm 1 requires computing all conditional probabilities $\mathbb{P}(\mathbf{A}^{(1)}_{i}=\boldsymbol{\alpha}^{(1)},\mathbf{A}^{(2)}_{i}=\boldsymbol{\alpha}^{(2)}\mid\mathbf{Y};\boldsymbol{\Theta}^{[t]})$ for all $\boldsymbol{\alpha}^{(1)}\in\{0,1\}^{K^{(1)}}$ and $\boldsymbol{\alpha}^{(2)}\in\{0,1\}^{K^{(2)}}$.
+For DDEs with a large number of latent variables, the E-step in standard EM algorithms (see Supplement S.3.1) computes all conditional probabilities $\mathbb{P}(\mathbf{A}^{(1)}_{i}=\boldsymbol{\alpha}^{(1)},\mathbf{A}^{(2)}_{i}=\boldsymbol{\alpha}^{(2)}\mid\mathbf{Y};\boldsymbol{\Theta}^{[t]})$ for all $\boldsymbol{\alpha}^{(1)}\in\{0,1\}^{K^{(1)}}$ and $\boldsymbol{\alpha}^{(2)}\in\{0,1\}^{K^{(2)}}$.
 This requires computing and storing $O(N\times 2^{\sum_{d=1}^{D}K^{(d)}})$ terms.
@@
 Therefore, we propose a penalized Stochastic Approximate EM (SAEM; see Delyon et al., , 1999; Kuhn and Lavielle, , 2004) by modifying both the E-step and M-step to more scalable versions using approximate sampling.
-As we illustrate below, this is a method with linear dependence of $\sum_{d=1}^{D}K^{(d)}$ on the computing time as well as memory, and is a scalable alternative to Algorithm 1.
-We elaborate on the details on how to modify the PEM to derive the SAEM.
+As we illustrate below, this is a method with linear dependence of $\sum_{d=1}^{D}K^{(d)}$ on the computing time as well as memory.
+We elaborate on the details on deriving the SAEM.
 First, we replace the E-step to a simulation step, which consists of simulating only a small number (denoted as $C$) of posterior samples of the latent variables.
 As exact sampling from the joint distribution $\mathbb{P}(\mathbf{A}_{i}^{(1)},\mathbf{A}_{i}^{(2)}\mid\mathbf{Y};\boldsymbol{\Theta}^{[t]})$ is expensive, we sample each latent variable from their complete conditionals.
-That is, we sample each $A_{i,k}^{(1),[t+1]}$ from $\mathbb{P}(A_{i,k}^{(1)}\mid(-),\boldsymbol{\Theta}^{[t]})$, where ()(-)( - ) denotes the estimates of all other latent variables from the $t$th iteration.
+That is, we sample each $A_{i,k}^{(1),[t+1]}$ from $\mathbb{P}(A_{i,k}^{(1)}\mid(-),\boldsymbol{\Theta}^{[t]})$, where ()(-) denotes the estimates of all other latent variables from the $t$th iteration.
 Since the latent variables are binary, the conditional distribution is Bernoulli and easy to evaluate.
 Consequently, the computationally expensive E-step is replaced by the simulation step, which computes and stores only $O(N\times\sum_{d=1}^{D}K^{(d)})$ terms.
-In terms of choosing the number of samples $C$ in each iteration, we empirically find that taking $C=1$ is computationally efficient without sacrificing much accuracy.
-Additional simulation studies in Supplementary Material S.4.2 demonstrate that taking a larger $C$ results in only little improvement in accuracy, but becomes significantly slower.
+In terms of choosing the number of samples $C$ in each iteration, we empirically find that taking $C=1$ is computationally efficient without sacrificing much accuracy (see Supplement S.4.4).
 This choice of $C=1$ was also suggested in the original paper that proposed the SAEM (Delyon et al., , 1999).
-Second, we modify the M-step objective function by replacing the exact conditional probability values (denoted as $\varphi^{[t+1]}$ in Algorithm 1) to sample-based quantities, and also stochastically averaging the objective functions.
-For example, we modify (9) as follows:
+Second, we modify the standard M-step objective function (expected complete data log-likelihood) by (i) replacing the exact conditional probability values to sample-based quantities, (ii) stochastically averaging the objective functions, and (iii) including the sparsity-inducing penalties in (6).
+The M-step objective for updating $\mathbf{B}^{(2)}$ is
 $Q^{(2),[t+1]}(\mathbf{B}^{(2)})$
@@
  // Simulation-step
-      for $(i,k)\in[N]\times[K^{(1)}]$ do
-             Sample each $A_{i,k}^{(1),[t+1]}$ from the complete conditional $\mathbb{P}(A_{i,k}^{(1)}\mid(-))$ using the previous parameter estimates $\boldsymbol{\Theta}^{[t]},\mathbf{A}^{[t]}$
-      for $(i,l)\in[N]\times[K^{(2)}]$ do
-             Sample each $A_{i,l}^{(2),[t+1]}$ from the complete conditional $\mathbb{P}(A_{i,l}^{(2)}\mid(-))$ using the previous parameter estimates $\boldsymbol{\Theta}^{[t]},\mathbf{A}^{[t]}$
+ Sample each $A_{i,k}^{(1),[t+1]},A_{i,l}^{(2),[t+1]}$ from the complete conditionals using the previous parameter estimates $\boldsymbol{\Theta}^{[t]},\mathbf{A}^{[t]}$
   // Stochastic approximation M-step
@@
 Algorithm 2 Penalized SAEM algorithm for the two-latent-layer DDE
-We similarly modify the optimizations in (8)–(10).
-Algorithm 3 summarizes our final SAEM algorithm with $C=1$, where exact formulas are deferred to Supplement S.3.1.
-Similar to the PEM algorithm, the M-step updates can be written in terms of low-dimensional maximizations for each row of the coefficient matrices.
-Selecting the latent dimension $\mathcal{K}$.
-To apply the above computation pipeline to real data, one additionally needs to specify the number of latent variables, $\mathcal{K}$.
-For this purpose, we propose a layer-wise estimation strategy, which can be incorporated into our initialization procedure in Algorithm 2.
+Data: $\widehat{\mathbf{A}}^{(0)}:=\mathbf{Y}_{N\times J},\widehat{K}^{(0)}:=J$, the number of latent layers $D$
+For each $1\leq d\leq D$, repeat:
+  1. 1.
+    Let $\mathfrak{K}^{(d)}:=\{\lceil\widehat{K}^{(d-1)}/4\rceil,\ldots,\lfloor\widehat{K}^{(d-1)}/2\rfloor\}$ be the candidate grid for $\widehat{K}^{(d)}$.
+  2. 2.
+    Define $\widehat{K}^{(d)}$ based on the largest spectral ratio of the denoised/linearized $\widehat{\mathbf{A}}^{(d-1)}$ (see step 1 in Algorithm 1):
+    $\widehat{K}^{(d)}:=\operatorname*{argmax}_{k\in\mathfrak{K}^{(d)}}\sigma_{k}/\sigma_{k+1}-1.$
+  3. 3.
+    Given $\widehat{K}^{(d)}$, estimate the $d$-th layer latent variables $\widehat{\mathbf{A}}^{(d)}_{N\times\widehat{K}^{(d)}}$ using Algorithm 1.
+Output : Estimated latent dimensions for all layers $\widehat{K}^{(1)},\ldots,\widehat{K}^{(D)}$.
+Algorithm 3 Practical methods for selecting the latent dimensions $\mathcal{K}$
+Algorithm 2 summarizes our proposed SAEM algorithm with $C=1$, where detailed formulas are deferred to Supplement S.3.2.
+Here, the detailed M-step updates can be written in terms of low-dimensional maximizations for each row of the coefficient matrices.
+Selecting the latent dimensions.
+To apply the above computational pipeline to real data, one also needs to specify the number of latent variables, $\mathcal{K}$.
+We propose a layer-wise estimation strategy in Algorithm 3, which can be incorporated into our initialization procedure in Algorithm 1.
 Recall the denoised data matrix $\widehat{\mathbf{Z}}$ from Step 1 of Algorithm 1 and let $\sigma_{1},\sigma_{2},\ldots$ be its singular values in the descending order.
-Given a candidate grid $\mathfrak{K}$, we estimate the number of latent variables in the first layer based on the largest spectral ratio: $\widehat{K}^{(1)}:=\operatorname*{argmax}_{k\in\mathfrak{K}}({\sigma_{k}}/{\sigma_{k+1}})-1$.
+Given a candidate grid $\mathfrak{K}^{(1)}$, we estimate the size of the first latent layer based on the largest spectral ratio: $\widehat{K}^{(1)}:=\operatorname*{argmax}_{k\in\mathfrak{K}}({\sigma_{k}}/{\sigma_{k+1}})-1$.
 Now, given $\widehat{K}^{(1)}$, we proceed with the remaining steps of Algorithm 1 to estimate the first-layer latent variables $\mathbf{A}^{(1)}$.
-By treating the estimated $\mathbf{A}^{(1)}$ as the observed variables of a one-latent-layer DDE, we can repeat the above procedure to estimate $\widehat{K}^{(2)}$.
-See Supplement S.3.2 for alternative selection criteria for $\mathcal{K}$.
+Inductively treating the estimated $\widehat{\mathbf{A}}^{(d)}$ as the observed variables of a one-latent-layer DDE, we can repeat the above procedure to estimate $\widehat{K}^{(d+1)}$.
+See Supplement S.3.3 for alternative selection criteria for $\mathcal{K}$.
 5 Simulation Studies
-We conduct extensive simulation studies in various settings to assess the performance of the proposed computation pipeline (Algorithms 2 and 3) and validate our identifiability conditions (in Section 3.1).
-We generate data from two-latent-layer DDEs, exploring a total of 90 true settings by varying the following:
+We conduct extensive simulation studies in various settings to assess the performance of the proposed computation pipeline (Algorithms 1 and 2) and validate our identifiability conditions (in Section 3.1).
+Two-latent-layer DDEs with general response types.
+First, we generate data from two-latent-layer DDEs, exploring a total of 90 true settings by varying the following:
   1. (a)
@@
   3. (c)
-    two parameter values: see (13) below,
+    two parameter values: see Supplement S.4.1,
   4. (d)
     five varying sample sizes: $N=500,1000,2000,4000,8000$.
-Given a value of $K^{(2)}$, set $K^{(1)}=3K^{(2)}$ and $J=9K^{(2)}$, which allows a large latent dimension with as many as $K^{(1)}=30$ binary latent variables in the shallowest latent layer.
-In terms of the true parameter values, we consider two sets of values based on the strict and generic identifiability conditions in Theorems 1 and 2, respectively.
-In particular, we define $\mathcal{B}_{s}=\{\mathbf{B}_{s}^{(d)}\}_{d=1,2}$ that satisfy Theorem 1 as follows:
-$\mathbf{B}_{s}^{(d)}=\begin{pmatrix}-2\mathbf{1}_{K^{(d)}}&4\mathbf{I}_{K^{(d)}}\\ -4\mathbf{1}_{K^{(d)}}&4\mathbf{I}_{K^{(d)}}\\ -2\mathbf{1}_{K^{(d)}}&\mathbf{B}_{1}^{(d)}\end{pmatrix},\quad\text{where}\quad\beta_{1;j,k}^{(d)}:=\begin{cases}4&\text{if }k=j,\\ 4/3&\text{ else if }|k-j|=K^{(d)}/2,\\ 0&\text{otherwise}.\end{cases}$ (13)
-We also consider $\mathcal{B}_{g}=\{\mathbf{B}_{g}^{(d)}\}_{d=1,2}$ that satisfy Theorem 2, whose exact values are displayed in Supplement S.4.1.
-We set the top-layer proportion parameters as $p_{k}=0.5$ for all $k\in[K^{(2)}]$, and set the dispersion parameters for the Normal distribution as $\gamma_{j}=\sigma_{j}^{2}=1$ for all $j\in[J]$.
-Figure 3: Estimation error for $\mathcal{G}$ and $\boldsymbol{\Theta}$ under the two-latent-layer DDE with parameters $\mathcal{B}_{s}$ and various observed-layer parametric families.
+Here, we consider two sets of parameter values that each satisfy the strict and generic identifiability conditions in Theorems 1 and 2.
+Regarding the parameter dimensions, given a value of $K^{(2)}$, we set $K^{(1)}=3K^{(2)}$ and $J=9K^{(2)}$.
+This allows a large latent dimension with as many as $K^{(1)}=30$ binary latent variables in the shallowest latent layer.
+Figure 4: Estimation error for $\mathcal{G}$ and $\boldsymbol{\Theta}$ under the two-latent-layer DDE with strictly identifiable true parameters and various observed-layer parametric families.
 For each scenario, we run $100$ independent simulations, and display the estimation results in Figure 4.
-The results under the generic identifiable parameters $\mathcal{B}_{g}$ and computation times are included in Supplementary Material S.4.1.
+The results under the generic identifiable parameters and computation times are included in Supplement S.4.3.
 We measure the estimation accuracy of the graphical matrices $\mathcal{G}$ by computing the average entrywise accuracy.
@@
 Additionally, by comparing the estimation accuracy across different parametric families, we observe that the Bernoulli is the most challenging to estimate, and the Normal is the easiest.
-Regarding the true parameters, the sparse model with coefficients $\mathcal{B}_{s}$ is easier to estimate.
-This is because sparsity helps the initialization and reduces the effective parameter dimension.
-Finally, under coefficients $\mathcal{B}_{s}$, estimation accuracy improves as the latent and observed dimensions increase, which is a consequence of the decreasing percentage of nonzero entries in $\mathcal{B}$.
-In Supplementary Material S.4.3, we present simulation results where the latent dimension $\mathcal{K}$ is unknown.
-We illustrate that the spectral-gap estimator has near-perfect selection accuracy for a large $N$ (e.g., larger than $4000$), and is superior in terms of both accuracy and computation time compared to alternative estimation strategies.
+Simulation studies under deeper models.
+We assess the scalability of our proposed method for deeper models (with $D=3,4,5$ latent layers) with potentially high-dimensional responses (with the observed data dimension set to $J=54,162,486$, respectively).
+For all settings, we consider Normal observations with $K^{(D)}=2$ latent variables for the top (deepest) layer and $K^{(d-1)}=3K^{(d)}$ variables for the $d$th layer for each $d=D,\ldots,2$.
+We summarize the estimation accuracy of the graph structure in each layer and the average runtime for DDEs with $D=5$ latent layers in Table 1.
+The results for $D=3,4$ are displayed in Supplement S.4.6.
+The computation time illustrates that our proposed method is scalable to deeper models.
+More specifically, the $D=3,4$ case only took a few minutes, and the most challenging case with $D=5$ and $N=16,000$ also took less than an hour on a personal laptop.
+In terms of estimation accuracy of graph structures $\mathcal{G}$, the accuracy is higher for shallower layers than for deeper layers.
+This results from the accumulation of uncertainty for deeper layers, which is inevitable as each layer consists of stochastic latent variables.
+Note that even for such deeper models with $D=4,5$ latent layers, the shallower structures $\mathbf{G}^{(1)},\mathbf{G}^{(2)}$ are recovered with high accuracy.
+This indicates that one can include additional latent layers to increase the models’ representational power, without sacrificing the accuracy of learning graph structures closer to the data layer.
+[…]
+$\mathbf{G}^{(1)}$ 1.00 1.00 1.00 1.00 1.00 1.00
+$\mathbf{G}^{(2)}$ 0.81 0.90 0.95 0.97 0.98 0.98
+$\mathbf{G}^{(3)}$ 0.77 0.83 0.86 0.88 0.90 0.92
+$\mathbf{G}^{(4)}$ 0.61 0.62 0.65 0.67 0.66 0.74
+$\mathbf{G}^{(5)}$ 0.60 0.59 0.62 0.62 0.65 0.64
+runtime (s) 220 259 326 563 893 2925
+Table 1: Average entrywise-accuracy of estimating the graphical matrices and runtime in seconds for DDEs with $D=5$ latent layers.
+Ablation studies.
+As our model closely resemble DBNs, we mainly compare DDEs with DBNs.
+For fair comparison, we considered two settings: (a) data generated from a Berounlli-DDE, and (b) data generated from a DBN with identical coefficients.
+The results in Figure 5 demonstrate that our proposed algorithm has better estimation accuracy for both the graphical structure and the continuous parameters under both well-specified and mis-specified settings, even when data are generated from a DBN with undirected edges in the top layer.
+This is because the DBN algorithm does not learn sparse coefficients, and also fundamentally suffers from local optima due to a random initialization.
+Note that our identifiability results in Propositions S.1-2 immediately guarantee the identifiability of DBNs, and theoretically justify the above positive results of our algorithm.
+The ablation studies also indicate that the current implementations of DDEs are slower than DBNs.
+We believe that our algorithm can be further scaled up by replacing the M-step in the SAEM algorithm to first-order optimization methods (e.g.
+gradient ascent or stochastic gradient ascent), which we leave for future work.
+Figure 5: Comparison of Bernoulli-DDE vs DBN.
+True data are generated from a DDE in (a)–(b), and from a DBN in (c)–(d).
+Red and blue lines show results obtained by the DBN algorithm and our DDE algorithm, respectively.
+(a) and (c): estimation error for the graphs $\mathcal{G}$.
+(b) and (d): estimation error for continuous parameters $\Theta$.
+Smaller values are better.
+Additional simulation results.
+In Supplement S.4.4 and S.4.6, we present additional simulation results where the latent dimension $\mathcal{K}$ is unknown.
+We illustrate that the spectral-gap estimator has near-perfect selection accuracy for a large $N$ (e.g., larger than $4000$), and is superior in terms of both accuracy and computation time compared to alternatives.
 6 Real Data Applications
@@
 Negative part [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Table 1: From left to right: Estimated basis images ${\boldsymbol{\beta}}_{0,0}^{(1)}$, ${\boldsymbol{\beta}}_{0,1}^{(1)}$, ${\boldsymbol{\beta}}_{0,2}^{(1)}$, ${\boldsymbol{\beta}}_{0,3}^{(1)}$, ${\boldsymbol{\beta}}_{0,4}^{(1)}$, ${\boldsymbol{\beta}}_{0,5}^{(1)}$ from the MNIST data, which define sparse subregions in the $28\times 28$ image.
-The second and third row shows the most significant portion of the basis images by thresholding at the value ±$\pm 1.5$.
-We fit the two-latent-layer DDE using Algorithm 1 with the latent dimensions set to $K^{(1)}=5$ and $K^{(2)}=2$ (see the Supplementary Material S.5.2 for the rationale for this choice).
+${\boldsymbol{\beta}}_{0,0}^{(1)}$ ${\boldsymbol{\beta}}_{0,1}^{(1)}$ ${\boldsymbol{\beta}}_{0,2}^{(1)}$ ${\boldsymbol{\beta}}_{0,3}^{(1)}$ ${\boldsymbol{\beta}}_{0,4}^{(1)}$ ${\boldsymbol{\beta}}_{0,5}^{(1)}$
+Table 2: From left to right: Estimated basis images (reshaped from the estimated coefficients ${\boldsymbol{\beta}}_{0,k}^{(1)}$) from the MNIST data, which define sparse subregions in the $28\times 28$ image.
+The second and third row shows the most significant positive and negative parts of the basis images by thresholding at the value $\pm 1.5$.
+We fit the two-latent-layer DDE with the latent dimensions set to $K^{(1)}=5$ and $K^{(2)}=2$ (see the Supplementary Material S.5.2 for the rationale for this choice).
 Table 2 displays the first-layer coefficients $\mathbf{B}^{(1)}$ by rearranging the each column into the original $28\times 28$ grid.
@@
 We first estimate the latent variables using the $\varphi$ matrix in the EM algorithm:
-$(\widehat{\mathbf{A}}^{(1)}_{i},\widehat{\mathbf{A}}^{(2)}_{i})=\operatorname*{argmax}_{\boldsymbol{\alpha}^{(1)}\in\{0,1\}^{K_{1}},~{}\boldsymbol{\alpha}^{(2)}\in\{0,1\}^{K_{2}}}\varphi_{i,\boldsymbol{\alpha}^{(1)},\boldsymbol{\alpha}^{(2)}};\quad\boldsymbol{\alpha}^{(\ell)}\in\{0,1\}^{K_{\ell}},~{}\ell=1,2.$ (14)
+$(\widehat{\mathbf{A}}^{(1)}_{i},\widehat{\mathbf{A}}^{(2)}_{i})=\operatorname*{argmax}_{\boldsymbol{\alpha}^{(1)}\in\{0,1\}^{K_{1}},~\boldsymbol{\alpha}^{(2)}\in\{0,1\}^{K_{2}}}\varphi_{i,\boldsymbol{\alpha}^{(1)},\boldsymbol{\alpha}^{(2)}};\quad\boldsymbol{\alpha}^{(d)}\in\{0,1\}^{K^{(d)}},~d=1,2.$ (10)
 Then, a decision tree that classifies the binary latent representations to the categorical label is built by using the misclassification error as the splitting criteria; see the left panel of Figure 6.
@@
 Although the state-of-art machine learning methods can achieve an accuracy as high as $97\%$ (Monnier et al., , 2020), we point out that the main goal here is not classification, but on interpretability and parsimony; indeed, the DDE uses more limited information and is a less complex but more interpretable model that provides the generative process for the image.
-In Table 3, we compare the classification accuracy and pixel-wise reconstruction accuracy of the two-latent-layer DDE with alternative interpretable models, such as the latent class model (LCM) and the one-latent-layer DDE.
-The results show that the two-latent-layer DDE performs the best.
-We also display example generated images alongside their latent configurations in Table 2, which illustrates various handwriting styles for each digit.
-Supplementary Material S.5.4 additionally displays sample reconstructed images based on these three methods, and illustrates that the LCM and the one-latent-layer DDE generate noisier and less flexible images than the two-latent-layer DDE.
-$[1,0,1,0,1]$ $[0,0,0,0,0]$ $[0,1,0,1,1]$ $[0,0,1,1,1]$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] $[1,0,1,1,1]$ $[0,0,0,1,0]$ $[0,1,1,0,1]$ $[0,0,1,1,0]$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Table 2: Example generated images from the two-latent-layer DDE and their latent representations.
-Accuracy LCM 1-DDE 2-DDE Train classif.
-(%) 89.5 84.8 92.0 Test classif.
-(%) 90.9 86.2 92.6 Train recon.
-(%) 48.2 79.5 79.6 Test recon.
-(%) 48.5 79.1 79.9 Table 3: Classification accuracy and pixel-wise reconstruction accuracy for the MNIST dataset.
+In Table 4, we compare the classification accuracy and pixel-wise reconstruction accuracy of the two-latent-layer DDE with alternative interpretable models (latent class model (LCM) and the one-latent-layer DDE), as well as popular unsupervised machine learning models (two-latent-layer DBN and VAE).
+The results show that the two-latent-layer DDE performs well for both measures.
+We also display example images generated from DDE alongside their latent representations in Table 3, which illustrate various handwriting styles for each digit.
+We provide implementation details, detailed comparison with iVAEs, and additional visualization in Supplementary Material S.5.3.
+$[1,0,1,0,1]$ $[0,0,0,0,0]$ $[0,1,0,1,1]$ $[0,0,1,1,1]$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] $[1,0,1,1,1]$ $[0,0,0,1,0]$ $[0,1,1,0,1]$ $[0,0,1,1,0]$ [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Table 3: Example images generated from DDE and their latent representations.
+Accuracy LCM 1-DDE 2-DDE 2-DBN VAE Train classif.
+(%) 89.5 84.8 92.0 89.3 97.1 Test classif.
+(%) 90.9 86.2 92.6 90.6 92.0 Train recon.
+(%) 48.2 79.5 79.6 76.4 82.5 Test recon.
+(%) 48.5 79.1 79.9 77.0 82.7 Table 4: Classification accuracy and pixel-wise reconstruction accuracy for the MNIST dataset.
 6.2 Count Data: Poisson-DDE Hierarchical Topic Modeling
@@
 While many of the existing works assume a tree-structured hierarchy, DDEs flexibly allow multiple parents for each variable.
-We analyze the text corpus from the 20 newsgroups dataset (Lang,, 1995), which was previously analyzed by other topic models with binary latent variables (Srivastava et al.,, 2013; Gan et al.,, 2015).
-The training set of the corpus consists of approximately 11,000 documents and 2,000 words, and is based on the email messages across 20 newsgroups.
-After preprocessing by focusing on 12 newsgroups and removing stop words and infrequent words, the dataset consists of $N=5,883$ documents and $J=653$ words.
-While not used for fitting the DDE, the label (newsgroup) of each document is also provided in the dataset, alongside three larger categories: recreation, computer, and science.
-Y1.0e0]ptrecreationtechnologycarssportsmotorcyclesoftwarelocation/nameshardwaregraphicsspacegeb bike banks riding pitt dod bmw gordonhockeynhlplayoffsteamsplayersleagueseasonbaseballenginecheappartsturnedmodelsdealermarketcarfilesdirectoryfilemscopydeletedsizecharlespeter ny chicagodevicesperiodyorkthomasusamotherboard dx CPU mhz controller board ram cardsmouse motif lines color display screen window imageorbit earth research science nasa development project institute
+Y1.0e0]ptrecreationcarssportsmotorcyclesoftwarelocation/nameshardwaregraphicsspacetechnologygeb bike banks riding pitt dod bmw gordonhockeynhlplayoffsteamsplayersleagueseasonbaseballenginecheappartsturnedmodelsdealermarketcarfilesdirectoryfilemscopydeletedsizecharlespeter ny chicagodevicesperiodyorkthomasusamotherboard dx CPU mhz controller board ram cardsmouse motif lines color display screen window imageorbit earth research science nasa development project institute
 Figure 7: DDE estimated from the 20 newsgroups dataset.
@@
 The width of the upper layer arrows is proportional to the corresponding coefficients and the red arrow indicates negative values.
-We fit the two-latent-layer DDE with a Poisson-distributed data layer (Poisson-DDE) $Y_{j}\mid(\mathbf{A}^{(1)}=\boldsymbol{\alpha}^{(1)})\sim\text{Poi}\big{(}\exp\big{[}\beta_{j,0}^{(1)}+\sum_{k\in[K^{(1)}]}\beta_{j,k}^{(1)}\alpha_{k}^{(1)}\big{]}\big{)}$, and display the estimated latent structure in Figure 5.
-We set $K^{(1)}=8$ based on the spectral ratio estimator and $K^{(2)}=2$ based on the identifiability conditions and better interpretability.
+We analyze the text corpus from the 20 newsgroups dataset (Lang, , 1995), which was previously analyzed by other topic models with binary latent variables (Srivastava et al., , 2013; Gan et al., , 2015).
+After preprocessing and focusing on 12 newsgroups, the dataset consists of $N=5,883$ documents and $J=653$ words.
+We fit the two-latent-layer DDE with a Poisson-distributed data layer (Poisson-DDE) $Y_{j}\mid(\mathbf{A}^{(1)}=\boldsymbol{\alpha}^{(1)})\sim\text{Poi}\big{(}\exp\big{[}\beta_{j,0}^{(1)}+\sum_{k\in[K^{(1)}]}\beta_{j,k}^{(1)}\alpha_{k}^{(1)}\big{]}\big{)}$ and latent dimensions $K^{(1)}=8,K^{(2)}=2$, and display the estimated latent structure in Figure 7.
 Additional details behind this choice are given in Supplementary Material S.5.2.
-Our choice of $K^{(2)}=2$ is also consistent with the correlation structure discovered in Gan et al., (2015).
 To better interpret individual latent variables, we define representative words for each topic $k$ based on the discrepancy between $\beta_{j,k}$ and all other coefficients, $(\beta_{j,l})_{l\neq k}$.
@@
 Here, each latent variable is named based on the representative words and the held-out newsgroup categories.
-Compared to the held-out tree structure of the 12 newsgroup labels (see Figure S.7 in the Supplement), one can see that the DDE has discovered a lower-dimensional structure in Figure 5.
+Compared to the held-out tree structure of the 12 newsgroup labels (see Figure S.7 in the Supplement), the DDE discovered a lower-dimensional structure in Figure 7.
 The latent structure in DDEs allow multiple parents for each topic and effectively model the complex label dependence.
@@
 We consider the following three metrics widely used in topic modeling to measure different aspects of fit (Chen et al., , 2023).
-The first measure is the perplexity, measuring the predictive likelihood of the words in the held-out set.
-While there are multiple notions of perplexities, we choose the definition that was adopted in similar Poisson-based models (Zhou et al.,, 2012; Gan et al.,, 2015; Ranganath et al.,, 2015, see Supplementary Material S.5.3 for more details).
-In addition to the train perplexity, we also display the test perplexity.
-The second measure is the average negative coherence, measuring the quality within each topic by computing $-\frac{1}{K^{(1)}}\sum_{k=1}^{K^{(1)}}\sum_{v_{1},v_{2}\in V_{k}}\log\left({(\text{freq}(v_{1},v_{2})+1)}/{\text{freq}(v_{2})}\right),$ where $V_{k}$ is the top $15$ representative words for the $k$th topic, and the function “freq” counts the number of documents containing the words specified in the input argument.
-The third measure is the similarity, computing the number of overlapping words within representative words across different topics: $\sum_{1\leq k_{1}<k_{2}\leq K^{(1)}}\sum_{v_{1}\in V_{k_{1}},v_{2}\in V_{k_{2}}}I(v_{1}=v_{2}).$ For all three measures, smaller values are better.
+The first is the perplexity, measuring the predictive likelihood of the words in the held-out set.
+The second is the average negative coherence, measuring the quality within each topic by computing $\frac{1}{K^{(1)}}\sum_{k=1}^{K^{(1)}}\sum_{v_{1},v_{2}\in V_{k}}\log\left({(\text{freq}(v_{1},v_{2})+1)}/{\text{freq}(v_{2})}\right),$ where $V_{k}$ is the top $15$ representative words for the $k$th topic, and the function “freq” counts the number of documents containing the words specified in the input argument.
+The third is the similarity, computing the number of overlapping representative words across different topics: $\sum_{1\leq k_{1}<k_{2}\leq K^{(1)}}\sum_{v_{1}\in V_{k_{1}},v_{2}\in V_{k_{2}}}I(v_{1}=v_{2}).$ For all three measures, smaller values are better.
 Table 5 summarizes the results and shows the promising fit of DDE.
@@
 We have also fit models with larger dimensions for comparison, by considering LDA with $256$ latent variables, and DPFA with $K^{(1)}=128,K^{(2)}=64$ latent variables.
-The dimension for LDA is motivated by that the DDE has $2562^{8}=$ 256 mixture components; while for DPFA, this is the same latent dimension specified in the original paper (Gan et al.,, 2015).
+The dimension for LDA is motivated by that the DDE has $2^{8}=256$ mixture components; while for DPFA, this is the same latent dimension specified in the original paper (Gan et al., , 2015).
 We can see that while considering a larger latent dimension may help in terms of perplexity, this leads to a loss of the within-topic coherence as well as dilutes the boundary of each topic, and hence gives less interpretable results.
 Model Dimension $\mathcal{K}$ Train perplexity Test perplexity Neg. coherence Similarity
-LDA $8$ 499 512 276 43
+LDA 88 499 512 276 43
 LDA $256$ 269 515 321 (41450)
-DPFA $8-2$ 289 499 211 5
-DPFA $128-64$ 175 232 280 (3378)
-DDE $8-2$ 322 398 270 0
-DDE $8-3$ 322 399 275 0
+DPFA 828-2 289 499 211 5
+DPFA 12864128-64 175 232 280 (3378)
+DDE 828-2 322 398 270 0
+DDE 838-3 322 399 275 0
 Table 5: Train and test perplexity scores of different models on the 20 Newsgroups dataset.
@@
 We parenthesize similarity scores for the models with different dimensions, as the measure is not normalized.
-6.3 Continuous Data: Lognormal-DDE for Response Times
-We apply the DDE to a response time dataset from the Trends in International Mathematics and Science Study (TIMSS) (Fishbein et al.,, 2021).
-We analyze the eighth-grade students’ response time for an internet-based mathematics assessment.
-After preprocessing, the dataset consists of $N=526$ students’ response times for $J=29$ questions.
-Here, we fit the two-latent-layer DDE where a lognormal distribution is used to model the continuous positive response times (Lognormal-DDE): $Y_{j}\mid(\mathbf{A}^{(1)}=\boldsymbol{\alpha}^{(1)})\sim\text{lognormal}\big{(}\beta_{j,0}^{(1)}+\sum_{k\in[K^{(1)}]}\beta_{j,k}^{(1)}\alpha_{k}^{(1)},\gamma_{j}\big{)}$.
-This dataset includes additional information regarding the latent structure $\mathbf{G}^{(1)}$, which is the so-called $Q$-matrix in the cognitive diagnostic modeling literature (von Davier,, 2008; Lee and Gu,, 2024).
-The provisional $\mathbf{G}^{(1)}$ matrix specifies $K^{(1)}=7$ latent cognitive skills (Number, Algebra, Geometry, Data and Probability, Knowing, Applying, and Reasoning) as well as the skills that are required to solve each item.
-That is, $g^{(1)}_{j,k}=1$ if item $j$ requires latent skill $k$ to solve it.
-Hence, following the confirmatory latent variable modeling convention in psychometrics, we estimate the DDE parameters by fixing $\mathbf{G}^{(1)}$ to this given structure.
+6.3 Multimodal Educational Data: Bernoulli-Lognormal-DDE
+We apply the DDE to an educational assessment dataset from the Trends in International Mathematics and Science Study (TIMSS) (Fishbein et al., , 2021).
+We analyze the eighth-grade students’ responses for an internet-based mathematics assessment.
+As the assessment is electronically conducted, multiple modalities of information are recorded.
+Here, we focus on two important modalities: binary response accuracy and continuous response time.
+For each individual student, our data consists of response accuracy (whether the student gave a correct answer) and response time (how long the student took) for each of the $J/2=29$ items.
+We use the same latent variables to model both data modalities and model the binary response accuracy via Bernoulli distributions, and model the continuous positive response times via Lognormal distributions: $Y_{j}\mid(\mathbf{A}^{(1)}=\boldsymbol{\alpha}^{(1)})\sim\text{lognormal}\big{(}\beta_{j,0}^{(1)}+\sum_{k\in[K^{(1)}]}\beta_{j,k}^{(1)}\alpha_{k}^{(1)},\gamma_{j}\big{)}$.
 Response \ Latent skill $A_{1}^{(2)}$ $A_{1}^{(1)}$: Number $A_{2}^{(1)}$: Algebra $A_{3}^{(1)}$: Geometry $A_{4}^{(1)}$: Data and Prob
-Agree a lot 0.83 0.72 0.88 0.86 0.87
-Agree a little 0.80 0.72 0.85 0.84 0.85
-Disagree a little 0.76 0.74 0.85 0.84 0.84
-Disagree a lot 0.69 0.67 0.76 0.75 0.78
+Agree a lot 0.62 0.59 0.62 0.42 0.57
+Agree a little 0.50 0.47 0.52 0.34 0.43
+Disagree a little 0.37 0.33 0.38 0.31 0.37
+Disagree a lot 0.29 0.29 0.32 0.22 0.29
 Table 6: Average latent variable estimate for each response category for the question “Mathematics is one of my favorite subjects”.
-We fit the two-latent-layer DDE with $K^{(1)}=7,~{}K^{(2)}=1$, and estimate the latent skills based on the posterior probability using (14).
+We fit the two-latent-layer DDE with $K^{(1)}=7,~K^{(2)}=1$, and estimate the latent skills based on the posterior probability using (10) (see Supplement S.5.2 for details).
 We compare the estimated latent variables with the held-out information of each student’s categorical response to a survey question: “Mathematics is one of my favorite subjects”, and display the results in Table 6.
-The first column shows that the higher-order latent variable, $A_{1}^{(2)}$, is highly correlated with the magnitude that students like math.
-This suggests that $A_{1}^{(2)}$ can be interpreted as a general indicator for the students’ interest in math, while the fine-grained latent variables $\mathbf{A}^{(1)}$ represent more specific skill mastery for each domain.
+The first column shows that the higher-order latent variable, $A_{1}^{(2)}$, is highly correlated with the extent that students like math.
+This suggests that $A_{1}^{(2)}$ can be interpreted as a general indicator of the students’ interest in math, while the fine-grained latent variables $\mathbf{A}^{(1)}$ represent students’ specific skill mastery profiles.
 In addition, we observe that the students who enjoy math tend to have a higher probability of mastering specific skills as well.
-This is coherent with the fact that the estimated $\mathbf{B}^{(1)}$-coefficients are all positive.
+This is coherent with the fact that the estimated $\mathbf{B}^{(1)}$-coefficients are nonnegative for both modes.
+In Supplement S.5.5, we also illustrate a strong correlation of the estimated intercepts for both data modalities, compared to another held-out information (whether or not the items are multiple choice questions).
 7 Discussion
-This paper makes contributions to core AI problems from statisticians’ perspectives by proposing a broad family of interpretable DGMs with solid identifiability guarantees, scalable computation pipelines, and promising application potential.
+This paper makes significant contributions to core AI problems from statisticians’ perspective by proposing a broad family of interpretable DGMs with solid identifiability guarantees, scalable computational pipelines, and promising application potential.
 It also opens up interesting directions for future research.
-Theoretically, it would be desirable to extend our identifiability and consistency results to the challenging setting with an unknown number of latent variables.
-We conjecture that one may be able to identify $\mathcal{K}$ under additional restrictions on the parameter space or considering specific parametric families.
-Second, it would be interesting to theoretically justify the good performance of the spectral-gap estimator for estimating the latent dimensions under general nonlinear link functions.
+First, the current formulation of DDEs mainly focuses on binary latent variables and a multi-layer graphical structure, but it provides foundations for understanding more complex discrete latent variable models.
+We believe both the identifiability theory and estimation methods are extendable to general categorical/polytomous latent variables.
+Additionally, our theoretical identifiability guarantees extend to “generalized DDEs” with cross-level edges, which we illustrate in Supplement S.1.5.
+It would be interesting to propose suitable estimation methods for these more flexible model settings.
 Another interesting problem pertains to high-dimensional settings, where the number of observed responses, $J$, may grow with the sample size $N$.
@@
 As modern datasets often comes with a large number of observed features, it would be interesting to explore whether our identifiability and estimability results can be generalized to such settings.
-In terms of the methodology and applications, an immediate future direction is working to effectively estimate and deploy deeper DDEs with $D\geq 3$ latent layers.
-The current theory and computational pipeline conceptually readily generalizes to deeper architectures than considered in this initial paper.
-It remains to test and refine these deeper models’ performance in larger-scale real-world data.
-Also, it would be interesting to extend DDEs to datasets with additional covariates.
+In terms of the methodology and applications, it would be interesting to extend DDEs to datasets with additional covariates.
 For example, the MNIST dataset comes with the actual digit labels as well as the spatial structure of the pixels in the image.
-In addition, many real datasets contain multiple modalities of data.
-For example, one may consider a dataset consisting of images uploaded in social media alongside text data such as tags.
-It is worthwhile to generalize DDEs to model such modern multimodal data.
 Finally, it would be interesting to extend DDEs for identifiable causal representation learning (Schölkopf et al., , 2021) to uncover causal structures among the higher-order latent variables.
@@
 The Supplement contains all technical proofs, additional theoretical results, and additional details about algorithms, simulations, and real data analyses.
-Supplementary Material
+Acknowldgement.
+The authors are partially supported by NSF Grant DMS-2210796.
+Supplement to “Deep Discrete Encoders: Identifiable Deep Generative Models for Rich Data with Discrete Latent Layers”
 This Supplementary Material is organized as follows.
-Section S.1 proves all main theorems and provides additional identifiability results under one-layer saturated models and variants.
+Section S.1 proves all main theorems and provides additional identifiability results under one-layer saturated models and generalized DDEs that goes beyond multi-layer structures.
 Section S.2 provides details regarding the spectral initialization algorithm.
-Section S.3 gives details regarding the EM algorithms (Algorithms 1 and 3) such as M-step update formulas, implementation details, and selection of the number of latent variables.
+Section S.3 gives details regarding the EM algorithm) such as the penalized EM algorithm, M-step update formulas, implementation details, and selection of the number of latent variables.
 Section S.4 provides various additional simulation results under (a) generically identifiable true parameters, (b) varying numbers of Monte Carlo samples in the SAEM algorithm, (c) unknown latent dimensions.
-Finally, Section S.5 gives additional data analysis details such as preprocessing, latent dimension selection, and additional visualizations.
+Section S.5 gives additional data analysis details such as preprocessing, latent dimension selection, and additional visualizations.
+Finally, Section S.6 discusses additional related works.
 S.1 Proof of Theorems
@@
 We prove the claims stated for the one-latent-layer models in Section S.1.3, and Theorem 3 in Section S.1.4.
-Additional identifiability results under related models with interaction effects (instead of the main-effect DDEs introduced in the main paper) are presented in Section S.1.5.
+Section S.1.5 establishes identifiability for generalized DDEs that allow additional edges compared.
+Additional identifiability results for selecting the latent dimension and results for related models with interaction effects (instead of the main-effect DDEs introduced in the main paper) are presented in Section S.1.6.
 S.1.1 Identifiability Under One-latent-layer Saturated Models
@@
 Definition S.2 (Equivalence relation).
-For the one-latent-layer saturated model, define an equivalence relationship “$\sim_{K^{(1)}}$” by setting $(\boldsymbol{\Theta}^{(1)},\mathbf{G}^{(1)})\sim_{K^{(1)}}(\widetilde{\boldsymbol{\Theta}}^{(1)},\widetilde{\mathbf{G}}^{(1)})$ if and only if $\boldsymbol{\gamma}=\widetilde{\boldsymbol{\gamma}}$ and there exist a permutation $\sigma^{(1)}\in S_{[K^{(1)}]}$ such that the following conditions hold:
+For the one-latent-layer saturated model, define an equivalence relationship $\sim_{K^{(1)}}$” by setting $(\boldsymbol{\Theta}^{(1)},\mathbf{G}^{(1)})\sim_{K^{(1)}}(\widetilde{\boldsymbol{\Theta}}^{(1)},\widetilde{\mathbf{G}}^{(1)})$ if and only if $\boldsymbol{\gamma}=\widetilde{\boldsymbol{\gamma}}$ and there exist a permutation $\sigma^{(1)}\in S_{[K^{(1)}]}$ such that the following conditions hold:
   • $\boldsymbol{\pi}_{(\alpha_{\sigma(1)},\ldots,\alpha_{\sigma(K^{(1)})})}=\widetilde{\boldsymbol{\pi}}_{\boldsymbol{\alpha}}$ for all $\boldsymbol{\alpha}\in\{0,1\}^{K^{(1)}}$
@@
 Consequently, the parameters $\mathbf{B}^{(d+1)},\mathbf{G}^{(d+1)},\boldsymbol{\pi}^{(d+1)}$ between the $d$th and $(d+1)$th layers are identified up to $\sim_{K^{(d+1)}}$.
-In particular, when $d=D-1$, it remains to determine ${\boldsymbol{p}}$ from the unstructured proportion parameter vector $\boldsymbol{\pi}^{(D)}$.
+In particular, when d=D1d=D-1, it remains to determine $\boldsymbol{p}$} from the unstructured proportion parameter vector $\boldsymbol{\pi}^{(D)}$.
 Since we already have identified the $D$th layer labels up to $\sim_{K^{(D)}}$, this simply follows by marginalizing out the irrelevant coordinates:
@@
 Now, assuming $\mathbf{B}^{(1)}\not\in N^{(1)}$, we can use a similar argument for deeper layers inductively.
-For $12\leq d\leq D-$ 1, let $N^{(d)}$ be the non-identifiable measure-zero subset of the $d$th layer coefficient space $\Omega(\mathbf{B}^{(d)};\mathbf{G}^{(d)})$.
+For $2\leq d\leq D-1$, let $N^{(d)}$ be the non-identifiable measure-zero subset of the $d$th layer coefficient space $\Omega(\mathbf{B}^{(d)};\mathbf{G}^{(d)})$.
 Note that we can still apply Proposition 2 since the conditional distribution of $\mathbf{A}^{(d)}\mid\mathbf{A}^{(d+1)}$ is modeled as a Bernoulli distribution with a logistic link $g_{\text{logistic}}$, which is indeed analytic and satisfy Assumption 2.
-Consequently, as long as $\mathbf{B}^{(d)}\not\in N^{(d)}$ for all $11\leq d\leq D-$ 1, the $D$-layer DDE is identifiable up to permutations of the latent variables within each layer.
+Consequently, as long as $\mathbf{B}^{(d)}\not\in N^{(d)}$ for all $1\leq d\leq D-1$, the $D$-layer DDE is identifiable up to permutations of the latent variables within each layer.
 The proof is complete since $\cup_{d=1}^{D}(N^{(d)})^{c}$ is a union of a finite number of measure-zero sets, and hence again measure-zero.
@@
 Most of these notations are consistent with Lee and Gu, (2024).
-First, we omit the superscript “${(1)}$” that indicates the first latent layer when dealing with the one-latent-layer saturated model.
+First, we omit the superscript “(1){(1)}” that indicates the first latent layer when dealing with the one-latent-layer saturated model.
 Let $j,k,\boldsymbol{\alpha}$ denote typical indices for $j\in[J],k\in[K],\boldsymbol{\alpha}\in\{0,1\}^{K}$.
@@
 Given $j$ and $\boldsymbol{\alpha}$, define the measure $\mathbb{P}_{j,\boldsymbol{\alpha}}$ on $\mathcal{Y}_{j}$ by setting
-[…]
+$\mathbb{P}_{j,\boldsymbol{\alpha}}(S):=\mathbb{P}(Y_{j}\in S\mid\mathbf{A}=\boldsymbol{\alpha})=\int_{S}p_{j}(y;\beta_{j,0}+\sum_{k}\beta_{j,k}\alpha_{k},\gamma_{j})dm_{j}(y).$ (S.3)
 In other words, $\mathbb{P}_{j,\boldsymbol{\alpha}}$ denotes the conditional distribution of $Y_{j}\mid\mathbf{A}=\boldsymbol{\alpha}$ in (4).
@@
 Then, we define the tensor product of $\mathbf{M}_{1},\mathbf{M}_{2},\mathbf{M}_{3}$ as
-$[\mathbf{M}_{1},\mathbf{M}_{2},\mathbf{M}_{3}]:=\sum_{l=1}^{r}\mathbf{m}_{1,l}\circ\mathbf{m}_{2,l}\circ\mathbf{m}_{3,l}.$
+$\mathbf{M}_{1},\mathbf{M}_{2},\mathbf{M}_{3}]:=\sum_{l=1}^{r}\mathbf{m}_{1,l}\circ\mathbf{m}_{2,l}\circ\mathbf{m}_{3,l}.$
 Using this notation, we can write $\mathbf{P}_{0}$ as follows:
@@
 Here, $\mathbf{M}_{a}$s are $\upsilon_{a}\times 2^{K}$ matrices, whose last rows are the all-one vector $\mathbf{1}_{2^{K}}^{\top}$ for $a=2,3$.
-Additionally, assuming that the graphical matrix $\mathbf{G}$ is known, the model parameters $(\boldsymbol{\Theta},\mathbf{G})$ are identifiable up to sign-flipping.
+Additionally, assuming that the graphical matrix $\mathbf{G}$ is known, the model parameters $\boldsymbol{\Theta}$ are identifiable up to sign-flipping.
+While Lee and Gu, (2024) establishes identifiability of the conditional distributions $\{\mathbb{P}_{j,\boldsymbol{\alpha}}\}$ (see (S.3)), recovering $\mathbf{B},\boldsymbol{\gamma}$ from $\{\mathbb{P}_{j,\boldsymbol{\alpha}}\}$ is straightforward since each $\mathbb{P}_{j,\boldsymbol{\alpha}}$ is the distribution of an identifiable parametric family.
 S.1.3.2 Proof of Proposition 1
-The following Lemma will be crucially utilized to identify and recover $\widetilde{\mathbf{G}}$. We present its proof after proving the main Proposition.
+The following Lemma will be crucially utilized to identify and recover $\widetilde{\mathbf{G}}$.
+Lemma S.2 generalizes the fact that when $Y_{j}$ is a pure child of $A_{k}^{(1)}$ (in other words $|H_{j}|=1$), we have
+$|\mathcal{S}_{j}|=|\{\mathbb{P}(Y_{j}\mid\mathbf{A}^{(1)}=\boldsymbol{\alpha}^{(1)}):\boldsymbol{\alpha}^{(1)}\in\{0,1\}^{K^{(1)}}\}|=|\{\mathbb{P}(Y_{j}\mid A_{k}^{(1)}=\alpha_{k}^{(1)}):\alpha_{k}^{(1)}=0,1\}|=2.$
+Thus, by partitioning the set of conditional distributions corresponding to $Y_{j}$ which are pure children, the binary indices $\{\boldsymbol{\alpha}^{(1)}:\alpha_{k}^{(1)}=1\}$ and $\widetilde{\mathbf{G}}$ can be recovered (up to the permutation $\sigma\in S_{[K^{(1)}]}$).
+We present its proof after proving the main Proposition.
 Lemma S.2.
@@
 Our proof builds upon Lemma S.1, which proved a more general notion of nonparametric identifiability of the one-layer model, but under a given graphical matrix $\mathbf{G}$.
-The cited Theorem 1 proves that the continuous parameters are identifiable up to sign flipping, given $\mathbf{G}$.
+Lemma S.1 proves that the continuous parameters are identifiable up to sign flipping, given $\mathbf{G}$.
 In our setting, Assumption 1(c) resolves the sign flipping issue and the continuous parameters can be uniquely determined.
@@
 We separate this proof into two steps.
-Step 1: Tensor decomposition and setup.
-Consider a one-latent-layer saturated model with true parameters $(\boldsymbol{\Theta},\mathbf{G})$ that satisfies conditions A, B.
-Suppose there exists an alternative set of parameters $(\widetilde{\boldsymbol{\Theta}},\widetilde{\mathbf{G}})$ that define a same marginal distribution of $\mathbf{Y}$.
-Define the notation ~$\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}$ similar as ${\mathbb{P}}_{j,{\boldsymbol{\alpha}}}$ in (S.3), and also define $\widetilde{\mathbf{N}}_{1},\widetilde{\mathbf{N}}_{2},\widetilde{\mathbf{N}}_{3}$ as the conditional probability matrices that specify the value of
+[…]
 $\mathbb{P}\big{(}(Y_{1},\ldots,Y_{K})\mid\mathbf{A}=\widetilde{\boldsymbol{\alpha}}\big{)},\quad\mathbb{P}\big{(}(Y_{K+1}\ldots,Y_{2K})\mid\mathbf{A}=\widetilde{\boldsymbol{\alpha}}\big{)},\quad\mathbb{P}\big{(}(Y_{2K+1},\ldots,Y_{J})\mid\mathbf{A}=\widetilde{\boldsymbol{\alpha}}\big{)}$
@@
 $\mathbf{P}_{0}=[{\mathbf{N}}_{1}\text{{diag}}({\boldsymbol{\pi}}),{\mathbf{N}}_{2},{\mathbf{N}}_{3}]=[\widetilde{\mathbf{N}}_{1}\text{{diag}}(\widetilde{\boldsymbol{\pi}}),\widetilde{\mathbf{N}}_{2},\widetilde{\mathbf{N}}_{3}].$ (S.6)
-By applying Lemma S.1, the tensor decomposition in (S.6) is unique, so $\widetilde{\mathbf{N}}_{1}\text{{diag}}({\boldsymbol{\pi}}),\widetilde{\mathbf{N}}_{2},\widetilde{\mathbf{N}}_{3}$ and $\mathbf{N}_{1}\text{{diag}}({\boldsymbol{\pi}}),\mathbf{N}_{2},\mathbf{N}_{3}$ are identical up to a common column permutation, say $\mathfrak{S}\in S_{\{0,1\}^{K}}$.
-In particular, as the last row of $\widetilde{\mathbf{N}}_{1}\text{{diag}}(\widetilde{\boldsymbol{\pi}})$ and $\mathbf{N}_{1}\text{{diag}}({\boldsymbol{\pi}})$ is exactly $\widetilde{\boldsymbol{\pi}}^{\top}$ and $\boldsymbol{\pi}^{\top}$, $\big{(}\widetilde{\boldsymbol{\pi}},\widetilde{\mathbf{N}}_{1}\big{)}$ and $\big{(}\boldsymbol{\pi},\mathbf{N}_{1}\big{)}$ are also identical up to the same permutation $\mathfrak{S}$.
+[…]
 We use this observation to prove that $\widetilde{\mathbf{G}}$ is equivalent to $\mathbf{G}$ under $\underset{K^{(1)}}{\sim}$.
@@
 $=$ $\Big{\{}\big{(}\widetilde{\mathbf{N}}_{a_{j}}((\kappa_{1},\ldots,\kappa_{j-1},{1},\kappa_{j+1},\ldots,\kappa_{J}),\widetilde{\boldsymbol{\alpha}}),\cdots,\widetilde{\mathbf{N}}_{a_{j}}((\kappa_{1},\ldots,\kappa_{j-1},{\kappa_{j}},\kappa_{j+1},\ldots,\kappa_{J}),\widetilde{\boldsymbol{\alpha}})\big{)}:\widetilde{\boldsymbol{\alpha}}\Big{\}}$
-[…]
+$=$ $\Big{\{}\big{(}\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{1,j}),\ldots,\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{\kappa_{j},j})\big{)}:{\widetilde{\boldsymbol{\alpha}}}\Big{\}}$, (S.8)
 and provides enough information about $\widetilde{\mathbf{g}}_{j}$. Here, the index $a_{j}=1,2,3$ can be understood from the context, for example $a_{j}=1$ when $j\leq K$, $a_{j}=2$ when $K<j\leq 2K$.
@@
 But this is immediate by noting that the cardinality of the pmf vector of $\{\mathbb{P}(Y_{k}\in S_{1,k},Y_{l}\in S_{1,l}\mid\mathbf{A}=\boldsymbol{\alpha}):\boldsymbol{\alpha}\}$ is 2 if and only if $k=l$.
-For future purposes, let us partition the $2^{K}$ row indices $\widetilde{\boldsymbol{\alpha}}$ in (S.8) into two groups $T_{k}$ and $T_{k}^{c}$ based on the value of (~$\big{(}\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{1,j}),\ldots,\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{\kappa_{j},j})\big{)}$.
+For future purposes, let us partition the $2^{K}$ row indices $\widetilde{\boldsymbol{\alpha}}$ in (S.8) into two groups $T_{k}$ and $T_{k}^{c}$ based on the value of $\big{(}\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{1,j}),\ldots,\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{\kappa_{j},j})\big{)}$.
 Then, the columns of $\widetilde{\mathbf{N}}$ that correspond to the $\widetilde{\boldsymbol{\alpha}}\in T_{k}$ must be identical to the columns of $\mathbf{N}$} indexed by $\boldsymbol{\alpha}$s such that $\alpha_{\sigma(k)}=0$ (or 1).
-Now, for each $j>K$ and $k\in[K]$, we show that $\widetilde{g}_{j,k}=g_{j,\sigma(k)}$ and complete the proof. By part (b) of Lemma S.2, we have $\widetilde{g}_{j,k}$ = 0 if and only if
-{(~$\big{\{}\big{(}\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{1,j}),\ldots,\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{\kappa_{j},j})\big{)}:\widetilde{\boldsymbol{\alpha}}\in T_{k}\big{\}}=\big{\{}\big{(}\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{1,j}),\ldots,\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{\kappa_{j},j})\big{)}:\widetilde{\boldsymbol{\alpha}}\not\in T_{k}\big{\}}.$ (S.9)
+Now, for each $j>K$ and $k\in[K]$, we show that $\widetilde{g}_{j,k}=g_{j,\sigma(k)}$. By part (b) of Lemma S.2, we have $\widetilde{g}_{j,k}$ = 0 if and only if
+$\big{\{}\big{(}\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{1,j}),\ldots,\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{\kappa_{j},j})\big{)}:\widetilde{\boldsymbol{\alpha}}\in T_{k}\big{\}}=\big{\{}\big{(}\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{1,j}),\ldots,\widetilde{\mathbb{P}}_{j,\widetilde{\boldsymbol{\alpha}}}(S_{\kappa_{j},j})\big{)}:\widetilde{\boldsymbol{\alpha}}\not\in T_{k}\big{\}}.$ (S.9)
 For notational simplicity, let $\boldsymbol{\alpha}_{\sigma}:=(\alpha_{\sigma(1)},\ldots,\alpha_{\sigma(k)})$. Then, by the construction of $T_{k}$, (S.9) simplifies to
 $\Big{\{}\big{(}\mathbb{P}_{j,\boldsymbol{\alpha}_{\sigma}}(S_{1,j}),\ldots,\mathbb{P}_{j,\boldsymbol{\alpha}_{\sigma}}(S_{\kappa_{j},j})\big{)}:\alpha_{\sigma(k)}=1\Big{\}}=\Big{\{}\big{(}\mathbb{P}_{j,\boldsymbol{\alpha}_{\sigma}}(S_{1,j}),\ldots,\mathbb{P}_{j,\boldsymbol{\alpha}_{\sigma}}(S_{\kappa_{j},j})\big{)}:\alpha_{\sigma(k)}=0\Big{\}}.$
-Another application of part (b) of Lemma S.2 shows that this is equivalent to $g_{j,\sigma(k)}=0$. Hence, $\widetilde{g}_{j,k}=g_{j,\sigma(k)}$ and the proof is complete. ∎
+Another application of part (b) of Lemma S.2 shows that this is equivalent to $g_{j,\sigma(k)}=0$.
+Hence, $\widetilde{g}_{j,k}=g_{j,\sigma(k)}$ and we have shown $\mathbf{G}\underset{K^{(1)}}{\sim}\mathbf{G}$.
+This completes the proof.
+∎
 Below, we provide a short proof of Lemma S.2.
@@
     The proof is immediate from the faithfulness of the graphical matrix $\mathbf{G}$.
-For example, the third bullet point follows from noting that $|H_{j}|\geq 2$ implies |$|\{\sum_{k\in H_{j}}\beta_{j,k}\alpha_{k}:\boldsymbol{\alpha}\}|\geq 3$, and that |$|\{\sum_{k\in H_{j}}\beta_{j,k}\alpha_{k}:\boldsymbol{\alpha}\}|\leq 2$ implies $|H_{j}|\leq 1$.
+For example, the third bullet point follows from noting that $|H_{j}|\geq 2$ implies $|\{\sum_{k\in H_{j}}\beta_{j,k}\alpha_{k}:\boldsymbol{\alpha}\}|\geq 3$, and that $|\{\sum_{k\in H_{j}}\beta_{j,k}\alpha_{k}:\boldsymbol{\alpha}\}|\leq 2$ implies $|H_{j}|\leq 1$.
   2. (b)
@@
 This is a nontrivial problem as we wish to identify the parameters up to a label switching of $K$ binary latent variables.
+Note that we can no longer utilize the pure children to address this problem, which was the strategy under the setting of strict identifiability.
 Here, to formally state the label switching, recall the notation of $\sim_{K}$ Definition S.2 and write $\mathbf{B}\sim_{K}\widetilde{\mathbf{B}},\mathbf{G}\sim_{K}\widetilde{\mathbf{G}}$, $\boldsymbol{\alpha}\sim_{K}\widetilde{\boldsymbol{\alpha}}$ when there exists a permutation $\sigma\in S_{[K]}$ such that $\beta_{j,l}=\widetilde{\beta}_{j,\sigma(l)}$, $g_{j,l}=\widetilde{g}_{j,\sigma(l)}$, and $\boldsymbol{\alpha}=(\widetilde{\alpha}_{\sigma(1)},\ldots,\widetilde{\alpha}_{\sigma(K)}).$
 Lemma S.4.
-Suppose that there exist two sets of parameters $(\mathbf{B},\mathbf{G})$, $(\widetilde{\mathbf{B}},\widetilde{\mathbf{G}})$ that satisfies Assumption 1 and defines an identical $\eta_{j,\boldsymbol{\alpha}}$ in the following sense: there exists a permutation $\mathfrak{S}\in S_{\{0,1\}^{K}}$ such that
+[…]
 $\eta_{j,\boldsymbol{\alpha}}=\widetilde{\eta}_{j,\mathfrak{S}(\boldsymbol{\alpha})},$ (S.11)
@@
 We assume that $\mathcal{D}_{j}\subset\mathcal{C}_{j}$, where $\mathcal{C}_{j}$ is a countable separating class whose values determine probability measure on $\mathcal{Y}_{j}$.
-The existence of such a separating class is a consequence of $\mathcal{Y}_{j}$ being a separating metric space; see Step 1 in the proof of Theorem 1 in Lee and Gu, (2024) for a proof.
-Suppose that $\mathbf{G}$ satisfy condition C, $\boldsymbol{\Theta}\in\Omega_{\mathcal{K}}(\boldsymbol{\Theta};\mathbf{G})$, and that there exists an alternate parameter $\widetilde{\boldsymbol{\Theta}},\widetilde{\mathbf{G}}$ that defines the same marginal likelihood.
-Here, we are dropping all superscripts in $\mathbf{G}=\mathbf{G}^{(1)\star}$ for simplicity.
-We show that $(\boldsymbol{\Theta},\mathbf{G})\sim_{\mathcal{K}}(\widetilde{\boldsymbol{\Theta}},\widetilde{\mathbf{G}})$ for $\boldsymbol{\Theta}\in\Omega_{\mathcal{K}}(\boldsymbol{\Theta};\mathbf{G})\setminus(\mathcal{N}_{1}\cup\mathcal{N}_{2})$, where $\mathcal{N}_{1},\mathcal{N}_{2}$ are measure-zero sets that will be defined later.
-Recall that for $a=1,2,3$, $\mathbf{N}_{a}$ are the $\upsilon_{a}\times 2^{K}$ conditional probability matrices that describe $\mathbf{Y}\mid(\mathbf{A}=\boldsymbol{\alpha})$ under the true parameters $\boldsymbol{\Theta},\mathbf{G}$.
-Similarly, define $\widetilde{\mathbf{N}}_{a}$ as the corresponding matrices under the alternative parameters $\widetilde{\boldsymbol{\Theta}},\widetilde{\mathbf{G}}$.
-Similar to the proof of Proposition 1, we start from the rank $2^{K}$ decomposition of the marginal probability tensor in (S.5).
-By applying Lemma S.3, the decomposition
+The existence of such a separating class is a consequence of $\mathcal{Y}_{j}$ being a separable metric space; see Step 1 in the proof of Theorem 1 in Lee and Gu, (2024) for a proof.
+[…]
 $\mathbf{P}_{0}=[\mathbf{N}_{1}\text{{diag}}(\boldsymbol{\pi}),\mathbf{N}_{2},\mathbf{N}_{3}]=[\widetilde{\mathbf{N}}_{1}\text{{diag}}(\widetilde{\boldsymbol{\pi}}),\widetilde{\mathbf{N}}_{2},\widetilde{\mathbf{N}}_{3}]$ (S.12)
@@
 Recalling that $\mathbb{P}_{j,\boldsymbol{\alpha}}=\text{ParFam}_{j}(\eta_{j,\boldsymbol{\alpha}},\gamma_{j})$ for some identifiable parametric family (see (S.3)), we must have $\eta_{j,\boldsymbol{\alpha}}=\widetilde{\eta}_{j,\mathfrak{S}(\boldsymbol{\alpha})}$ and $\gamma_{j}=\widetilde{\gamma}_{j}$ for all $j,\boldsymbol{\alpha}$.
-Now, additionally assuming that $\boldsymbol{\Theta}\not\in\mathcal{N}_{2}$, we can apply Lemma S.4.
-Here, $\mathcal{N}_{2}$ is the null set defined in Lemma S.4.
-Then, we have $(\mathbf{B},\mathbf{G})\sim_{K}(\widetilde{\mathbf{B}},\widetilde{\mathbf{G}})$.
-Also, because $\boldsymbol{\alpha}\sim_{K}\mathfrak{S}(\boldsymbol{\alpha})$, $\widetilde{\pi}_{\mathfrak{S}(\boldsymbol{\alpha})}=\pi_{\boldsymbol{\alpha}}$ implies $\widetilde{\boldsymbol{\pi}}\sim_{K}\boldsymbol{\pi}$ and the proof is complete.
-∎
+[…]
 We finally present the postponed proof of Lemmas S.3 and S.4.
@@
 We claim that it suffices to show that
-$rk_{k}(\mathbf{N}_{1})=2^{K},~{}rk_{k}(\mathbf{N}_{2})=2^{K},~{}rk_{k}(\mathbf{N}_{3})\geq 2$
+$rk_{k}(\mathbf{N}_{1})=2^{K},~rk_{k}(\mathbf{N}_{2})=2^{K},~rk_{k}(\mathbf{N}_{3})\geq 2$ (S.14)
 for generic parameters in $\Omega(\boldsymbol{\Theta};\mathbf{G})\setminus\mathcal{N}$.
@@
 We first assume that the true parameter $\mathbf{B}$ belongs in the coefficient space
-$\Omega^{g}(\mathbf{B};\mathbf{G}):=\{\mathbf{B}:\text{for each $j$, all elements of the form }\sum_{k\in H_{j}}\beta_{j,k}a_{k},a_{k}=0,1,\text{ are distinct}\}.$
+[…]
 It is clear that the complement $\Omega(\mathbf{B};\mathbf{G})\setminus\Omega^{g}(\mathbf{B};\mathbf{G})$ can be spelled out as a finite union of lower dimensional subspaces, and has measure-zero with respect to $\Omega(\mathbf{B};\mathbf{G})$.
@@
 Consequently, we can construct the permutation $\mathfrak{T}_{j}$ by setting
-$(\mathfrak{T}_{j}(\boldsymbol{\alpha}))_{k}:=\begin{cases}\alpha_{l_{k}}&\text{if }\beta_{j,l_{k}}>0\\ 1-\alpha_{l_{k}}&\text{if }\beta_{j,l_{k}}<0,\end{cases}$
+[…]
 for $k\leq I_{j}$ and arbitrarily defining the remaining coordinates. By plugging in these definitions, we get
@@
 Inductively for $t>1$, let
-$\mathcal{T}_{j,k}^{[t]}$ :=$:=\{\boldsymbol{\alpha}\not\in T_{j,k}^{(t-1)},V_{j,k}^{(t-1)}:\eta_{j,\boldsymbol{\alpha}}=\min_{\boldsymbol{\alpha}^{\prime}\not\in T_{j,k}^{(t-1)},V_{j,k}^{(t-1)}}\eta_{j,\boldsymbol{\alpha}^{\prime}}\},$ (S.16)
-$\mathcal{V}_{j,k}^{[t]}$ :=$:=\{\boldsymbol{\alpha}\not\in T_{j,k}^{(t-1)},V_{j,k}^{(t-1)}:\eta_{j,\boldsymbol{\alpha}}=\min_{\boldsymbol{\alpha}^{\prime}\not\in T_{j,k}^{(t-1)},V_{j,k}^{(t-1)}}\eta_{j,\boldsymbol{\alpha}^{\prime}}+c_{j,k}\},$ (S.17)
+$\mathcal{T}_{j,k}^{[t]}$ $:=\{\boldsymbol{\alpha}\not\in T_{j,k}^{(t-1)},V_{j,k}^{(t-1)}:\eta_{j,\boldsymbol{\alpha}}=\min_{\boldsymbol{\alpha}^{\prime}\not\in T_{j,k}^{(t-1)},V_{j,k}^{(t-1)}}\eta_{j,\boldsymbol{\alpha}^{\prime}}\},$ (S.17)
+$\mathcal{V}_{j,k}^{[t]}$ $:=\{\boldsymbol{\alpha}\not\in T_{j,k}^{(t-1)},V_{j,k}^{(t-1)}:\eta_{j,\boldsymbol{\alpha}}=\min_{\boldsymbol{\alpha}^{\prime}\not\in T_{j,k}^{(t-1)},V_{j,k}^{(t-1)}}\eta_{j,\boldsymbol{\alpha}^{\prime}}+c_{j,k}\},$ (S.18)
 and define $T_{j,k}^{[t]}=T_{j,k}^{(t-1)}\cup\mathcal{T}_{j,k}^{[t]}$ and $V_{j,k}^{[t]}=V_{j,k}^{(t-1)}\cup\mathcal{V}_{j,k}^{[t]}$.
@@
 Hence, for some permutation $\widetilde{\tau}\in S_{[K]}$, we must have
-$\widetilde{T}_{l}$ =$=\{\widetilde{\boldsymbol{\alpha}}:\widetilde{\alpha}_{\widetilde{\tau}(l)}=0\}=\mathfrak{S}(\{\boldsymbol{\alpha}:\alpha_{\tau(l)}=0\})=\mathfrak{S}(T_{l}),$ (S.22)
-$\quad\widetilde{V}_{l}$ =$=\{\widetilde{\boldsymbol{\alpha}}:\widetilde{\alpha}_{\widetilde{\tau}(l)}=1\}=\mathfrak{S}(\{\boldsymbol{\alpha}:\alpha_{{\tau}(l)}=1\})=\mathfrak{S}(V_{l}).$ (S.23)
+$\widetilde{T}_{l}$ $=\{\widetilde{\boldsymbol{\alpha}}:\widetilde{\alpha}_{\widetilde{\tau}(l)}=0\}=\mathfrak{S}(\{\boldsymbol{\alpha}:\alpha_{\tau(l)}=0\})=\mathfrak{S}(T_{l}),$ (S.23)
+$\quad\widetilde{V}_{l}$ $=\{\widetilde{\boldsymbol{\alpha}}:\widetilde{\alpha}_{\widetilde{\tau}(l)}=1\}=\mathfrak{S}(\{\boldsymbol{\alpha}:\alpha_{{\tau}(l)}=1\})=\mathfrak{S}(V_{l}).$ (S.24)
 Now, we finish the proof by proving $\mathbf{B}\sim\widetilde{\mathbf{B}}$ and $\boldsymbol{\alpha}\sim\mathfrak{S}(\boldsymbol{\alpha})$.
@@
 ∎
-S.1.4 Proof of Theorem 3
+S.1.4 Technical Conditions and Proof of Theorem 3
+We first state the technical assumptions regrading the penalty function $p_{\lambda_{N},\tau_{N}}$ and tuning parameters $(\lambda_{N},\tau_{N})$ that are imposed for Theorem 3.
+For some tuning parameters $\lambda_{N},\tau_{N}>0$, $p_{\lambda_{N},\tau_{N}}:\mathbb{R}\to[0,\infty)$ is a sparsity-inducing symmetric penalty that is nondecreasing on $0,\infty)$, nondifferentiable at 0, differentiable at $(0,\tau_{N})$, $p_{\lambda_{N},\tau_{N}}\propto{\lambda_{N}}/{\tau_{N}}$ around 0 and satisfy
+$p_{\lambda_{N},\tau_{N}}(b)=0,~\text{if}~b=0;\quad p_{\lambda_{N},\tau_{N}}^{\prime}(b)\leq\frac{C\lambda_{N}}{\tau_{N}},~\text{if}~|b|\leq\tau_{N};\quad p_{\lambda_{N},\tau_{N}}(b)=\lambda_{N},\quad\text{if}~|b|\geq\tau_{N}.$
+Note that $\lambda_{N}$ is the magnitude of the penalty, and $\tau_{N}$ is the point of truncation.
+We also assume that $\lambda_{N}$ and $\tau_{N}$ depend on $N$ such that ${1}/{\sqrt{N}}\ll\tau_{N}\ll{\lambda_{N}}/{\sqrt{N}}\ll 1.$ Here, for nonnegative sequences $\{(a_{N},b_{N})\}_{N\geq 1}$, we write $a_{N}\ll b_{N}$ when ${a_{N}}/{b_{N}}\to 0$ as $N\to\infty$.
 Proof.
@@
 We first re-write the inequality $Q_{N}(\widetilde{\boldsymbol{\Theta}})\leq Q_{N}(\boldsymbol{\Theta}^{\star})$ as
-$-\ell(\widetilde{\boldsymbol{\Theta}})+\ell(\boldsymbol{\Theta}^{\star})\leq p_{\lambda_{N},\tau_{N}}(\mathbf{B}_{0})-p_{\lambda_{N},\tau_{N}}(\widetilde{\mathbf{B}}).$ (S.26)
+[…]
 By a Taylor expansion, we can bound the LHS of (S.27) as
@@
 Thus, $\|\widetilde{\boldsymbol{\Theta}}-\boldsymbol{\Theta}^{\star}\|_{2}=O_{p}\left(\frac{1}{\sqrt{N}}\right)$.
-[…]
+Finally, we prove the estimation consistency for the discrete graph structures $\mathbf{G}^{(d)}$s.
+It suffices to show that for any fixed $d,l,k$, $\widetilde{g}_{l,k}^{(d)}=g_{0;l,k}^{(d)}$ with high probability.
+As a first case, suppose that $g_{0;l,k}^{(d)}=1$.
+Then, the consistency result implies $\widetilde{\beta}_{l,k}^{(d)}\xrightarrow{p}\beta_{0;l,k}^{(d)}\neq 0$.
+Hence, $\widetilde{\beta}_{l,k}^{(d)}\neq 0$ with high probability, so $\widetilde{g}_{l,k}^{(d)}=1$.
+Next, consider the case when $g_{0;l,k}^{(d)}=1$.
+Assume the converse, and suppose $\widetilde{\beta}_{l,k}^{(d)}\neq 0$.
+By the $\sqrt{N}$-consistency and the assumption that $\tau_{N}\gg\frac{1}{\sqrt{N}}$, we must have $|\widetilde{\beta}_{l,k}^{(d)}|\ll\tau_{N}$ with high probability.
+Then, the first-order conditions (KKT conditions) give that
 $\partial_{\beta_{l,k}^{(d)}}\ell(\widetilde{\boldsymbol{\Theta}}):=\frac{\partial\ell(\boldsymbol{\Theta})}{\partial\beta_{l,k}^{(d)}}\mid_{\boldsymbol{\Theta}=\widetilde{\boldsymbol{\Theta}}}=p^{\prime}_{\lambda_{N},\tau_{N}}(\widetilde{\beta}_{l,k}^{(d)})=\Theta_{p}\left(\frac{\lambda_{N}}{\tau_{N}}\right).$
@@
 and $\sqrt{N}\ll{\lambda_{N}}/{\tau_{N}}.$ Hence, we must have $\widetilde{g}_{l,k}^{(d)}=0$, and the proof is complete. ∎
-S.1.5 Detour: Identifiability of One-latent-layer Saturated Models With Interaction Effects of Binary Latent Variables
+Remark S.1.
+One natural question is to whether our estimator would still be consistent when the number of latent variables are unknown.
+This extension is not straightforward since the number of the top-layer latent variable, $K^{(D)}$, determines the number of deepest mixture components of DDEs.
+Estimating the number of mixture components is a challenging problem even in simple parametric models, and often leads to a slower (than ${1}/{\sqrt{N}}$) rate of convergence in parameter estimation (Goffinet et al., , 1992; Ho and Nguyen, , 2016).
+Note that for such cases, the Fisher information becomes singular, and Theorem 3 cannot be applied.
+S.1.5 Identifiability of Generalized DDEs with More Complex Latent Structures
+As pointed out by a reviewer, the current architecture of DDEs do not allow cross-level edges, which may limit the models’ representational power.
+Here, we illustrate that DDEs can still be identified, even under the presence of cross-level dependencies.
+Our main idea is to re-formulate multi-layer graphical structures with cross-level edges into a structure with within-layer edges.
+See Figure S.1 for a visual illustration; where the latent variable $A_{1}$ with cross-level edges (in the left panel) is moved to the lower latent layer (in the right panel).
+Based on this re-formulation, we instead establish identifiability of multi-layer latent structures where within-layer arrows are permitted.
+In other words, layer-wise local dependence is allowed.
+$A_{1}$$A_{2}$$A_{3}$$A_{4}$$A_{5}$$Y_{1}$$Y_{2}$$Y_{3}$$Y_{4}$$Y_{5}$$Y_{6}$$Y_{7}$$Y_{8}$
+[…]
+Figure S.1: Left: Example graphical structure without within-level edges, but allowing cross-level edges.
+Right: Equivalent formulation by moving $A_{1}$ to the middle layer; now we have within-level edges $\Lambda^{(1)}=\{A_{1}\to A_{3}\}$ instead of cross-level edges.
+Generalized DDEs.
+To formalize each layer in such generalized DDEs, we assume that this multi-layer latent structure is ordered in a manner such that
+$\text{{pa}}(\mathbf{A}_{1:K^{(d)}}^{(d)})\setminus\mathbf{A}_{1:K^{(d)}}^{(d)}=\mathbf{A}_{1:K^{(d+1)}}^{(d+1)},\quad\forall 0\leq d\leq D-1,$
+where we write $\mathbf{A}^{(0)}=\mathbf{Y}$ for notational simplicity.
+That is, the parents of variables in the $d$th layer must belong either in the current ($d$th) or directly upper ($d+1$th) layer.
+Assume that there are no edges $Y_{j}\to Y_{j^{\prime}}$ in the observed (bottom) layer, in other words, assume that local dependence occurs only among the latent variables.
+For each latent layer (indexed by $d$), assume without loss of generality that $\mathbf{A}^{(d)}$ are indexed in a topological manner, so that there exist no arrows of the form $A_{k}^{(d)}\to A_{k^{\prime}}^{(d)}$ for $k\geq k^{\prime}$.
+Let $\Lambda^{(d)}$ denote the graph on the vertex set $\mathbf{A}^{(d)}$; for example, $\Lambda^{(1)}=\{A_{1}\to A_{3}\}$ in Figure S.1.
+Also, for each $A_{k}^{(d)}$, extend the conditional probabilities in (3) by including $\delta$-parameters to model the dependence of the same ($d$th) layer variables:
+$A_{k}^{(d)}\mid\mathbf{A}_{1:(k-1)}^{(d)},\mathbf{A}^{(d+1)}\sim\text{Ber}\Big{(}\operatorname{\emph{g}_{\text{logistic}}}(\beta_{k,0}^{(d+1)}+\sum_{l=1}^{K^{(d+1)}}\beta_{k,l}^{(d+1)}A_{l}^{(d+1)}+\sum_{k^{\prime}=1}^{k-1}\delta_{k^{\prime},k}^{(d)}A_{k^{\prime}}^{(d)})\Big{)}.$ (S.28)
+For simplicity, assume monotonicity of the conditional distributions in the sense that all coefficients $\beta_{k,l}^{(d+1)},\delta_{k^{\prime},k}^{(d)}$ are nonnegative.
+In the following, we establish identifiability of generalized DDEs under the class of “topologically double triangular” models.
+This notion follows from a recent work Lee and Gu, (2025), where a single latent layer was considered and no arrows between the observed variables were allowed.
+Here, we extend the “double triangular” graphical matrices by additionally incorporating the edges $\Lambda^{(d)}$ in each layer as follows.
+Definition S.3 (Triangular graphical matrix).
+A $L\times K^{(d)}$ matrix $\mathbf{G}_{1}^{(d)}$ with binary entries is “triangular” when it takes the following form:
+[…]
+Here, the starred entries are allowed to take any values 0 or 1, and each $\mathbf{G}_{1,k}^{(d)}$ can be a row-vector.
+For each $d\geq 2$, we say that the $K^{(d-1)}\times K^{(d)}$ binary matrix $\mathbf{G}^{(d)}$ is “topologically double-triangular” when there exists a topological ordering of $K^{(d-1)}]$ (with respect to the graph $\Lambda^{(d-1)}$) such that:
+$\mathbf{G}^{(d)}=\begin{pmatrix}\mathbf{G}_{1}^{(d)}\\ \mathbf{G}_{2}^{(d)}\\ \mathbf{G}_{3}^{(d)}\end{pmatrix}$,
+where
+  1. (i)
+    the matrices $\mathbf{G}_{1}^{(d)},\mathbf{G}_{2}^{(d)}$ are triangular after individual column permutations,
+  2. (ii)
+    $\mathbf{G}_{3}^{(d)}$ does not have empty columns,
+  3. (iii)
+    there exists no arrows between $\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}$ in $\Lambda^{(d-1)}$ (i.e.
+$\mathcal{I}_{1},\mathcal{I}_{2},\mathcal{I}_{3}$ are $d$-separated given $\mathbf{A}$).
+Here, each $\mathcal{I}_{a}$ denotes the set of row indices corresponding to $\mathbf{G}_{a}^{(d)}$.
+We elaborate on the details of Definition S.3.
+For simplicity, assume that each $\mathbf{G}_{1,k}^{(d)}$ is a row vector $(\underbrace{0,\ldots,0}_{k-1},1,\ldots,*)$.
+Then, $\mathbf{G}_{1}^{(d)}$ simplifies to the upper triangular matrix $\begin{pmatrix}1&*&\cdots&*\\ 0&1&\cdots&*\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&1\\ \end{pmatrix}$.
+Thus, the topologically double-triangular condition is requiring two such triangular matrices.
+Here, note that condition (i) requires $\mathbf{G}_{1}^{(d)},\mathbf{G}_{2}^{(d)}$ to be a triangular matrix after arbitrary column permutations.
+For example, in Figure S.2, by letting
+$\mathbf{G}_{1}^{(d)}=\begin{pmatrix}1&1\\ 0&1\end{pmatrix},\quad\mathbf{G}_{2}^{(d)}=\begin{pmatrix}1&1\\ 1&0\end{pmatrix}$
+respectively denote the (1,2)(1,2)nd and $(3,4)$th rows of $\mathbf{G}^{(d)}$, both matrices are triangular.
+We also have $\mathbf{G}_{3}^{(d)}=(1,1)$, so condition (ii) as well as (iii) are also satisfied, and Figure S.2 defines an topologically double-triangular graphical model.
+$A_{1}^{(d-1)}$$A_{2}^{(d-1)}$$A_{3}^{(d-1)}$$A_{4}^{(d-1)}$$A_{5}^{(d-1)}$$A_{1}^{(d)}$$A_{2}^{(d)}$
+Figure S.2: Graphical illustration of a topologically double-triangular structure, with $\mathcal{I}_{1}=\{1,2\},\mathcal{I}_{2}=\{3,4\},\mathcal{I}_{3}=\{5\}$.
+Definition S.4.
+A $D$-latent-layer generalized DDE is a statistical model defined by the conditional distributions in (1), (S.28), and (4).
+Now, we establish identifiability of generalized DDEs, under the class of topologically double triangular models.
+Theorem S.1.
+Consider a $D$-latent-layer generalized DDE where the number of latent variables per layer, $\mathcal{K}$, is known.
+Among the class of topologically double-triangular generalized DDEs, the model is identifiable up to (i) Markov equivalence for tree-structures within each layer and (ii) label-switching, when the observed (bottom) layer satisfy conditions A and B (see Theorem 1).
+Here, the restriction to Markov equivalence is a fundamental ambiguity for identifying DAG models.
+For example, suppose that the true graph is $A_{1}^{(1)}\to A_{2}^{(1)}\to A_{3}^{(1)}\leftarrow A_{1}^{(2)}$, where the graph in the first latent layer is a tree.
+Then, this structure is indistinguishable from $A_{1}^{(1)}\leftarrow A_{2}^{(1)}\to A_{3}^{(1)}\leftarrow A_{1}^{(2)}$.
+Note that colliders with multiple parents (e.g.
+$A_{3}^{(1)}$) do not suffer from this ambiguity, thanks to the parametric assumption in (S.28).
+Compared to identifiability results in the main text, we additionally impose structural restrictions on the class of alternative graphical models.
+In other words, both the true and alternative models are required to be topologically double-triangular.
+Proof.
+To prove Theorem S.1, it suffices to show the following proposition for the generalized DDE with one-latent layer.
+Then, Theorem S.1 follows by the exact same layer-wise identifiability argument as that in the proof of Theorem 1.
+Proposition 3.
+Consider a saturated generalized DDE with only one latent layer, and known $K$.
+Among the class of topologically double-triangular saturated generalized DDEs, the model is identifiable up to Markov equivalence.
+We prove Proposition 3 using the following key property for matrix ranks, whose proof is deferred to the end of the section.
+Lemma S.6.
+Suppose that $\mathbf{G}_{\mathcal{I}}$ is triangular. Then, $\mathbb{P}(Y_{\mathcal{I}}\mid\mathbf{A})$ has full column-rank.
+Proof of Proposition 3.
+We use the notations from Section S.1.3, and denote the latent variables as $\mathbf{A}$ and observed variables as $\mathbf{Y}$.
+Let $\cup_{a=1}^{3}I_{a}=[J]$ denote the partition that corresponds to the double triangular structure in Definition S.3.
+As we assume that each $\mathbf{A}_{\mathcal{I}_{a}}^{(d)}$ are disconnected (or d-separated) by $\Lambda^{(d)}$, we have the conditional independence
+$\mathbf{Y}_{I_{1}}\perp\mathbf{Y}_{I_{2}}\perp\mathbf{Y}_{I_{3}}\mid\mathbf{A}.$
+Thus, the tensor decomposition in (S.5) still holds. We separate the proof into two steps below.
+Step 1: Kruskal’s theorem.
+Lemma S.6 ensures that $\mathbf{N}_{1},\mathbf{N}_{2}$ in the tensor decomposition (S.5) are full rank.
+Additionally, all $2^{K}$ columns in $\mathbf{N}_{3}$ are distinct.
+To see this, note that for any $\boldsymbol{\alpha}\neq\boldsymbol{\alpha}^{\prime}\in\{0,1\}^{K}$, there exists some $k$ such that $\alpha_{k}\neq\alpha_{k}^{\prime}$.
+By the assumption that $\mathbf{G}_{3}$ (corresponding to the rows indexed by $\mathcal{I}_{3}$) does not have empty columns, there exists some $j\in\mathcal{I}_{3}$ such that $A_{k}\to Y_{j}$.
+Then, as $\beta_{j,k}\neq 0$, $\mathbb{P}(Y_{j}\mid\mathbf{A}=\boldsymbol{\alpha})\neq\mathbb{P}(Y_{j}\mid\mathbf{A}=\boldsymbol{\alpha}^{\prime})$.
+Hence, the corresponding columns in $\mathbf{N}_{3}$ are distinct.
+Hence, (S.14) holds.
+Applying Kruskal’s theorem shows that the decomposition (S.5) is unique up to column permutations.
+In other words, $(\mathbf{N}_{a},\boldsymbol{\pi})$ can be recovered up to a common column permutation $\mathfrak{S}\in S_{\{0,1\}^{K}}$.
+For more details, see Step 1 in the proof of Proposition 1.
+Step 2: structuring the permutation $\mathfrak{S}$.
+The above argument recovers $\mathbb{P}(\mathbf{Y}\mid\mathbf{A})$ and $\mathbb{P}(\mathbf{A})$ up to a common column permutation $\mathfrak{S}$ for the latent configurations $\mathbf{A}$.
+Now, it suffices to structure $\mathfrak{S}$ up to the trivial label permutation ambiguities.
+Then, as the probability distribution $\mathbb{P}(\mathbf{Y},\mathbf{A})$ is faithful to the underlying DAG $\mathbf{G}\cup\Lambda$, we can identify all components up to Markov equivalence.
+This allows us to construct the graphical structure up to the completed partially DAG (CPDAG).
+Directions can be assigned to the undirected edges using (i) the assumption that there are no directed edges of the type $A_{k}\to Y_{j}$, (ii) the linearity of the conditional distributions in (S.28), excluding the fundamental ambiguities regarding tree structures among $\mathbf{Y}$ in the observed graph $\Lambda$.
+Without loss of generality, let $Y_{1},\ldots,Y_{J_{1}}$ be the observations corresponding to $\mathcal{I}_{1}$, that are indexed via topological ordering under true graph $\Lambda$.
+Index the latent variables $\mathbf{A}$ according to the triangular graphical structure in Definition S.3, and let ${j_{k}}:=\min\{j\in\mathcal{I}_{1}:A_{k}\to Y_{j}\}$ denote the child of $A_{k}$ with the largest index.
+Note that we assume $\text{{pa}}(\mathbf{Y}_{\mathcal{I}_{1}})\supseteq\mathbf{A}$, so $j_{k}$ is well-defined.
+We prove by backwards induction and show that we can recover the set $T_{k}=\{\boldsymbol{\alpha}:\alpha_{k}=1\}$ only from the conditional probability matrix $\mathbf{N}_{1}$ and $\boldsymbol{\pi}$ (without the correct column labels).
+For the base case with $k=K$, $Y_{j_{K}}$ has an unique latent parent $A_{K}$ (excluding the observed parents in $\mathbf{Y}_{1:j_{K}-1}$).
+Hence, for any $\mathbf{y}_{1:j_{K}}\in\{0,1\}^{j_{K}}$,
+$\mathbb{P}(Y_{j_{K}}=y_{j_{K}}\mid\mathbf{A}=\boldsymbol{\alpha},\mathbf{Y}_{1:j_{K}-1}=\mathbf{y}_{1:j_{K}-1})=\mathbb{P}(Y_{j_{K}}=y_{j_{K}}\mid A_{K}=\alpha_{K},\mathbf{Y}_{1:j_{K}-1}=\mathbf{y}_{1:j_{K}-1}).$
+By fixing $\mathbf{y}=\mathbf{0}$ and considering $\boldsymbol{\alpha}\in\{0,1\}^{K}$, the above takes exactly two values, depending on $\alpha_{K}=0/1$.
+As the LHS of the above display can be computed from $\mathbf{N}_{1}$ and $\boldsymbol{\pi}$, $T_{K}$ must correspond to the column indices of $\mathbf{N}_{1}$ with the larger value.
+Here, we have used the monotonicity assumption that $\beta_{j_{K},K}>0$.
+Next, we recover $T_{k}$, assuming that we are given $T_{l}$ for all $k<l\leq K$.
+We similarly proceed by noting that $Y_{j_{k}}$ must have $A_{k}$ as a latent parent, and other latent parents are a subset of $\mathbf{A}_{k+1:K}.$ This implies that for any $\mathbf{y}_{1:j_{k}}\in\{0,1\}^{j_{k}}$,
+$\mathbb{P}(Y_{j_{k}}=y_{j_{k}}\mid\mathbf{A}=\boldsymbol{\alpha},\mathbf{Y}_{1:j_{k}-1}=\mathbf{y}_{1:j_{k}-1})=\mathbb{P}(Y_{j_{k}}=y_{j_{k}}\mid\mathbf{A}_{k:K}=\boldsymbol{\alpha}_{k:K},\mathbf{Y}_{1:j_{k}-1}=\mathbf{y}_{1:j_{k}-1}).$
+By fixing $\mathbf{y}_{1:j_{k}}=\mathbf{0}$ and considering any $\boldsymbol{\alpha}\in\{0,1\}^{K}$ with $\boldsymbol{\alpha}_{k+1:K}=\mathbf{0}$, the above takes exactly two values, depending on $\alpha_{k}=0/1$.
+Similar as above, the LHS can be computed from $\mathbf{N}_{1},\boldsymbol{\pi}$, and hence $T_{k}$ is recovered by selecting the column indices of $\mathbf{N}_{1}$ corresponding to a larger value.
+This completes the induction, and we have recovered all $T_{k}$s.
+This completes the proof.
+∎
+We finally prove Lemma S.6 using the following key property, which allows us to ignore redundant observables within $\mathbf{Y}_{\mathcal{I}}$.
+Lemma S.7.
+For any $\mathcal{I}\subset\mathcal{I}^{\prime}\subseteq[J]$ we have $\text{{rank}}(\mathbb{P}(\mathbf{Y}_{\mathcal{I}}\mid\mathbf{A}))\leq\text{{rank}}(\mathbb{P}(\mathbf{Y}_{\mathcal{I}^{\prime}}\mid\mathbf{A}))$.
+The proof directly follows by the fact that $\text{{rank}}(\mathbf{M}\mathbf{N})\leq\text{{rank}}(\mathbf{N})$ for compatible matrices, as one can recover $\mathbb{P}(\mathbf{Y}_{\mathcal{I}}\mid\mathbf{A})$ by appropriately marginalizing out rows in $\mathbb{P}(\mathbf{Y}_{\mathcal{I}^{\prime}}\mid\mathbf{A})$.
+Proof of Lemma S.6.
+For notational convenience, let $\mathcal{I}=\{1,\ldots,J\}$ and suppose that $\mathbf{Y}_{\mathcal{I}}$ is indexed via topological ordering so that $Y_{j}\to Y_{j^{\prime}}$ if and only if $j\leq j^{\prime}$.
+We proceed in an induction on $K$.
+First, the case when $K=1$ is straightforward.
+Assuming that the statement holds when there are K1K-1 latent variables, we show the claim for models with $K$ latent variables.
+Suppose that we have observed variables $Y_{1},\ldots,Y_{J}$ and latent variables $A_{1},\ldots,A_{K}$.
+For each $k$, let $Y_{j_{k}}$ denote the minimum element in $\text{{ch}}(A_{k})$.
+Index the latent variables $\mathbf{A}$ so that $j_{k}<j_{l}$ for $k<l$.
+By Lemma S.7, we can assume without loss of generality that $j_{1}=1$, since ignoring the observations $Y_{j}$s for $j<j_{1}$ do not increase the rank of $\mathbb{P}(\mathbf{Y}\mid\mathbf{A})$.
+We show that the $2^{J}\times 2^{K}$ matrix $\mathbf{P}:=\mathbb{P}(Y_{1},\ldots,Y_{J}\mid A_{1},\ldots,A_{K})$ has full column rank (of $2^{K}$).
+Fix the row/column ordering of $\mathbf{P}$ by mapping each binary configuration $(a_{1},\ldots,a_{K})$ to an integer $\sum_{k=1}^{K}a_{k}2^{k-1}$ and assume that this integer is increasing (e.g.
+$(0,0)<(1,0)<(0,1)<(1,1)$).
+Let $\theta_{\boldsymbol{\alpha}}:=\mathbb{P}(Y_{1}=1\mid\mathbf{A}=\boldsymbol{\alpha})$ denote the conditional probability of $Y_{1}=1$ given its parents.
+Also, define conditional probability matrices
+$\mathbf{M}_{0}:=\mathbb{P}(\mathbf{Y}_{2:J}\mid\mathbf{A}_{2:K}=\boldsymbol{\alpha}_{2:K},Y_{1}=0),\quad\mathbf{M}_{1}:=\mathbb{P}(\mathbf{Y}_{2:J}\mid\mathbf{A}_{2:K}=\boldsymbol{\alpha}_{2:K},Y_{1}=1).$
+Our main observation is that, under the topologically triangular assumption, we have the conditional independence $A_{1}\perp\mathbf{Y}_{2:J}\mid Y_{1},\mathbf{A}_{2:K}$. Using this, we have
+$\mathbb{P}(\mathbf{Y}\mid\mathbf{A})=\mathbb{P}(Y_{1}\mid\mathbf{A})\mathbb{P}(\mathbf{Y}_{2:J}\mid Y_{1},\mathbf{A}_{2:K}),$
+which gives the following decomposition of $\mathbf{P}$:
+$\mathbf{P}\Big{(}(0,\mathbf{y}),(0,\boldsymbol{\alpha})\Big{)}=\mathbf{M}_{0}(\mathbf{y},\boldsymbol{\alpha})(1-\theta_{0,\boldsymbol{\alpha}}),\quad$ $\mathbf{P}\Big{(}(0,\mathbf{y}),(1,\boldsymbol{\alpha})\Big{)}=\mathbf{M}_{0}(\mathbf{y},\boldsymbol{\alpha})(1-\theta_{1,\boldsymbol{\alpha}}),$
+$\mathbf{P}\Big{(}(1,\mathbf{y}),(0,\boldsymbol{\alpha})\Big{)}=\mathbf{M}_{1}(\mathbf{y},\boldsymbol{\alpha})\theta_{0,\boldsymbol{\alpha}},\quad$ $\mathbf{P}\Big{(}(1,\mathbf{y}),(1,\boldsymbol{\alpha})\Big{)}=\mathbf{M}_{1}(\mathbf{y},\boldsymbol{\alpha})\theta_{1,\boldsymbol{\alpha}}.$
+Let $\mathbf{p}_{(1,\boldsymbol{\alpha})}$ and $\mathbf{p}_{(0,\boldsymbol{\alpha})}$ denote the corresponding columns in $\mathbf{P}$.
+Let $\bar{\mathbf{P}}$ denote the matrix by performing the following column operation on $\mathbf{P}$: for each $\boldsymbol{\alpha}\in\{0,1\}^{K-1}$, subtract $\frac{1-\theta_{1,\boldsymbol{\alpha}}}{1-\theta_{0,\boldsymbol{\alpha}}}\mathbf{p}_{(0,\boldsymbol{\alpha})}$ from $\mathbf{p}_{(1,\boldsymbol{\alpha})}$.
+Then, we have
+$\bar{\mathbf{P}}\Big{(}(0,\mathbf{y}),(0,\boldsymbol{\alpha})\Big{)}=\mathbf{M}_{0}(\mathbf{y},\boldsymbol{\alpha})(1-\theta_{0,\boldsymbol{\alpha}}),\quad$ $\bar{\mathbf{P}}\Big{(}(0,\mathbf{y}),(1,\boldsymbol{\alpha})\Big{)}=0,$
+$\bar{\mathbf{P}}\Big{(}(1,\mathbf{y}),(0,\boldsymbol{\alpha})\Big{)}=\mathbf{M}_{1}(\mathbf{y},\boldsymbol{\alpha})\theta_{0,\boldsymbol{\alpha}},\quad$ $\bar{\mathbf{P}}\Big{(}(1,\mathbf{y}),(1,\boldsymbol{\alpha})\Big{)}=\mathbf{M}_{1}(\mathbf{y},\boldsymbol{\alpha})\frac{\theta_{1,\boldsymbol{\alpha}}-\theta_{0,\boldsymbol{\alpha}}}{1-\theta_{0,\boldsymbol{\alpha}}}.$
+Since $\text{{rank}}(\mathbf{P})=\text{{rank}}(\bar{\mathbf{P}})$ and viewing $\bar{\mathbf{P}}$ as a $2\times 2$ block matrix, it suffices to show that the conditional probability matrices $\mathbf{M}_{0},\mathbf{M}_{1}$ have full column rank.
+Here, we use the fact that $\theta_{1,\boldsymbol{\alpha}}-\theta_{0,\boldsymbol{\alpha}}\neq 0$ for all $\boldsymbol{\alpha}\in\{0,1\}^{K-1}$ under our parametrization for $Y_{1}\mid\mathbf{A}$ (see (S.28)), since the triangular assumption gives $g_{1,1}=1$ (in other words, $A_{1}\to Y_{1}$).
+Finally, the full rankness of $\mathbf{M}_{0},\mathbf{M}_{1}$ follows from the induction hypothesis.
+To see this, use (S.28) to spell out the conditional distribution of $Y_{2}\mid\mathbf{A}_{\mathbf{A}_{2:K}},Y_{1}$:
+$\mathbb{P}(Y_{2}=1\mid\mathbf{A}_{{2:K}},Y_{1}=0)$ $=\operatorname{\emph{g}_{\text{logistic}}}(\beta_{2,0}+\sum_{k=2}^{K}\beta_{2,k}A_{k}),$
+$\mathbb{P}(Y_{2}=1\mid\mathbf{A}_{{2:K}},Y_{1}=1)$ $=\operatorname{\emph{g}_{\text{logistic}}}(\beta_{2,0}+\delta_{1,2}+\sum_{k=2}^{K}\beta_{2,k}A_{k}).$
+Each conditional distributions above can be viewed as that arising from separate reduced models with K1K-1 latent variables $\mathbf{A}_{2:K}$, where the second equation considers a combined intercept parameter $\beta_{2,0}+\delta_{1,2}$.
+The same logic applies to all conditional distributions $Y_{k}\mid\mathbf{A}_{k:K},Y_{2:k-1},Y_{1}$, allowing us to use the induction hypothesis.
+∎
+S.1.6 Additional Identifiability Results
+Identifying the latent dimension.
+Our main identifiability results in the main theorems have assumed that the latent dimension $\mathcal{K}$ is known.
+Here, we illustrate that one can additionally establish the identifiability of $\mathcal{K}$ under a weaker notion of identifiability.
+To elaborate, we identify $\mathcal{K}$ under the class of DDEs that satisfy the two-pure-children condition A.
+Theorem S.2 (Modification of Theorem 1 in Lee and Gu, (2025)).
+Assume a $D$-latent layer DDEs satisfying the two-pure-children condition A (see Theorem 3.1), where $D$ is given but $\mathcal{K}$ is unknown.
+Then, the number of latent variables $\mathcal{K}$ is identifiable.
+The above result follows directly from a more general claim from Lee and Gu, (2025) alongside the layerwise identifiability in the proof of Theorem 1.
+Detour: Identifiability of one-latent-layer saturated models with interaction effects of binary latent variables
 One may ask whether the linear/additive parametrization $\beta_{j,0}+\sum_{k\in[K]}\beta_{j,k}A_{k}$ in the one-latent-layer saturated model is necessary for its identifiability.
@@
 For Binary responses, we set
-$\widehat{y}_{K^{(1)}}(i,j)=\begin{cases}\epsilon,&\text{ if }y_{K^{(1)}}(i,j)=0,\\ 1-\epsilon,&\text{ if }y_{K^{(1)}}(i,j)=1.\end{cases}$
+[…]
 For Poisson responses, we set
@@
 Here, we consider the same three parametric families as the main paper, the identifiable true parameter values $\mathcal{B}_{s}$ (see eq.
-(13)), and two sample sizes $N=1000,4000$.
+(S.36)), and two sample sizes $N=1000,4000$.
 We set the random initialization as follows:
@@
 $B^{(1)}_{j,0},B^{(2)}_{k,0}$ $\sim\text{Unif}(-1,0),\quad\gamma_{j}\sim\text{Unif}(0.5,1.5),$
-where Unif$(a,b)$ denotes the uniform distribution on the interval $(a,b)$.
-ParFam Accuracy($\mathcal{G}$) RMSE$(\boldsymbol{\Theta})$ Time (s) # iterations
-Initialization \$N$ 1000 4000 1000 4000 1000 4000 1000 4000
+where Unif(a,b)(a,b) denotes the uniform distribution on the interval $(a,b)$.
+[…]
+[…]
 Bernoulli Random 0.617 0.547 1.37 1.32 23.6 48.4 20.1 27.11
@@
 Spectral 0.996 1 0.13 0.06 1.2 3.0 4.1 4.4
-Table S.1: Accuracy measures for $\mathcal{G}$ and $\boldsymbol{\Theta}$, computation time and iterations for 2-layer DDE estimates under different initializations.
-For the Accuracy$(\mathcal{G})$ column, larger is better.
-For the other columns, smaller is better.
+[…]
 The results in Table S.1 clearly illustrates that random initialization does not effectively converge to the true parameter values.
@@
 S.3 Additional Algorithm Details
-S.3.1 Detailed Update Formulas
-Simplified M-step for the PEM Algorithm.
-We display more explicit expressions for the M-step in Algorithm 1.
-Recall that we need to solve three optimizations detailed in equations (8)-(10).
-By plugging-in the distributional formulas from Section 2, we can simplify each equation as follows.
-The computations are straightforward and we just display the simplified formulas.
-First, the maximization in (8) simplifies to updating
-$p_{l}^{[t+1]}=\frac{\sum_{i,\operatorname{\boldsymbol{\alpha}^{(1)}},\operatorname{\boldsymbol{\alpha}^{(2)}}2}\alpha_{l}^{(2)}\varphi^{[t+1]}_{i,\operatorname{\boldsymbol{\alpha}^{(1)}},\operatorname{\boldsymbol{\alpha}^{(2)}}2}}{N},$
-for $l\in[K^{(2)}]$. The maximization in (9) simplifies to maximizing the $k$th row of $\mathbf{B}^{(2)}$ for each $k\in[K^{(1)}]$:
-${\boldsymbol{\beta}}_{k}^{(2),[t+1]}$
-$=$ $\underset{{\boldsymbol{\beta}}^{(2)}\in\mathbb{R}^{K^{(2)}+1}}{\operatorname*{argmax}}\sum_{i,\operatorname{\boldsymbol{\alpha}^{(1)}},\operatorname{\boldsymbol{\alpha}^{(2)}}2}\log\mathbb{P}(\mathbf{A}_{i,k}=\alpha_{k}^{(1)};g_{\text{logistic}}(\beta_{0}^{(2)}+\sum_{l=1}^{K^{(2)}}\alpha_{l}^{(2)}\beta_{l}^{(2)}))\varphi_{i,\operatorname{\boldsymbol{\alpha}^{(1)}},\operatorname{\boldsymbol{\alpha}^{(2)}}2}^{[t+1]}-p_{\lambda_{2},\tau}({\boldsymbol{\beta}}^{(2)}).$
-Finally, the maximization in (10) boils down to maximizing the $j$th row of $\mathbf{B}^{(1)}$ and $\gamma_{j}$ for each $j\in[J]$:
-$({\boldsymbol{\beta}}_{j}^{(1),[t+1]},\gamma_{j}^{[t+1]})$
-$=$ $\underset{{\boldsymbol{\beta}}^{(1)}\in\mathbb{R}^{K^{(1)}+1},\gamma>0}{\operatorname*{argmax}}~{}\sum_{i,\operatorname{\boldsymbol{\alpha}^{(1)}},\operatorname{\boldsymbol{\alpha}^{(2)}}2}\log\mathbb{P}(Y_{i,j};\beta_{0}^{(1)}+\sum_{k=1}^{K^{(1)}}\alpha_{k}^{(1)}\beta_{k}^{(1)},\gamma)\varphi_{i,\operatorname{\boldsymbol{\alpha}^{(1)}},\operatorname{\boldsymbol{\alpha}^{(2)}}2}^{[t+1]}-p_{\lambda_{1,N},\tau}({\boldsymbol{\beta}}^{(1)}),$
+S.3.1 Penalized EM Algorithm
+This supplement presents the details of the standard penalized EM algorithm (PEM) that was mentioned in the main text, which does not use stochastic approximation.
+The PEM computes the penalized maximum likelihood estimator in (6) to estimate the potentially sparse coefficients $\mathbf{B}^{(1)}$ and $\mathbf{B}^{(2)}$.
+The PEM is an iterative procedure that consists of an expectation step followed by a penalized maximization step (Green, , 1990; Chen et al., , 2015).
+In the $(t+1)$th iteration, the E-step computes the expectation of the complete data penalized-log-likelihood
+$\ell_{c}(\mathbf{Y},\mathbf{A}^{(1)},\mathbf{A}^{(2)};\boldsymbol{\Theta})$- $\sum_{d=1}^{2}p_{\lambda_{N},\tau_{N}}(\mathbf{B}^{(d)})=~\sum_{i=1}^{N}\Bigg{[}\log\mathbb{P}(\mathbf{A}_{i}^{(2)};{\boldsymbol{p}})+\log\mathbb{P}(\mathbf{A}_{i}^{(1)}\mid\mathbf{A}_{i}^{(2)};\mathbf{B}^{(2)})$
+$+\log\mathbb{P}(\mathbf{Y}_{i}\mid\mathbf{A}_{i}^{(1)},\mathbf{A}_{i}^{(2)};\mathbf{B}^{(1)},\boldsymbol{\gamma})\Bigg{]}-\sum_{d=1}^{2}p_{\lambda_{N},\tau_{N}}(\mathbf{B}^{(d)}).$
+This requires calculating the conditional probability for each latent configuration using the previous parameter estimates; that is, calculating $\mathbb{P}(\mathbf{A}_{i}^{(1)}=\boldsymbol{\alpha}^{(1)},\mathbf{A}_{i}^{(2)}=\boldsymbol{\alpha}^{(2)}\mid\mathbf{Y};\boldsymbol{\Theta}^{[t]})$ for all $i\in[N]$, $\boldsymbol{\alpha}^{(1)}\in\{0,1\}^{K_{1}}$, and $\boldsymbol{\alpha}^{(2)}\in\{0,1\}^{K_{2}}$.
+Here, the superscript $[t]$ denotes the $t$th iteration estimates.
+In the M-step, we update the parameters by maximizing the expectation computed in the E step, which boils down to solving the following three maximizations:
+${\boldsymbol{p}}^{[t+1]}:=$ $\operatorname*{argmax}_{{\boldsymbol{p}}}\sum_{i=1}^{N}\mathbb{E}\left[\log\mathbb{P}(\mathbf{A}_{i}^{(2)};{\boldsymbol{p}});{\boldsymbol{p}}^{[t]}\right],$ (S.32)
+$\mathbf{B}^{(2),[t+1]}:=$ $\operatorname*{argmax}_{\mathbf{B}^{(2)}}\sum_{i=1}^{N}\mathbb{E}\left[\log\mathbb{P}(\mathbf{A}_{i}^{(1)}\mid\mathbf{A}_{i}^{(2)};\mathbf{B}^{(2)});\mathbf{B}^{(2),[t]}\right]-p_{\lambda_{N},\tau_{N}}(\mathbf{B}^{(2)}),$ (S.33)
+$(\mathbf{B}^{(1),[t+1]},\boldsymbol{\gamma}^{[t+1]}):=$ $\operatorname*{argmax}_{\mathbf{B}^{(1)},\boldsymbol{\gamma}}\sum_{i=1}^{N}\mathbb{E}\left[\log\mathbb{P}(\mathbf{Y}_{i}\mid\mathbf{A}_{i}^{(1)},\mathbf{A}_{i}^{(2)});\mathbf{B}^{(1),[t]},\boldsymbol{\gamma}^{[t]})\right]-p_{\lambda_{N},\tau_{N}}(\mathbf{B}^{(1)}).$ (S.34)
+Here, the optimizations for each layer are separated; we update the top-layer proportion parameters $\boldsymbol{p}$} in (S.32), the middle latent layer coefficients $\mathbf{B}^{(2)}$ in (S.33), and the bottom layer coefficients $(\mathbf{B}^{(1)},\boldsymbol{\gamma})$ in (S.34).
+Additionally, due to the conditional independence assumption in each layer, the maximizations can be further simplified into low-dimensional optimizations over each row of $\mathbf{B}$.
+Algorithm 5 summarizes the PEM algorithm.
+Data: $\mathbf{Y},\mathcal{K}$, tuning parameters $\lambda_{N},\tau_{N}$.
+Initialize $\boldsymbol{\Theta}^{[0]}$ using the layerwise initialization in Algorithm 1.
+while log-likelihood has not converged do
+  In the $[t+1]$th iteration,
+ // E-step
+ for $(i,\boldsymbol{\alpha}^{(1)},\operatorname{\boldsymbol{\alpha}^{(2)}}2)\in[N]\times\{0,1\}^{K^{(1)}}\times\{0,1\}^{K^{(2)}}$ do
+$\varphi^{[t+1]}_{i,\operatorname{\boldsymbol{\alpha}^{(1)}},\operatorname{\boldsymbol{\alpha}^{(2)}}2}=\mathbb{P}(\mathbf{A}_{i}^{(1)}=\boldsymbol{\alpha}^{(1)},\mathbf{A}_{i}^{(2)}=\boldsymbol{\alpha}^{(2)}\mid\mathbf{Y};\boldsymbol{\Theta}^{[t]})$
+  end for
+ // M-step
+  update $\boldsymbol{\Theta}^{[t+1]}=({\boldsymbol{p}}^{[t+1]},\mathbf{B}^{(1),[t+1]},\mathbf{B}^{(2),[t+1]},\boldsymbol{\gamma}^{[t+1]})$ by solving (S.32)-(S.34).
+end while
+Estimate $\mathbf{G}$ based on the sparsity structure of $\widehat{\mathbf{B}}$ according to (7).
+Output : Estimated continuous parameters $\widehat{\boldsymbol{\Theta}}$ and graphical matrices $\widehat{\mathbf{G}}^{(1)}$, $\widehat{\mathbf{G}}^{(2)}$.
+Algorithm 5 Penalized EM (PEM) algorithm for the two-latent-layer DDE
+S.3.2 Detailed Update Formulas for Algorithm 2
+Continuing from Section 4, we describe Algorithm 2 under $D=2$ latent layers. The extension to more latent layers is straightforward and we omit the detailed updates for simplicity.
 Simplified simulation step for the SAEM Algorithm.
@@
 Simplified M-step for the SAEM Algorithm.
-In the main paper, we have motivated the SAEM M-step in terms of the parameter $\mathbf{B}^{(2)}$ (see (12)).
-Here, for completeness, we present the fully expanded formulas for updating each parameter.
+In the main paper, we have motivated the SAEM M-step in terms of the parameter $\mathbf{B}^{(2)}$ (see (9)). Here, for completeness, we present the fully expanded formulas for updating each parameter.
 First, for each $l\in[K^{(2)}]$, we update $p_{l}$ in closed form as follows:
@@
 $({\boldsymbol{\beta}}_{j}^{(1),[t+1]},\gamma_{j}^{[t+1]})$ $:=\operatorname*{argmax}_{{\boldsymbol{\beta}}_{j}^{(1)},\gamma_{j}}\left[Q_{j}^{(1),[t+1]}({\boldsymbol{\beta}}_{j}^{(1)},\gamma_{j})-p_{\lambda_{N},\tau_{N}}({\boldsymbol{\beta}}_{j}^{(1)})\right].$ (S.35)
-S.3.2 Estimating the Number of Latent Variables
+S.3.3 Alternatives for Estimating the Number of Latent Variables
 Recall that we have proposed a spectral-ratio estimator to select the latent dimension $\mathcal{K}$ in Section 4.2.
@@
     $K^{(1)}:=\operatorname*{argmin}_{k\in\mathfrak{K}}\text{EBIC}(k).$
-    We postpone the details of the formula of EBIC$(k)$ to Section S.3.4.
+    We postpone the detailed formula of EBIC(k)(k) to the end of this subsection.
 Here, we consider EBIC instead of more traditional information criteria such as AIC and BIC since it is designed to handle large parameter spaces.
@@
 For analyzing real data, we recommend making the final selection of $\mathcal{K}$ by also incorporating qualitative aspects of the data such as domain knowledge and interpretability.
-S.3.3 Implementation Details
-We elaborate on the choices made to implement the EM Algorithms 1 and 3.
+Definition of EBIC.
+[…]
+$EBIC(\mathcal{K}):=-2\ell_{\mathcal{K}}(\widehat{\boldsymbol{\Theta}}_{\mathcal{K}};\mathbf{Y})+\text{df}(\mathcal{K})\log N+2{|\boldsymbol{\Theta}|\choose\text{df}(\mathcal{K})}$.
+S.3.4 Implementation Details
+We elaborate on the choices made to implement the EM Algorithms 5 and 2.
 Tuning parameter selection.
@@
 Preliminary simulations revealed that the results are very similar under other penalties such as SCAD.
-Based on the consistency result in Theorem 3, we consider $\lambda_{N}=N^{1/4},\tau_{N}=3N^{-0.3}$ for the simulation studies.
+Based on the consistency result in Theorem 3, we consider $\lambda_{N}=N^{1/4},\tau_{N}=\max(3N^{-0.3},0.3)$ for the simulation studies.
 For real data analysis, tuning parameters were selected from the following grid:
@@
 Under the spectral initialization, the PEM and SAEM generally took less than 10 and 5 iterations until convergence, respectively.
-In terms of implementing the PEM algorithm to select the number of latent variables, $K^{(1)}$ or $K^{(2)}$, we consider a more generous threshold of $5003\times N/$ 500.
+In terms of implementing the PEM algorithm to select the number of latent variables, $K^{(1)}$ or $K^{(2)}$, we consider a more generous threshold of $3\times N/500$.
 This is for the sake of faster implementation, as we only use the resulting likelihood to implement the EBIC and LRT method.
@@
 This is because there are already other approximations being made in the SAEM, and we did not want to increase the randomness for our simulation studies.
-For practical implementation for a larger dataset, we recommend the user to modify the M-step to a faster but approximate optimization procedure.
+For the simulation studies under deeper models with $D\geq 3$ (see Section 5 in the main paper), we have slightly modified the M-step of the Normal observed-layer (see (S.35)) by considering the hard-thresholding penalty $p_{\tau}(b):=\frac{\tau^{2}}{2}I(|b|\neq 0)$.
+This leads to a closed-form simplification of (S.35), as spelled-out below.
+Letting $\mathbf{y}_{j}$ denote the $j$th column of $\mathbf{Y}$, $\mathbf{A}^{[t+1]}_{\text{long}}:=(\mathbf{1}_{N},\mathbf{A}^{(1),[t+1]})$, and $\text{thres}_{\tau}(b):=bI(|b|>\tau)$, we can write the updates for each $\beta_{j},\gamma_{j}$ as
+$\widehat{{\boldsymbol{\beta}}}_{j}^{(1),[t+1]}=\text{thres}_{\tau}\Big{(}(\mathbf{A}^{[t+1]\top}_{\text{long}}\mathbf{A}^{[t+1]}_{\text{long}})^{-1}\mathbf{A}^{[t+1]\top}_{\text{long}}\mathbf{y}_{j}\Big{)},\quad\widehat{\gamma}_{j}^{[t+1]}=\sqrt{\frac{\sum_{i}(Y_{i,j}-\eta_{j,\mathbf{A}_{i}^{(1),[t+1]}})^{2}}{N}},$
+which speeds up the SAEM algorithm.
+For practical implementation for a larger dataset with general observed-layer distributions, we recommend the user to modify the M-step to a faster but approximate optimization procedure.
 For example, one may choose to do a one-step gradient ascent in each iteration of the M-step.
-Additional latent variable permutation for simulations.
+Addressing latent variable permutation for simulations.
 In our simulations, resolving the latent variable permutation is necessary to accurately compute the errors.
@@
 This method ensures computational efficiency while accurately resolving the latent variable permutation.
-S.3.4 Definition of EBIC
-Consider a 2-latent layer DDE with $\mathcal{K}=(K^{(1)},K^{(2)})$ latent variables.
-This is a parametric model with $|\boldsymbol{\Theta}|=J(K^{(1)}+1)+K^{(1)}(K^{(2)}+1)+K^{(2)}+J$ (if there exists a dispersion parameter) or $J(K^{(1)}+1)+K^{(1)}(K^{(2)}+1)+K^{(2)}$ (otherwise) parameters.
-Here, we do not count the discrete parameters $\mathcal{G}$, as they are implied by the coefficients $\mathcal{B}$.
-Let $\widehat{\boldsymbol{\Theta}}_{\mathcal{K}}$ and $\ell_{\mathcal{K}}(\widehat{\boldsymbol{\Theta}}_{\mathcal{K}};\mathbf{Y})$ respectively denote the MLE and the log-likelihood under the 2-layer DDE with $\mathcal{K}$ parameters.
-Also, let df$(\mathcal{K})$ be the number of non-zero parameters in $\widehat{\boldsymbol{\Theta}}_{\mathcal{K}}$.
-Then, following Chen and Chen, (2008), the EBIC objective in Section S.3.2 is defined as follows:
-$EBIC(\mathcal{K}):=-2\ell_{\mathcal{K}}(\widehat{\boldsymbol{\Theta}}_{\mathcal{K}};\mathbf{Y})+\text{df}(\mathcal{K})\log N+2{|\boldsymbol{\Theta}|\choose\text{df}(\mathcal{K})}.$
 S.4 Additional Simulation Results
-S.4.1 Simulation Results Under the Generic Identifiable Parameters $\mathcal{B}_{g}$ and Computation Time
+S.4.1 True Parameter Values
+In terms of the true parameter values, we consider two sets of values based on the strict and generic identifiability conditions in Theorems 1 and 2, respectively.
+First, we define $\mathcal{B}_{s}=\{\mathbf{B}_{s}^{(d)}\}_{d\in[D]}$ that satisfy Theorem 1 as follows:
+$\mathbf{B}_{s}^{(d)}=\begin{pmatrix}-2\mathbf{1}_{K^{(d)}}&4\mathbf{I}_{K^{(d)}}\\ -4\mathbf{1}_{K^{(d)}}&4\mathbf{I}_{K^{(d)}}\\ -2\mathbf{1}_{K^{(d)}}&\mathbf{B}_{1}^{(d)}\end{pmatrix},\quad\text{where}\quad\beta_{1;j,k}^{(d)}:=\begin{cases}4&\text{if }k=j,\\ 4/3&\text{ else if }|k-j|=K^{(d)}/2,\\ 0&\text{otherwise}.\end{cases}$ (S.36)
+We also consider $\mathcal{B}_{g}=\{\mathbf{B}_{g}^{(d)}\}_{d\in[D]}$ that satisfy Theorem 2, defined as:
+[…]
+and $\mathbf{B}_{1}^{(d)}$ is the matrix defined in Equation S.36.
+Note that $\mathbf{B}_{g}^{(d)}$ is a less sparse version of $\mathbf{B}_{s}^{(d)}$, where the identity matrices are modified to $\mathbf{B}_{2}^{(d)}$.
+While $\mathbf{B}_{2}^{(d)}$ and $\mathcal{B}_{g}$ have the same value of 4 on the main diagonals, $\mathcal{B}_{s}$ is sparser.
+In particular, $\mathcal{B}_{s}$ has two pure children per latent variable whereas $\mathcal{B}_{g}$ has none.
+Regarding the remaining parameters, we set the top-layer proportion parameters as $p_{k}=0.5$ for all $k\in[K^{(D)}]$, and set the dispersion parameters for the Normal distribution as $\gamma_{j}=\sigma_{j}^{2}=1$ for all $j\in[J]$.
+S.4.2 Additional Details on Ablation Studies
+We provide additional details regarding ablation studies.
+We separately considered data generated from a DDE and a DBN with the same coefficients $\mathcal{B}_{s}$ in (S.36).
+We implemented the DBN using the Deep Neural Network toolbox in MATLAB (Tanaka and Okutomi, , 2014).
+In addition to the analysis in the main text, we compare the RMSE values for the continuous parameters under sparse, identifiable models (panel (d) in Figure 5) versus that under fully-connected, non-identifiable models (Figure S.3 below).
+Interestingly, the RMSE values returned by the DBN algorithm are similar.
+This indicates the unreliability of parameter estimation via estimation algorithms for learning DBNs.
+In contrast, our proposed estimation method for DDEs encourages sparsity as well as adopts a wiser initialization strategy that is close to the true parameters, which leads to better parameter estimation and structure recovery.
+Figure S.3: Parameter estimation accuracy for fully-connected DBNs with $J=K^{(1)}=K^{(2)}=18$. The true coefficients are generated iid from $N(0,2)$.
+S.4.3 Simulation Results Under the Generic Identifiable Parameters $\mathcal{B}_{g}$ and Computation Time
 Here, we present the omitted details from Section 5 in the main paper, regarding (a) the estimation accuracy under generically identifiable parameters, (b) performance of the PEM algorithm, and (c) computation time.
 Estimation accuracy under generically identifiable parameters.
-We present the analogs of Figure 3 under the generically identifiable true parameters $\mathcal{B}_{g}$ defined as:
-$\mathbf{B}_{g}^{(d)}=\begin{pmatrix}-2\mathbf{1}_{K^{(d)}}&\mathbf{B}_{2}^{(d)}\\ -4\mathbf{1}_{K^{(d)}}&\mathbf{B}_{2}^{(d)\top}\\ -2\mathbf{1}_{K^{(d)}}&\mathbf{B}_{1}^{(d)}\end{pmatrix},\quad\text{where}\quad\beta_{2;j,k}^{(d)}$ :=$:=\begin{cases}4&\text{if }k=j,\\ 4/3&\text{ else if }0<k-j\leq\lceil K^{(d)}/3\rceil,\\ 0&\text{otherwise},\end{cases}$ (S.31)
-and $\mathbf{B}_{1}^{(d)}$ is the matrix defined in Equation 13.
-Note that $\mathbf{B}_{g}^{(d)}$ is a less sparse version of $\mathbf{B}_{s}^{(d)}$, where the identity matrices are modified to $\mathbf{B}_{2}^{(d)}$.
-While $\mathbf{B}_{2}^{(d)}$ and $\mathcal{B}_{g}$ have the same value of 4 on the main diagonals, $\mathcal{B}_{s}$ is sparser.
-In particular, $\mathcal{B}_{s}$ has two exclusive children per latent variable whereas $\mathcal{B}_{g}$ has none.
+We present the analogs of Figure 4 under the generically identifiable true parameters $\mathcal{B}_{g}$.
 While all other simulation settings are identical to that described in the main text, we make the following two changes.
@@
 This is because the layerwise initialization turned out to be less effective in this high-dimensional but less-sparse scenario.
-Second, for Poisson-based-DDEs, we modify the intercept values to be smaller compared to the values in (13) as follows.
+Second, for Poisson-based-DDEs, we modify the intercept values to be smaller compared to the values in (S.36) as follows.
 When $K^{(2)}=2$, we consider a smaller intercept parameter for $\mathbf{B}_{g}^{(1)}$ as follows:
@@
 As a result, PEM was only applied in the low-dimensional scenario with with $(J,K^{(1)},K^{(2)})=(18,6,2)$.
-The results, presented in Tables S.3-S.5 (error of estimating $\mathcal{G}$) and S.6-S.8 (error of estimating $\boldsymbol{\Theta}$) in Section S.4.4, show that the PEM estimates exhibit a faster rate of convergence as $N$ increases.
+The results, presented in Tables S.6-S.8 (error of estimating $\mathcal{G}$) and S.9-S.11 (error of estimating $\boldsymbol{\Theta}$) in Section S.4.7, show that the PEM estimates exhibit a faster rate of convergence as $N$ increases.
 In particular, the RMSE values decrease at the parametric rate ${1}/{\sqrt{N}}$, as established in Theorem 3.
@@
 The legends indicate the value of $K^{(1)}$ and estimation algorithm.
-S.4.2 Experiments With Varying Numbers of Gibbs Samples in Algorithm 3
+S.4.4 Experiments With Varying Numbers of Gibbs Samples in Algorithm 2
 Here, we conduct additional experiments to assess the effect of the number of Gibbs samples, $C$, on both estimation accuracy and computation time.
@@
 Based on these findings, we selected $C=1$ as a baseline for the SAEM procedure.
-ParFam Accuracy($\mathcal{G}$) RMSE$(\boldsymbol{\Theta})$ time (s) # iterations
-# Gibbs $C$$N$ 1000 4000 1000 4000 1000 4000 1000 4000
+[…]
+[…]
 Bernoulli 1 0.916 0.962 0.43 0.36 3.1 10.0 3.0 3.6
@@
 For the other columns, smaller is better.
-S.4.3 Simulations for Selecting the Number of Latent Variables
+S.4.5 Simulations for Selecting the Number of Latent Variables
 We first evaluate the performance of the three estimators (denoted as EBIC, LRT, Spectral) for selecting $K^{(1)}$, as introduced in Section S.3.3, assuming the knowledge of $K^{(2)}$.
@@
 The equality in $2K^{(2)}\leq K^{(1)}$ is allowed to consider an equal number of underfitted/overfitted models.
-Figure S.3: Selection accuracy of $K^{(1)}$ under the two-latent-layer DDE with true parameters $\mathcal{B}_{s}$. The spectral ratio estimator shows near-perfect accuracy for large $N$.
+Figure S.6: Selection accuracy of $K^{(1)}$ under the two-latent-layer DDE with true parameters $\mathcal{B}_{s}$. The spectral ratio estimator shows near-perfect accuracy for large NN.
 Figure S.7: Simulation results for selecting $K^{(1)}$ under the 2-layer DDE with true parameters $\mathcal{B}_{g}$.
-We fit the three estimators $400$ times for each scenario, and display the correct selection percentage for the Bernoulli, Poisson, and Normal-based DDEs with true parameters $\mathcal{B}_{s}$/$\mathcal{B}_{g}$ in Figure S.3/Figure S.4.
+We fit the three estimators $400$ times for each scenario, and display the correct selection percentage for the Bernoulli, Poisson, and Normal-based DDEs with true parameters $\mathcal{B}_{s}$$\mathcal{B}_{g}$ in Figure S.6/Figure S.7.
 Comparing the three estimators, the spectral ratio-based estimator has near-perfect accuracy when $N$ is large enough, say $4000$.
@@
 The candidate set for $K^{(2)}$ is $\mathfrak{K}:=\{1,2,3\}$.
-Figure S.5/Figure S.6 displays the correct selection percentage under the true parameters $\mathcal{B}_{s}$/$\mathcal{B}_{g}$.
+Figure S.8/Figure S.9 displays the correct selection percentage under the true parameters $\mathcal{B}_{s}$$\mathcal{B}_{g}$.
 Here, we implement the LRT estimator by naively assuming that Wilk’s theorem holds.
@@
 Then, $K^{(2)}$ can be determined using either the EBIC or the spectral ratio estimator, incorporating domain knowledge if needed.
-S.4.4 Exact Values of Estimation Errors
+S.4.6 Simulations for Deeper Models with $D\geq 3$ Latent Layers
+Estimation accuracy of graphical structures
+We display the postponed tables from the the main text (Section 5) in Tables S.3 and S.4, which correspond to experiments for deeper models with Normal responses and $D=3,4$ latent layers.
+[…]
+$\mathbf{G}^{(1)}$ 1.00 1.00 1.00 1.00 1.00 1.00
+$\mathbf{G}^{(2)}$ 0.89 0.89 0.90 0.93 0.94 0.98
+$\mathbf{G}^{(3)}$ 0.87 0.89 0.88 0.89 0.90 0.91
+runtime (s) 3 3 3 5 8 15
+Table S.3: Average entrywise-accuracy of estimating graphical matrices and runtime when $D=3$.
+[…]
+$\mathbf{G}^{(1)}$ 1.00 1.00 1.00 1.00 1.00 1.00
+$\mathbf{G}^{(2)}$ 0.83 0.89 0.93 0.95 0.96 0.97
+$\mathbf{G}^{(3)}$ 0.72 0.74 0.75 0.77 0.78 0.80
+$\mathbf{G}^{(4)}$ 0.64 0.65 0.70 0.70 0.71 0.73
+runtime (s) 75 96 74 89 186 312
+Table S.4: Average entrywise-accuracy of estimating graphical matrices and runtime when $D=4$.
+Selecting the number of latent variables
+To assess our spectral-ratio estimator’s performance, we have conducted additional experiments for deeper models with $D=3$.
+Table S.5 illustrates the satisfactory performance of Algorithm 3 for all layers, where each entry reports the accuracy across 400 replications.We have also evaluated the spectral-ratio estimator for more challenging cases with $D\geq 4$ latent layers, where the accuracy degrades due to accumulation of uncertainty and larger grid size.
+[…]
+$K^{(1)}$ 100 100 100 100 100 100 14
+$K^{(2)}$ 28 44 68 79 87 92 5
+$K^{(3)}$ 65 71 82 89 92 95 2
+Table S.5: Correct selection percentage for $\mathcal{K}$, under a true model with Normal responses and $D=3,K^{(D)}=2$. The last column reports the typical grid size for each $K^{(d)}$.
+S.4.7 Exact Values of Estimation Errors
 We display the complete results of the simulation outputs corresponding to plots in Section 5, Section S.4.3, and Section S.4.5.
 Accuracy of Estimating $\mathcal{G}$.
-Tables S.3-S.5 display the average entrywise estimation accuracy for $\mathcal{G}$ under each parametric family. These values correspond to the upper rows of Figures 3 and S.1.
+Tables S.6-S.8 display the average entrywise estimation accuracy for $\mathcal{G}$ under each parametric family. These values correspond to the upper rows of Figures 4 and S.4.
 […]
@@
 Accuracy of Estimating $\boldsymbol{\Theta}$.
-Tables S.6S.8 reports the RMSE values for estimating continuous parameters $\boldsymbol{\Theta}$. These values correspond to the bottom rows of Figures 3 and S.1.
+Tables S.9S.11 reports the RMSE values for estimating continuous parameters $\boldsymbol{\Theta}$. These values correspond to the bottom rows of Figures 4 and S.4.
 […]
@@
 Accuracy of selecting $\mathcal{K}$.
-Tables S.9-S.10 reports the accuracy of selecting $\mathcal{K}$ under Bernoulli and Poisson-based DDEs, which correspond to Figure S.3 and Figure S.4.
+Tables S.12-S.13 reports the accuracy of selecting $\mathcal{K}$ under Bernoulli and Poisson-based DDEs, which correspond to Figure S.6 and Figure S.7.
 Accuracy values for Normal-based DDEs are omitted, as all methods demonstrated near-perfect accuracy in this setting.
-Similarly, Tables S.11-S.12 display the results for selecting $K^{(2)}$, which corresponds to Figures S.5 and S.6.
+Similarly, Tables S.14-S.15 display the results for selecting $K^{(2)}$, which corresponds to Figures S.8 and S.9.
 These tables provide additional insights beyond those presented in figures, which only display the correct selection probability for the events $\widehat{K}^{(1)}=6$ and $\widehat{K}^{(2)}=2$.
 The percentage of incorrect estimates in the tables reveal systematic tendencies of the estimators: the EBIC and LRT estimators frequently overselect ($\widehat{K}^{(1)}\geq K^{(1)}$) whereas the spectral estimator sometimes underselects ($\widehat{K}^{(1)}\leq K^{(1)}$).
-ParFam, value $N$ Method \ $\widehat{K}^{(1)}$ 4 5 6 7 8
-Bernoulli, $\mathcal{B}_{s}$ 500 EBIC 0.25 96.25 3.5
-LRT 0.5 98.75 0.75
-Spectral 1.75 30.0 67.5 0.75
-1000 EBIC 94.5 5.5
-LRT 96.5 3.5
-Spectral 0.25 18.25 81.5
-2000 EBIC 88.25 11.75
-LRT 93.5 6.25 0.25
-Spectral 13.0 87.0
-4000 EBIC 86.0 12.25 1.75
-LRT 90.25 9.25 0.5
-Spectral 3.75 96.25
-6000 EBIC 86.0 13.5 1.5
-LRT 87.75 12.25
-Spectral 0.5 99.5
-8000 EBIC 87.75 11.0 1.25
-LRT 88.25 10.75 1.00
-Spectral 0.75 99.25
-Bernoulli, $\mathcal{B}_{g}$ 500 EBIC 84.5 13.75 1.75
-LRT 0.75 1.0 86.25 12.0
-Spectral 8.5 4.25 87.25
-1000 EBIC 87.25 10.5 2.5
-LRT 80.5 18.5 1.0
-Spectral 0.5 99.5
-2000 EBIC 85.25 11.5 3.25
-LRT 82.75 16.25 1.0
-Spectral 100.0
-4000 EBIC 97.0 3.0
-LRT 92.25 7.75
-Spectral 100.0
-6000 EBIC 97.5 2.5
-LRT 91.5 8.5
-Spectral 100.0
-8000 EBIC 98.0 2.0
-LRT 90.75 9.25
-Spectral 100.0
+ParFam, value NN Method \ $\widehat{K}^{(1)}$ 4 5 6 7 8
+Bernoulli, $\mathcal{B}_{s}$ 500 EBIC 0 0.25 96.25 3.5 0
+LRT 0.5 0 98.75 0.75 0
+Spectral 1.75 30.0 67.5 0.75 0
+1000 EBIC 0 0 94.5 5.5 0
+LRT 0 0 96.5 3.5 0
+Spectral 0.25 18.25 81.5 0 0
+2000 EBIC 0 0 88.25 11.75 0
+LRT 0 0 93.5 6.25 0.25
+Spectral 0 13.0 87.0 0 0
+4000 EBIC 0 0 86.0 12.25 1.75
+LRT 0 0 90.25 9.25 0.5
+Spectral 0 3.75 96.25 0 0
+6000 EBIC 0 0 86.0 13.5 1.5
+LRT 0 0 87.75 12.25 0
+Spectral 0 0.5 99.5 0 0
+8000 EBIC 0 0 87.75 11.0 1.25
+LRT 0 0 88.25 0 10.75 1.00
+Spectral 0 0.75 99.25 0 0
+Bernoulli, $\mathcal{B}_{g}$ 500 EBIC 0 0 84.5 13.75 1.75
+LRT 0.75 1.0 86.25 12.0 0
+Spectral 8.5 4.25 87.25 0 0
+1000 EBIC 0 0 87.25 10.5 2.5
+LRT 0 0 80.5 18.5 1.0
+Spectral 0.5 0 99.5 0 0
+2000 EBIC 0 0 85.25 11.5 3.25
+LRT 0 0 82.75 16.25 1.0
+Spectral 0 0 100.0 0 0
+4000 EBIC 0 0 97.0 3.0 0
+LRT 0 0 92.25 7.75 0
+Spectral 0 0 100.0 0 0
+6000 EBIC 0 0 97.5 2.5 0
+LRT 0 0 91.5 8.5 0
+Spectral 0 0 100.0 0 0
+8000 EBIC 0 0 98.0 2.0 0
+LRT 0 0 90.75 9.25 0
+Spectral 0 0 100.0 0 0
 Table S.12: Empirical distribution of the estimated $\widehat{K}^{(1)}$ values under the Bernoulli DDE with true $K^{(1)}=6$ and parameters $\mathcal{B}_{s},\mathcal{B}_{g}$.
@@
 “ParFam” is short for “parametric family”.
-ParFam, value $N$ Method \ $\widehat{K}^{(1)}$ 4 5 6 7 8
-Poisson, $\mathcal{B}_{s}$ 500 EBIC 99.5 0.5
-LRT 99.5 0.5
-Spectral 100.0
-1000 EBIC 99.0 1.0
-LRT 99.0 1.0
-Spectral 100.0
-2000 EBIC 97.5 2.5
-LRT 96.25 3.75
-Spectral 100.0
-4000 EBIC 98.25 1.5 0.25
-LRT 97.25 2.75
-Spectral 100.0
-6000 EBIC 99.0 1.0
-LRT 97.25 2.75
-Spectral 100.0
-8000 EBIC 98.75 0.75 0.5
-LRT 97.75 1.75 0.5
-Spectral 100.0
-Poisson, $\mathcal{B}_{g}$ 500 EBIC 92.0 7.75 0.25
-LRT 1.0 91.0 8.0
-Spectral 8.5 91.5
-1000 EBIC 90.0 9.5 0.5
-LRT 88.5 10.75 0.75
-Spectral 2.75 97.25
-2000 EBIC 90.5 8.5 1.0
-LRT 0.25 87.25 11.5 1.0
-Spectral 0.5 99.5
-4000 EBIC 91.25 7.25 1.5
-LRT 89.0 9.25 1.75
-Spectral 100.0
-6000 EBIC 95.0 4.5 0.5
-LRT 91.25 7.75 1.0
-Spectral 100.0
-8000 EBIC 94.0 5.0 1.0
-LRT 90.75 8.25 1.0
-Spectral 100.0
+ParFam, value NN Method \ $\widehat{K}^{(1)}$ 4 5 6 7 8
+Poisson, $\mathcal{B}_{s}$ 500 EBIC 0 0 99.5 0.5 0
+LRT 0 0 99.5 0.5 0
+Spectral 0 0 100.0 0 0
+1000 EBIC 0 0 99.0 1.0 0
+LRT 0 0 99.0 1.0 0
+Spectral 0 0 100.0 0 0
+2000 EBIC 0 0 97.5 2.5 0
+LRT 0 0 96.25 3.75 0
+Spectral 0 0 100.0 0 0
+4000 EBIC 0 0 98.25 1.5 0.25
+LRT 0 0 97.25 2.75 0
+Spectral 0 0 100.0 0 0
+6000 EBIC 0 0 99.0 1.0 0
+LRT 0 0 97.25 2.75 0
+Spectral 0 0 100.0 0 0
+8000 EBIC 0 0 98.75 0.75 0.5
+LRT 0 0 97.75 1.75 0.5
+Spectral 0 0 100.0 0 0
+Poisson, $\mathcal{B}_{g}$ 500 EBIC 0 0 92.0 7.75 0.25
+LRT 0 1.0 91.0 8.0 0
+Spectral 8.5 0 91.5 0 0
+1000 EBIC 0 0 90.0 9.5 0.5
+LRT 0 0 88.5 10.75 0.75
+Spectral 2.75 0 97.25 0 0
+2000 EBIC 0 0 90.5 8.5 1.0
+LRT 0 0.25 87.25 11.5 1.0
+Spectral 0.5 0 99.5 0 0
+4000 EBIC 0 0 91.25 7.25 1.5
+LRT 0 0 89.0 9.25 1.75
+Spectral 0 0 100.0 0 0
+6000 EBIC 0 0 95.0 4.5 0.5
+LRT 0 0 91.25 7.75 1.0
+Spectral 0 0 100.0 0 0
+8000 EBIC 0 0 94.0 5.0 1.0
+LRT 0 0 90.75 8.25 1.0
+Spectral 0 0 100.0 0 0
 Table S.13: Empirical distribution of the estimated $\widehat{K}^{(1)}$ values under the Poisson DDE with true $K^{(1)}=6$ and parameters $\mathcal{B}_{s},\mathcal{B}_{g}$.
@@
 “ParFam” is short for “parametric family”.
-ParFam, value $N$ Method \ $\widehat{K}^{(2)}$ 1 2 3
-Bernoulli, $\mathcal{B}_{s}$ 500 EBIC 0.5 99.5
+ParFam, value NN Method \ $\widehat{K}^{(2)}$ 1 2 3
+Bernoulli, $\mathcal{B}_{s}$ 500 EBIC 0.5 99.5 0
 LRT 16.75 83.0 0.25
 Spectral 12.0 85.5 2.5
-1000 EBIC 100.0
+1000 EBIC 0 100.0 0
 LRT 0.5 98.75 0.75
 Spectral 6.5 91.5 2.0
-2000 EBIC 100.0
-LRT 98.0 2.0
+2000 EBIC 0 100.0 0
+LRT 0 98.0 2.0
 Spectral 3.25 96.5 0.25
-4000 EBIC 99.5 0.5
-LRT 96.0 4.0
-Spectral 2.75 97.25
-6000 EBIC 98.0 2.0
-LRT 93.0 7.0
-Spectral 2.0 98.0
-8000 EBIC 97.25 2.75
-LRT 94.25 5.75
-Spectral 2.0 98.0
-Bernoulli, $\mathcal{B}_{g}$ 500 EBIC 0.75 99.25
+4000 EBIC 0 99.5 0.5
+LRT 0 96.0 4.0
+Spectral 2.75 97.25 0
+6000 EBIC 0 98.0 2.0
+LRT 0 93.0 7.0
+Spectral 2.0 98.0 0
+8000 EBIC 0 97.25 2.75
+LRT 0 94.25 5.75
+Spectral 2.0 98.0 0
+Bernoulli, $\mathcal{B}_{g}$ 500 EBIC 0.75 99.25 0
 LRT 13.5 86.0 0.5
 Spectral 7.0 91.75 1.25
-1000 EBIC 99.25 0.75
+1000 EBIC 0 99.25 0.75
 LRT 1.5 96.25 2.25
 Spectral 3.5 95.25 1.25
-2000 EBIC 99.25 0.75
+2000 EBIC 0 99.25 0.75
 LRT 0.5 94.0 5.5
 Spectral 8.75 90.75 0.5
-4000 EBIC 98.25 1.75
-LRT 98.25 1.75
+4000 EBIC 0 98.25 1.75
+LRT 0 98.25 1.75
 Spectral 2.75 96.25 1.0
-6000 EBIC 97.5 2.5
-LRT 96.5 3.5
+6000 EBIC 0 97.5 2.5
+LRT 0 96.5 3.5
 Spectral 1.75 97.75 0.5
-8000 EBIC 99.25 0.75
-LRT 99.25 0.75
+8000 EBIC 0 99.25 0.75
+LRT 0 99.25 0.75
 Spectral 0.5 99.0 0.5
@@
 “ParFam” is short for “parametric family”.
-Parfam, value $N$ Method \ $\widehat{K}^{(2)}$ 1 2 3
-Poisson, $\mathcal{B}_{s}$ 500 EBIC 99.75 0.25
-LRT 99.5 0.5
-Spectral 0.25 99.75
-1000 EBIC 99.75 0.25
-LRT 99.75 0.25
-Spectral 100.0
-2000 EBIC 99.25 0.75
-LRT 99.0 1.0
-Spectral 100.0
-4000 EBIC 98.25 1.75
-LRT 98.25 1.75
-Spectral 100.0
-6000 EBIC 99.0 1.0
-LRT 98.25 1.75
-Spectral 100.0
-8000 EBIC 98.5 1.5
-LRT 98.5 1.5
-Spectral 100.0
+Parfam, value NN Method \ $\widehat{K}^{(2)}$ 1 2 3
+Poisson, $\mathcal{B}_{s}$ 500 EBIC 0 99.75 0.25
+LRT 0 99.5 0.5
+Spectral 0.25 99.75 0
+1000 EBIC 0 99.75 0.25
+LRT 0 99.75 0.25
+Spectral 0 100.0 0
+2000 EBIC 0 99.25 0.75
+LRT 0 99.0 1.0
+Spectral 0 100.0 0
+4000 EBIC 0 98.25 1.75
+LRT 0 98.25 1.75
+Spectral 0 100.0 0
+6000 EBIC 0 99.0 1.0
+LRT 0 98.25 1.75
+Spectral 0 100.0 0
+8000 EBIC 0 98.5 1.5
+LRT 0 98.5 1.5
+Spectral 0 100.0 0
 Poisson, $\mathcal{B}_{g}$ 500 EBIC 1.0 97.5 1.5
@@
 Spectral 13.75 84.75 1.5
-2000 EBIC 95.75 4.25
+2000 EBIC 0 95.75 4.25
 LRT 0.5 94.0 5.5
@@
 Spectral 5.5 94.0 0.5
-6000 EBIC 97.25 2.75
-LRT 97.0 3.0
+6000 EBIC 0 97.25 2.75
+LRT 0 97.0 3.0
 Spectral 3.5 96.0 0.5
@@
 We work with the preprocessed version of the default training set, which consists of $60,000$ images, each containing information on the $28^{2}$ pixel values.
-Initially, each pixel takes integer values between 00 and $255$.
+Initially, each pixel takes integer values between 0 and $255$.
 As the data values are highly concentrated around values near 0 or larger than 200, we transform the data into binary responses by thresholding at a value of 128.
@@
 After filtering out the short/long documents, the resulting data matrix consists of $N=3320$ documents.
-TIMSS Response Time Data.
+TIMSS Assessment Data.
 We use the preprocessed dataset provided in Lee and Gu, (2024).
-Additionally, we remove all rows of the data matrix with missing values.
+Additionally, we remove all rows of the data matrix with missing entries, which resulted in a total of $N=435$ observations.
 While our estimation procedure can be modified to handle missing data (missing at random) by modifying the likelihood function, we take this additional preprocessing step for the sake of consistency with other parts of the paper.
+This dataset includes additional information regarding the latent structure $\mathbf{G}^{(1)}$, which is the so-called $Q$-matrix in the cognitive diagnostic modeling literature (von Davier, , 2008; Lee and Gu, , 2024).
+The provisional $\mathbf{G}^{(1)}$ matrix specifies $K^{(1)}=7$ latent cognitive skills (Number, Algebra, Geometry, Data and Probability, Knowing, Applying, and Reasoning) as well as the skills that are required to solve each item.
+That is, $g^{(1)}_{j,k}=1$ if item $j$ requires latent skill $k$ to solve it.
+Hence, following the confirmatory latent variable modeling convention in psychometrics, we estimate the DDE parameters by fixing $\mathbf{G}^{(1)}$ to this given structure.
 S.5.2 Selecting the Number of Latent Variables
@@
 The width of the upper layer arrows is proportional to the corresponding coefficients and the red arrow indicates negative values.
-TIMSS Response Time Data.
-The dataset comes with a rough classification of the seven latent skills into $K^{(2)}=2$ categories: content skills and cognitive skills.
+TIMSS Assessment Data.
+This dataset includes additional information regarding the latent structure $\mathbf{G}^{(1)}$, which is the so-called $Q$-matrix in the cognitive diagnostic modeling literature (von Davier, , 2008; Lee and Gu, , 2024).
+The provisional $\mathbf{G}^{(1)}$ matrix specifies $K^{(1)}=7$ latent cognitive skills (Number, Algebra, Geometry, Data and Probability, Knowing, Applying, and Reasoning) as well as the skills that are required to solve each item.
+That is, $g^{(1)}_{j,k}=1$ if item $j$ requires latent skill $k$ to solve it.
+Hence, following the confirmatory latent variable modeling convention in psychometrics, we estimate the DDE parameters by fixing $\mathbf{G}^{(1)}$ to this given structure.
+Additionally, the dataset comes with a rough classification of the seven latent skills into $K^{(2)}=2$ categories: content skills and cognitive skills.
 However, we observed that fitting $K^{(2)}=2$ results in one of the proportion parameters exhibiting an unusually large value of $p_{2}=0.96$.
@@
 The EBIC for the two cases were also comparable and justified this choice.
-S.5.3 Performance Evaluation for the 20 Newsgroups Dataset
+S.5.3 Additional Visualization and Performance Evaluation for the MNIST Dataset
+S.5.3.1 Additional visualization
+[…]
+Figure S.13: Graphical model representation of the learned latent structure.
+The edge widths are proportional to coefficients’ absolute values.
+The edge colors (black/red) imply positive/negative coefficients, respectively.
+In Figure S.13, we visualize the learned latent structure of the MNIST dataset.
+The corresponding graphical matrix $\mathbf{G}^{(2)}$ was used to interpret the top layer latent variables in the main text.
+For example, $A_{1}^{(2)}$ is connected to all lower-layer variables, and we interpret this as an indicator for the pixel density of the image.
+Label True LCM 1-layer DDE Spectral init 2-layer DDE
+0 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+1 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+2 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+3 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Table S.16: True (first column) and reconstructed (other columns) images under various estimators. Here, each pixel takes values in $[0,1]$, where a darker shade indicates a larger value.
+We additionally visualize examples of reconstructed images under various models in Table S.16, which have been used to compute the pixel-wise reconstruction accuracy in Table 4 in the main paper.
+The reconstructions are based on the corresponding conditional mean $\mathbb{E}\left(\mathbf{Y}\mid\mathbf{A}^{(1)}\right)$ for each latent representation $\mathbf{A}^{(1)}$, which is rearranged in the original $28\times 28$ grid.
+The visualization illustrates that the reconstructions from the 2-layer DDE exhibit greater clarity compared to those from alternative models.
+S.5.3.2 Comparison with VAEs and iVAEs
+We provide additional numerical comparison against VAEs and identifiable VAEs (iVAE), alongside implementation details.
+We use the standard VAE and iVAE in Khemakhem et al., (2020) with varying latent dimensions $K=2,5$ (these numbers are motivated by the latent dimensions $K^{(1)}=5,K^{(2)}=2$ used for DDEs).
+The values presented in Table 4 in the main text correspond to the VAE with latent dimension $K=2$.
+While we use identical neural network architectures (two-layer perceptrons with 50 hidden variables in the middle layer) and factorized priors for both VAE and iVAE, the latter requires auxiliary information to construct “label priors” while the former does not.
+Hence, we have provided the true digit labels as auxiliary data for iVAEs.
+Accuracy 2-DDE VAE ($K=2$) VAE ($K=5$) iVAE ($K=2$) iVAE ($K=5$)
+Train classif. (%) 92.0 97.1 97.5 97.5 98.2
+Test classif. (%) 92.6 92.0 92.1 92.5 93.2
+Train recon. (%) 79.6 82.5 83.4 82.5 86.6
+Test recon. (%) 79.9 82.7 83.6 82.8 86.6
+Table S.17: Performance comparison of DDEs versus VAEs and iVAEs with varying latent dimensions $K$ on the MNIST dataset.
+Note that the iVAEs are trained in a supervised manner, where the true labels are incorporated as auxiliary information.
+We report the evaluation metrics in Table S.17.
+Compared to VAEs, DDEs have better test classification accuracy but lower reconstruction accuracy.
+We believe that the discrepancy in reconstruction accuracy results from the continuous latent variables in VAEs, which provide more detailed information.
+In contrast, for classification, the lack of a clear threshold in continuous latent variables seem to result in potential overfitting and lower test accuracy.
+Compared to iVAEs, DDEs have lower performance in both metrics but this is not a fair comparison as iVAEs are weakly supervised (requiring the true label information), whereas DDEs are fully unsupervised.
+Focusing on the classification error, we note that iVAEs with $K=2$ latent dimensions have similar performance compared to DDEs.
+This suggests that DDEs are able to compete with benchmark iVAEs as well.
+DDE [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+iVAE [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+${\boldsymbol{\beta}}_{0,0}^{(1)}$ ${\boldsymbol{\beta}}_{0,1}^{(1)}$ ${\boldsymbol{\beta}}_{0,2}^{(1)}$ ${\boldsymbol{\beta}}_{0,3}^{(1)}$ ${\boldsymbol{\beta}}_{0,4}^{(1)}$ ${\boldsymbol{\beta}}_{0,5}^{(1)}$
+Table S.18: From left to right: Estimated basis images (reshaped last-layer coefficients) from the MNIST data.
+Brown indicates positive values and green indicates negative values.
+The basis images under DDEs are distinct, smoother, and more interpretable whereas that under the iVAE exhibit overlaps.
+Our most interesting finding is the qualitative comparison of the basis images learned via DDEs versus iVAEs (with $K=2$) in Table 2.
+Here, the iVAE basis images are the last-layer perceptron coefficients.
+We see that the third, fourth, fifth column of the iVAE panel are almost identical.
+This strongly indicates that the estimated perceptron (neural network) parameters suffer from overparametrization.
+Consequently, it is challenging to interpret the shallow layer in the perceptron.
+In contrast, the basis images under DDEs are distinct, smoother, as well as highly interpretable (e.g.
+the basis images in the first two columns directly resemble the digits 1 and 0).
+S.5.3.3 Implementation details
+The VAE and iVAE was implemented by the Python codes provided in Khemakhem et al., (2020), where no auxiliary information was used for VAEs and the true digit labels were used as auxiliary information for iVAEs.
+The DBN is implemented similar to the earlier ablation studies in Supplement 4.2.
+For both models, we have used default tuning parameters and architectures provided in the original codes.
+To compute the classification error from latent representations, we have fitted a decision tree analogous to that for DDEs as described in the main text.
+S.5.4 Performance Evaluation for the 20 Newsgroups Dataset
 Perplexity.
@@
 Notably, our preprocessing steps enhanced the signal-to-noise ratio and leads to a smaller perplexity.
-S.5.4 Additional Visualization of MNIST Dataset
-[…]
-Figure S.10: Graphical model representation of the learned latent structure.
-The edge widths are proportional to coefficients’ absolute values.
-The edge colors (black/red) imply positive/negative coefficients, respectively.
-In Figure S.10, we visualize the learned latent structure of the MNIST dataset.
-The corresponding graphical matrix $\mathbf{G}^{(2)}$ was used to interpret the top layer latent variables in the main text.
-For example, $A_{1}^{(2)}$ is connected to all lower-layer variables, and we interpret this as an indicator for the pixel density of the image.
-Label True LCM 1-layer DDE Spectral init 2-layer DDE
-0 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-1 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-2 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-3 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Table S.13: True (first column) and reconstructed (other columns) images under various estimators. Here, each pixel takes values in $[0,1]$, where a darker shade indicates a larger value.
-We additionally visualize examples of reconstructed images under various models in Table S.13, which have been used to compute the pixel-wise reconstruction accuracy in Table 3 in the main paper.
-The reconstructions are based on the corresponding conditional mean $\mathbb{E}\left(\mathbf{Y}\mid\mathbf{A}^{(1)}\right)$ for each latent representation $\mathbf{A}^{(1)}$, which is rearranged in the original $28\times 28$ grid.
-The visualization illustrates that the reconstructions from the 2-layer DDE exhibit greater clarity compared to those from alternative models.
+S.5.5 Additional Analysis of the TIMSS Assessment Dataset
+We display the estimated coefficients for each modailty in Figure S.14, focusing on the first four skills (Number, Algebra, Geometry, Data&Probability).
+The magnitude of the coefficients are larger for the response accuracy, indicating a larger effect of the latent variables on the responses.
+Additionally, we observe an interesting pattern learned from the intercept values, compared to the held-out information regarding the items (constructed-response versus multiple-choice).
+To be specific, the constructed-response items (indexed by the rows 2, 4, 7, 15-17, 19-20, 24-26, 29) have smaller intercept values for the left panel (response accuracy) of Figure S.14 and larger intercept values for the right panel (response time).
+This is analogous to the previous analysis of the same dataset in (Lee and Gu, , 2024), where each mode (response accuracy and time) were analyzed separately.
+Figure S.14: Estimated coefficients for (left) response accuracy, (right) response time.
+Additionally, we visualize the analog of Table 6 solely-based on the response time in Table S.19, as opposed to the analysis using both modes.
+Here, the same lognormal-DDE was used with identical latent dimensions.
+While we observe a similar trend that students enjoying math are more likely to master each skill, the discrepancy between the students enjoying/not enjoying math is lower.
+For instance, the discrepancy between the category ‘agree a lot’ versus ‘disagree a lot’ in Table 6 of the main paper is 0.35 versus the value of 0.14 in Table S.19.
+This illustrates considering both data modalities help to better learn the latent variables in a more interpretable manner.
+Response \ Latent skill $A_{1}^{(2)}$ $A_{1}^{(1)}$: Number $A_{2}^{(1)}$: Algebra $A_{3}^{(1)}$: Geometry $A_{4}^{(1)}$: Data and Prob
+Agree a lot 0.83 0.72 0.88 0.86 0.87
+Agree a little 0.80 0.72 0.85 0.84 0.85
+Disagree a little 0.76 0.74 0.85 0.84 0.84
+Disagree a lot 0.69 0.67 0.76 0.75 0.78
+Table S.19: Average latent variable estimate for each response category for the question “Mathematics is one of my favorite subjects”, solely based on response times.
+S.6 Additional Literature Review
+iVAEs.
+We compare DDEs versus iVAEs and its extensions.
+One distinction is that iVAEs essentially have only one latent layer of random variables transformed by deterministic deep neural networks, while DDEs allow multiple latent layers.
+Thus, DDEs are able to model the hierarchical nature of uncertainty and are more suitable for tasks like hierarchical topic modeling (Ranganath et al., , 2015).
+Also, iVAEs typically require continuous latent variables, additional auxiliary variables, and its notion of identifiability still suffers from rotational ambiguity.
+Most iVAEs are used to model continuous observed responses (Khemakhem et al., , 2020; Moran et al., , 2022; Kivva et al., , 2022), although there exist a few variants exclusively proposed for discrete data (binary, count) as well (Hyttinen et al., , 2022).
+In particular, up to our knowledge, identifiability guarantees for discrete data has been only established in Hyttinen et al., (2022) under strong parametric restrictions on the encoder.
+[^4] The original paper Khemakhem et al., (2020) only provided identifiability theory for continuous data, although their simulation studies suggest potential extensions to discrete data.
+In contrast, DDEs with discrete latent variables do not require auxiliary variables, do not suffer from rotational ambiguity, and the identifiability of DDEs holds regardless of the observed data types.
+Psychometrics.
+Multidimensional binary latent variables are also popular for modeling students’ item response data in educational measurement (Junker and Sijtsma, , 2001; von Davier, , 2008; de la Torre, , 2011).
+These psychometric models are known as cognitive diagnostic models (CDMs).
+A CDM uses a single latent layer of cognitive skills to model a student’s binary responses to many questions in a test, with a sparse loading graph between the observed and latent layers.
+Here, each binary latent variable represents a student’s mastery or deficiency of a cognitive skill, and the sparse graph between the observed item responses and latent skills encodes which skills each item is designed to measure in the test.
+The proposed DDEs substantially generalize this modeling idea by allowing (a) multilayer latent variables, which can model one’s knowledge structure at multiple resolutions ranging from fine-grained details to general concepts (Gu, , 2024), and (b) modeling a rich class of general responses, going beyond the typical binary correct/incorrect responses in educational tests.